A Bibliography of Publications in Theoretical Computer Science: 2005–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 April 2025 Version 1.17

Title word cross-reference

#CSP [2264]. #P [1788]. #P-completeness [1788].

(1 + 1) [1683]. (1, λ) [115]. (18/11)n [2213]. (α, β) [2215]. (n, k) [1557, 2392]. (n, t) [1049]. (p, k) [467]. (r, p) [2357]. $-calculus [1275]. 1 [220, 503]. 2 [1425, 2304, 1890, 2253, 2333, 1287, 1302]. 2 < α ≤ 7/3 [2163]. 3 [1628, 1771]. 3.4713 [2387]. 4 [2260]. 5 + ϵ [1968]. 5/2^{+} [90]. 5/4 [778]. [0, 1] [503]. ^{∗} [58]. ^{,} [109]. _{∞} [1543]. _{p} [196]. α [2163, 132]. b [1319]. β [1456]. c [469]. C^{bω} [2338]. C^{∞} [1502]. d [1245, 912, 1690, 817, 105]. Δ [439]. D⌊log_{m} D⌋ [648]. ℓ_{1} [2296]. ℓ_{∞} [2296, 1543]. ℓ_{p} [782]. ϵ [1183, 729]. F_{0} [2005]. F S [1741].

H [1823]. k [1116, 1805, 796, 803, 173, 781, 1961, 1609, 775, 1989, 178, 1310, 630, 1311, 951, 62, 1823, 1809, 2253, 2061, 2003, 1624, 953, 1226, 1306, 1290, 2007]. k > 2 [173]. K_{max}, K_{min} [541]. κ [1836]. L(2, 1) [2243, 1635]. L(d, s) [440]. l_{∞} [961]. λ [2281, 400, 1588, 562]. λμ [1159]. log_{d} N [547]. LU [1217]. m [1760, 1324, 2258, 1745, 1226, 1306]. N^{d} [1444]. R [417]. R^{d} [2062]. Z [1859]. A [1200]. C [641]. H [1585]. Q [1034]. T^{2} [1320]. μ [187, 400, 589, 359, 101]. n [1909, 958, 2079, 1451]. n ≥ 30 [2169]. ≠ [789]. O(n^{1.75}) [2243]. ∨(n^{2}polylogn) [1915]. o(n log n) [1313]. Ω [1372, 974, 1655, 663, 826]. ω(1) [1752]. P [1527, 433]. ϕ [115]. π [35, 56, 1436, 143, 1440]. Π_{2}^{p} [441]. R [974]. σ [1233]. t [893]. Õ(n^{ω}) [1553]. w̃ × w [1502, 2338]. × [445]. U [1768]. x^{k} = z_{1}^{k_{1}}z_{2}^{k_{2}} ⋯ z_{n}^{k_{n}} [12].

* [400]. *- [400].

-achievement [1881]. -antipodal [1049]. -approximation [2387, 1968, 1287, 778]. -ary [953]. -automata [826]. -avoidable [1890]. -batch [1989]. -bit [1313]. -calculi [101]. -calculus [35, 2281, 187, 56, 589, 359, 143, 562, 1440, 1436]. -categories [1372, 1034]. -center [775, 62]. -centroid [2357]. -cliques [469]. -collapsing [1451]. -coloring [467, 439]. -complete [441]. -connected [1311, 2253, 1226, 1306]. -context-free [1624]. -contextual [893]. -counter [220]. -cube [958]. -dimensional [1752, 817, 105]. -domains [1741]. -dominating [2253, 1306]. -EA [1683]. -edge [1319]. -estimation [2005]. -facility [1290]. -flat [1760]. -free [433, 1823]. -game [1233]. -iet [1628]. -in-a-row [1310]. -induced [1823]. -integers [1456]. -irreducible [115]. -labeling [2243, 1635]. -labelings [440]. -language [663]. -limited [2079]. -longest [1809]. -maximum-sums [796]. -Median [781]. -nearest-neighbor [2061]. -panconnectivity [1324]. -path [630]. -points [729]. -polygons [1768]. -power-free [2163]. -reducibilities [1745]. -regular [1245, 1624]. -SA-ES [115]. -SAT [178]. -semirings [400]. -spanner [173]. -spanners [2215]. -splittable [951]. -star [1557, 2392]. -Steiner [2007]. -strict [1200]. -sub-recursive [417]. -subshifts [1444]. -tape [1909]. -terms [974]. -theories [1588]. -tonic [2003]. -transitions [1183]. -trees [1961]. -tuple [1226]. -valued [912, 1690]. -varieties [641]. -variety [1655]. -words [1502, 2338].

.NET [858].

1 [385]. 1/2 [2399].

2 [386]. 2-approximate [1915].

2-approximation [1963]. 2-categories [102]. 2-coloring [2312]. 2-constraint [431]. 2-cyclic [2362, 1991]. 2-interval [1322, 1541, 2123]. 2-letter [1493]. 2-median [1979]. 2-radiian [1821]. 2-radius [1821]. 2005 [987, 1434]. 2006 [1513, 2302]. 2nd [1554]. 2nd-order [1554]. 2SAT [74].

3-COL [1716]. 3-coloring [1554].

3-dimensional [1818]. 3-regular [1289]. 3-SAT [1025, 1810]. 3-tier [835]. 3-vertex-connectivity [1012]. 380 [2399]. 3D [689]. 3SAT [74].

4-coloring [1414]. 4-cycles [944]. 4-way [2040]. 400 [2398].

5 [898].

6-man [446]. 60th [2270, 1531, 2287].

64-bit [503].

7/8 [2346]. 7/8-approximation [2346].

8-approximation [2346].

abelian [2260, 1184, 1890, 1911]. absence [1139]. Absolute [1577, 317, 1296]. absoluteness [1514]. absorption [372]. Abstract [154, 2325, 672, 1685, 278, 356, 856, 1069, 163, 1084, 1926, 186, 159, 1677, 1864]. abstraction [593, 1729, 1784, 1436, 1695, 1029]. abstractions [1698]. Accelerating [2231, 2245]. acceptance [24, 72]. Accepting [993, 1135, 2199, 1941, 1251]. Access [1596, 1242, 1597, 1003, 159]. accessible [1214]. achievable [2109]. achievement [1881]. achieving [1016]. achromatic [1974, 773]. acid [1932]. Acknowledged [1103]. acknowledgment [1983]. acknowledgments [947]. across [1793]. action [753, 59]. action-based [753, 59]. Actions [641, 916]. actor [1863]. Actors [1925]. Acyclic [2364, 1565, 1179, 413, 1059, 2154, 971]. ad [1665, 1663, 157, 1664, 1103]. ad-hoc [1665, 1664]. Adams [2089]. Adaptive [1367, 1663, 2212, 1952, 2020, 909, 1098]. adaptiveness [107]. Addendum [2399]. Adding [164, 1038]. Addition [1127]. Additive [381, 2331, 1323, 546]. Additivity [1772]. address [2378]. addressable [610]. addressing [1857, 67]. AdHoc [2136]. adic [1193]. Admissible [501, 374]. ADMs [457, 1301]. advances [1682]. Adversarial [1960, 1231]. adversaries [224, 227]. advice [1676, 2018]. Affine [2055, 1389, 569, 1540]. affix [1410]. after [1040]. again [1104]. against [574]. agent [805]. agents [288, 1461, 1305, 1926, 1487, 515]. aggregate [981]. aggregation [737, 310, 2022, 1875]. AGh [921]. agnostic [1110]. agreeable [1054]. Agreement [1300, 133, 2020, 1954, 985, 1050]. agreements [473]. al. [1909]. Alberto [1531, 363]. Alder [43]. algebra [881, 141, 137, 140, 300, 882, 1469, 1275, 1074, 509, 360, 210, 2053, 579, 1440]. Algebraic [939, 238, 682, 1475, 2321, 338, 957, 485, 1870, 1067, 579, 788, 497]. algebraicity [2278]. algebraization [584]. Algebras [1377, 514, 2366, 406, 2058, 1830, 584, 1100]. Algol [1066, 1438]. algorithm [2113, 811, 553, 1168, 1109, 1963, 2376, 2162, 2387, 1700, 859, 640, 2295, 469, 128, 2308, 2362, 960, 711, 1143, 1426, 1663, 1783, 1869, 1968, 778, 1551, 1951, 2000, 1182, 1139, 329, 79, 1055, 1169, 1330, 1955, 1140, 2356, 966, 387, 1321, 1647, 2147, 723, 815, 1111, 1404, 2263, 909, 1721, 2021, 1267, 727, 2230, 1428, 265, 468, 728, 1976, 1771, 1132, 1451, 1991, 2293, 1454, 2138, 1287, 1503, 2110, 431, 1688, 462, 627, 1290, 2007, 693]. Algorithmic [1161, 1423, 1276, 720, 1169, 2192, 949, 253, 408]. Algorithms [1886, 1531, 1539, 1809, 1399, 130, 1856, 1226, 1101, 1973, 1129, 1777, 937, 386, 1148, 1906, 2313, 458, 2391, 433, 1683, 2019, 86, 2052, 859, 3, 402, 1979, 2062, 268, 1216, 2351, 796, 1648, 1801, 1606, 1179, 1607, 1114, 2242, 2023, 1219, 421, 781, 549, 836, 1309, 1609, 1853, 1173, 1676, 2309, 217, 833, 1908, 1284, 269, 511, 1629, 1311, 624, 2049, 773, 1499, 1025, 1982, 297, 2289, 1042, 1128, 84, 1699, 1927, 539, 542, 1631, 1221, 1532, 530, 1237, 1050, 694, 44, 2132, 2005, 1011, 2374, 1306, 710, 2063, 1774, 1997, 2368, 1547, 2397]. Aliased [1800]. Alice&Bob [899]. alignment [926, 966]. all- [132]. all-ones [1307]. all-optical [109]. All-pairs [1915, 1553]. all-shortest-path [1408]. all-to-all [109]. allocation [950, 2100, 1180, 549, 1781, 1800]. allowing [388]. almost [2366, 2025, 1110]. almost-complete [2025]. alpha [1193]. alpha-adic [1193]. Alphabet [1313, 245, 2037, 2200, 1446]. Alphabet-independent [1313]. alphabet-size [2200]. alphabets [262, 1493, 1314, 2168, 1909, 75]. alternating [194, 251, 170, 567]. Alternating-time [567]. Alternative [889, 1728, 691]. Amalgamating [1923]. Amalgamation [51]. amazons [451]. amb} [35]. Ambainis [214]. ambient [666, 883, 37, 1596, 667]. ambient-like [666]. ambients [1030, 97, 1597, 2016]. ambiguity [959, 352]. Ambiguous [101]. amiable [1430]. among [2002, 1487]. amplification [984, 2075]. analog [1063, 995, 996]. analyser [274]. analyses [857]. Analysis [1652, 693, 447, 598, 1161, 1423, 1693, 897, 2322, 2192, 591, 1468, 197, 154, 898, 734, 268, 931, 2189, 2274, 595, 1373, 157, 222, 192, 1558, 1555, 1169, 1071, 1150, 1362, 1637, 723, 1534, 1404, 2253, 373, 1151, 568, 449, 903, 1132, 540, 1677, 2333, 225, 1463, 905, 1410, 332, 885, 462, 1811, 551, 1922, 2191]. analytical [1760, 1323]. Analyzer [902]. Analyzing [1072, 637, 1421, 1839]. anarchy [934]. ancestor [1553]. ancestors [1179]. AND/OR [1647]. Angel [1401, 465]. animation [1738]. Anisotropic [1850]. annealing [1330]. annotated [1620]. anonymity [2271]. anonymous [1814, 1611]. answer [1884]. answering [1001, 149]. Ant [322]. antichain [1147]. antimagic [1047]. antimagicness [1965]. antipodal [1049]. antiport [1021, 692]. Antonio [230]. Any [284, 612, 2183]. Any-world [284]. Apartness [1734]. Apostolico [1531]. Application [735, 34, 804, 1421, 581, 689, 1175, 1131]. Applications [2322, 1255, 870, 1291, 602, 1584, 1468, 1197, 979, 365, 1061, 1932, 1279, 1924, 2299, 251, 559, 1311, 2298, 1721, 601, 1410, 1500]. Applied [271]. approach [404, 974, 1850, 757, 919, 1626, 737, 949, 967, 883, 2377, 1156, 58, 705, 295, 328, 748, 1323, 60, 247, 2009, 1621, 2288, 152, 2239, 466, 962, 841]. approaches [957, 1933]. approaching [909]. appropriateness [480]. Approximability [1319, 288, 951, 68, 267]. approximants [358]. Approximate [2378, 548, 1871, 1149, 1915, 414, 2047, 311, 1452, 550, 1556]. Approximately [999, 1999]. Approximating [1053, 1484, 2312, 1541, 173, 1232, 1491, 1221, 2373, 818, 754, 2308, 175, 1057, 1543, 1426, 597, 2049, 1995]. Approximation [2106, 1648, 732, 2000, 836, 1309, 2309, 1989, 182, 624, 390, 1267, 1631, 1248, 536, 2248, 2353, 216, 85, 1481, 1963, 2101, 2387, 1987, 2308, 1216, 2351, 249, 1968, 778, 2150, 1114, 730, 1548, 1182, 1646, 1472, 2059, 942, 1298, 2346, 773, 1152, 1428, 1110, 1287, 1306, 1290, 2007, 2109, 1050]. approximations [1353, 618, 409]. APX [134, 216, 730]. APX- [216]. APX-hard [134]. Arbitrage [1793, 832]. Arbitrary [1814, 262, 205, 1967, 1298, 2099, 1756, 1431, 77]. arboricity [1224]. architectural [141]. architecture [55]. Area [1848, 2392]. Area-time [1848]. areas [2004]. arising [532]. Arithmetic [1127, 918, 502, 505, 506, 503, 1453, 1403, 138, 679]. arithmetical [1198, 207, 1907]. Arity [1704]. arm [1122, 627]. armed [2069]. arrangements [474]. array [1410]. arrays [2386, 548, 1313, 1537]. arrivals [2151]. Artemis [2105]. artificial [686, 1682]. ary [953]. ASM [161]. AsmL [303]. Aspects [64, 1240, 2271, 47, 1235, 1516, 312, 790, 2006]. assembling [1936]. assembly [2039, 1127, 1522, 1524, 2034, 1048, 1715, 721, 2030, 1523, 1940, 528]. assertion [54, 1830]. assertion-based [54]. assignment [288, 1168, 558, 1796, 1594, 438, 1267, 1556, 2104, 1590]. assignments [1501]. associated [1790, 1229, 767, 716, 1799]. associativity [1639, 1207]. assumption [227]. assumptions [284, 1966]. asymmetric [1959]. Asymmetry [775]. Asymptotic [2092, 723, 1481, 1347]. Asynchronous [2118, 628, 700, 1339, 2279, 1261, 139, 786, 1924, 168, 1814, 1424, 1052, 1957, 2333, 2130]. Atomic [2210]. atoms [1070]. ATSP [978]. attack [1900]. attacks [910, 905]. Attila [1759]. attribute [1344, 752]. attribute-efficiently [1344]. attribute-element [752]. Attributed [1096]. auction [1140]. auction-based [1140]. auctions [1480, 182]. augmentation [734, 1648, 2080]. authentication [226, 901, 10, 551]. Author [40, 61, 91, 112, 125, 144, 162, 183, 198, 221, 493, 759, 785]. autocorrelation [2114]. Automata [2155, 1328, 594, 2241, 974, 1350, 829, 2073, 5, 1085, 713, 2301, 2193, 1214, 636, 1727, 1813, 339, 1441, 244, 637, 456, 2057, 483, 698, 2199, 1492, 71, 1489, 823, 1208, 2168, 1909, 50, 2222, 1714, 695, 1178, 2202, 264, 9, 1888, 2331, 2265, 891, 1181, 2223, 2077, 787, 786, 56, 595, 1574, 2074, 194, 1794, 2076, 648, 972, 2176, 2229, 24, 650, 2225, 929, 825, 1210, 801, 114, 826, 1802, 1052, 1091, 2220, 1437, 1229, 1902, 1684, 827, 100, 328, 28, 30, 1358, 546, 1528, 893, 2228, 1343, 2206, 1115, 1010, 2181, 1687, 1093, 2333]. automata [1621, 323, 2396, 830, 538, 936, 72, 1130, 2226, 1277, 2075, 124, 922, 659, 1696]. automata-theoretic [974]. Automated [1097, 727]. Automatic [1731, 573, 566, 2160, 724, 1729, 948, 2232, 1043]. automaticity [1162]. Automaton [2348, 344, 1033, 2230, 1686, 266]. automaton-based [1033]. Automorphism [2090]. Autonomous [2031, 1606, 2023, 1299, 1681]. Autopoietic [1277]. auxiliary [825]. availability [2145, 793, 2334, 2349]. Average [1150, 931, 1262, 1356, 2001, 439]. Averages [724]. avoid [940]. avoidability [1990]. avoidable [1890]. avoidance [202, 2179]. Avoiding [203, 1911, 2386, 1546, 1453]. aware [1030, 1369]. Axiomatisations [619, 758]. axiomatization [33, 1509, 187, 567]. Axiomatizations [1032, 1005]. axiomatized [745]. Axiomatizing [1006]. axis [1767].

back [885]. back-and-forth [885]. Backtracking [484]. backup [106]. Backward [2229]. Balanced [2150, 1180, 311, 928]. balancing [782, 952, 2022]. ball [367]. balls [1898, 1749]. balls-into-bins [1898]. bandits [2069, 1575]. bandwidth [811, 709, 1818, 1781, 1117, 834, 930]. barrier [1331, 2395]. bars [1400]. base [949, 2214]. based [54, 880, 1149, 1242, 821, 34, 753, 1001, 911, 976, 1461, 128, 212, 1147, 255, 1143, 47, 1597, 2325, 664, 343, 902, 610, 720, 1373, 2324, 1925, 479, 1396, 2178, 1730, 910, 1955, 2033, 1362, 1140, 2354, 846, 994, 1033, 1695, 1266, 1004, 449, 312, 1328, 411, 1732, 1454, 59, 1879, 1966, 779, 2255, 123, 1811, 319, 920, 551, 1144]. bases [1193, 1845, 1655]. basic [882, 2135, 1169, 746]. basis [586, 1879]. batch [2151, 1058, 1989, 2153, 1999, 390, 1054, 1567, 1975, 1045, 2110]. batching [2237, 1225, 2361]. Baumslag [220]. Bayes [1342]. Bayesian [1835, 1282, 1755]. be [483, 97, 612, 1084, 1686, 1308]. behavior [2162, 456, 786, 1732]. behavioral [821, 1527, 1672, 1155]. behaviors [1032, 1098]. Behavioural [301, 49, 754]. being [1080]. Belnap [1386]. benefits [1251]. Benenson [2396, 323]. Bernstein [1105]. Besicovitch [2235]. best [1503, 2110]. beta [1722, 2337, 2191]. beta-binders [1722, 2191]. beta-shifts [2337]. better [1548]. between [278, 336, 1513, 1069, 2367, 1001, 978, 1068, 1561, 382, 2214, 690, 2233, 1638, 1976, 788, 231, 779, 563]. Beyond [43, 730, 1363, 1439]. Bezoutian [1105]. bi [399]. bi-infinite [399]. Bias [453]. bibliography [1620]. bicliques [172]. biclusters [927]. biconnectivity [1556]. Bicriteria [1225]. bifinite [870]. bijective [1178]. billiard [240, 1197]. bimachines [1633]. bimatrix [1828, 2049]. bimorphisms [2224]. bin [2353, 1481, 794, 1899, 1431, 1500]. Binary [2114, 2363, 977, 1367, 1885, 2163, 1764, 1766, 1890, 1911, 749, 2174, 210, 392, 25, 2231, 1770, 372, 982, 1499, 1151, 2003, 203, 629, 627, 1774]. Bincoloring [1797]. binders [2191, 1722]. binding [1849, 2194]. binoid [1006]. binomial [540]. bins [1898, 1180]. bio [1126, 2189]. bio-operations [1126]. Bio-PEPA [2189]. biological [1464, 2190, 1710, 2189, 1998, 1841, 2297, 369, 1935, 1462, 2188]. biology [1685]. bipanconnectedness [1972]. bipanconnectivity [1324]. bipanpositionable [1972]. bipartite [1227, 1228, 2088, 81, 2307, 1212]. biprefix [2117]. biRFSA [652]. Birthday [1531, 2270, 2287]. bisecting [265]. bisection [1245, 1246, 1121]. bisimilar [1696]. Bisimilarity [880, 38]. Bisimulation [1489, 281, 748, 661, 1340, 187, 188, 2229, 1187, 1677, 1073]. bit [2378, 544, 911, 1296, 503, 1313, 1185]. bit-probe [1185]. Bitonal [1710]. bits [2344]. biunitary [535]. bivariate [375]. black [425, 1298]. black-box [425]. BlenX [1841]. blind [1888, 650, 801]. Block [1217, 1286, 956, 780, 1963, 803, 290, 166, 905, 726]. block-synchronized [166]. blocks [1201, 973]. blockwise [1364]. Board [270, 279, 298, 403, 443, 561, 575, 738, 761, 769, 807, 847, 868, 886, 906, 913, 923, 964, 986, 997, 1017, 1027, 1035, 1064, 1081, 1094, 1124, 1136, 1165, 1190, 1234, 1325, 1348, 1387, 1415, 1433, 1458, 1505, 1512, 1570, 1600, 1636, 1660, 1832, 1866, 1919, 2158, 2184, 17, 52, 92, 103, 126, 135, 153, 184, 199, 211, 228]. Bodlaender [1404]. bond [832, 119]. bond-free [119]. boolean [1330, 471, 1129, 2264, 1815, 1382, 1549, 1501, 406, 1561, 302, 1284, 752, 1854, 277, 742, 68, 291]. Boost [505]. bordered [236]. Borel [2103, 492]. both [2347]. bottom [389, 213, 2107]. bottom-up [389, 213, 2107]. bound [1242, 43, 791, 1271, 1294, 1543, 2213, 1178, 1796, 87, 1907, 75, 953, 710, 428]. boundaries [601]. boundary [1406, 71, 321]. bounded [327, 988, 434, 381, 1648, 597, 1230, 1295, 9, 1263, 2380, 648, 719, 1402, 1730, 1171, 1247, 348, 1119, 121, 813, 383, 498, 634]. bounded-hop [1295]. Bounding [1520]. Bounds [1245, 1442, 1560, 2019, 1987, 1122, 2089, 1296, 715, 933, 1016, 421, 1371, 1189, 706, 2177, 1820, 2129, 2253, 2230, 729, 1808, 65, 1635, 1817, 779, 214]. box [425]. boxed [1597]. boxes [685, 1171]. Boyer [2301, 2231]. BPA [880]. bracketing [747]. Brainstorming [1018]. branches [1650]. Branching [661, 791, 1163, 2279, 2314, 706, 1894, 285, 121]. branching-time [1163, 285]. brane [994, 1711, 1714]. break [1535]. Breaking [1131, 694]. breakpoint [638]. breakpoints [1534]. bridges [1711]. brief [137]. broadcast [1261, 1604]. Broadcasting [607, 558, 2216, 107, 2025, 983, 1103]. broadcasts [1558, 1643]. Brownian [1516, 1937]. Bruijn [610]. brushes [1617]. Brzozowski [1630]. Büchi [829, 194]. budget [1480]. buffering [1602]. bundles [2311]. Burrows [89, 1361, 1363, 1362, 1368, 2182, 2293]. Busi [1930]. busing [65]. butterfly [46]. BWT [1364]. Byzantine [111].

C [502]. C# [156, 914, 306]. cache [2143, 1678]. cache-friendly [2143]. cache-oblivious [1678]. caches [47]. cacti [630]. cactus [1148]. cage [2032]. calculations [1143]. Calculi [1584, 1861, 662, 1069, 96, 1671, 664, 994, 37, 1596, 1711, 101, 1384]. calculus [57, 275, 849, 1630, 1591, 35, 1598, 283, 2281, 1836, 187, 1592, 56, 666, 174, 883, 589, 1436, 866, 1586, 586, 1587, 489, 1720, 359, 568, 860, 143, 678, 305, 562, 1440, 1002, 1159, 429, 751]. calibrated [839]. calibration [838, 1347, 2063]. call [733, 296]. call-by-name [296]. camp [2214]. Can [988, 1084, 1651, 1686]. Cancellation [1447]. candidate [2388]. cannot [483, 97]. Canonical [969, 1393, 639, 1147, 672]. Cantelli [2103]. capacitated [1319, 2387, 1428]. capacities [1211]. capacity [166]. capercaillie [2185]. cardiac [2193]. cares [2291, 803]. carrier [1663]. carrying [856]. Cartesian [1965]. case [155, 776, 2162, 931, 1141, 1639, 1150, 411, 1747, 812]. case-based [411]. cases [134, 165]. CASL [579, 51]. catalysts [23]. catalytic [1019, 824]. categorial [1379, 1706]. Categorical [849, 562, 488]. categories [874, 573, 584, 1372, 742, 1743, 102, 1034]. category [879, 1707]. category-theoretic [879]. caterpillar [2219]. caterpillars [834].

Caucal [2398, 1622]. Causal [1748, 2276]. Causality [1337, 1739]. Cavity [1849]. Cayley [395]. CCP [39]. CCS [33]. cell [2315, 1716, 1174]. cells [1720, 28]. cellular [2241, 1464, 339, 456, 2331, 2265, 786, 648, 114, 1052, 28, 546, 1528, 1115, 1010, 2333, 1621, 538, 936, 1686, 1130]. center [1148, 775, 62]. central [233]. Centralized [1261, 1264]. centroid [2357]. CERES [1693]. certain [504, 2245, 1042, 231, 800]. certification [586, 330]. certified [856, 499]. chain [1228, 835, 2214, 2226]. chains [115, 1948, 1470, 2030, 767]. Chaitin [1318]. challenges [1932]. Chang [960]. Change [1470, 364]. changeover [950]. Changing [2217]. channel [11, 2104]. channels [1170]. chaotic [456]. Characteristic [2165, 2166, 169, 533]. Characteristics [1943, 661]. characterization [1197, 2342, 2297, 1446, 399, 948, 1359, 557, 745]. Characterizations [1015, 692, 1060, 644]. Characterizing [702, 1592, 646, 2233, 261]. Charts [2276]. chasing [1603]. cheating [1853]. checkers [1020]. checking [356, 2192, 753, 1163, 1238, 2285, 1462, 1726, 1730, 572, 703, 2280, 1927, 1237, 1732, 1002, 917, 333]. Chemical [1837]. Chess [446]. Choffrut [656]. choice [143]. Chomsky [583]. Choosing [504]. chord [401, 457]. chordal [1227, 1411, 2352, 437, 527, 1212]. Chordless [524]. Chosen [910, 2354]. chosen-ciphertext [2354]. Christoffel [2251]. chromatic [1798, 436, 82]. chromosomal [1535]. Chudnovsky [553]. Chvátal [2156]. Ciancaglini [1584]. CiE [987, 1513]. ciliates [638]. ciphertext [910, 2354]. circle [1551]. Circuit [1270, 255, 1563, 84, 398]. circuits [1848, 2131, 414, 718, 1854, 1907, 912, 1690, 1403, 2268, 1732, 941, 1817]. circulant [1801, 1427, 2084]. Circular [2295, 90, 2342, 1638, 1450, 2081, 82]. clarifications [401]. class [2291, 1389, 753, 639, 196, 836, 1525, 1200, 2076, 532, 559, 317, 1155, 1248, 477, 332, 1308, 1500]. Classes [1107, 1406, 1256, 336, 216, 1254, 2168, 1546, 730, 474, 970, 371, 1284, 692, 1283, 2064, 1936, 1247, 475, 2002, 213, 374, 1026, 542, 1578, 101, 936, 2394, 745, 545]. classical [849, 2129, 2326, 325]. Classification [1320, 1255, 438, 1577]. Classifying [2099, 1056]. clause [2172]. clauses [1640]. Claw [2374, 1554]. claw-free [1554]. Cleaning [1617]. cleavage [1838]. clique [2091, 1053, 1411, 979, 1216, 436, 719, 2384, 1321, 1568, 1119]. clique-width [719, 1119]. cliques [608, 469, 1825, 1332, 812]. clock [328]. Clos [1014]. Close [427]. closed [1254, 580, 2126]. closest [2054, 1504]. Closure [581, 1155, 538, 150, 580, 806]. cluster [2391, 1964]. clustering [774, 1787, 1955, 1810]. clusters [2061]. CNF [393]. CNN [710]. co [1838]. co-transcriptional [1838]. coagnostic [473]. coalgebra [881, 756]. Coalgebraic [56, 862, 1375, 883, 1439]. coalgebras [2273, 739, 517, 756]. coarse [768]. Cobham [2057, 1444]. cocomplete [1034]. code [1417, 856]. codeletable [1770]. Codes [243, 636, 233, 888, 2300, 2262, 1170, 2117, 1049, 1126, 10, 1546, 234, 117, 1407, 239, 1229, 1638, 654, 1450, 735, 736]. Coding [956, 1243, 10, 2247]. coefficients [1656, 375, 1876]. coherent [2097]. coinductive [757, 305]. coiteration [94]. COL [1716]. Collage [242]. collapse [1316]. collapses [340, 1865]. Collapsing [2097, 2203, 1451]. Collatz [1429, 1831]. Collatz-like [1429]. collisions [2155]. colonies [1951]. colony [322]. color [1062]. colorability [45, 1642]. Colored [1062, 1897, 1847]. Coloring [810, 435, 2105, 811, 2312, 467, 1795, 2000, 1627, 439, 2246, 441, 1414, 1554, 527, 1131, 2364]. colorings [434, 2097, 944, 432, 2081, 1785]. colors [176, 2246]. colouring [1294, 2026, 1933, 631, 1395]. colourings [2367, 1791, 442]. colours [1791]. column [1366]. columns [2100, 394]. combinations [2059]. Combinatorial [458, 554, 1546, 1784, 2006, 8, 388, 1182, 1363, 2377, 374, 779, 466]. Combinatorics [44, 1441, 2368, 1886, 720]. combinatory [2282]. combined [2204, 1272]. Combining [155, 673, 1067, 590]. comma [1638]. comma-free [1638]. commands [357]. Commitment [1855]. commitments [1061]. commodity [1967]. common [669, 338, 1804, 1179, 1539, 2379, 942, 2382, 1982, 965, 788, 2154, 2008, 1553, 962]. Commons [287]. communicating [1705, 695, 1354, 1026, 1723]. Communication [1222, 780, 1782, 821, 2136, 1339, 952, 1265, 1646, 2177, 1820, 2129, 2022, 28]. Communications [1297, 109]. Commutation [239, 165]. commutative [1630, 2011]. commutativity [1639]. commuting [651]. Compact [1033, 766, 1896, 2164, 1803]. compactness [1734]. companion [1217]. comparability [1559]. comparative [1596]. comparator [2024]. compare [1926]. Comparing [1683, 1183, 705]. Comparison [2128, 1561, 1242, 1917, 1345, 1330, 622, 161, 2384]. comparison-based [1242]. comparisons [918]. compartments [2191, 2188]. Compatibility [519]. Compatible [804]. Competitive [1485, 600, 1558, 2237, 1480, 1647, 1757]. competitiveness [258]. competitivity [710]. complementary [1942]. Complementation [1244, 330, 1092]. Complementing [30]. Complete [567, 1372, 885, 2197, 1524, 835, 1375, 438, 2088, 1510, 2111, 1830, 1610, 1807, 441, 2025, 1800, 1359, 678, 813, 2394]. completely [535]. Completeness [216, 730, 1227, 2367, 174, 290, 1830, 1788]. Completing [2300, 654, 1450, 736]. completion [390, 2258, 1910, 267, 1942]. Completions [8, 1908]. complex [608, 852, 1462, 1627, 2370, 609, 611]. complex-value [852]. complexities [541]. Complexity [1884, 1256, 1163, 597, 1646, 1246, 2254, 1936, 951, 1152, 1333, 1135, 1642, 809, 1878, 2091, 2113, 616, 1978, 1195, 1389, 2388, 1665, 1649, 1560, 804, 1762, 1307, 1494, 181, 2200, 418, 2264, 1815, 1276, 1271, 1198, 20, 1657, 212, 521, 2304, 2086, 2142, 1432, 179, 1852, 1997, 2202, 1296, 1355, 2120, 1828, 1552, 2204, 1917, 1912, 438, 1141, 2208, 207, 293, 1174, 2092, 1790, 2261, 1356, 1173, 2176, 2135, 460, 192, 1654, 24, 1595, 2177, 1310, 2338, 2129, 25, 1392, 707, 165, 1278, 1781, 475, 1403, 1184, 107, 425, 1739, 2122, 1422, 675, 1499, 1523, 1414, 2152, 827]. complexity [2371, 151, 842, 1918, 1413, 2232, 1285, 296, 2090, 349, 1185, 983, 2206, 2125, 1722, 735, 1272, 121, 492, 1547, 1258, 2337, 1274, 936, 812, 941, 1477, 1817, 31, 1569, 398, 1995, 129, 1342, 1944, 1289, 552, 565, 556, 841, 1212, 1509, 993, 1277]. Component [821, 1383, 1236]. component-oriented [1236]. Components [517, 1761, 19, 660, 1354, 1153, 60, 862]. composable [2008]. composition [57, 2294, 821, 300, 725, 865, 264, 301, 1889, 1807, 1386]. Compositional [897, 1527, 591, 1097, 58, 1041]. Compositionality [604]. Compositions [892]. Compound [2087]. Compressed [2161, 1988, 1369, 2380, 2299, 548, 827, 1982, 1313]. Compressing [1366]. compression [856, 1361, 75, 1362, 839, 766]. Comput [1529, 2398, 402, 1690, 560, 1135, 1707, 2396, 1123, 2399, 922]. Computability [1595, 996, 256, 674, 1154, 1513, 1461]. computable [417, 541, 840, 1864, 708, 498, 497]. Computation [832, 292, 1538, 404, 1291, 327, 1127, 1329, 1279, 1997, 384, 290, 532, 358, 1517, 1939, 1157, 1063, 449, 679, 1913, 291, 1994, 122, 319, 614]. Computational [871, 602, 1815, 1931, 1932, 1715, 1935, 1739, 2390, 2090, 1434, 1289, 790, 278, 2227, 1494, 179, 926, 681, 1275, 1717, 1510, 224, 1737, 2384, 884, 2152, 37, 854, 1050, 2337, 1475, 323, 2396, 812]. Computationally [23, 2197]. Computations [1868, 500, 2296, 6, 2322, 338, 660, 688, 1853, 1816, 1519, 321, 1880]. compute [988, 2376, 1982]. Computer [286, 1745]. computers [1689, 1518]. Computing [987, 133, 620, 178, 651, 1540, 1956, 1876, 2154, 1504, 813, 77, 1717, 1139, 251, 1539, 1938, 2147, 908, 177, 375, 1528, 1958, 2289, 1042, 2090, 1856, 1711, 1416, 995, 430, 754, 1018]. concatenation [957, 641, 350]. concept [474, 475, 602]. concepts [2384, 2238, 409, 496]. conceptual [1063]. concerning [564, 1830]. concise [908]. conclusive [2024]. concordance [450]. Concurrency [32, 1435, 704, 1672, 682, 1337, 487, 700, 1077, 1922]. Concurrent [662, 1380, 1338, 1731, 1840, 1076, 108, 186, 863, 513, 860, 59]. condensation [466]. condensing [1374]. condition [372, 1621, 1308]. Conditional [2141, 1806, 1554, 1285, 1386, 1892, 1969]. conditions [1731, 2097, 24, 492, 2359, 72, 31]. conference [733]. confidence [838]. Configuration [2277, 1152]. configuring [510]. confinement [223]. confirmation [1282]. conflict [405]. conflict-free [405]. conflicts [1555]. confluence [2282]. congestion [2044, 2210, 2345]. congruence [2222, 1722]. congruences [142]. conjecture [2167, 1314, 2169, 206, 209, 309, 442]. Conjugacy [2117, 1206]. Conjunction [1029]. Conjunctive [1000, 413]. connected [506, 2332, 1311, 1959, 2253, 2397, 1226, 1306, 531, 1973, 2385]. connection [1519, 1221]. Connectionist [688, 990]. Connections [231, 1104, 445]. connectives [36]. connectivity [1012, 1486, 215, 1223, 312, 2139]. connectivity-based [312]. connectors [882]. coNP [345, 1316]. coNP-hard [345]. conquer [532]. consecutive [2082]. consensus [1533]. consequence [2011]. consequences [364, 816]. Conservation [2241, 1486]. considerations [1518]. considering [2140]. Consistency [1755, 47, 2327]. consistent [980, 419]. consolidation [1989]. constant [1180, 1791, 1907, 1817]. constant-depth [1817]. constants [116]. constrained [1614, 634, 1322, 311, 1646, 2261, 1388, 1980, 1809, 1976, 1500]. constraint [2056, 419, 1839, 150, 290, 793, 2257, 2053, 1695, 2349, 359, 431, 1044]. constraints [1304, 2252, 243, 1764, 1766, 1480, 1515, 1865, 50, 1443, 743, 1859, 148, 1003, 1795, 1483, 2145, 752, 2266, 1463, 332, 1819]. Constructibility [1514, 1518]. Constructing [248, 1984, 1311, 727, 272, 2103]. Construction [1959, 553, 823, 859, 1375, 1506, 1305, 10, 872, 2058, 2061, 2230, 1313, 2180, 1803, 816]. constructions [2073, 1796, 2074, 117, 633, 1820]. Constructive [953, 274, 2103, 1478]. consumption [1952, 1251]. contact [406, 427, 1857]. contain [1651]. Containers [272]. containment [147, 148]. contains [90, 1823]. content [681, 610]. content-addressable [610]. contention [108, 1952]. Context [1091, 29, 385, 386, 583, 1904, 2198, 1069, 1020, 1230, 643, 2360, 220, 692, 581, 1918, 38, 166, 1393, 2079, 170, 1624, 2228, 1538, 1379, 1706, 31, 1098]. Context-dependent [1091]. Context-free [29, 385, 386, 583, 1904, 2198, 1230, 643, 2360, 220, 581, 38, 166, 170, 2228, 1538, 1379, 1706]. context-sensitive [1069, 692]. contextual [959, 893]. Continuant [2162]. continuants [395]. continuations [1067]. continued [1192, 1878, 1477, 2239]. Continuity [256, 1218, 2284]. continuous [327, 2286, 1718, 874, 1595, 1169, 1666, 1747, 1034]. continuous-time [327, 2286]. continuum [1010]. contours [376]. contract [897]. contract-signing [897]. contraction [1186, 2399]. contracts [821]. Control [1381, 701, 2388, 1672, 1597, 487, 281, 1508, 2152, 1596, 2191]. controllable [574]. controlled [2203]. Convergence [115, 2030, 1637, 1132, 1551, 840, 511, 909]. converse [1635]. conversion [1454, 1879]. converting [393]. Convex [1768, 1228, 397, 1765, 394, 640, 63, 371, 376, 1400]. Conway [1903, 465]. cooperating [19]. cooperative [701, 291]. Coordinating [1118]. Coordination [2209, 2048, 355, 1926, 2334, 60, 160, 2138]. cop [1616]. cop-robber [1616]. copies [511]. Copower [2010]. Coppo [1584, 1592]. copulas [1143]. copy [1480]. Coquand [1474]. Core [1074]. corecursive [499]. cores [1801]. corona [1662]. coroutines [489]. correction [599, 2359]. correctness [1921, 1334, 514, 291]. Correlation [774, 1829, 1249]. correspondence [278, 1069, 1852, 2175]. corresponding [908]. Corrigendum [1707, 1706]. corrupt [1344]. Cost [960, 946, 380, 1645, 108, 1641, 784, 1952, 1142, 836, 934, 1889, 1012, 2290, 1499, 2140]. costs [2022, 1575]. Could [612]. count [100]. countable [1506, 1575]. counter [9, 220, 124]. Counterexample [593]. Counterexample-guided [593]. counterexamples [1405, 2137]. counters [27]. Counting [74, 1146, 236, 1537, 2055, 859, 643, 2360, 2254, 574, 1452, 1289]. cover [1177, 2101, 1228, 1604, 823, 978, 175, 249, 423, 1916, 630, 1059, 2372, 1221, 2397, 818, 1973, 2385, 85]. coverage [2395, 62]. Covering [2386, 2088, 1544, 608, 835, 2350, 66, 1114, 732, 217, 2259, 1631, 1011]. CPO [1863]. cpos [874]. CPS [665, 99]. Cramer [976]. Cramer-Damgård [976]. crash [1957]. creation [1382, 991]. Creol [863]. Crick [695, 2202]. criteria [480]. Criterion [838, 1276, 1525]. criterium [208]. critical [1317, 368, 1089, 1220]. Crochemore [2287]. crossed [1789]. Crossing [963, 81]. crossover [911, 694]. crowdedness [1619]. Cryptographic [1507, 95, 50, 568]. Cryptographically [1670]. Cryptographically-masked [1670]. cryptography [938, 1062, 251, 910]. CSC [1383]. CSC-solving [1383]. CSL [1237]. CSP [579]. CSP-CASL [579]. CSPs [2327]. CTL [570, 646, 58, 359]. CTL-like [646]. cube [958, 817]. cubes [1806, 1984, 1789, 1653]. cubic [442]. Cubical [872]. cubically [872]. Curien [1592]. Curtis [1634, 2235]. curve [1469, 1105]. Curves [1476, 1149, 1218, 430]. cut [1102, 2376, 2150, 1309, 555, 893, 565, 1121]. cutsets [1827]. cycle [978, 381]. Cycles [1026, 2358, 109, 1806, 944, 1475, 2081, 1224, 1785]. Cyclic [2146, 2362, 1911, 528, 965, 1991, 1819].

D

[216, 1425, 259, 366, 2304, 1771, 2333, 1302]. D. [553]. D0L [13]. D2B [610]. Damgård [976]. Daniell [708]. darkening [1062]. Data [149, 2022, 1304, 1390, 1005, 446, 1804, 274, 1805, 459, 1674, 1839, 1924, 311, 1174, 276, 324, 1189, 1369, 1694, 1405, 1171, 1581, 857, 1323, 1251, 992, 1579, 1578, 225, 161, 1410, 476, 1643, 1366, 2, 1123, 2065]. data-aware [1369]. data-flow [857]. data-independence [225]. data-oriented [1924]. databases [150, 852, 564, 285]. datalog [1697]. dataspace [55]. datatypes [94]. date [942]. dates [390, 1054, 1488, 1567]. Deadline [1895, 2106]. deadline-TSP [2106]. deadlock [1059]. deadlock-free [1059]. Deadlocks [876]. debugging [586]. decay [2274]. decentralized [427]. Decidability [2113, 343, 1508, 746, 39, 661, 1514, 259, 747, 851, 150, 567, 174, 1201, 525, 496, 1722, 1455].

Decidable [147, 1028, 165, 1438, 894]. decide [1019, 2076]. decide-and-halt [2076]. Deciding [896, 2279, 852, 2219, 1828, 1552, 1954]. Decimations [2122]. Decision [1149, 1713, 15, 327, 1442, 193, 590, 454, 1575, 1092]. Decision-making [1149]. declarative [2321]. declustering [715]. decoders [1857]. Decoding [1170]. Decomposition [138, 979, 1506, 1768, 912, 1690, 1206, 1383]. decompositions [634, 1088, 1618, 1699, 1876]. dedicated [1441, 2302]. Dedication [1260]. deductive [285]. Deeper [447]. Default [853, 189]. Defect [1115]. deficiency [1085, 1815]. definability [1962, 294]. defined [1640, 1126, 888, 359, 1186, 2399]. definedness [1002]. Defining [2224]. definite [765]. Definition [1506, 156, 691]. definitions [164]. deflation [714]. defying [1639]. Degree [634, 1167, 943, 2308, 1395, 1870, 1961, 648, 1023, 1916, 2356, 439, 2228, 777, 1906]. Degree-constrained [634]. Degrees [929, 1798, 81, 813]. Dejean [1314, 2169]. Del [363]. delay [1602, 1178, 1558, 2001, 2022]. delays [1646, 648, 317, 1681]. delegation [264, 1889]. deletable [1770]. deleted [76]. deletion [1641, 64, 29, 2140]. deletions [1908]. delivery [2334, 1488, 2138, 1045, 1503]. Della [1584]. demand [1602, 1643, 1500]. denotational [757]. Dense [1812, 1987, 814, 2238]. Density [1702]. dependence [1756]. Dependences [392]. dependencies [619, 852, 758, 564, 1013]. dependency [585, 1366]. dependent [1496, 1091]. deployment [1667]. depth [602, 1352, 2198, 1988, 1055, 1907, 2003, 351, 1817]. depth-first [1055]. Derek [1490]. derivated [232]. derivation [1630, 2198, 354, 625, 391]. derivations [2079]. Derivatives [69, 391]. descendent [985]. description [338, 157, 152]. descriptional [2202, 24, 27, 1918, 31]. descriptions [1285]. descriptive [877]. design [2045, 154, 505, 222, 511, 1666, 1811]. Designing [80, 2211, 83]. detect [1853]. detecting [128]. Detection [1533, 1849, 1824, 1534, 815, 1957, 1701, 225, 727]. deteriorating [793, 2249, 1975]. determinacy [543]. determinants [1875]. Determination [2199, 63, 690]. determined [535]. Determining [1684, 2356, 1739]. determinism [2272, 318, 107, 2219]. Deterministic [1777, 2037, 798, 824, 2346, 2107, 823, 628, 1607, 389, 801, 826, 2040, 213, 1615, 2268, 1359, 1343, 1132, 1520, 124]. determinization [829]. determinized [483]. development [2014]. Developments [1082, 1579]. devices [932, 1836, 2031, 1144]. deviousness [1844]. Dezani [1584, 1592]. Dezani-Ciancaglini [1584]. DFA [26, 2124]. DFAs [21]. DFCA [2124]. diagonal [1009]. diagrams [1850, 213]. diameter [1827, 1914]. dichotomy [435]. dictionaries [1242, 1180, 1369]. dictionary [246]. difference [1656]. differences [2002]. different [1482, 985, 26, 1739]. Differentiable [768]. Differential [850, 216, 85, 1548, 1022, 1835, 1159]. diffusion [1637, 1681]. Digest [459, 401]. digit [251, 501, 1902]. Digital [1112, 1746, 2239]. digits [76]. Digraph [1618, 2388, 1823, 1822]. digraphs [2358, 2081]. dihomotopy [876]. Dilemma [1118]. dimension [1576, 339, 461, 1470, 1215, 474, 633, 1937, 475, 2099, 842, 517, 1623]. dimensional [925, 245, 2227, 2037, 1752, 1818, 242, 179, 2148, 320, 388, 963, 370, 2377, 2225, 707, 817, 105, 1115, 1457, 377, 378, 2006, 922, 659]. dimensionality [2052, 992]. dimensions [718, 465]. diophantine [559, 1057, 2363]. Direct [2354, 2282]. directed [2358, 1565, 1179, 555, 1372, 1566, 2154, 971, 2224, 2139].

Directional [1621, 2331]. discipline [1589]. Discontinuities [1580]. discounted [333]. discounting [2223]. discovery [1845, 2291, 846, 130]. discrepancy [44]. Discrete [1196, 1763, 1022, 884, 1188, 1760, 1761, 1327, 1529, 365, 1449, 2304, 1767, 2096, 396, 1718, 370, 1907, 1171, 1381, 572, 1940, 1394, 1132, 1755, 391, 1966, 1943]. discrete-event [1381]. discrete-time [572]. discreteness [995]. Disjoint [2358, 1761, 1107, 1648, 969, 958, 719, 1984, 1610, 1553]. Disk [2026, 1604, 1294, 1968, 1121, 1306, 531, 2385]. displacement [1792]. displacements [2108]. dissimilarities [1391]. Distance [2352, 1609, 729, 2376, 171, 2119, 179, 269, 1545, 2253, 177, 1976, 1343, 1829, 935, 556]. Distance- [1609]. distance-hereditary [269]. distances [2296, 2367, 1122, 1784, 1538, 2339, 754]. distillation [1249]. distinct [1948, 1146]. Distributed [1603, 2023, 19, 1672, 859, 865, 47, 1783, 1607, 1296, 1951, 1552, 292, 1924, 384, 157, 55, 863, 1311, 1381, 1422, 373, 1957, 1958, 1221, 1677, 1913, 321]. Distribution [532, 471, 943, 1167, 960, 1961, 623]. distributions [1761, 2062, 1171]. Divergence [498, 1019]. diversity [2128]. divide [532]. Divisibility [201, 799]. division [2015, 980, 1716]. divisions [248, 536]. DLS [731]. DLS-trees [731]. DNA [22, 1717, 117, 2032, 2033, 119, 689, 1942, 2031]. DNA-based [2033]. DNAzymes [2031]. DNF [393, 537]. do [1531, 111]. Do-All [111]. documents [1367]. does [1939]. Dolev [224, 227]. Domain [1378, 98, 1514, 915, 1877, 877]. domain-like [877]. Domain-theoretical [1378]. domains [870, 919, 179, 1595, 1156, 277, 1741, 1188, 1034, 699]. dominance [2242]. dominating [1319, 1307, 1987, 1968, 2356, 999, 1321, 1959, 2253, 2390, 1226, 1306, 531]. domination [2352, 2351, 445, 1992, 726, 545, 534]. domino [1786]. don’t [2291, 803]. dot [351]. dot-depth [351]. double [2376, 711, 786, 1213]. double-loop [711]. double-quiescent [786]. doubling [2006]. down [151]. Dpi [1038]. Dr [2303]. Drawing [1847, 2004]. driven [791, 582]. Dual [1171, 1445, 1139, 350, 1556, 751]. Dual-bounded [1171]. duality [1445]. dualization [1247]. duals [530]. due [942, 1211, 1054]. duplex [2032]. duplication [943, 380, 1977]. duplications [1312]. duration [174]. Dutch [268]. Duval [206]. Dyck [1790]. Dynamic [2045, 1899, 952, 2275, 1983, 2298, 1242, 740, 2151, 1805, 1783, 1674, 1608, 914, 1483, 276, 2299, 1185, 1034, 1073]. dynamical [2241, 1327, 1529, 281, 1394, 313, 2337]. dynamically [142]. Dynamics [1516, 1196, 2331, 1561, 1595, 1026, 1621, 1681].

E-pattern [477]. e-services [264]. EA [1683]. early [1954, 2361]. early-deciding [1954]. EAs [767]. Easiness [577]. economies [1852]. EDAs [766]. Eddy [1525]. Edge [631, 1395, 930, 1892, 1897, 2340, 709, 1319, 1987, 2312, 543, 2026, 961, 1262, 2000, 958, 1908, 944, 1984, 1247, 1610, 2246, 1427, 1556, 2364]. Edge-bandwidth [930, 709]. edge-biconnectivity [1556]. edge-colored [1897]. edge-coloring [2246]. Edge-colouring [631, 1395]. edge-disjoint [1984, 1610]. Edge-fault-tolerant [1892]. edge-intersections [1247]. edge-labeled [543]. edge-modification [2340]. edge-pancyclicity [1427]. edges [2386, 772, 1631]. edit [171, 1976]. editing [1351, 2391, 1964]. edition [1075]. Editor [1946, 136, 1341]. Editorial [230, 17, 52, 92, 103, 126, 135, 153, 184, 199, 211, 228, 270, 279, 298, 403, 443, 587, 561, 575, 738, 761, 769, 807, 847, 868, 886, 906, 913, 923, 964, 986, 997, 1017, 1027, 1035, 1064, 1081, 1094, 1124, 1136, 1165, 1190, 1234, 1325, 1348, 1387, 1415, 1479, 1433, 1458, 1505, 1512, 1570, 1600, 1636, 1660, 1832, 1866, 1919, 2158, 2184, 1466, 518, 820, 697, 576]. EF [702]. Effective [1937, 1964, 874, 500]. effects [278, 673, 884, 1067, 854]. efficiency [1953, 1278, 2065]. Efficient [2221, 2296, 327, 1148, 1662, 2101, 1979, 931, 1219, 549, 384, 1824, 1730, 773, 572, 374, 1175, 1982, 2180, 2280, 1666, 580, 1994, 2075, 2136, 711, 558, 737, 204, 1264, 329, 966, 815, 1699, 981, 391, 2005, 1144, 976, 1173, 1171]. Efficiently [608, 430, 1344, 2238, 971]. Ehrenfeucht [294]. eigenvalues [908]. Eigenvariables [747]. eigenvectors [908]. Eilenberg [1909]. Einstein [1513]. Election [1605, 960, 723, 667, 1384]. elections [2388, 460]. electromagnetic [1774]. element [752, 2290, 1312]. element-cost [2290]. Elementarily [417]. Elementary [2002, 786]. elements [1156, 2108, 1113, 2126]. eligibility [2266]. eliminating [1701]. Elimination [704, 36, 968, 1702, 904, 252, 1880, 565, 1906]. embedded [1712, 665, 1938, 472]. Embedding [1727, 180, 1653, 1789]. emergence [611]. emerging [129]. Empire [1951]. empirical [1143, 1278]. EMSO [698]. encapsulation [2354, 1099]. encoded [1367, 1540, 2298, 2263]. encoding [665, 366, 1210, 143]. encryption [910, 227]. endgame [447]. endgames [446]. Energy [2136, 1486, 988, 558, 311, 1817, 1604]. energy-constrained [311]. enforces [95]. Enforcing [1639]. enhanced [234, 321]. enquiries [1579]. Enriched [1743, 1379, 1706]. ensuring [1648]. entanglement [1390]. entrant [1629]. Entropy [461, 490, 1016, 2380, 423, 1536]. entropy-bounded [2380]. enumerability [340]. enumerating [2238]. Enumeration [1350, 1214, 394, 371, 945, 2221, 522, 2384]. environment [910, 1072]. environments [1737, 1575, 1756]. epichristoffel [2251]. epistemic [569]. episturmian [1482, 2165, 1200, 1905, 261]. eponymous [1956]. equal [2168, 1644]. equal-processing-time [1644]. equality [1031, 152]. equation [342, 787, 302, 12, 77, 1837]. Equational [1698, 896, 33, 2015, 1031, 2058]. equations [2055, 2170, 343, 1022, 2174, 1835, 1202, 559, 1656, 349, 454, 1092, 1057]. equidistribution [2247]. equilibria [2044, 776, 1649, 864, 2046, 1852, 2047, 1483, 1141, 2208, 293, 1882, 2345, 2261, 2049, 1138]. equilibrium [1139, 360, 182, 1140]. Equitable [434, 1785, 439]. Equivalence [1086, 1028, 822, 13, 746, 690, 1684, 349, 2181, 2239]. equivalences [1527]. equivalent [337, 698, 857]. erasing [1089, 410, 844]. Errata [1529]. Erratum [2398, 402, 758, 1690, 560, 1135, 2396, 1123, 922]. erroneous [459]. error [84, 2063]. error-free [84]. errors [2378, 813, 88]. Essays [1584]. essence [303]. establishment [1948]. estimates [2069]. Estimation [2204, 1927, 2005]. EU [1416]. EU-FET [1416]. Euclidean [1295, 2006]. Euler [1692]. Eulerian [1874]. evaluating [1647]. evaluation [1000, 250, 452, 150, 450, 451]. evaluators [278]. event [1925, 1381, 699, 1819, 2277]. event-based [1925]. Events [1083]. Every [1220, 90]. everywhere [1110]. evolution [1535, 764, 1713, 1841]. evolutionary [2197, 1683, 20, 731, 1169, 1330, 1128, 2132, 1688, 693]. Evolving [1841, 1743, 1277]. EX [702]. Exact [311, 131, 502, 499, 1321, 1157, 626, 86]. examples [1252, 412, 1345, 2137, 1738, 844, 626]. exception [537]. Exchange [1138, 149, 1144, 1421]. Excluded [479]. exclusion [2259, 876]. exclusivity [291]. Executability [2013]. execution [354]. execution-time [354]. exercise [862]. exercises [1079]. existence [2094, 1327, 1529, 2256, 1353, 1088, 2033, 816, 872].

Existential [2267, 589]. Expander [2372]. Expanders [1894]. expansion [938, 1703, 251, 2228]. expansions [1193, 251, 76]. expected [468, 314]. Experiments [826, 2384, 1050, 812]. Explicit [16, 1417, 96, 670]. exploitation [2069]. Exploration [2069, 344, 600, 1658, 160, 1910]. explorer [2214]. Exploring [606]. explosion [328]. exponent [2363, 87, 1220]. Exponential [2145, 1817, 791, 2142, 2274, 317, 1316, 1342]. exponents [1089]. EXPRESS [32]. expression [1896]. expressions [245, 1727, 1630, 2205, 972, 69, 2219]. expressive [1865, 664, 663, 1511, 491, 454]. expressively [698]. Expressiveness [32, 1931, 922, 659, 1699]. Expressivity [1439]. EXPTIME [1807, 38]. EXPTIME-complete [1807]. EXPTIME-hard [38]. extended [663, 206, 1511, 2327, 1370]. Extending [1909, 1599, 516, 1582, 1592]. extensible [128]. Extension [1630, 1201, 901, 1368]. extensions [970]. external [1189, 1055]. external-memory [1055]. Extracting [1322, 274]. extraction [846]. extractions [1413]. extremal [1117]. extreme [293, 2345].

face [2004]. facility [836, 1290, 781]. Factor [1195, 2157, 1389, 1262]. factorial [2174]. factoring [1522, 976]. factorisation [1862]. factorization [1317, 368, 1966]. factorizations [649]. factorizing [234]. Factors [169, 1217, 240, 1197, 464, 2233, 1625, 557]. Fagin [564]. fail [1966]. fail-stop [1966]. failing [1262]. failures [34, 111, 983]. failures-based [34]. Fairness [660, 1310]. families [1581, 2334, 2079, 1623, 84, 1342, 816]. family [1716]. Fano [442]. farthest [1504]. Fast [1896, 1607, 2083, 1981, 1364, 2289, 1042, 728, 1532, 599, 1481, 1658, 1872, 2263, 1941, 797]. Faster [925, 1179, 1114, 2380, 1365, 46, 1182]. fat [1949]. Fault [1806, 1789, 1913, 1969, 180, 2397, 1892, 1973]. Fault-free [1806, 1969, 180]. Fault-tolerant [1789, 1913, 2397, 1973]. faults [255, 1608, 2305, 1806, 180, 1300, 1969]. faulty [2070, 1676, 2025, 1113, 1427]. Feasibility [983]. FeatherTrait [1599]. featured [1729]. Feedback [1266, 838, 530]. feedforward [414]. Festschrift [2270]. FET [1416]. few [2344, 1949, 2088, 178, 15, 1079]. fewer [1400]. FFT [2143]. Fibonacci [181, 464, 828]. Fibonacci-like [181]. fibrational [1066]. Fibrations [1376]. field [2370]. fields [553, 2109]. FIFO [1104]. figures [1210]. fill [445]. fill-in [445]. filter [1593]. filters [727]. final [739, 2347, 756, 191, 448]. find [711, 2345]. Finding [2367, 2164, 1767, 2111, 985, 927, 1332, 2034, 1804, 1565, 1626, 1179, 523, 480, 1278, 1809, 2263, 2374, 129]. Fine [1745, 205, 1446, 59, 2183]. fine-grained [59]. Finitary [1735, 191]. Finite [71, 1909, 1736, 2225, 1210, 625, 799, 33, 553, 1327, 1529, 1727, 2244, 2199, 461, 823, 1244, 2117, 246, 9, 1162, 1888, 1032, 2077, 646, 344, 1574, 1446, 1356, 972, 24, 358, 650, 1916, 1807, 705, 2147, 209, 1684, 349, 1343, 2181, 1093, 2075, 843, 2109, 644]. finite-equational [33]. finite-state [461, 1032, 972, 358, 705]. finitely [880]. finiteness [894]. first [741, 1363, 1055, 1558, 39, 294]. first-order [39]. Fisher [1139]. Fitness [1130]. five [2356]. Fixed [2313, 1764, 1815, 1657, 2304, 484, 595, 2092, 236, 1089, 1422, 209, 28, 490, 940, 491, 777]. Fixed-parameter [2313]. fixed-priority [595]. fixpoint [856]. fixpoints [516]. FKP [1167]. flag [268]. flat [1760, 976]. flat-tree [976]. flexible [914, 1936]. flips [1194]. Flow [2191, 1071, 274, 370, 555, 1629, 857, 305, 1374, 962]. flows [1670, 422, 555, 1173, 951]. Floyd [387]. fly [329, 2285]. Focusing [2326]. folded [1324, 257]. following [1474]. follows [1659]. forbidden [2386, 721]. forecasting [2063]. forecasts [992]. forest [1050, 2007]. forests [1648]. Foreword [481, 1750, 1583, 127, 1250, 1929, 1709, 2270, 1571, 1095, 470, 2317, 1360, 2375, 2310, 837, 1531, 2287, 869, 1733, 2365, 444, 136, 1341]. Forewords [1613]. form [583, 1904, 822, 1502, 2338, 1943]. Formal [898, 859, 157, 511, 1927, 306, 905, 637, 757, 1420, 1735, 890, 1659, 375, 449, 1732, 1744, 1749, 844, 920]. Formalising [1417]. Formalization [1239]. formalized [1692]. formation [1551, 1814, 1942]. formats [1040]. formed [1705]. former [230]. forms [749, 1024, 765, 1873]. formula [1692]. formulae [74, 1501]. formulas [1815, 2172, 826, 1810]. forth [885]. forward [2229, 161, 1144]. FOSSACS [1434]. foundations [740, 755, 222, 1416, 1434]. four [1443, 2293]. four-stage [2293]. Fourier [1637]. fourth [1911, 1018]. FPT [519]. FPTAS [2145]. fraction [2239]. fractional [1608]. fractions [1192, 1899, 1878, 1477]. Fraenkel [2012]. fragment [680]. fragments [146, 589, 1090]. Frame [2283, 1072]. framework [356, 1037, 1761, 1419, 1421, 2189, 1066, 1063, 903, 622, 67, 1811]. Free [1440, 90, 385, 386, 583, 1904, 237, 1630, 2198, 2163, 433, 1230, 1562, 643, 2360, 1028, 220, 310, 2058, 1806, 1373, 1412, 2135, 12, 180, 2240, 119, 1893, 581, 2260, 405, 1059, 621, 1554, 1823, 29, 38, 166, 1638, 2316, 170, 1624, 2228, 1538, 84, 1205, 1379, 1706, 16, 314, 535, 1969, 806]. Free-algebra [1440]. Freedom [1138]. Freeness [1409, 1318, 1971, 2324, 1986]. Frege [671]. frequencies [2157, 1043]. frequency [6, 1509, 1205]. frequent [790]. freshness [273]. frictional [832]. friendly [2143]. Fulkerson [442]. Full [1436, 1029].

Fully [786, 1926, 186, 942, 1771]. function [725, 259, 643, 1284, 1229, 451, 2233]. Functional [1194, 278, 619, 852, 758, 192, 1071, 908, 917]. functionalities [1782]. functionally [1830]. functionals [708]. Functions [195, 471, 1086, 988, 2256, 417, 1560, 418, 179, 1549, 2360, 1429, 452, 2128, 836, 450, 1790, 1573, 789, 1639, 1330, 1967, 1133, 1448, 1807, 297, 1874, 472, 291, 1476, 1253, 1286, 1994, 1572]. functor [191]. functors [2010, 756]. fundamental [138]. Fürstenberg [1693]. future [1711]. futures [860]. Fuzzy [385, 386, 920].

G [553, 1427, 2084]. Game

[286, 1798, 99, 592, 1594, 2208, 1616, 451, 1490, 294, 1233, 1056, 2339, 448, 1906]. Games [289, 864, 1882, 296, 492, 1884, 2044, 1898, 458, 2210, 1380, 78, 1852, 543, 484, 1828, 1881, 2345, 2261, 1310, 1618, 1967, 2049, 1985, 700, 686, 1338]. gamma [1874]. gap [1540]. gapped [2381, 2292]. gaps [2257, 1110]. gas [1686]. gates [123]. Gathering [1949, 168, 1424, 311, 1299]. GBL [2011]. GCD [1543]. GCDM [1543]. GCDs [1871]. gem [433]. GemCell [1464]. gene [380, 638, 1837, 1523]. General [2230, 774, 253, 1055, 1345, 475, 705, 767, 1780, 496, 2337, 1747, 745]. generalisation [138]. Generalised [205]. Generalising [1640, 1162]. generalization [1971, 2057, 1634, 2308, 346, 2251, 710]. generalizations [1482, 161]. Generalized [1891, 2269, 915, 1452, 1723, 1217, 5, 1194, 943, 755, 2165, 1979, 1219, 1790, 1310, 840, 463, 1043, 1342, 385]. generated [1655, 1106, 2342, 975, 2079, 2130]. Generating [583, 1904, 1172, 172, 1761, 932, 1171, 1448, 812]. Generation [341, 1350, 155, 1947, 1214, 1731, 397, 21, 566]. generative [166]. generator [1900]. generators [1393]. Generic [854, 1464, 1516, 1156, 1120]. generics [1923]. genes [911]. genetic [911, 694, 1931]. genome [1977, 1534, 1312]. genomes [1534]. genomic [2376]. genus [1692]. geodesic [220]. geometric [1177, 2150, 948, 1210, 309]. Geometrical [680, 2039, 1197, 2377]. geometry [365, 486, 488, 1738, 1744, 405]. geotomography [1774]. GF [263, 552]. Giorgio [2318, 2319]. Girard [1075]. girth [2364]. given [1798, 711, 167, 2363, 474, 2144, 1117, 1914, 992, 1877]. Glance [1950]. Global [909, 1681, 645, 1416]. Go [449, 448]. Gödel [1513]. Goemans [598]. Going [2391]. good [1626]. gossip [384]. gossiping [1264, 46, 1487, 1103, 1302]. GPs [767]. grafting [1885]. grained [59]. graining [768]. grammar [19, 1354, 827]. grammar-compressed [827]. grammars [385, 583, 1904, 822, 2198, 1020, 642, 585, 1392, 1564, 959, 1918, 166, 2079, 170, 893, 1379, 1706, 235, 2, 1123, 2065]. granularities [1033]. Graph [1492, 366, 344, 834, 432, 2018, 2091, 1406, 1614, 2157, 709, 364, 577, 791, 2367, 71, 804, 638, 1485, 1565, 1933, 172, 1962, 612, 420, 1912, 1620, 1613, 324, 648, 1658, 2033, 2027, 1117, 795, 526, 1615, 81, 265, 2154, 2392, 1910, 1131, 1308, 1096, 2090]. graph-driven [791]. Graphical [466, 1525, 2261, 689]. graphs [1897, 1406, 2113, 48, 592, 1227, 1177, 1847, 1818, 2330, 634, 1813, 1798, 1468, 1622, 2398, 1172, 433, 434, 1827, 1411, 1228, 2389, 804, 1804, 1294, 1987, 1559, 175, 1216, 1557, 2312, 2352, 2351, 381, 1779, 280, 1506, 1179, 1305, 628, 631, 1395, 774, 467, 436, 1245, 1121, 2150, 1263, 2369, 1791, 437, 1498, 2088, 970, 719, 1015, 1908, 1055, 525, 2332, 944, 180, 269, 1916, 2393, 1637, 2004, 2356, 1610, 439, 1059, 1816, 1414, 2105, 1486, 1554, 215, 1563, 2253, 442, 2061, 527, 2316, 1568, 1780, 2372, 973, 1119, 305, 828, 971, 1635, 1332, 1306, 1892, 294]. graphs [1569, 2255, 531, 1224, 1289, 726, 466, 1887, 2084, 2364, 1101, 1965, 2385, 2006, 1785, 1212, 2094, 2026]. Gravitational [1874]. Gray [1049, 1546]. greater [1220]. greed [2382]. greedoid [1978]. Greedy [76, 251, 1267]. Green [365]. Greibach [1904]. grid [1604, 1853, 1055]. grids [1047, 525, 930]. Gröbner [1879]. grooming [2248, 1293, 1301]. gross [1139]. ground [2107]. Group [1645, 441, 1979, 952, 220, 2092, 1820, 1893, 1999, 1209, 2140]. Group-strategyproof [1645]. groups [713, 248, 1346, 349, 16, 736]. Growth [237, 1087, 259]. growths [13]. GTP [2194]. GTP-binding [2194]. guaranteed [1620, 1895]. guarded [357]. Guest [136, 1341]. guided [593, 1351, 22].

H [435, 321]. H-Coloring [435].

Hadamard [1801, 532]. Hadwiger [1995]. Haines [1357]. hairpin [1942]. Hajós [429]. halt [2076]. halting [1520, 1120, 1943]. Hamiltonian [1806, 269, 267]. Hamiltonicity [1892]. Hamming [1818, 1545, 177]. handling [1275, 1491, 88]. Hanoi [954]. haplotype [1152]. haplotyping [134]. Harary [2094]. Hard [2257, 1406, 134, 1990, 670, 80, 459, 1767, 1121, 1275, 2269, 2345, 1373, 345, 1650, 38, 779, 626]. Hardness [2340, 2248, 1057, 2351, 1298, 2106, 1565, 1498, 1391, 413, 2053, 984, 1408, 1852]. hardware [2322]. HasCasl [562, 2014]. hash [816]. hashes [1510]. HASSE [213]. Hausdorff [633]. having [1493, 13]. HD [56]. HD-automata [56]. HD0L [13]. HDTP [582]. Head [967]. heads [27]. heart [1590]. Hedlund [2235, 1634]. helix [132]. Hennessy [33, 604]. Herbelin [1592]. hereditary [970, 269, 545]. heritage [1592]. Herman [468]. Hermite [1872]. Heuristic [1801, 1629, 582, 2138]. heuristic-driven [582]. heuristics [611, 1332]. Hex [445]. hidden [1536, 1184]. hierarchical [1796, 301, 2328]. Hierarchies [542, 2200, 2203, 893, 120, 101, 1745, 2234, 1865].

Hierarchy [2181, 337, 1352, 1622, 2398, 34, 520, 901, 118, 2079, 1316, 1704, 351]. High [1782, 154, 156, 492]. high-level [154, 156]. higher [94, 96, 620, 1071, 1072, 2014, 922, 659, 665]. higher-order [94, 96, 620, 1071, 1072, 2014, 665]. highly [601]. Higman [1357]. histories [1799]. history [137]. hitting [2311, 1916]. Hoare [1041]. hoc [1665, 1663, 157, 1664, 1103]. HOL [357]. Holant [1281]. hold [1739]. holds [2169]. Hole [1298]. holes [2288]. Holographic [2052]. holonomic [375, 2236]. homeomorphism [1053]. homogeneous [2104]. homomorphism [1695]. homomorphisms [535]. honeycomb [371]. honor [1531, 2287]. honour [1584]. Hop [1948, 1295, 312]. Hopcroft [2162, 2295, 2124]. Hopfield [324]. Horn [1005, 1640, 1031]. Horspool [728]. hospitals [1626]. hospitals/residents [1626]. hotlink [1267]. Huffman [2262]. Hybrid [1061, 2197, 593, 1161, 1423, 2193, 140, 20, 2046, 281, 1726, 912, 1690]. hybridization [88]. hyperball [1762]. hyperbolic [1786, 1749]. hypercomputation [1515, 687]. hyperconcepts [809]. hypercube [1324, 1223, 1113, 2141]. hypercube-like [1223, 1113, 2141]. hypercubes [2309, 1972, 257]. hypergraph [45]. hypergraphs [1247, 2139]. hypermap [1692]. hypermap-formalized [1692]. hypersurfaces [729]. hypotheses [1402]. Hypothesis [480].

I/O [2193]. ICALP [2302]. icosahedral [2032]. ideal [1469, 2147]. Idealized [1438]. ideals [1478]. idempotency [975]. identical [778, 1156, 1108]. Identification [652, 1129, 1811, 2, 1123]. identifying [2061]. identities [1948, 1655, 1006]. Identity [1396, 1862, 2354, 735, 1887, 1811, 1144].

Identity-based [1396, 2354, 1811, 1144]. IEEE [503, 551]. iet [1628]. If [789]. II [1423, 222, 1154, 2183]. im [1249]. image [1774]. images [1770]. imbalance [982]. immune [1682]. immunity [383]. impact [2132]. imperative [857, 867]. imperfect [522]. implementation [1860, 499, 571]. Implementing [502]. implements [1939]. implication [852]. implications [1301, 431]. Implicit [369]. implicitization [1105]. Impossibility [1299]. Improved [1129, 2073, 1987, 796, 249, 715, 2074, 833, 2246, 984, 1635, 2113, 598, 811, 1543, 255, 1428, 120, 1110, 2005]. improvement [553, 2312]. Improving [2001]. Inapproximability [2095, 2145, 1992, 1779]. including [132]. inclusion [752, 2259, 282, 2394]. incomparable [1605]. increase [1211]. increased [1251]. increases [2068]. increasing [1147]. Incremental [1845, 823, 781, 859, 479, 449]. Independence [800, 1514, 2324, 225, 735]. independence-based [2324]. Independent [545, 2369, 1313, 1780, 1992, 531, 2084, 534]. index [40, 61, 91, 112, 125, 144, 162, 183, 198, 221, 41, 493, 759, 785, 760, 494, 1798, 1917, 550]. indexed [388, 1581]. indexes [1371]. indicating [800]. individual [134]. induce [1882]. Induced [973, 1352, 1228, 381, 1414, 1823]. Induced-path [973]. Inductive [409, 1574, 1573, 479]. inductively [476]. inefficient [1171]. Inequalities [261, 3, 402, 426]. inequationally [1385]. inertia [1763]. inexact [128, 1108]. inexact-suffix-tree-based [128]. inferable [1578, 476]. inference [1589, 902, 1574, 1835, 1573, 479, 409, 1323, 1580]. Inferring [814, 1391]. infinitary [2055, 890]. Infinite [1207, 1195, 1389, 1352, 458, 2056, 554, 2168, 1909, 464, 1790, 1986, 399, 1996, 1632, 705, 2079, 1205, 203, 1274, 39]. infinite-state [705, 39]. infinite-valued [2056]. infinity [1693]. inflationary [484]. Influence [992]. Information [1531, 1747, 336, 355, 164, 292, 2212, 963, 2324, 873, 1916, 1616, 348, 1742, 1374, 1108, 614]. Ingo [2303, 2302]. inheritance [1096]. inhibitors [1719]. initial [2347, 1621]. initialization [1663]. initiative [1416]. Injective [901, 745]. innocence [900, 700]. input [459, 953, 1895]. input-queued [1895]. insecurity [193]. insertion [1351, 585, 1564, 29, 2140]. insights [1906]. inspired [1717, 1942]. instability [1211]. instances [1295, 1201, 178]. institutions [1160]. integer [791, 1762, 743, 1769, 373, 68, 962]. integers [1456, 941]. integral [1772]. integrands [503]. Integrated [2014]. Integration [1675, 579, 708]. integrity [1003]. intelligence [686]. intelligent [515]. intensional [756]. interaction [850, 488]. interactions [513, 1710]. Interactive [486, 1863, 1316, 1249]. Interchange [2290]. intercluster [935]. interconnection [1113, 2141]. Interfaces [756, 1599]. Interference [1687, 1665, 1953, 1664, 1508]. interleaved [1170]. intermediate [2190]. Internal [2373, 1232, 959]. International [32]. interpolating [331]. Interpolation [2144, 1869, 2009]. interpretation [856, 2325, 1685, 1510, 163, 224, 99]. interpretations [192]. interrupt [880]. Intersecting [1046]. Intersection [615, 881, 1591, 1589, 1254, 2116, 1780, 2372, 1590]. intersection-closed [1254]. intersections [1247]. Interval [110, 1519, 1322, 505, 1216, 1541, 506, 1952, 1795, 1912, 2123, 2316, 1568, 332, 1910, 1408]. Interval-valued [1519]. intervals [811]. intractability [177]. intrinsic [841]. Introducing [1934]. introduction [396]. intruders [1603]. intrusion [225]. intuitionistic [1860, 1692, 688, 2326]. Invariance [916]. invariant [1204, 795]. invariants [867]. Inverse [1545, 345, 1218, 1448]. inverter [2003]. Invertible [1283, 1009, 789]. Investigating [2094]. Investigations [751]. irrational [169, 1112, 2239]. irreducible [115]. irredundant [1845]. irrefutable [622]. irregular [1593]. irrevocable [622]. Ising [320, 2030]. isolated [714]. Isolation [2384, 2238, 1878, 1477, 2339]. isomorphic [985]. isomorphism [804, 349, 1075]. isoperimetrically [1773]. isotopic [871]. isotropic [765]. Issue [2302, 334, 1441, 1613, 1531, 2287, 326, 588, 93]. issues [1277]. items [1899, 1146]. Itemset [1509]. Iterated [932, 1447, 975, 16, 1118]. Iteration [94, 174, 374, 1438, 1994]. iterations [504]. iterative [2068]. iterativity [1377].

Java [54, 1417, 914, 1599]. job

[2151, 2153, 2334, 1488, 2138, 542]. jobs [805, 793, 2249, 2361, 2334, 1054, 1975, 1013]. John [1080]. join [2376, 631]. joining [1783, 2083]. joint [2232]. Jones [1042]. Jordan [1218]. Jug [1547]. Juhani [2185]. jump [2002]. juntas [1280, 2341, 2269]. Justified [669].

k-Partitioning [1044]. Karhumaki [2185]. Katriel [1404]. Kayles [2087]. Kelly [1618]. Kemeny [2313, 460]. Kemp [632]. Kerberos [898]. Kernel [1753, 1864]. kernelization [1964]. kernels [2067]. key [1641, 1796, 910, 2354, 1999, 2140, 1144]. keyboards [80]. KGS [1144]. kinematics [1115]. KLAIM [664]. KLAIM-based [664]. Kleene [1659]. knapsack [693, 536]. kneading [181]. knock [1494, 2019]. knock-out [1494, 2019]. know [1531]. knowledge [896, 669, 670, 1061, 1265, 1609, 846, 2008, 1133]. Kolmogorov [616, 541, 1285, 735, 1274]. Korteweg [77]. Kuba [1759].

L [394, 640, 711]. L-convex [394, 640].

L-shape [711]. Label [940]. labeled [1172, 1243, 543]. labeling [2243, 1635]. labelings [440]. Labelle [1850]. labelled [98]. labelling [1150]. labels [1605]. laceability [215, 1223]. Lambda [1586, 615, 275, 1585, 1591, 1592, 1703, 1587, 860]. lambda-terms [1703]. Lambek [678]. landscape [1130]. Langevin [1837]. Language [2205, 1092, 2190, 2039, 932, 1355, 220, 663, 1373, 1084, 510, 13, 692, 1082, 1345, 426, 1157, 727, 2079, 454, 2328, 917, 614]. language-based [1373]. Languages [1315, 975, 2221, 278, 2321, 1240, 245, 2227, 2037, 385, 386, 70, 603, 244, 1252, 2244, 2056, 241, 116, 2199, 71, 2200, 2116, 1923, 8, 642, 1230, 1048, 643, 2360, 2203, 2148, 2342, 1235, 1006, 238, 2174, 2343, 2076, 1790, 1356, 2175, 1412, 2135, 692, 1926, 1405, 1346, 25, 853, 26, 165, 119, 581, 1753, 1448, 2122, 1581, 652, 857, 1358, 1072, 2228, 1538, 410, 1579, 379, 235, 477, 1041, 1100, 2337, 476, 894, 1056, 2130]. Laplacian [1914]. Large [2369, 817, 1780, 1435, 1792, 7, 1314, 1395, 1646, 1816, 81, 203, 1332, 2364]. Larger [1010]. largest [381, 474, 87, 908, 788]. last [2168]. latency [1967]. lateness [1225]. Lattice [1047, 1885, 1765, 371, 910, 2236, 1686]. lattice-based [910]. Lattices [2148]. law [1435, 1498, 1816]. Lawson [1749, 681]. Lawvere [884]. layered [414]. lazy [1071]. LCS [1891, 2289]. lead [1740]. Leader [667, 960, 723, 1384]. Leading [1757]. leads [1138]. leaf [2314, 985, 2394]. leaf-agreement [985]. learnability [2341, 1579, 1343]. learnable [477]. learners [2137]. Learning [1280, 1752, 1344, 1252, 1254, 845, 412, 1405, 1346, 1582, 1581, 410, 411, 844, 2065, 1572, 448, 471, 473, 2068, 250, 2062, 414, 2070, 1794, 324, 1284, 1345, 2064, 475, 1753, 909, 2066, 478, 1754, 282, 1132, 1755, 1756, 472, 2359, 408, 843, 841, 1531, 479]. least [1255, 491]. left [251, 982, 627]. left-arm [627]. left-right-imbalance [982]. left-to-right [251]. Leiss [954]. lemma [2103, 2288]. lemmata [1520]. length [90, 218, 2200, 381, 1648, 1263, 1540, 2298, 2361, 1809, 2263]. length-constrained [1809]. lengths [219]. Leontief [2046]. Leontief’s [1138]. less [1511]. Letter [1205, 1085, 245, 2037, 1493, 2168, 2175, 2181]. letters [2260, 1043]. level [1782, 154, 156, 510, 814, 1041]. level-1 [814]. levels [26]. Levi [2318, 2319]. lexicalized [2228]. lexicographic [172]. liberal [1028]. library [505]. Life [1010]. light [1233]. lightweight [1950]. like [181, 2026, 606, 1429, 646, 666, 1223, 1113, 2141, 877]. limit [1010, 843]. limited [168, 1309, 1405, 2079, 2003, 838]. Limiting [2268]. limits [2177, 1439, 1678]. line [1846, 167, 219, 467, 66, 73, 2144, 1824, 1800, 528, 1803, 1045, 1503, 839, 1757, 1013]. Linear [421, 389, 117, 269, 569, 2003, 2139, 2376, 1792, 71, 741, 1244, 2046, 2096, 1178, 1028, 558, 1716, 474, 207, 1964, 382, 1908, 793, 2249, 581, 826, 99, 2326, 2267, 1313, 285, 1875, 680, 1374, 1287, 708, 627, 1224, 552, 1057]. linear-Leontief [2046]. Linear-size [2003]. Linear-time [269, 1908, 1313, 1287]. linearization [275]. lines [1112, 2239]. link [1806, 1211, 1969]. linking [914]. links [1042]. lion [1619]. Lipschitz [1256]. liquid [1998]. List [437, 944, 440, 364, 2237, 1642, 1175, 1785]. lists [619, 758, 540]. lit [1233]. lit-only [1233]. literals [2172]. lived [1265]. liveness [1359]. LMNtal [2328]. Load [782, 952]. loading [1602]. Local [1606, 2026, 1199, 2327, 611, 685, 2312, 2148, 645, 872, 851, 1916, 1012, 195, 1128, 1077, 763]. localization [314]. Locally [419, 277, 1655, 2343, 1984]. Locating [464, 102]. location [836, 157, 1012, 2257, 1290, 781]. Löf [1257]. log [2003]. log-depth [2003]. logarithm [1966]. Logarithmic [2094]. Logic [677, 614, 1037, 1418, 2055, 1860, 1513, 698, 2323, 2011, 1076, 741, 274, 1031, 1506, 1238, 747, 2077, 933, 1839, 2286, 604, 990, 567, 666, 1675, 578, 290, 358, 916, 912, 1690, 826, 675, 676, 1090, 621, 284, 478, 1039, 159, 750, 304, 1386, 1328, 1041, 189, 1439, 491, 516, 680, 152, 2328, 123, 1078, 319, 920]. logical [919, 1962, 902, 1577, 1416, 1463, 1507]. logics [1236, 702, 704, 1375, 283, 645, 891, 1181, 2223, 646, 150, 620, 569, 2326, 36, 1704]. Long [1890, 1265, 1853, 1414, 427, 1101]. Long-lived [1265]. long-range [427]. long-term [1853]. longer [88]. Longest [1147, 803, 965, 180, 1539, 2379, 1809, 1982, 2154, 1969]. look [2323]. lookahead [1846, 264, 2361]. Looking [529]. Loop [1701, 711, 1213, 1202]. Losing [1164]. loss [1977]. losslessness [1001]. Low [936, 108, 1311, 439, 1041, 44, 1144]. low-contention [108]. low-discrepancy [44]. low-level [1041]. low-power [1144]. low-weight [1311]. Lower [706, 1820, 791, 1271, 1543, 2089, 1371, 1189, 2177, 1907, 75, 1817, 779, 710, 214, 428]. lowest [1179]. LP [598, 779]. LP-based [779]. LP-relaxation [598]. LPT [462]. LTL [2192, 743, 329, 2280, 39]. Lucas [364]. Ludics [486]. Lungo [363]. Lyapunov [87]. Lyndon [889, 1905, 2235, 1206].

M. [1584]. M2M [798]. Machine [212, 1858, 2350, 805, 324, 217, 833, 1225, 574, 1694, 793, 2249, 1629, 2153, 942, 624, 1496, 1587, 2361, 390, 1054, 1488, 1567, 838, 1975, 2138, 1011, 1045, 1503, 1943]. Machine-based [212]. machines [278, 455, 1777, 937, 1069, 154, 1705, 2350, 1888, 732, 1182, 324, 1084, 217, 14, 254, 2266, 690, 2133, 2093, 2258, 159, 797, 1970, 539, 687, 1677, 1633, 1108, 2110, 77, 1520, 1944, 2234, 967, 1864]. Magnus [1490]. Main [1816]. Main-memory [1816]. maintaining [2214]. Majority [123, 1130]. Majority-based [123]. makespan [1645, 1058, 1225, 2153, 2093, 1488, 1567]. Making [670, 1478, 1149]. malicious [1953]. malleable [778]. man [446, 1619]. management [1641, 1999, 2140]. Manhattan [1426]. manipulating [2283]. manufacturing [83]. Many [626, 1031, 1651]. many-sorted [1031]. Map [1305, 1893]. mappings [231]. maps [956, 1851, 1736]. Marcus [893]. margin [474]. marked [2342, 1201]. Market [2046, 292, 1140, 1138]. markets [1793, 832]. Markoff [2167]. Markov [1435, 115, 327, 1331, 2030, 1536, 767, 98]. Markovian [87, 360]. Marriage [1219]. Martin [1257]. Martin-Löf [1257]. Martingale [1623]. masked [1670, 966]. Masking [2291]. mass [988]. massive [1804]. Master [41, 760, 494, 1837]. matchgate [1281]. Matching [1827, 925, 1542, 2378, 2161, 1896, 1849, 1228, 175, 388, 73, 1412, 79, 2377, 548, 1855, 2021, 130, 1398, 550, 728, 2141, 1532, 1829, 2307, 131]. matchings [1626, 1887]. math [2303]. Mathematical [2077, 287]. Mathematics [479]. Matrices [1628, 1430, 1009, 1868, 1764, 1766, 1801, 977, 1769, 87, 647, 2240, 1105, 580, 1799, 1772]. Matrix [2100, 1217, 1442, 2329, 437, 2275, 908, 625, 231, 2307]. matroid [1978, 2306, 1044]. matroids [2308, 525, 1563]. matters [258]. Maude [1927]. max [555, 2178, 1545, 462, 1819, 598, 1121, 2109]. MAX-BISECTION [1121]. MAX-CUT [1121]. max-flow [555]. MAX-MQ [2109]. Maximal [2056, 707, 2292, 1815, 469, 1825, 2297, 204, 985, 2021, 1398, 2180, 1803, 812, 531, 790, 1478]. Maxime [2287]. maximize [1673, 2361]. Maximizing [473, 450]. Maximum [2307, 2104, 1053, 1116, 794, 2210, 2312, 796, 133, 2329, 2000, 1309, 1225, 2356, 1610, 2346, 1809, 1398, 1050, 1301, 1411, 979, 2373]. maximum-sum [1809]. may [186, 1957]. Mazurkiewicz [645, 655]. MCA [1132]. McCarthy [35]. MCS [2086]. Meadows [2015]. mean [1939]. meaning [584]. means [3, 402, 1769]. measure [8, 1402, 2228]. measures [1236, 1369, 1618]. Measuring [1576, 139, 223, 1619, 1547]. mechanical [2239]. mechanism [2045, 920]. mechanisms [2211, 1645, 2209, 1981, 2128, 2048, 2050, 31, 2102]. mechanized [357]. media [1602]. media-on-demand [1602]. medial [1767]. Median [781, 1979]. medians [1051]. Meet [2284, 1802, 1611]. Meets [286]. Mem [1714]. membership [1009, 1252, 1031, 941]. Membrane [1712, 1938, 1018, 1711, 1037, 1710, 1713, 1023, 991, 118, 1719, 1026]. membranes [1712, 1710]. memetic [2132]. Memoriam [1759, 2303, 363]. memory [50, 1952, 2305, 2302, 1189, 1055, 1816, 159]. Memoryless [1676, 710]. Menger [218]. merge [33]. Mergible [7]. merging [2245]. Merkle [1008]. mesh [1266, 1013]. mesh-based [1266]. meshes [105, 1653, 1302]. Message [698, 2276]. Message-passing [698]. messages [2149]. meta [902, 1040]. meta-logical [902]. meta-theory [1040]. metabelian [16]. metabolic [1022, 1721]. Metaphors [582]. method [197, 255, 961, 1246, 1637, 714, 1132, 1879, 816, 1880]. methods [2054, 1923, 212, 1753, 1932]. Metric [1741, 1652, 2119, 2346, 550, 935, 1186, 2399]. metrical [258]. metrics [875]. Michaux [208]. microscope [977]. midbit [472]. middle [479]. Milner [604]. min [2101, 555, 2178, 1545, 462, 1819, 85]. min-cut [555]. min/max [2178]. minded [182]. miniaturized [521]. minima [2054]. Minimal [364, 1770, 795, 1910, 2109, 2366, 1763, 823, 1134, 747, 933, 721, 970, 1908, 2177, 1914, 2180, 2292, 2126]. Minimality [1518]. minimax [453]. Minimization [1301, 2162, 457, 56, 2229, 81, 2124, 777]. minimize [805, 1058, 1225, 2349, 2093, 2258, 1488, 1567]. minimizes [765]. Minimizing [1664, 1665, 1262, 2153, 624, 390, 1408]. Minimum [1889, 2314, 1012, 772, 1233, 531, 2311, 1307, 2308, 175, 1543, 1426, 249, 1767, 1968, 423, 1321, 1303, 1980, 62, 1152, 1128, 1221, 1226, 557, 818, 1906, 1604, 1057, 962]. Minimum-cost [1889]. minimum-degree [2308]. Minimum-Energy [1604]. minimum-entropy [423]. mining [790]. minor [1053]. minority [2333]. Mirror [1397]. mismatch [728]. mismatches [2222, 1315]. mitotic [2085]. mix [424]. Mixed [70, 1483, 2264, 1114, 2074, 143, 1732]. mixed-signal [1732]. Mixing [1236, 1177]. mixture [1342]. mixtures [2062]. Mobile [2016, 2395, 1671, 665, 1949, 1238, 1551, 485, 2023, 1667, 157, 1424, 512, 1596, 903, 60, 1299, 1487]. mobility [1417]. Möbius [725, 1653]. modal [2272, 283, 1162, 739, 990, 514, 1439, 920]. modalities [990]. mode [391]. Model [753, 1238, 703, 333, 1593, 356, 447, 2227, 2215, 2192, 943, 1331, 1167, 1231, 380, 1977, 1163, 355, 1127, 1522, 1524, 2034, 1804, 2194, 1461, 255, 926, 914, 420, 1924, 320, 1717, 1510, 1373, 731, 2285, 488, 290, 1462, 1726, 1730, 863, 2290, 1140, 858, 1837, 572, 313, 287, 1778, 1267, 1958, 2280, 1927, 1237, 940, 306, 1892, 122, 2008, 745, 497, 452]. model-based [255]. modeled [264]. Modeling [2193, 1738, 1098, 132, 1526, 1464, 1836, 846, 1518, 1927, 1471]. Modelling [1838, 1550, 276, 513, 2189, 2188, 1351]. models [1435, 615, 1291, 1652, 577, 1378, 34, 835, 1279, 1997, 164, 74, 2274, 2070, 1718, 2030, 1736, 55, 1694, 1926, 2027, 1739, 1422, 1155, 449, 1743, 854, 1844, 876, 1093, 160, 1440, 839, 1943, 563]. modern [910]. modification [2340]. modified [24, 2349]. Modular [592, 1375, 190, 156, 1506, 1381]. modules [1083, 867, 712]. moduli [1218]. modulus [908]. Mogens [2270]. molecular [1026]. molecules [369]. moles [783]. moment [1763]. momentary [259]. monadic [278, 1506, 1068, 853, 491, 273, 1697]. monads [742]. monitoring [1666]. monochromatic [1303]. monoid [801]. monoids [201, 1630, 799, 16, 1901, 535, 806]. monomials [1110]. Monotone [373, 471, 1777, 1854]. monotonic [5, 2175, 1908, 1273]. Monotonicity [1615, 1822, 929, 893]. Moore [2301, 2231]. morphic [1996]. Morphically [2098]. morphism [1911]. morphisms [1628, 2165, 1447, 1089, 209, 1007, 1206, 16]. morphological [1595]. Morse [90, 1413]. most [2170, 1558]. most-request-first [1558]. Mostowksi [2012]. motes [1953]. Motif [1425, 1845, 2291]. motifs [2164, 1533, 2297, 2180, 2292]. motion [1516, 1937]. Move [1487]. Move-optimal [1487]. movement [2395]. Moves [2108]. Moving [528, 1469, 1666]. MQ [2109]. MSC [48]. MSO [525, 2267]. MSOL [1119]. MST [2308, 1295]. MTLA [512]. mu [283]. mu-calculus [283]. Muller [72]. Multi [1535, 805, 1464, 2069, 1102, 547, 414, 389, 510, 2259, 912, 1690, 1437, 727, 312, 2181, 378, 1922, 2086]. multi- [547]. Multi-agent [805]. multi-armed [2069]. Multi-break [1535]. multi-cellular [1464]. multi-covering [2259]. multi-dimensional [378]. multi-hop [312]. multi-layered [414]. multi-letter [2181]. multi-level [510]. Multi-MCS [2086]. multi-multiway [1102]. multi-priced [1437]. multi-qudit [912, 1690]. multi-regular [727]. multi-threaded [1922]. multicast [946, 2387, 1641, 798, 1428]. multicolored [1303]. multicoloring [772]. Multicommodity [1173, 370, 555]. multicounter [1888, 14]. multicover [1484]. multicuts [955]. multidimensional [1197, 368, 1171, 936]. multihead [1888, 650]. multihop [1959]. multilayer [992]. Multimodal [750, 1379, 1706]. multimorphisms [1640]. Multiobjective [693]. multiparty [2059]. multiperiodic [1358]. Multiple [2062, 1443, 1312, 716, 1168, 1305, 926, 888, 1912, 1726, 1346, 317, 2188]. multiple-interval [1912]. multiples [263]. multiplication [553, 791, 1127, 2307]. multiplicative [1560, 680]. multiplicity [69]. multiplier [1543, 1473]. multiprocessor [734, 167, 2335]. multiprocessors [1373]. multirate [1014]. multirecombination [764]. Multiset [1020, 1461, 1625]. multiset-based [1461]. multisets [619, 758]. multitape [2168, 830]. multithreaded [54, 190, 917]. Multitree [100]. multivalued [564]. multivariate [1869]. multiway [1102]. mutable [867]. Mutation [320]. mutations [765]. mutual [876]. mutually [2283].

Nadia [1930]. Name [666, 665, 296]. Name-passing [666, 665]. nano [1836]. nanorobotic [2031]. Nanowire [1857]. Narayana [377]. narrations [1420]. narrowing [2320, 885]. Nash [2044, 1649, 2047, 1483, 2208, 293, 2345, 2261, 2049]. national [268]. natural [987, 1041, 835]. naturally [1740]. navigability [2080]. Navigable [2344]. navigation [1550]. NC [340]. NCA [539]. near [1472, 1643]. near-singular [1472]. nearest [2137, 425, 2061, 1877, 2006]. nearest-neighbor [425]. nearly [1915, 1411, 971, 2006]. nearly-Euclidean [2006]. nearness [1734]. Necessary [2359, 401]. negation [2003]. negation-limited [2003]. negations [2268]. negative [2137]. negotiation [1552]. Nei [961]. neighbor [2083, 425, 2061, 2006]. neighborhood [1764]. neighborhoods [1804]. Nerode [2222]. nested [94, 2176, 2206, 1002]. net [1153, 712]. nets [1435, 662, 196, 1329, 850, 663, 324, 744, 2285, 1511, 2013, 2277, 1696]. network [1167, 1603, 711, 1784, 610, 2080, 2345, 1664, 1676, 814, 1650, 1611, 908, 1617, 1666, 1927, 962]. networks [780, 2197, 1782, 288, 607, 544, 608, 1148, 2136, 1662, 946, 1665, 1604, 20, 547, 1426, 1261, 133, 1663, 1783, 2344, 606, 558, 1841, 1950, 1014, 310, 733, 2024, 421, 311, 990, 798, 1264, 110, 1953, 157, 1561, 1213, 324, 1297, 1967, 1298, 1781, 2022, 1211, 2025, 1422, 2095, 427, 313, 2245, 1486, 1959, 1223, 1863, 1941, 317, 1266, 993, 1135, 1428, 903, 1113, 2141, 981, 1300, 609, 46, 1463, 599, 611, 314, 1969, 996, 1103, 1681, 2104, 1526]. neural [1329, 2118, 250, 990, 1024, 908, 317, 1681, 1526, 2130]. News [1, 560]. Newton [504, 714]. Newtonian [988]. NFA [402, 7, 3, 1794, 2177]. NFAs [1183, 2347]. Nielsen [2270]. NL [616]. NL-printable [616]. NLC [382, 1015]. NLC-width [382, 1015]. no [33, 391, 1224]. node [1096, 2087]. noise [1280]. noisy [1170, 2061, 1118]. nominal [1671, 1700, 2012]. Non [1551, 2085, 2324, 1285, 291, 1347, 1209, 685, 1432, 342, 2030, 574, 910, 929, 1517, 1770, 1089, 372, 909, 1615, 2236, 2268, 477, 68, 2292, 1520, 1400, 1249, 1681]. Non-asymptotic [1347]. non-autonomous [1681]. non-Boolean [68]. non-codeletable [1770]. non-controllable [574]. Non-cooperative [291]. non-deletable [1770]. non-deterministic [1615, 2268, 1520]. non-erasing [1089]. non-gapped [2292]. non-holonomic [2236]. non-interactive [1249]. non-learnable [477]. non-local [685]. Non-mitotic [2085]. non-monotonicity [929]. non-parametrizability [342]. non-quantum [910]. Non-reducible [1285]. non-saturated [2030]. non-standard [1400]. Non-strict [2324]. Non-uniform [1551]. Non-unique [1209, 1432]. non-uniqueness [372]. non-unitary [1517]. non-zero-approaching [909]. nonblocking [547]. nonclairvoyant [802]. nondeterminism [318, 1091, 2220, 2228]. Nondeterministic [2256, 1522, 2176, 616, 2271, 639, 9, 290, 621, 30]. nonidentical [2153]. noninvertibility [1639]. noninvertible [789]. nonlinear [1907]. nonmonotonic [189]. nonnegative [87]. nonparametric [735]. Nonstochastic [1575]. nonuniform [648, 121]. norm [782, 1543]. Normal [749, 1024, 583, 1904, 822, 1109, 1873, 1943]. Normalisation [96, 1590]. normalization [1589, 1592]. note [2157, 1825, 2362, 89, 2047, 718, 1203, 2153, 2027, 959, 2133, 981, 2397, 995, 398, 2008]. notions [842]. NP [1406, 1227, 1107, 1524, 1565, 193, 1767, 1121, 2269, 969, 789, 2111, 1830, 1800, 984, 678, 813, 779]. NP-complete [1524, 2111, 1800, 678, 813]. NP-completeness [1227, 1830]. NP-hard [1406, 1767, 1121, 2269, 779]. NP-hardness [1565]. NP-pairs [969]. NRL [902]. nucleic [1932]. number [1947, 1441, 2163, 19, 1974, 805, 205, 1354, 1791, 1502, 2178, 1203, 1405, 1220, 773, 2361, 28, 1157, 1301, 557, 1995, 1233, 1400, 1887, 82]. numbers [1435, 1109, 2057, 417, 2019, 2167, 169, 499, 436, 1870, 1658, 2250, 1627, 257, 953, 377, 498, 500]. numeration [716]. numerical [1839, 1880, 324]. numerous [1156].

O [1845, 2193]. obfuscation [1569]. object [1760, 1923, 301, 866, 863, 304, 1922]. object-oriented [1923, 863, 304, 1922]. objectives [1380]. objects [186, 1666, 59]. obligatory [2355]. Oblivious [250, 1551, 1814, 1424, 105, 1678]. obliviousness [107]. observability [486]. Observational [142, 1099]. Observations [829]. obstacles [940]. obstructions [1772]. Obtaining [972]. Occurrences [533, 2170, 178, 800]. occurring [557].

Ockham [1278]. octrees [810]. off [1008, 430]. Offline [1619, 1499]. offspring [1277]. offspring-producing [1277]. OJAn [1132]. Oja’s [909]. Old [1711]. Olympus [1130]. on-demand [1643]. On-line [66, 1803, 1045, 1013, 1846, 219, 1824, 1503, 839, 1757]. On-the-fly [2285, 329]. once [791, 706]. One [336, 967, 2137, 201, 245, 2037, 50, 2170, 9, 1888, 320, 1639, 254, 2129, 1133, 347, 1684, 81, 2050, 1115]. one-dimensional [320, 1115]. one-letter [245, 2037]. one-pebble [254]. one-reversal [1888]. One-shot [2137]. one-sided [81]. one-way [1639, 2129, 1133, 347, 1684]. ones [1307, 2082]. Online [784, 1795, 1787, 1058, 73, 1644, 2266, 2361, 2349, 2093, 2258, 1431, 2234, 1497, 2211, 811, 1484, 1231, 1294, 2350, 217, 802, 1, 560, 815, 1404, 1970, 1011, 1108, 2110]. only [1233]. onto [180, 535]. Open [780, 1340, 1419, 1812, 188, 2123]. open-shop [1812]. opened [97]. Operational [1153, 2206, 1037, 604, 511, 1385]. Operations [603, 2241, 245, 1449, 1126, 2171, 1355, 2204, 2205, 2135, 752, 25, 994, 1942, 2125, 1272]. operators [569, 514, 806]. Opponent [452]. Opponent-Model [452]. opportunities [1793]. optical [780, 218, 109, 122]. Optimal [1602, 1331, 1008, 2395, 2024, 2305, 1371, 1213, 217, 1517, 2214, 1437, 2061, 1004, 105, 1011, 1302, 2140, 1168, 2256, 938, 2114, 928, 2254, 427, 1999, 728, 1487, 1643, 1773, 1556, 431]. Optimality [1374, 581]. Optimally [2237, 727, 816]. Optimization [1860, 1271, 597, 1640, 322, 1498, 1363, 1169, 2306]. Optimizing [1641]. optimum [1967]. oracle [633]. Oracles [252, 1129, 1344, 2354]. orbits [2160, 2103]. Orc [1679]. order [94, 96, 741, 665, 2281, 1506, 1470, 172, 620, 1071, 1554, 328, 1072, 296, 1438, 2280, 2014, 39, 294]. Ordered [476, 541, 463, 1976, 2154]. Ordering [189, 1549, 1404]. orderings [1244, 1618, 252]. orders [1106, 872, 426, 138, 2267, 2226]. ordinal [1518, 842]. ordinary [1703]. organizing [1951]. orientation [1059]. oriented [1236, 1923, 1924, 1694, 863, 304, 1922]. orthogonal [2309, 376]. other [1645, 290, 692, 1247]. out-branching [2314]. outline [304]. outputs [2245]. Overlap [1986, 2240]. Overlap-free [2240]. Overlap-freeness [1986]. overlay [607]. overloaded [1895].

P [989, 340, 1019, 2118, 1134, 1021, 1716, 1022, 23, 11, 789, 318, 118, 692, 1024, 994, 1721, 1623, 120, 1992, 1723, 2130]. PA [1511]. PAC [1284, 1343]. PAC-learnability [1343]. packed [1180]. packet [1059, 1895]. packing [2252, 2353, 1481, 794, 1899, 1883, 1114, 1795, 2309, 1431, 1500]. PAFAS [139]. page [1550]. Paging [1104, 2242]. pair [1640, 1829]. paired [2352, 2351]. paired-domination [2352, 2351]. pairing [132]. Pairs [1501, 1915, 969, 1553]. Pairwise [958]. Palindrome [1562]. palindromes [1192, 533, 2381, 1982, 2179]. Palindromic [240, 1195, 2182, 1397]. palindromization [1893]. pancake [215, 1892]. Panconnectivity [1113, 1427, 1324]. pancyclicity [1557, 1113, 1427]. paper [1529]. papers [32, 1434, 1416]. paradigmatic [1275]. Parallel [1329, 2249, 1496, 1808, 2113, 937, 1906, 1494, 2019, 2151, 695, 1354, 787, 1058, 1182, 1561, 1523, 746, 1567, 1771, 539, 1975, 2110, 321, 2234, 865]. parallel-batch [1975]. Parallel-machine [2249, 1496]. Parallelism [2068, 1134, 1023]. parallelization [2324]. Parallelizing [2131]. parameter [2313]. Parameterised [302, 1729]. Parameterized [2341, 2388, 522, 2152, 527, 526, 530, 591, 212, 521, 2086, 2142, 1912, 2379, 2306, 1856, 2391]. parameters [1331]. Parametric [1671, 1588, 1378, 1738, 478]. parametricity [1385]. parametrizability [342]. parametrization [2132]. parametrized [1377]. parentheses [928]. Pareto [1149]. Parikh [1430, 2360]. Parity [791, 1493, 706]. parses [1393]. Parsing [246, 386]. part [1161, 1423, 385, 386, 534]. Partial [1317, 459, 158, 584, 875, 955, 336, 862, 1971, 1990, 872, 1986, 328, 1409, 514, 2090, 2280, 562, 1108, 2226, 1421, 401]. partially [1630, 1888, 541, 650, 463, 1956, 1631, 1901]. participation [1673]. particle [1115]. particles [2155]. particular [2108]. partition [1978, 418, 1147, 1216, 1303, 973, 462, 1044, 962]. Partitioned [1559]. Partitioning [2332, 1044, 2023, 1568, 1119, 1131]. partitions [437, 1394]. partizan [78]. partners [1168]. parts [2332]. pass [2062]. passing [698, 665, 666]. password [551]. password-based [551]. Path [1304, 2034, 1731, 1293, 506, 1619, 106, 630, 973, 1913, 1408, 267]. path-connected [506]. pathology [453]. Paths [1897, 1563, 2081, 2358, 1915, 1449, 1648, 1262, 109, 958, 2087, 719, 524, 180, 1388, 1610, 1414, 1809, 1789, 703, 971, 2357, 1969, 1553, 1101, 2367]. pathways [1718, 1462]. Pattern [202, 647, 2179, 1829, 1256, 925, 1255, 1541, 1546, 1814, 1412, 2377, 707, 130, 727, 410, 1579, 1874, 477, 1580]. patterns [1425, 2291, 638, 845, 128, 1890, 1003, 463, 1586, 2123, 282, 129, 131, 790]. payment [1981]. PCA [909]. PCP [1201]. pebble [254]. pebbles [27]. pebbling [1175]. Pedagogical [2281]. peer [607, 1947]. peer-to-peer [607, 1947]. penalty [950, 1540]. pencils [1217]. people [1531]. PEPA [2189, 1718]. perceptron [992]. percolation [609]. perfect [2330, 175, 1887, 227]. Perfectly [2038]. Performance [47, 1236, 139, 746, 1927]. Periodic [2355, 1046, 599]. periodical [396]. Periodicity [2160, 743, 399, 1447, 1457, 2288]. periods [205, 546, 2183]. peripheral [1713]. permutation [544, 1766, 2097, 2247]. permutations [583, 1904, 1455, 1546, 424, 347, 1678, 429]. Permuting [455]. Personal [2319]. perspective [1652, 396]. PEsort [931]. Petri [1696, 712, 196, 663, 1153, 744, 2285, 1511, 2013, 2277]. Phase [1423, 694, 626]. phylogenetic [519, 133, 961, 985, 814, 813]. phylogenies [1391]. phylogeny [522]. Physical [1515, 1939]. physics [1513]. pi [1598]. pi-calculus [1598]. Pict [1921]. Picture [642, 164, 1060, 235, 894]. Pictures [2039]. pigeonhole [1270, 2156]. PIN [905]. PIOA [865]. Pipelined [1853]. Pisa [2318]. Pisot [1193, 716]. pixel [938]. plan [515]. planar [1851, 2395, 1172, 944, 2390, 1569, 1308, 1224, 1289, 2364, 1785]. planarity [634, 1525]. plane [1949, 1664, 1786, 2236, 2155, 466, 1887]. planning [1619, 1156]. plateau [2128]. platform [1464]. player [2044, 78]. player-specific [2044]. players [1118]. Playing [1903]. plurality [176]. point [1657, 2304, 1388, 1422, 490, 940, 491, 1475]. point-shaped [940]. pointer [539]. pointers [2283]. points [1847, 1762, 484, 2092, 1171, 1937, 1089, 209, 729, 716, 1504]. polarization [2326]. Polarized [300, 295]. policies [528]. policy [1142]. Poly [216]. Poly- [216]. polyforms [1773]. polygon [1400]. polygonal [1161, 1423]. polygons [1768, 600, 1388, 1287]. polygraphic [744]. Polyhedra [1692]. polyhedral [2322, 1762]. Polylogarithmic [1316]. Polylogarithmic-round [1316]. polymorphic [56, 1436]. polymorphism [1378]. Polynomial [134, 1168, 671, 2049, 2067, 2, 1123, 1522, 1700, 845, 250, 1028, 1870, 739, 934, 150, 2205, 718, 942, 1218, 475, 714, 568, 1413, 1873, 2050, 679, 1877, 1476, 476, 1880].

Polynomial-size [671]. Polynomial-time [2, 1123, 568, 2050, 679]. polynomially [332]. Polynomials [1090, 1868, 2162, 1869, 2360, 2144, 1871, 263, 1042, 777]. Polyomino [1627, 1881]. polyominoes [394, 640, 371, 376]. Population [1688]. port [1658]. portrait [1423, 2319]. poset [725, 980]. posets [720, 2284]. positional [543]. positioning [312]. positions [1562, 2263, 448]. Positive [379, 272, 1252, 1470, 412, 747, 765, 1405, 2137, 1581, 1579, 1578, 2067, 476, 844, 2, 1123, 2065]. possibility [1756, 1249]. possible [1160, 474, 1066]. possible-world [1066]. possibly [1909]. post [2175]. postponed [516]. potential [1550]. Power [726, 90, 1846, 248, 2163, 757, 685, 2052, 2068, 1865, 695, 664, 2120, 890, 1917, 1498, 663, 14, 1659, 27, 1816, 454, 491, 323, 2396, 2394, 1556, 214, 1144]. power-index [1917]. power-law [1498, 1816]. powerful [2260]. Powers [1200, 968, 2389, 1911, 2233]. PP [398]. pre [1602, 1838]. pre-buffering [1602]. pre-loading [1602]. pre-rRNA [1838]. Precedence [282, 2252, 1646]. precision [77]. preclusion [2141]. Predecessor [1327, 1529]. predicate [593, 968, 2168, 747]. Predictable [2201]. predicting [1257, 132]. Prediction [413, 1932, 1282, 1757]. Prediction-hardness [413]. predictions [839]. preemptive [217]. Preface [1843, 315, 361, 415, 507, 613, 657, 683, 696, 770, 819, 1240, 1291, 1268, 1335, 1465, 1530, 1612, 1668, 1691, 1724, 1775, 1842, 1945, 2042, 2078, 2112, 2134, 2159, 2195, 2051, 407, 668, 658, 1166, 1336, 1758, 1441, 635, 1867, 1920, 1191, 1269, 1834, 2196, 2187, 353, 1145, 808, 1279, 1997, 1241, 18, 1065, 495, 895, 1137, 1235, 416, 998, 887, 878, 1601, 2302, 1467, 93, 200, 2072, 855, 2218, 2060, 1176, 771, 2029, 2036, 1776, 4, 335, 1259, 605, 145, 1349, 596, 1751, 508, 1725, 2207, 831, 1669, 1036, 308, 1661, 42, 1292, 104, 362, 2017, 1459, 113, 316, 684]. Preface [907, 1125, 1326, 1521, 1680, 1708, 1833, 1928, 2028, 2035, 2071, 2127, 2186, 924, 482, 2043, 762, 2383, 53, 299, 861, 229, 848, 1946, 588]. preferential [563]. Prefix [1412, 1318, 2213, 654]. Prefix-free [1412]. prefix-freeness [1318]. prefixing [662]. preliminary [67]. Presburger [1902]. Prescribed [2064]. presence [918, 288, 1280, 2305, 619, 758, 2188]. presentations [672, 744]. preserving [1628, 821, 603, 2171, 83, 2226, 1297]. presheaves [2273]. price [934, 1967]. priced [1647, 1437]. pricing [1138]. primal [1139]. primal-dual [1139]. Prime [822, 2278, 970, 1088]. primes [1693]. primitive [236, 649, 263, 2098, 844]. primitivity [2171]. principle [2274, 1907, 2156]. principled [160]. principles [1363, 1270, 1361, 1721]. printable [616]. priorities [23]. priority [455, 595, 542]. priors [840]. prisms [1047]. Prisoner [1118]. Private [2059]. Probabilistic [2271, 268, 283, 1462, 357, 1555, 2253, 699, 1435, 661, 338, 196, 1158, 1421, 1728, 1032, 2274, 2076, 358, 1577, 571, 373, 568, 468, 1538, 1343, 223, 1093, 49, 754, 2016]. probabilities [327, 360, 728]. Probability [623, 1171, 2075]. Probable [900]. Probe [2330, 1559, 1432, 1174, 1185, 1209]. probes [606]. problem [201, 176, 974, 288, 2311, 1978, 1102, 954, 134, 1319, 794, 979, 804, 1307, 1494, 2387, 1825, 1979, 978, 2308, 2362, 268, 1216, 2304, 1974, 2086, 1779, 1295, 1541, 401, 737, 715, 1619, 1275, 733, 947, 111, 438, 949, 977, 370, 2175, 423, 564, 1989, 13, 1539, 1012, 1903, 630, 1392, 165, 942, 1184, 1610, 1448, 1807, 1321, 693, 1980, 1111, 1152, 1545, 328, 2040, 2382, 1786, 2390, 537, 1568, 1491, 1128, 965, 2082, 539, 2397, 1050, 1991, 2373, 1808, 1301, 1226, 2138, 1131, 1287, 1130, 710, 267, 1500, 1290, 2007, 1973, 367, 1411, 401, 1401, 1120, 962, 2397]. problems [1406, 1846, 1053, 2340, 1222, 1227, 1327, 1529, 2100, 1442, 1148, 1813, 1484, 2388, 171, 341, 1649, 1228, 1645, 1307, 1524, 8, 1713, 1979, 960, 521, 835, 2352, 796, 597, 2329, 1640, 467, 1114, 1619, 419, 1275, 173, 420, 1912, 1483, 1219, 549, 836, 1646, 2314, 2309, 833, 290, 269, 2257, 801, 1855, 1171, 1403, 2053, 1831, 1303, 2123, 62, 2152, 527, 526, 2306, 15, 2090, 1970, 454, 1092, 1248, 530, 1119, 68, 2357, 1131, 1108, 2139, 941, 779, 1384, 1821, 462, 1044, 536, 1289, 1822, 2104, 1212]. procedure [193]. procedures [590, 2283]. Process [140, 1861, 141, 137, 662, 300, 753, 1460, 2194, 1275, 509, 360, 1536, 568, 579, 1384]. Processes [757, 1719, 327, 1419, 1030, 485, 746, 1029, 1754, 38, 1957, 667, 39, 1277, 98]. processing [1001, 167, 1058, 1189, 1644, 1694, 2153, 1496, 1054, 1079]. processing-oriented [1694]. processor [111]. processors [2197, 1497, 20, 1941, 993, 1135]. producing [1277]. product [957, 641, 835, 87, 1635, 1965]. production [2041, 1140]. products [709, 641, 1962, 263]. Prof. [2303, 2287]. Professor [656, 1531]. profile [1532]. program [1418, 740, 1028, 2324, 1675, 1175, 879, 1707, 2014]. programmability [1689]. programmable [2031]. programming [2321, 1240, 2039, 2323, 1235, 1924, 1925, 858, 284, 449, 750, 304, 68, 2328]. programs [791, 1158, 2279, 1382, 1841, 1334, 190, 2144, 706, 1071, 1894, 1800, 354, 756, 296, 189, 121, 1922]. Progresses [2333]. progressions [1453]. Progressive [787]. project [1039]. projection [582, 2147]. projections [1765, 927, 376]. Projective [870]. prologue [562]. prolongation [2147]. promoters [1719]. Proof [856, 1334, 54, 1693, 849, 1495, 2256, 741, 259, 1061, 2089, 2282, 889, 1692, 739, 480, 969, 666, 2111, 206, 304, 1677]. Proof-carrying [856]. proofs [671, 1097, 675, 676, 1316, 1288, 565, 1590]. properties [2241, 146, 753, 554, 2279, 164, 1471, 1162, 939, 1552, 902, 851, 2103, 1447, 254, 119, 2121, 581, 2284, 1155, 1359, 2372, 538, 333]. Property [1542, 83, 1256, 1585, 339, 2166, 280, 234, 1139, 1732, 736]. Property-preserving [83]. proposed [551]. propositional [1860, 849, 933, 618, 969, 920]. Protean [2393]. proteins [1712, 2194, 1713, 1938, 1720, 132]. Protocol [902, 1420, 193, 1783, 551, 1144]. protocols [896, 897, 1662, 95, 197, 899, 1421, 50, 1952, 568, 904, 314]. provably [1796]. provably-secure [1796]. prove [506]. prover [331]. proving [2177]. pruning [1885]. pruning-grafting [1885]. pseudo [1974, 1330]. pseudo-achromatic [1974]. pseudo-boolean [1330]. pseudoknots [1525]. pseudometric [49]. pseudometrics [1740]. Pseudopalindrome [806]. PSPACE [2367, 290, 1519]. PSPACE-completeness [2367, 290]. PTAS [216, 2151, 2385, 777]. PTAS-completeness [216]. public [910]. Pure [2044, 37, 539]. purity [1099]. push [2308]. Pushdown [1215, 1163, 1091, 2220, 1328]. puzzles [290]. pyramids [1772]. python [502].

Q [397, 63]. Q-convex [397, 63]. QBDs [1237]. quadratic [1963, 2362, 765, 1991, 1887]. Quadrature [503]. qualitative [1859]. quality [992]. quantified [1815]. Quantifier [968, 621, 904]. Quantifier-free [621]. quantifiers [704, 187]. Quantitative [355, 1235, 359, 915, 1240]. quantities [800]. Quantization [686]. Quantum [1933, 1574, 1358, 121, 325, 124, 1129, 2073, 338, 244, 637, 2131, 2077, 1334, 1689, 2074, 910, 2129, 1133, 347, 1517, 912, 1690, 1184, 1025, 690, 1684, 2133, 686, 84, 1328, 2181, 2374, 2075, 319]. quarter [2236]. quartic [1765]. Quasi [2096, 1549, 1106, 2360, 188, 426, 875, 1740, 1211]. Quasi-linear [2096]. quasi-metrics [875]. quasi-open [188]. quasi-ordering [1549]. quasi-orders [1106, 426]. quasi-polynomials [2360]. quasi-pseudometrics [1740]. Quasi-Static [1211]. Quasiperiodic [1905, 1007]. quasiseparable [1868]. qudit [912, 1690]. queries [918, 1252, 147, 1805, 148, 1003, 620, 413, 1345, 410, 981, 67, 1913, 2359]. Query [1273, 1388, 1000, 336, 1001, 149, 150, 1189, 1184, 1504]. Query-monotonic [1273]. Query-point [1388]. querying [1950]. question [1884, 954]. question/answer [1884]. questions [1624]. queued [1895]. queueing [1231]. queues [455]. queuing [219, 1960]. Quicksort [2108]. quiescent [786]. quilted [2038]. Quorum [989]. quotient [1746]. quotient-universal [1746].

r.e [2064]. Rabin [257]. radiian [1821]. radio [288, 1261, 1663, 558, 798, 1264, 1781, 107, 1103]. radius [961, 1914, 1821]. Rainer [632]. Rambo [1265]. Random [1318, 397, 21, 1791, 1177, 1947, 1214, 1172, 1977, 1020, 1295, 2344, 1501, 436, 1245, 2150, 2369, 1961, 1239, 1257, 1916, 2354, 1985, 927, 373, 1810, 1188, 1780, 2372, 983, 1332, 2255]. randomised [2136]. randomization [107]. Randomized [1111, 1128, 1484, 458, 1783, 2216, 14, 2356, 1311, 1647, 143, 1857, 65, 1332, 2063]. randomized-contact [1857]. randomness [253]. Range [558, 288, 1585, 427, 67, 314, 332, 2005, 2102]. range-efficient [2005]. range-free [314]. Rank [2299, 1370, 2366, 1792, 1244, 2089, 737, 2275, 1371, 2298, 1398, 2156]. rank-maximal [1398]. Rank/select [2299, 2298]. Ranking [1116, 2393]. rankings [2313]. Ranks [1059]. Raphson [504]. Rapid [2025]. rate [1460, 909]. rates [461, 2030]. ratio [1087, 1548, 555, 462, 2109]. Rational [1901, 1630, 1244, 1087, 1230, 69, 1876, 953, 2125, 1678]. Rationality [725, 1138]. Rauzy [2157, 378]. rays [63]. razor [1278]. RBIBDs [816]. rCOS [866]. re [1629, 307]. re-entrant [1629]. re-use [307]. reachability [327, 1552, 1437, 885, 1338, 1161]. reachable [256]. reaction [1083, 1934, 102, 1681]. reaction-diffusion [1681]. Reactive [2273, 300, 1575, 1756, 712]. Reactors [1924]. read [791, 706]. Read&Modify&Write [108]. read-once [791, 706]. ready [1029]. ready-tree [1029]. Real [1870, 2057, 417, 502, 499, 1097, 1373, 1220, 2147, 354, 908, 1157, 1958, 44, 1878, 1477, 498, 500, 1927]. real-number [1157]. real-time [1097, 1373, 354, 1958, 1927]. Realizability [48]. realizes [711]. RealLife [1010]. reals [968, 1318, 2366]. Rearrangeability [544]. rearrangeable [1014]. rearrangement [2296, 1977, 2290, 1534, 1312]. rearrangements [1535]. reasonable [1587]. Reasoning [1158, 933, 57, 688, 1077, 411, 152, 920, 867]. reciprocal [1874]. recognisable [801]. recognition [2091, 1849, 1874, 1451, 386]. recognizability [280, 238, 1164]. recognizable [2227, 2076, 1060]. recognizing [244]. recombinant [1152]. recombination [22, 320]. Reconciling [380]. reconfigurable [142]. reconstructed [369]. Reconstruction [1764, 1765, 376, 1625, 2255, 871, 1766, 640, 961, 522, 1774]. records [619, 758]. recover [1957]. rectangles [1752]. Rectangular [1881, 2038, 1769]. rectilinear [600]. recurrence [2233]. recurrences [532]. recurrent [1195, 317]. recursion [1038]. recursive [592, 417, 1705, 147, 324, 1573, 409, 1581, 879, 1707, 1427, 2283, 2084, 1572, 1516]. Recursively [1186, 2399]. Rédei [736]. reduced [1941]. reducibilities [336, 1745]. Reducibility [638, 351]. reducible [1285]. Reducing [1398]. Reduction [370, 3, 402, 2282, 509, 1942, 830, 405]. Reductions [1854, 629, 1273, 83]. Reeb [1468]. Reed [1170]. Rees [1469, 625]. reference [859, 2008]. Referential [676]. Refined [1122]. Refinement [586, 862, 866, 512, 517, 1383, 161, 59, 1374]. refinements [1926]. reflected [675]. Reflection [1031]. reflections [1361]. Reflective [1573]. reflexive [116]. reformulation [148]. refutations [523]. regarding [954, 1624]. regex [2116]. regimes [1837]. region [2395]. regions [1471]. register [1800]. regression [1728]. Regular [6, 241, 2041, 1448, 426, 2221, 245, 1727, 603, 244, 1896, 1649, 116, 753, 1565, 2116, 845, 3, 402, 1353, 1470, 2342, 1395, 1245, 2369, 238, 1373, 2205, 972, 1412, 2135, 25, 727, 1624, 1538, 1450, 1359, 1913, 1289, 1056]. regularity [2094, 38, 195]. regularized [1255]. Regulated [1351, 2220, 31]. regulation [2192]. regulatory [1463]. rejection [1481, 2151, 1567]. relabel [2308]. related [1777, 2388, 171, 1649, 1307, 2167, 8, 960, 2329, 2314, 392, 1392, 1247, 309]. Relating [1718]. Relation [2233, 2011, 1638, 788, 411]. Relational [1407, 1078, 1736, 564, 282, 1002]. relations [1367, 1489, 1087, 1230, 749, 1068, 1788, 194, 281, 392, 571, 2239, 1507]. relationship [1001, 978, 382, 779]. relative [2242]. relator [201]. relaxation [598]. relaxations [1871]. relays [2214]. release [2252, 1812, 390, 1488, 1567, 1431]. reliability [110, 26]. reliable [218]. remarks [2173]. remote [947]. renaming [2020]. rendezvous [1605, 628, 2099]. renewal [2146]. Reoptimization [2217, 2106]. repairman [1, 560]. repeats [803, 1803]. repetition [237, 1824, 346, 1205]. repetition-free [237, 1205]. Repetitions [2368, 2160, 204, 1453]. reporter [2033]. reporting [469, 1825]. representable [1107]. Representation [639, 1350, 35, 1735, 406, 928, 689]. representations [1851, 26, 2027, 1033, 2180, 766, 2292]. representatives [999]. represented [1476]. representing [474, 2225, 990]. request [1558, 1499]. required [2103]. requirements [1012]. residents [1626]. resilient [2305]. resolution [671, 1565, 2325, 523, 2111, 1288, 1347]. resolution-based [2325]. resolved [2052]. Resource [701, 348, 383, 82, 950, 794, 1672, 734, 549, 1402, 265]. resource-bounded [1402]. Resource-sharing [82]. Resources [1077, 1418, 665, 83]. respect [1445]. Responsiveness [1861]. Restarting [825, 929, 893]. restarts [1058]. restricted [1015, 706, 1894, 825, 703, 904, 68]. restriction [742]. restrictions [1798]. result [921, 321]. Results [263, 2113, 1227, 115, 456, 1163, 1804, 2351, 1779, 1852, 1298, 994, 2053, 993, 1135, 767, 67, 1384]. retrospective [2323]. Reversal [1051, 1654, 9, 1888, 360]. reversal-bounded [9]. Reversals [1192, 2213]. reverse [391]. Reversibilities [1328]. Reversible [1528, 404, 2343, 2076, 1901, 123]. review [237]. Revising [1253]. Revision [622, 515]. revisited [2320, 1000, 1921, 1317, 1850, 1340, 435, 976, 900, 1074, 1654, 348, 1452, 1370, 1747, 1590]. rewards [1236]. rewrite [753, 755, 2107]. Rewriting [918, 1003, 120, 1037, 1351, 1001, 642, 1031, 1106, 902, 1039, 1701, 1578, 235, 31]. rewriting-based [902]. rewritten [1686]. Reynolds [1080, 1075]. Rho [2194]. Rich [1826, 2166, 1284]. richness [2182]. right [1122, 251, 982, 627]. right-arm [1122, 627]. ring [1667, 1396, 1424, 1956]. rings [1474, 667, 1301]. Rivas [1525]. Rivest [387]. RMI [1417]. RNA [1351, 926, 1525, 130, 1410, 131]. robber [1616]. Roberts [960]. robot [1606]. robotic [2362, 1991]. robots [1949, 1551, 2023, 168, 1814, 1424, 1299]. Robust [1947, 1955, 280, 384, 2377, 265]. Robustness [2286, 1461, 2216, 549]. Rocca [1584]. Role [1597, 438, 1616]. Role-based [1597]. rollback [1175]. Ronchi [1584]. Roommates [1219]. root [953, 1878, 1477]. rooted [814]. roots [2121, 813]. rotation [1122, 2377]. rotations [925, 1320, 1196, 388, 627]. round [1316]. rounding [1426]. roundings [44]. route [1287]. routing [1846, 218, 1948, 1231, 2210, 2387, 547, 1142, 1483, 2212, 109, 2208, 110, 1213, 106, 1059, 427, 1778, 105, 309, 1408]. routings [1428]. row [1310]. rows [394]. rRNA [1838]. RRWW [2228]. RRWW-automata [2228]. RSA [1396]. rule [1199, 2283, 1819]. rules [1334, 15, 745]. run [1540, 2298, 2263]. run-length [1540, 2298, 2263]. runs [373, 1651]. runtime [155, 2216, 1688].

S [2323]. S-semantics [2323]. S. [1584]. S4 [1379, 1706]. SA [115]. Safe [1957, 1619, 863]. safely [1070]. safety [1392]. SAIL [1531]. Saitou [961]. same [1493]. sample [931]. Sampling [2054, 1791, 2059, 408]. Sand [2265, 1208]. sandpiles [1208, 963]. sandwich [1212]. SAT [598, 1501, 991, 178, 1730, 1025, 1810, 1732]. SAT-assignments [1501]. SAT-based [1730, 1732]. Satgraphs [534]. satisfaction [419, 2257, 2053, 431]. satisfiability [1509, 2172, 1810, 68, 428, 86]. saturated [2273, 2030]. say [1939]. Scala [1925]. Scalable [609]. scalar [1009]. Scale [310, 2316]. Scale-free [310]. scaled [436]. scaling [2315, 2377]. scan [2104]. Scanning [1769]. scattered [1244, 1918, 1393, 2079]. scenarios [731, 2013]. Schedulability [595]. scheduler [1643]. schedules [1812, 802]. Scheduling [594, 950, 1673, 793, 2335, 2334, 1054, 780, 1777, 937, 1497, 2211, 1645, 2151, 734, 2362, 167, 805, 865, 778, 732, 549, 1058, 2145, 1182, 1646, 833, 1225, 1558, 1644, 2048, 1555, 2249, 1496, 2266, 2361, 2349, 2093, 2258, 1488, 1567, 528, 1491, 1970, 1975, 1991, 2355, 1895, 2138, 1108, 1045, 1503, 2110, 779, 1013, 82, 2102, 2234, 542]. schemas [1028]. scheme [2353, 1481, 1016, 942, 1311, 1566, 1966]. schemes [94, 938, 1796, 715, 1552, 732, 1981, 2080, 345, 1820, 106, 2393, 1150, 390, 879, 1707, 1248, 1408, 536, 1811]. Schröder [377]. Schur [1875]. Schützenberger [116, 889, 1659]. Sci [1529, 2398, 402, 1690, 560, 1135, 1707, 2396, 1123, 2399, 922]. Science [286, 1434, 1745]. sciences [987]. scientific [632]. score [448]. sealing [1073]. search [1000, 1331, 741, 1793, 452, 733, 204, 1676, 1055, 210, 425, 982, 1499, 1151, 1128, 453, 629, 609, 1332, 1131, 766, 763, 1822, 1298]. search-tree [1000]. searches [1485, 2231]. Searching [2381, 1614, 1367, 980, 784, 2305, 1620, 1613, 1615, 2018]. second [2281, 1506, 296, 1075]. second-order [2281, 1506, 296]. secondary [1932, 1525, 130, 1410, 131]. secrecy [1144]. secret [1297]. Secrets [1421]. sector [436]. secure [864, 1641, 1796, 1820, 2354, 159, 1948]. Securing [858]. Security [1674, 551, 896, 197, 899, 222, 1675, 903, 904]. Seeing [1650]. seeking [763]. segment [1504]. Segmental [463]. segments [1809]. select [1371, 2299, 2298, 1370, 387]. Selected [32, 271, 1434, 1232]. selected-internal [1232]. Selecting [452, 2329]. selection [1432, 1111, 1248, 1258, 1209]. Selective [1385]. Self [1048, 1667, 1445, 1110, 2039, 1715, 1951, 1609, 2030, 1955, 1936, 1940, 2021, 468, 1951]. self-assembling [1936]. Self-assembly [1048, 2039, 1715, 2030, 1940]. Self-deployment [1667]. Self-dual [1445]. Self-improved [1110]. self-organizing [1951]. self-stabilizing [1609, 1955, 2021, 468, 1951]. Selfish [422, 288, 937, 1483, 2208, 2048, 1778]. Semantic [755, 1598, 303, 55, 1416, 356, 1419, 2009, 1002]. Semantics [1157, 515, 1037, 1527, 226, 2273, 757, 156, 2323, 1420, 1076, 899, 1460, 2325, 1594, 1160, 149, 1068, 604, 163, 1153, 883, 1066, 1858, 186, 99, 295, 1863, 328, 1029, 1039, 271, 1041, 51, 562, 273, 680, 1385, 1679]. Semenov [1444]. Semi [1497, 2350, 167, 1732, 1108, 2273, 420, 217, 15, 1970, 1976, 1332, 1011]. semi- [2273]. Semi-formal [1732]. Semi-on-line [167]. Semi-online [1497, 2350, 1108, 217, 1970, 1011]. semi-ordered [1976]. semi-random [1332]. semi-streaming [420]. semi-Thue [15]. semiautomata [639, 2201]. semidefinite [2067]. semigroup [12]. semigroups [2348, 1655, 948, 573, 625, 535]. semimeasures [1257]. Seminormal [1474]. semiretract [949]. semiring [1695]. semirings [400, 359, 580]. Semisimple [2366]. semistructured [1304]. sense [547]. sensing [989, 1663]. sensitive [339, 1069, 692]. sensitivity [1621, 1286]. sensor [1782, 1662, 1665, 1950, 310, 311, 2022, 313, 1486, 1666, 1927, 981, 314]. sensors [2395, 1667, 1953]. sentences [851, 1393]. separability [1230]. Separable [360]. Separating [2070]. Separation [1339, 2250, 1076, 526, 516, 1384, 1078]. separations [1402, 1819]. separator [979]. separators [1411, 294]. Sequence [2276, 2160, 2167, 1499, 1809, 44, 2154, 191]. Sequences [207, 724, 2291, 181, 2114, 911, 2297, 1988, 2299, 1908, 1257, 966, 1473, 689, 2003, 1624, 2232, 1043, 1397, 88, 627, 552]. Sequencing [88]. Sequential [1906, 1883, 388, 2178, 1523, 653, 297, 539, 408, 1005, 932, 1158, 690, 2133, 1157]. serial [1561, 735]. series [1630, 757, 890, 1839, 1659, 213, 892, 375, 1060, 735]. server [1846, 1231, 947]. service [1950, 1889, 157]. services [226, 264]. session [917]. sessions [1923]. SET [85, 1411, 2259, 546, 2311, 1652, 1319, 1484, 2217, 2101, 1307, 1987, 50, 1968, 506, 1881, 2247, 2020, 1954, 423, 814, 1807, 2390, 1699, 1575, 1299, 530, 2397, 877, 1226, 1258, 1748, 191, 1973, 979]. set-agreement [1954]. set-up [1652]. set-valued [2247]. sets [616, 1761, 1990, 1322, 1763, 397, 1765, 1244, 256, 280, 63, 888, 2369, 2269, 2024, 541, 396, 2012, 2085, 1357, 619, 758, 722, 1202, 501, 2356, 1770, 1959, 2253, 1266, 795, 703, 265, 1780, 76, 16, 1306, 941, 531, 2126]. Setting [1658, 1626, 1577]. setups [1960]. Severe [622]. shape [1468, 711, 1767]. shaped [940]. shared [83, 1922]. shared-variable [1922]. Sharing [946, 1645, 2324, 867, 82]. shelf [536]. Sherali [2089]. Shewchuk [1850]. shift [337, 2300, 1275]. shifts [337, 2337]. shop [1812, 1629]. Short [218, 1966, 523, 1405, 1224, 1785]. shorter [972]. Shortest [2262, 1915, 2054, 1262, 958, 106, 1388, 2382, 971, 1408]. shortest-path [106]. shot [2137, 2050]. shrinking [1900]. Shuffle [1449, 2244, 1106, 64]. shuffles [260, 1207]. Shyr [2336]. sided [81]. Sierpinski [1940]. Signal [846, 1874, 305, 1732]. signalling [1998, 1718]. signals [2286]. signature [1566, 1966]. signatures [1254, 976, 1396]. signed [2113]. signing [897]. signs [2264]. similar [608]. Similarity [563, 2119, 1828, 1540]. Simple [1455, 2174, 424, 822, 1086, 2312, 1016, 2128, 928, 1202, 1330, 2249, 1311, 1388, 1720, 1523, 1566, 1332, 1287, 535, 1688, 2, 1123, 2065, 623]. simpler [1114, 1362]. simplex [777]. Simplicial [2389]. simplicity [383]. simplification [1996]. Simplifying [520]. simulate [1841]. simulated [911, 1330]. Simulating [1888]. Simulation [194, 2190, 2193, 1836, 161, 98]. simulations [606]. simultaneous [2082]. Single [1908, 106, 1488, 1975, 805, 421, 833, 182, 793, 1967, 2153, 942, 624, 1054, 1045, 1503]. Single-edge [1908]. single-minded [182]. singular [1472]. singularities [714]. singularity [503]. singularization [1454]. site [1849]. six [2170]. size [2197, 2301, 244, 1764, 2200, 671, 20, 1180, 1357, 2205, 718, 1183, 2177, 1985, 1059, 2003, 830, 266, 1817, 1288, 1688]. sizes [2153, 2335]. SJF [734]. Skew [890, 908, 1635]. skip [540]. Slicing [307, 740, 720, 2004]. sliding [290, 1983]. sliding-block [290]. slopes [1112, 2239]. slow [1754]. Small [2215, 244, 797, 80, 1021, 612, 1870, 523, 1364, 1534, 1941, 1503, 1817, 1944, 2344]. small-world [612, 2344]. smallest [474, 788]. Smooth [262, 1493, 554]. Smoothed [1151, 1149, 258]. snake [2038]. snapshots [2280]. SNP [2178]. social [934]. sofic [243, 337, 636, 2300]. Soft [1695]. Software [1434, 57, 2322, 1730]. Soittola [1495]. solar [2315]. Solitaire [2371]. Solitar [220]. Solomon [1170]. solution [991, 1401, 879, 1707, 2090, 77, 367]. solutions [787, 1022, 2254, 426, 1810, 1057]. solvability [559]. solvable [332]. solve [459, 1025]. Solving [1792, 1524, 445, 1656, 971, 1716, 1275, 1839, 1873, 1383]. Some [637, 456, 1216, 2171, 401, 2173, 254, 767, 1578, 644, 2241, 1053, 1227, 1761, 364, 244, 1197, 365, 1354, 2030, 789, 1595, 25, 842, 1358, 1624, 1093, 67, 31, 1619]. SONET [457, 1301]. sort [931]. Sorted [585, 1031]. sorting [1963, 2213, 2305, 2001, 1364, 1365, 2245]. SOS [1871, 1040]. Sound [1510]. source [421, 1012]. Space [1030, 204, 988, 355, 1271, 253, 509, 254, 1364, 1739, 1740, 2230, 1313, 1808, 121, 1131, 1749, 1408]. Space-aware [1030]. space-bounded [121]. spaces [1736, 1737, 915, 1741, 2235, 496, 1743, 601, 2006, 1186, 2399, 872]. spanner [173]. Spanners [1263, 2215, 381, 1607]. Spanning [2373, 1984, 1311, 1980, 1223, 1545, 1128, 611, 2084]. sparse [381, 1230, 1869, 643, 2360, 1607, 1791, 1171, 1816, 1871, 580, 1332]. Spatial [1672, 704, 873, 666, 569, 36]. Spatial-behavioral [1672]. Special [334, 2302, 1531, 326, 1441, 836, 973, 1613, 588, 93, 2287]. species [380, 985]. specific [2044]. Specification [512, 571, 1236, 2015, 301, 579, 2014, 306]. specifications [899, 2286, 509, 510, 705]. specified [1882]. spectra [2240]. spectral [1914, 2006]. spectrum [1744]. Speed [2315, 748, 497]. spike [2178]. spiking [1329, 2118, 1024, 1526, 2130]. splicing [241, 116, 2342, 691, 1333, 1941, 993, 1135]. split [1056]. splittable [951]. Splitting [1070]. sponsored [1793]. Spreading [2149, 1606]. square [1562, 2260, 815, 1302]. square-free [1562, 2260]. squares [1255, 1203, 2263, 1876, 203, 1544]. SRPT [734]. st [671, 1856]. Stability [497, 317]. stabilization [1879, 1943]. stabilizers [2173, 1632]. stabilizing [1951, 1609, 1955, 2021, 468]. Stable [1411, 1873, 1168, 919, 1626, 1985, 297, 979, 1219]. Stackelberg [1967]. stage [1014, 1855, 2293]. stagnation [259]. stale [2212]. stamps [655]. standard [2161, 216, 1510, 1200, 261, 1206, 1400, 1239]. star [1557, 1445, 180, 215, 2392, 1969]. star-duality [1445]. stars [529]. start [1558]. start-up [1558]. starting [504]. State [2200, 2120, 509, 2135, 25, 2125, 1864, 1272, 154, 461, 1705, 1032, 2077, 2204, 2074, 1356, 1084, 972, 2176, 360, 358, 574, 2225, 1511, 2122, 705, 170, 195, 159, 867, 1343, 2206, 1732, 39]. state-alternating [170]. State-complexity [2200]. state-extended [1511]. stateless [882]. statements [1739]. states [7, 11, 2347, 1743]. Static [1740, 449, 2191, 36, 1211]. stationary [1754]. Statistical [1133, 2103, 846]. statistics [2247, 647, 2104]. steady [1732]. Steiner [2217, 1779, 1232, 1399, 2007]. stepped [1194, 1199]. STG [1383]. Stochastic [2188, 1342, 2190, 1840, 1238, 1855, 2333, 843]. Stone [1735, 708]. stop [1966]. storage [1016]. straight [2144, 1800]. straight-line [2144, 1800]. strands [2033]. Strassen [43]. Strategic [1900]. Strategies [590, 592, 764, 547, 1985, 2214, 1757]. strategy [446, 766]. strategyproof [1645]. stratification [285]. stream [1804, 574, 1694, 305]. streaming [2215, 2062, 420, 1189]. streams [1146]. strength [37]. stretch [2215, 734, 1262]. Strict [1940, 1394, 1200, 2324, 392]. strictly [272, 2267]. strictness [1385]. strikes [1104]. String [1531, 1350, 2378, 1845, 2294, 2161, 79, 548, 2263, 550, 728, 2180, 1651, 2292, 2008]. string-matching [2161]. strings [364, 2262, 2368, 1048, 1016, 1540, 1072, 965, 1537, 2288]. Strip [2252]. Strong [1156, 257, 1589, 2282, 1592, 1639]. Strongly [2050, 789, 1959, 297, 1120, 1212]. Structural [146, 1902, 404, 2164, 926, 604, 1722, 152, 1385]. Structure [293, 565, 2321, 458, 1622, 2398, 116, 1805, 1932, 643, 1141, 2208, 1736, 2111, 1381, 130, 1072, 828, 1410, 2394, 131]. structured [1367, 1792, 1019, 1028, 324]. Structures [1434, 1162, 1174, 1525, 2032, 1369, 2033, 2298, 877, 699, 2140, 2277]. structuring [741]. student [1080]. students [230]. study [662, 2012, 2390, 2333]. Sturmian [208, 232, 2161, 2167, 233, 1198, 2295, 533, 1007, 2233, 261, 1206, 378, 644, 1826]. stuttering [916]. sub [417, 2329]. sub-matrix [2329]. subadditive [836]. subclass [845]. subclasses [1227, 241, 2352]. subdivided [872]. subdivision [1980]. subdivision-constrained [1980]. subexponential [458]. subgraph [1053, 1649, 1228, 1303]. subgraphs [2026, 2088, 1311, 2238, 1823, 611]. subgroup [1184]. subject [2266]. Subjective [1142]. Subjective-cost [1142]. sublanguages [1582]. sublinear [2353, 1221]. sublinear-time [2353]. Sublogarithmic [352, 254]. submodularity [1640]. submonoids [654, 1450, 535]. subnet [83]. subregular [2199]. subregularly [2203]. subsequence [2294, 1886, 1539, 2379, 965, 2154, 266]. subsequences [1147]. subset [1248, 1808]. subset-sum [1808]. subsets [1322, 523, 496, 1901]. subshifts [1444, 2225]. subspace [1754]. subspace-learning [1754]. substitutability [1139]. substitution [939, 2012, 2260, 516]. substitutions [1194, 96, 1199, 2092, 1397, 378]. substrings [1982]. substructures [1455]. Subthreshold [763]. Subthreshold-seeking [763]. subtle [1533]. subtrees [985, 788, 945]. subtyping [1598, 677]. Subword [1799, 2092, 1413, 2232, 828, 800]. subwords [231]. successive [2054]. successors [2267]. Succinct [1851, 26, 1174, 430]. succinctly [1367]. suffice [421]. sufficient [23, 372, 2359, 1308]. Suffix [2297, 128, 2222, 2380, 2135, 548, 1364, 1365, 2230, 1313, 1803, 828, 1537]. suffix-free [2135]. suffixes [1845]. suites [1728]. sum [1111, 1809, 1545, 772, 1876, 1808, 673]. sum-type [1545]. summaries [2067]. summary [1323]. sums [1116, 796]. super [215]. supercover [1760]. superlinear [345]. superpolynomial [2367]. superstring [2382]. supertrees [788]. supply [835]. support [1985, 1476]. surface [871, 369, 1771, 2392]. surface-thinning [1771]. surfaces [1194, 1471, 1199, 1472, 1476]. survey [171, 322, 114, 1581, 1534, 1528, 1532, 1944, 1572]. survivable [2095]. suspicious [1953]. Swapping [1262]. swaps [1312]. swarm [2023]. Swept [1471]. Switched [865]. switches [1895]. switching [219]. symbol [2154]. symbol-wise [2154]. symbolic [1196, 197, 164, 967, 1510, 571, 2337, 885]. symbolic-probabilistic [571]. symbols [825]. symmetric [1560, 1592, 1548, 1286]. symmetries [2157]. Symmetrized [1456]. symmetry [371, 2032, 348, 694, 1131]. symport [692]. symport/antiport [692]. Synapses [1840]. synchronisation [901]. Synchronization [648, 1352, 1178]. synchronizations [1671]. Synchronized [260, 2198, 166]. Synchronizing [5, 1085, 713, 2226, 2262, 1461, 727]. synchronous [701, 1339, 1087, 1924, 1300, 894]. synchronous/asynchronous [1924]. synopses [1175, 1004]. Syntactic [487, 1069, 1508, 535]. Syntax [295, 562, 2224]. syntax-directed [2224]. Synthesis [912, 1690, 1838]. system [54, 154, 911, 660, 2089, 1594, 902, 1058, 1862, 1675, 1939, 1957, 1873, 82, 920]. systematic [10]. systems [593, 1464, 661, 1037, 1161, 1423, 2055, 1947, 2322, 2190, 1327, 1529, 591, 636, 988, 2272, 1731, 140, 989, 2256, 2273, 19, 185, 1792, 753, 1840, 1712, 142, 1019, 1672, 1729, 1710, 1061, 1713, 2118, 642, 1134, 2189, 1606, 139, 1021, 695, 2170, 1354, 2342, 1429, 1716, 1083, 1934, 1870, 1685, 595, 2058, 1022, 23, 11, 1097, 1935, 1373, 969, 2041, 302, 1023, 281, 1202, 2146, 13, 83, 1644, 318, 118, 692, 824, 1024, 2178, 1694, 866, 863, 1719, 512, 1381, 994, 705, 572, 857, 1394, 714, 2152, 1333, 571, 1721, 29, 15, 1026, 120, 349]. systems [454, 1875, 223, 1578, 225, 716, 258, 2337, 1747, 1682, 844, 2107, 321, 1723, 2130, 1819, 49, 754, 1527, 1931].

table [980, 106, 1366]. tabled [894].

TACAS [588]. Tag [1429]. tagging [95]. tail [1380]. tandem [1977]. tape [1909]. tardiness [942, 624, 2349]. tardy [805]. Tarjan [329]. Tarski [1657]. task [1644, 258]. tasks [937, 782, 778, 2335, 2099]. taut [742]. tautologies [623]. taxonomic [411]. Taylor [1703]. tccp [356]. TCP [947, 1983, 287]. TCTL [2285]. teachability [1576]. teachers [2070]. teaching [1576]. team [1118]. technique [1729, 2177, 2124, 1454]. techniques [1860, 1619, 2012, 1187, 1677, 1879, 830, 408, 307]. Template [22, 388]. Template-guided [22]. templates [1874]. Temporal [285, 1463, 283, 645, 1839, 2286, 567, 569, 916, 826, 1090, 1704, 67, 333]. temporary [782]. tennis [367]. tensor [673]. tensors [1281]. Tenth [32]. term [1853, 1701, 1578, 476, 2107]. terminal [2217, 1399]. Termination [2320, 95, 150, 1957]. terms [974, 1703, 692, 1343]. ternary [1562]. test [155, 255, 253, 752, 347, 16]. testable [1655]. Testing [1990, 45, 574, 1694, 1823, 1339, 186, 571, 735, 1209, 98]. tests [486, 285]. texts [2231, 2298]. their [1860, 1061, 840, 1650, 2121, 1448, 705, 908, 263, 1393, 1519, 376, 530]. theme [1787]. themes [202]. theorem [218, 1495, 1317, 2057, 1634, 1657, 1909, 1735, 205, 2336, 1692, 406, 1444, 368, 2370, 1659, 138, 331, 756, 1079, 2235, 708, 745, 365, 1281, 889]. theorems [1943]. Theoret [1529, 2398, 402, 1690, 560, 1135, 1707, 2396, 1123, 2399, 922]. theoretic [974, 364, 879, 1707]. Theoretical [286, 740, 222, 312, 1682, 1378, 480, 1745]. theories [896, 755, 884, 853, 621, 1588, 189]. Theory [286, 1279, 114, 1390, 115, 1585, 849, 1441, 34, 212, 521, 485, 10, 322, 739, 1084, 582, 1082, 581, 1960, 915, 354, 1333, 1742, 756, 1040, 1328, 877, 408, 319, 98, 1291, 606]. There [391]. Thierry [1474]. thinning [1771]. Third [1438]. Third-order [1438]. thousand [2039]. thread [1382, 1925, 306]. thread-based [1925]. threaded [1922]. threads [701]. Three [78, 1025, 496, 377, 176, 1443, 2170, 2172, 1014, 370, 524, 2246, 465]. three-dimensional [370]. Three-player [78]. three-stage [1014]. threshold [2330, 2172, 346, 1810, 1253, 1817]. thresholds [1608]. throughput [1301]. Thue [90, 1971, 15, 1413]. tier [835]. Tight [1560, 2089, 1189, 1294, 1178, 1404]. tightly [1180]. Tile [235, 1127, 1522, 1524, 2034, 142, 642]. tiles [2039, 1936]. Tiling [1813, 2227, 2040, 642]. tilings [2038, 2096, 1445, 2041, 716]. Time [1614, 1264, 107, 1819, 655, 1177, 1652, 327, 1168, 1915, 2353, 988, 1963, 2376, 1848, 1163, 1522, 1765, 845, 2279, 2142, 167, 249, 1028, 1796, 1716, 1934, 1839, 2286, 1097, 1373, 567, 648, 2285, 1173, 360, 1908, 269, 1644, 1738, 1894, 1739, 1496, 1033, 572, 354, 390, 2258, 568, 1958, 1313, 1491, 468, 2372, 285, 1221, 1132, 2050, 735, 1808, 2307, 679, 476, 1287, 1045, 812, 321, 1943, 627, 1681, 2, 1123, 1553, 428, 1696, 1927]. time-bound [1796]. time-bounded [327]. time-restricted [1894]. Time-stamps [655]. time-varying [321, 1681]. timed [594, 1731, 1696, 196, 595, 55, 1926, 572, 1437, 1863, 328, 703, 39, 1679, 1696]. timed-probabilistic [196]. times [2252, 1812, 1916, 1740, 1496, 1054, 1491, 1431, 1503, 1520, 1943]. timeslot [950]. timing [1819]. Tissue [11, 2193, 1716]. tolerant [1789, 2397, 1913, 1892, 1973]. tolls [421]. tomographic [1766, 1765]. tomographical [396, 977]. tomography [370, 372]. tonic [2003]. tools [1423, 637, 1926]. top [1805, 151]. top- [1805]. top-down [151]. topics [271]. topological [1954, 1736, 1737, 1404, 247, 2154]. topologies [804, 1735]. Topology [2115, 1742, 258, 870, 871, 681, 682, 1297, 1738, 478, 1746, 1748, 1749]. tori [930]. torus [1320, 1637]. total [734, 167, 805, 944, 942, 624, 390, 2349, 2258]. totality [1639]. totally [2154, 1034]. tournament [1640, 1118]. tournaments [530]. Tower [954]. TPAL [196]. traceroute [606]. traceroute-like [606]. traces [821, 645, 1090, 655]. tractable [1033, 1640]. trade [1008, 430]. trade-off [1008, 430]. tradeoff [2069]. Tradeoffs [776, 1848, 325]. Trading [1265]. Traffic [1293, 2248, 1301, 1895]. Tragedy [287]. trails [1897]. training [1662]. trajectories [939, 888, 64]. trajectory [343]. trajectory-based [343]. transcendental [1656]. transcription [1837]. transcriptional [2192, 1838]. transducers [932, 1020, 389, 799, 2107]. transductions [247]. transductive [838]. transfer [2370, 2066]. transfinite [1520, 1943]. transform [1363, 532, 1872, 540, 2182, 2293, 1368]. transformation [89, 511, 627, 1096]. transformations [2096, 213, 892, 151]. Transforming [163]. Transition [185, 1355, 2272, 1019, 1356, 49, 754]. transitions [185, 1183, 694, 626]. transitive [150, 1566, 580]. transitive-closure [150]. translation [196, 586]. translations [2224]. translocation [556]. transmembrane [132]. transmission [983]. transmissions [946]. transport [1713, 1938]. Transposing [862]. Transposition [1204, 1051]. transversals [522]. trapezoidal [1826]. traveling [1, 560]. traversal [1008, 1175]. treatment [1954]. Tree [483, 50, 1100, 1222, 1000, 1367, 1331, 1008, 1252, 171, 702, 380, 2387, 976, 128, 1779, 2203, 1262, 1263, 891, 238, 2297, 389, 2111, 2229, 1232, 358, 1150, 581, 1499, 1980, 1809, 1545, 827, 1029, 213, 892, 2289, 1128, 379, 2224, 476, 72, 2107, 1233, 267, 935, 627, 2140, 2373]. tree-length [1263]. tree-structured [1367]. tree-to-tree-series [213]. Tree-walking [483]. Trees [525, 1885, 2217, 519, 1243, 1979, 2097, 1784, 646, 2380, 1961, 731, 2243, 210, 985, 1984, 2250, 1650, 982, 1647, 955, 1999, 1914, 827, 1151, 1399, 772, 1566, 1976, 828, 629, 2357, 1821, 945, 2084]. Treewidth [1962, 634, 434, 1119]. triangle [2041, 1816]. triangles [1940, 817]. Triangulating [1400, 1906]. triangulations [289]. triconnected [265]. trinucleotide [1638]. trip [1534]. triplet [1391]. triplet-dissimilarities [1391]. triplets [814]. trivially [2330]. True [1435, 1337, 700]. True-concurrency [1435, 1337]. truncated [2143]. trust [2274, 1422]. truth [674, 1154, 1278]. truth-finding [1278]. Truthful [937, 2102, 2211, 1981, 2050]. TSP [2106, 1548, 2346]. TTL [47]. TTL-based [47]. Tuning [450]. tuple [355, 1226]. Turing [1109, 1273, 254, 797, 1633, 77, 1520, 1943, 1944]. turn [784]. turned [612]. Tutorial [1384, 305]. Twin [2121]. Twin-roots [2121]. twisted [1806, 1984]. Two [957, 2203, 1296, 744, 874, 1831, 1942, 2236, 1970, 2005, 2239, 925, 1497, 245, 2227, 2037, 968, 911, 1979, 2350, 1648, 1801, 242, 179, 9, 2148, 388, 23, 1769, 963, 217, 833, 2377, 1629, 2225, 707, 2335, 1403, 2266, 2123, 2093, 28, 2267, 265, 1457, 1093, 1108, 2110, 2075, 2102, 2234, 1855]. two-bit [911]. two-dimensional [925, 245, 2227, 2037, 242, 179, 2148, 388, 963, 2377, 2225, 707, 1457]. two-range-values [2102]. Two-stage [1855]. two-way [9, 28, 1093, 2075]. Type [857, 917, 1589, 1594, 1862, 1675, 863, 1545, 625, 1323, 756, 1002, 1096, 1590]. type-checking [1002]. type-safe [863]. typechecking [151]. typed [1591, 485]. types [272, 615, 1584, 1591, 1672, 1674, 1685, 56, 917].

ubiquitous [1300]. UET [2335]. ultrametric [915]. Ultraproducts [1160]. UML [512]. unambiguous [2037, 636]. unary [30]. Unavoidable [722, 1455]. unbalance [1309]. unbounded [1382, 1567, 1575]. uncertain [1805]. uncertainty [1907, 1156, 1855, 1251, 314]. Unconditional [1480]. uncountable [496]. undecidability [1442, 704, 921, 195, 1120]. undecidable [1831, 2040, 1786]. undirected [2358]. unfair [1207]. unfoldings [1701]. unification [1700]. unified [1242, 1626, 748]. Uniform [253, 1472, 471, 1497, 327, 1761, 1649, 2200, 2350, 1551, 1716, 1952, 991, 217, 2266, 2093, 2258, 1970, 841, 1239]. Uniformity [1703, 84, 1551]. uniformly [1195, 1172, 373]. Unifying [478, 622, 1925]. unions [1752]. unique [1087, 1432, 1201, 649, 1642, 2125, 1209]. Uniqueness [578, 2244, 2167, 372]. unit [1899, 1968, 1121, 1787, 817, 2361, 531, 2385, 2026]. unitary [1517]. unitation [1330]. UNITY [1098]. Universal [1788, 2080, 1794, 347, 2053, 1848, 1021, 23, 589, 840, 1282, 824, 1376, 1564, 1941, 2066, 797, 735, 1746, 816, 1944, 2008]. Universality [1689, 994]. universe [618]. unknown [2136, 1305, 1297, 75]. unknowns [2170]. unlabeled [412]. unoriented [965]. unpredictable [1844]. unpublished [1109]. unrelated [732, 1182, 2102]. Unresolved [454]. unrooted [519]. unsatisfiable [523, 178]. unsigned [556]. unsolvability [1276]. unsplittable [422]. unstructured [607]. Unsupervised [1754]. untranquil [1534]. up-to [1187]. updatability [1376]. update [1146]. updates [158, 1999]. updating [940, 2293]. Upper [2019, 2213, 1122, 933, 87, 2253, 1808, 1635, 2239]. upper-bounds [1808]. Upward [1402, 2126]. upward-closed [2126]. usability [141]. use [307, 1478]. useful [1084]. Using [1019, 506, 1677, 225, 918, 2069, 115, 1761, 1576, 2193, 1331, 638, 47, 1003, 56, 595, 1058, 2012, 2324, 666, 991, 503, 548, 1916, 574, 765, 2137, 1637, 1072, 1313, 728, 1701, 2031, 1878, 2374, 885, 1477, 627]. utilities [2046]. utility [1138]. Utilization [802, 825].

V [553, 792]. vacations [2355]. Valiant [1281]. validation [586]. valuation [1143]. value [852]. valued [2056, 1865, 2247, 912, 1690, 1519]. values [504, 2102]. Vandermonde [647]. VaR [1143]. variable [956, 1239, 178, 317, 1922]. variables [373, 1188]. variance [2069]. variant [1977, 728, 2102]. variants [1576, 957, 597, 1619, 775, 1539, 1831, 1270]. variation [1343]. variational [1342]. variations [1533, 202, 1093, 1787]. varieties [641, 1802, 379, 1100]. variety [1655, 2393, 1090, 2147]. various [547]. varying [321, 1681]. VC [842]. VC-dimension [842]. vector [1883, 2059, 1902]. vectors [2054, 1803]. vehicle [1491]. Verification [95, 1705, 1382, 1859, 59, 48, 155, 2322, 591, 1981, 190, 329, 511, 1730, 58, 566, 1732, 306, 307]. verifying [1729]. version [457, 440, 708]. versus [1514, 1195, 642, 320, 1337, 318, 1577, 1288, 1688, 124, 497]. Vertex [719, 175, 1557, 66, 180, 1012, 1266, 795, 1221, 252, 818, 2385]. vertex-covering [66]. vertex-weighted [795]. vertices [608, 524]. very [1816, 2, 1123, 2065]. via [2376, 1484, 116, 1339, 2312, 865, 283, 1852, 187, 1363, 360, 2259, 1907, 1029, 294, 1384]. video [1500]. video-on-demand [1500]. View [1001, 141, 873, 1376, 1596, 2306, 788, 1475, 1130]. View-based [1001]. views [918, 1003, 1726]. Villemaire [208]. virtual [445]. virtuality [1935, 995]. visibility [168, 1388]. visible [601]. visited [1937]. Visual [938, 1062]. visualization [369, 1738]. VLSI [1848]. volume [2329]. volumetric [1471]. voting [2152]. VPSPACE [2370]. Vries [77]. vs [107, 1523, 295, 2155].

W [835]. walk [2374]. walking [483]. walks [1916, 1517, 2236, 2255]. Walrasian [182]. Walsh [532]. watchman [1287]. Watson [695, 2202]. wave [1874]. wavebands [109]. wavelet [1004]. way [9, 1639, 2129, 1133, 347, 1684, 2040, 28, 1093, 2075]. WCET [1373]. Weak [275, 1557, 38, 72, 1881, 1862, 1587, 1270, 1187, 818]. Weak-vertex-pancyclicity [1557]. Weakening [227]. weaker [1402, 2267]. weakly [1696, 1286]. web [226, 1674, 1550, 276]. Week [1018]. weft [520]. Wegener [2303, 2302]. Weight [1411, 236, 1955, 1564, 1311, 1809, 1398, 2307, 818, 979]. weight-based [1955]. weight-constrained [1809]. Weighted [764, 891, 1181, 2223, 1052, 1697, 2044, 1542, 1898, 2391, 1489, 2264, 1216, 2351, 805, 1968, 774, 1637, 942, 624, 834, 390, 2258, 795, 1004, 1568, 267, 1545]. Well [1106, 1002, 839, 1705, 1019, 424, 426]. Well-calibrated [839]. Well-definedness [1002]. well-formed [1705]. well-structured [1019]. Weyl [2235]. whack [783]. Wheeler [1368, 89, 1361, 1363, 1362, 2182, 2293]. Wheeler-based [1362]. where [2347]. whether [1823]. which [1739, 476]. who [1531]. whose [1952, 1656]. Wide [547, 79]. Wide-sense [547]. widenings [332]. Width [1288, 1245, 382, 719, 1015, 1119]. Wiener [1271]. Wilf [205, 2183]. Williamson [598]. window [79, 1983]. windows [1147, 1769, 1491]. winning [492]. winnowing [1484]. wireless [1782, 946, 1604, 1950, 733, 1953, 1664, 313, 1486, 1959, 903, 1927, 2104]. wise [2154]. withdrawal [622]. within [1727, 1118]. without [1318, 2191, 185, 1062, 23, 944, 1183, 2354, 1414, 332, 1785, 1186, 2399]. Word [721, 801, 90, 974, 464, 342, 2170, 1106, 2176, 1202, 1203, 2228, 2206, 1625]. word-expansion [2228]. word-shuffle [1106]. Words [232, 1453, 90, 2221, 208, 1430, 1195, 2161, 1389, 237, 262, 2244, 1317, 1971, 1990, 2163, 240, 1197, 554, 1493, 639, 1482, 2165, 2166, 2115, 233, 1198, 2295, 2097, 242, 2168, 169, 1562, 1443, 2171, 2363, 721, 533, 1200, 1446, 1905, 1790, 1407, 1986, 236, 399, 649, 1996, 1502, 2338, 2240, 707, 2121, 1632, 1204, 1007, 1409, 2233, 1205, 2251, 1451, 261, 1457, 203, 2098, 1206, 828, 1046, 1274, 378, 2183, 557, 2239, 644, 1826]. Work [1531]. workflow [566]. working [1313, 1258]. workload [1004]. workload-based [1004]. works [632, 2382]. Workshop [32]. World [1699, 612, 1066, 1534, 284, 2344]. World-set [1699]. worlds [1160]. worst [776, 2162, 1141, 1639, 812]. worst-case [776, 2162, 1141, 1639, 812]. worth [2039]. wreath [641].

X [63, 574, 1694]. X-machine [574, 1694]. X-rays [63]. XML [148, 151, 152]. XOR [193]. XPath [146, 827]. xyz [342].

Yao [2026, 224, 227]. Yao-like [2026]. years [1040]. yeast [1838]. Yu [2336].

Zariski [1744]. Zeno [687]. zero [1061, 909, 1877, 2008, 1133]. zero-knowledge [1061, 2008]. zeta [1229].

References

Krumke:2003:NOT

[1] Sven O. Krumke, Willem E. de Paepe, Diana Poensgen, and Leen Stougie. News from the online traveling repairman. Theoretical Computer Science, 295(1–3):279–294, February 24, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [560].

Yokomori:2003:PTI

[2] Takashi Yokomori. Polynomial-time identification of very simple grammars from positive data. Theoretical Computer Science, 298(1):179–206, April 4, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [1123].

Champarnaud:2004:NRA

[3] J.-M. Champarnaud and F. Coulon. NFA reduction algorithms by means of regular inequalities. Theoretical Computer Science, 327(3):241–253, November 2, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [402].

Karhumaki:2005:P

[4] Juhani Karhumäki and Grzegorz Rozenberg. Preface. Theoretical Computer Science, 330(1):1, January 31, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ananichev:2005:SGM

[5] D. S. Ananichev and M. V. Volkov. Synchronizing generalized monotonic automata. Theoretical Computer Science, 330(1):3–13, January 31, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Austinat:2005:RFC

[6] Holger Austinat, Volker Diekert, Ulrich Hertrampf, and Holger Petersen. Regular frequency computations. Theoretical Computer Science, 330(1):15– 21, January 31, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Campeanu:2005:MSL

[7] Cezar Câmpeanu, Nicolae Sântean, and Sheng Yu. Mergible states in large NFA. Theoretical Computer Science, 330(1):23–34, January 31, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carpi:2005:CML

[8] Arturo Carpi and Aldo de Luca. Completions in measure of languages and related combinatorial problems. Theoretical Computer Science, 330(1):35– 57, January 31, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dang:2005:TWN

[9] Zhe Dang, Oscar H. Ibarra, and Zhi-Wei Sun. On two-way nondeterministic finite automata with one reversalbounded counter. Theoretical Computer Science, 330(1):59–79, January 31, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ding:2005:CTC

[10] Cunsheng Ding and Xuesong Wang. A coding theory construction of new systematic authentication codes. Theoretical Computer Science, 330(1):81– 99, January 31, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Freund:2005:TPS

[11] Rudolf Freund, Gheorghe Păun, and Mario J. Pérez-Jiménez. Tissue P systems with channel states. Theoretical Computer Science, 330(1):101– 116, January 31, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harju:2005:EFS

[12] Tero Harju and Dirk Nowotka. On the equation x^{k} = z_{1}^{k_{1}}z_{2}^{k_{2}} ⋯ z_{n}^{k_{n}} in a free semigroup. Theoretical Computer Science, 330(1):117–121, January 31, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Honkala:2005:LEP

[13] Juha Honkala. The language equivalence problem for HD0L systems having D0L growths. Theoretical Computer Science, 330(1):123–133, January 31, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hromkovic:2005:PRM

[14] Juraj Hromkovič and Georg Schnitger. On the power of randomized multicounter machines. Theoretical Computer Science, 330(1):135–144, January 31, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Matiyasevich:2005:DPS

[15] Yuri Matiyasevich and Géraud Sénizergues. Decision problems for semi-Thue systems with a few rules. Theoretical Computer Science, 330(1):145–169, January 31, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ruohonen:2005:ETS

[16] Keijo Ruohonen. Explicit test sets for iterated morphisms in free monoids and metabelian groups. Theoretical Computer Science, 330(1):171–191, January 31, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBa

[17] Anonymous. Editorial board. Theoretical Computer Science, 330(1):v– ix, January 31, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Csuhaj-Varju:2005:P

[18] E. Csuhaj-Varjú and D. Wotschke. Preface. Theoretical Computer Science, 330(2):193, February 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bordihn:2005:NCC

[19] Henning Bordihn. On the number of components in cooperating distributed grammar systems. Theoretical Computer Science, 330(2):195–204, February 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Castellanos:2005:SCH

[20] Juan Castellanos, Peter Leupold, and Victor Mitrana. On the size complexity of hybrid networks of evolutionary processors. Theoretical Computer Science, 330(2):205–220, February 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Champarnaud:2005:RGD

[21] Jean-Marc Champarnaud and Thomas Paranthoën. Random generation of DFAs. Theoretical Computer Science, 330(2):221–235, February 2, 2005. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daley:2005:TGD

[22] Mark Daley and Ian McQuillan. Template-guided DNA recombination. Theoretical Computer Science, 330(2): 237–250, February 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Freund:2005:CUP

[23] Rudolf Freund, Lila Kari, Marion Oswald, and Petr Sosík. Computationally universal P systems without priorities: two catalysts are sufficient. Theoretical Computer Science, 330(2):251– 266, February 2, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Holzer:2005:DCF

[24] Markus Holzer and Martin Kutrib. On the descriptional complexity of finite automata with modified acceptance conditions. Theoretical Computer Science, 330(2):267–285, February 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jiraskova:2005:SCS

[25] Galina Jirásková. State complexity of some operations on binary regular languages. Theoretical Computer Science, 330(2):287–298, February 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kappes:2005:SRL

[26] Martin Kappes and Frank Nießner. Succinct representations of languages by DFA with different levels of reliability. Theoretical Computer Science, 330(2):299–310, February 2, 2005. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kutrib:2005:DPH

[27] Martin Kutrib. On the descriptional power of heads, counters, and pebbles. Theoretical Computer Science, 330(2): 311–324, February 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Malcher:2005:TWC

[28] Andreas Malcher. On two-way communication in cellular automata with a fixed number of cells. Theoretical Computer Science, 330(2):325–338, February 2, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Margenstern:2005:CFI

[29] Maurice Margenstern, Gheorghe Păun, Yurii Rogozhin, and Sergey Verlan. Context-free insertion–deletion systems. Theoretical Computer Science, 330(2):339–348, February 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mera:2005:CUN

[30] Filippo Mera and Giovanni Pighizzini. Complementing unary nondeterministic automata. Theoretical Computer Science, 330(2):349–360, February 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vaszil:2005:DCS

[31] György Vaszil. On the descriptional complexity of some rewriting mechanisms regulated by context conditions. Theoretical Computer Science, 330(2): 361–373, February 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Corradini:2005:SPT

[32] Flavio Corradini and Uwe Nestmann. Selected papers of the Tenth International Workshop on Expressiveness in Concurrency (EXPRESS 2003). Theoretical Computer Science, 330(3):375– 376, February 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aceto:2005:CHM

[33] Luca Aceto, Wan Fokkink, Anna Ingólfsdóttir, and Bas Luttik. CCS with Hennessy’s merge has no finiteequational axiomatization. Theoretical Computer Science, 330(3):377–405, February 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bolton:2005:HFB

[34] Christie Bolton and Gavin Lowe. A hierarchy of failures-based models: theory and application. Theoretical Computer Science, 330(3):407–438, February 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carayol:2005:RMA

[35] Arnaud Carayol, Daniel Hirschkoff, and Davide Sangiorgi. On the representation of McCarthy’s amb in the π-calculus. Theoretical Computer Science, 330(3):439–473, February 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lozes:2005:ESC

[36] Étienne Lozes. Elimination of spatial connectives in static spatial logics. Theoretical Computer Science, 330(3):475– 499, February 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maffeis:2005:CSP

[37] Sergio Maffeis and Iain Phillips. On the computational strength of pure ambient calculi. Theoretical Computer Science, 330(3):501–551, February 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mayr:2005:WBR

[38] Richard Mayr. Weak bisimilarity and regularity of context-free processes is EXPTIME-hard. Theoretical Computer Science, 330(3):553–575, February 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Valencia:2005:DIS

[39] Frank D. Valencia. Decidability of infinite-state timed CCP processes and first-order LTL. Theoretical Computer Science, 330(3):577–607, February 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:AIa

[40] Anonymous. Author index. Theoretical Computer Science, 330(3):609– 610, February 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:MI

[41] Anonymous. Master index. Theoretical Computer Science, 330(3):611– 622, February 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Orejas:2005:P

[42] Fernando Orejas and Jan van Leeuwen. Preface. Theoretical Computer Science, 331(1):1–2, February 15, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blaser:2005:BAS

[43] Markus Bläser. Beyond the Alder– Strassen bound. Theoretical Computer Science, 331(1):3–21, February 15, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sadakane:2005:CAL

[44] Kunihiko Sadakane, Nadia Takki-Chebihi, and Takeshi Tokuyama. Combinatorics and algorithms for lowdiscrepancy roundings of a real sequence. Theoretical Computer Science, 331(1):23–36, February 15, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czumaj:2005:THC

[45] Artur Czumaj and Christian Sohler. Testing hypergraph colorability. Theoretical Computer Science, 331(1):37– 52, February 15, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sibeyn:2005:FGB

[46] Jop F. Sibeyn. Faster gossiping on butterfly networks. Theoretical Computer Science, 331(1):53–72, February 15, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cohen:2005:PAD

[47] Edith Cohen, Eran Halperin, and Haim Kaplan. Performance aspects of distributed caches using TTL-based consistency. Theoretical Computer Science, 331(1):73–96, February 15, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alur:2005:RVM

[48] Rajeev Alur, Kousha Etessami, and Mihalis Yannakakis. Realizability and verification of MSC graphs. Theoretical Computer Science, 331(1):97– 114, February 15, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanBreugel:2005:BPP

[49] Franck van Breugel and James Worrell. A behavioural pseudometric for probabilistic transition systems. Theoretical Computer Science, 331(1):115– 142, February 15, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Comon:2005:TAO

[50] Hubert Comon and Véronique Cortier. Tree automata with one memory set constraints and cryptographic protocols. Theoretical Computer Science, 331(1):143–214, February 15, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schroder:2005:ASC

[51] Lutz Schröder, Till Mossakowski, Andrzej Tarlecki, Bartek Klin, and Piotr Hoffman. Amalgamation in the semantics of CASL. Theoretical Computer Science, 331(1):215–247, February 15, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBb

[52] Anonymous. Editorial board. Theoretical Computer Science, 331(1):v– ix, February 15, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deBoer:2005:Pa

[53] Frank de Boer and Marcello Bonsangue. Preface. Theoretical Computer Science, 331(2–3):249–250, February 25, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abraham:2005:ABP

[54] Erika Ábrahám, Frank S. de Boer, Willem-Paul de Roever, and Martin Steffen. An assertion-based proof system for multithreaded Java. Theoretical Computer Science, 331(2–3):251– 290, February 25, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hooman:2005:SMT

[55] Jozef Hooman and Jaco van de Pol. Semantic models of a timed distributed dataspace architecture. Theoretical Computer Science, 331(2–3):291–323, February 25, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferrari:2005:CMH

[56] Gianluigi Ferrari, Ugo Montanari, and Emilio Tuosto. Coalgebraic minimization of HD-automata for the π-calculus using polymorphic types. Theoretical Computer Science, 331(2–3):325– 365, February 25, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Achermann:2005:CRA

[57] Franz Achermann and Oscar Nierstrasz. A calculus for reasoning about software composition. Theoretical Computer Science, 331(2–3):367– 396, February 25, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kesten:2005:CAC

[58] Yonit Kesten and Amir Pnueli. A compositional approach to CTL^{∗} verification. Theoretical Computer Science, 331(2–3):397–428, February 25, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sekerinski:2005:VRF

[59] Emil Sekerinski. Verification and refinement with fine-grained actionbased concurrent objects. Theoretical Computer Science, 331(2–3):429– 455, February 25, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pattinson:2005:CAM

[60] Dirk Pattinson and Martin Wirsing. A coordination approach to mobile components. Theoretical Computer Science, 331(2–3):457–483, February 25, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:AIb

[61] Anonymous. Author index. Theoretical Computer Science, 331(2–3):485– 486, February 25, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lim:2005:CPM

[62] Andrew Lim, Brian Rodrigues, Fan Wang, and Zhou Xu. k-Center problems with minimum coverage. Theoretical Computer Science, 332(1–3):1– 17, February 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daurat:2005:DQC

[63] Alain Daurat. Determination of Qconvex sets by X-rays. Theoretical Computer Science, 332(1–3):19–45, February 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kari:2005:ASD

[64] Lila Kari and Petr Sosík. Aspects of shuffle and deletion on trajectories. Theoretical Computer Science, 332(1– 3):47–61, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seiden:2005:NBR

[65] Steven S. Seiden, Peter P. Chen, R. F. Lax, J. Chen, and Guoli Ding. New bounds for randomized busing. Theoretical Computer Science, 332(1–3):63– 81, February 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demange:2005:LVC

[66] Marc Demange and Vangelis Th. Paschos. On-line vertex-covering. Theoretical Computer Science, 332(1–3): 83–108, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shi:2005:NFA

[67] Qingmin Shi and Joseph JáJá. A new framework for addressing temporal range queries and some preliminary results. Theoretical Computer Science, 332(1–3):109–121, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Serna:2005:ANB

[68] Maria Serna, Luca Trevisan, and Fatos Xhafa. The approximability of non-Boolean satisfiability problems and restricted integer programming. Theoretical Computer Science, 332(1–3):123– 139, February 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lombardy:2005:DRE

[69] Sylvain Lombardy and Jacques Sakarovitch. Derivatives of rational expressions with multiplicity. Theoretical Computer Science, 332(1–3):141–177, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berstel:2005:ML

[70] Jean Berstel, Luc Boasson, and Michel Latteux. Mixed languages. Theoretical Computer Science, 332(1–3):179– 198, February 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brandenburg:2005:FGA

[71] F. J. Brandenburg and K. Skodinis. Finite graph automata for linear and boundary graph languages. Theoretical Computer Science, 332(1–3):199– 232, February 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Torre:2005:WMA

[72] Salvatore La Torre, Aniello Murano, and Margherita Napoli. Weak Muller acceptance conditions for tree automata. Theoretical Computer Science, 332(1–3):233–250, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fuchs:2005:OML

[73] Bernhard Fuchs, Winfried Hochstättler, and Walter Kern. Online matching on a line. Theoretical Computer Science, 332(1–3):251–264, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dahllof:2005:CMF

[74] Vilhelm Dahllöf, Peter Jonsson, and Magnus Wahlström. Counting models for 2SAT and 3SAT formulae. Theoretical Computer Science, 332(1–3):265– 291, February 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jevtic:2005:LBC

[75] Nikola Jevtić, Alon Orlitsky, and Narayana P. Santhanam. A lower bound on compression of unknown alphabets. Theoretical Computer Science, 332(1–3):293–311, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pedicini:2005:GES

[76] Marco Pedicini. Greedy expansions and sets with deleted digits. Theoretical Computer Science, 332(1–3):313– 336, February 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Weihrauch:2005:CSK

[77] Klaus Weihrauch and Ning Zhong. Computing the solution of the Korteweg–de Vries equation with arbitrary precision on Turing machines. Theoretical Computer Science, 332 (1–3):337–366, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cincotti:2005:TPP

[78] A. Cincotti. Three-player partizan games. Theoretical Computer Science, 332(1–3):367–389, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

He:2005:WWS

[79] Longtao He, Binxing Fang, and Jie Sui. The wide window string matching algorithm. Theoretical Computer Science, 332(1–3):391–404, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cardinal:2005:DSK

[80] Jean Cardinal and Stefan Langerman. Designing small keyboards is hard. Theoretical Computer Science, 332(1–3):405–415, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nagamochi:2005:OSC

[81] Hiroshi Nagamochi. On the one-sided crossing minimization in a bipartite graph with large degrees. Theoretical Computer Science, 332(1–3):417– 446, February 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yeh:2005:RSS

[82] Hong-Gwa Yeh and Xuding Zhu. Resource-sharing system scheduling and circular chromatic number. Theoretical Computer Science, 332(1–3): 447–460, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huang:2005:PPS

[83] H. J. Huang, L. Jiao, and T. Y. Cheung. Property-preserving subnet reductions for designing manufacturing systems with shared resources. Theoretical Computer Science, 332(1–3):461– 485, February 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nishimura:2005:UQC

[84] Harumichi Nishimura and Masanao Ozawa. Uniformity of quantum circuit families for error-free algorithms. Theoretical Computer Science, 332(1–3): 487–496, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bazgan:2005:DAM

[85] Cristina Bazgan, Jérôme Monnot, Vangelis Th. Paschos, and Fabrice Serrière. On the differential approximation of MIN SET COVER. Theoretical Computer Science, 332(1–3):497–513, February 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Byskov:2005:NAE

[86] Jesper Makholm Byskov, Bolette Ammitzbøll Madsen, and Bjarke Skjernaa. New algorithms for Exact Satisfiability. Theoretical Computer Science, 332(1–3):515–541, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gharavi:2005:UBL

[87] Reza Gharavi and V. Anantharam. An upper bound for the largest Lyapunov exponent of a Markovian product of nonnegative matrices. Theoretical Computer Science, 332(1–3):543– 557, February 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tsur:2005:SHE

[88] Dekel Tsur. Sequencing by hybridization with errors: handling longer sequences. Theoretical Computer Science, 332(1–3):559–566, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crochemore:2005:NBW

[89] Maxime Crochemore, Jacques Désarménien, and Dominique Perrin. A note on the Burrows–Wheeler transformation. Theoretical Computer Science, 332 (1–3):567–572, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aberkane:2005:TMW

[90] Ali Aberkane and James D. Currie. The Thue–Morse word contains circular 5/2^{+} power free words of every length. Theoretical Computer Science, 332(1–3):573–581, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:AIc

[91] Anonymous. Author index. Theoretical Computer Science, 332(1–3):583– 584, February 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBc

[92] Anonymous. Editorial board. Theoretical Computer Science, 332(1–3):v– ix, February 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gordon:2005:PSI

[93] Andrew D. Gordon, Programme Chair FOSSACS 2003. Preface for the Special Issue. Theoretical Computer Science, 333(1–2):1, March 1, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abel:2005:ICS

[94] Andreas Abel, Ralph Matthes, and Tarmo Uustalu. Iteration and coiteration schemes for higher-order and nested datatypes. Theoretical Computer Science, 333(1–2):3–66, March 1, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanchet:2005:VCP

[95] Bruno Blanchet and Andreas Podelski. Verification of cryptographic protocols: tagging enforces termination. Theoretical Computer Science, 333(1– 2):67–90, March 1, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonelli:2005:NHO

[96] Eduardo Bonelli. Normalisation for higher-order calculi with explicit substitutions. Theoretical Computer Science, 333(1–2):91–125, March 1, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boneva:2005:WAC

[97] Iovka Boneva and Jean-Marc Talbot. When ambients cannot be opened. Theoretical Computer Science, 333(1– 2):127–169, March 1, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanBreugel:2005:DTT

[98] Franck van Breugel, Michael Mislove, Joël Ouaknine, and James Worrell. Domain theory, testing and simulation for labelled Markov processes. Theoretical Computer Science, 333(1–2):171– 197, March 1, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laird:2005:GSL

[99] J. Laird. Game semantics and linear CPS interpretation. Theoretical Computer Science, 333(1–2):199– 224, March 1, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lugiez:2005:MAC

[100] Denis Lugiez. Multitree automata that count. Theoretical Computer Science, 333(1–2):225–263, March 1, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Santocanale:2005:ACC

[101] Luigi Santocanale and André Arnold. Ambiguous classes in μ-calculi hierarchies. Theoretical Computer Science, 333(1–2):265–296, March 1, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sassone:2005:LRC

[102] Vladimiro Sassone and Paweł Sobociński. Locating reaction with 2-categories. Theoretical Computer Science, 333(1– 2):297–327, March 1, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBd

[103] Anonymous. Editorial board. Theoretical Computer Science, 333(1–2): v–ix, March 1, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Peleg:2005:P

[104] David Peleg and Jop Sibeyn. Preface. Theoretical Computer Science, 333(3): 329–330, March 3, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Osterloh:2005:OOR

[105] Andre Osterloh. Optimal oblivious routing on d-dimensional meshes. Theoretical Computer Science, 333(3):331– 346, March 3, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ito:2005:SBT

[106] Hiro Ito, Kazuo Iwama, Yasuo Okabe, and Takuya Yoshihiro. Single backup table schemes for shortest-path routing. Theoretical Computer Science, 333(3):347–353, March 3, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kowalski:2005:TCR

[107] Dariusz R. Kowalski and Andrzej Pelc. Time complexity of radio broadcasting: adaptiveness vs. obliviousness and randomization vs. determinism. Theoretical Computer Science, 333(3):355– 371, March 3, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Busch:2005:CCL

[108] Costas Busch, Marios Mavronicolas, and Paul Spirakis. The cost of concurrent, low-contention Read&Modify&Write. Theoretical Computer Science, 333(3):373–400, March 3, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flammini:2005:RWA

[109] Michele Flammini, Alfredo Navarra, and Andrzej Proskurowski. On routing of wavebands for all-to-all communications in all-optical paths and cycles ^{,}. Theoretical Computer Science, 333(3): 401–413, March 3, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gavoille:2005:IRR

[110] Cyril Gavoille and Martin Nehéz. Interval routing in reliability networks. Theoretical Computer Science, 333(3): 415–432, March 3, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fernandez:2005:DAP

[111] Antonio Fernández, Chryssis Georgiou, Alexander Russell, and Alex A. Shvartsman. The Do-All problem with Byzantine processor failures. Theoretical Computer Science, 333(3):433– 454, March 3, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:AId

[112] Anonymous. Author index. Theoretical Computer Science, 333(3):455– 456, March 3, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2005:Pa

[113] G. Rozenberg. Preface. Theoretical Computer Science, 334(1–3):1–2, April 15, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kari:2005:TCA

[114] Jarkko Kari. Theory of cellular automata: a survey. Theoretical Computer Science, 334(1–3):3–33, April 15, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Auger:2005:CRS

[115] Anne Auger. Convergence results for the (1, λ)-SA-ES using the theory of ϕ-irreducible Markov chains. Theoretical Computer Science, 334(1–3):35– 69, April 15, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonizzoni:2005:SRR

[116] Paola Bonizzoni, Clelia De Felice, and Rosalba Zizza. The structure of reflexive regular splicing languages via Schützenberger constants. Theoretical Computer Science, 334(1–3):71– 98, April 15, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gaborit:2005:LCD

[117] Philippe Gaborit and Oliver D. King. Linear constructions for DNA codes. Theoretical Computer Science, 334(1– 3):99–113, April 15, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2005:MHP

[118] Oscar H. Ibarra. On membrane hierarchy in P systems. Theoretical Computer Science, 334(1–3):115–129, April 15, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kari:2005:PBF

[119] Lila Kari, Stavros Konstantinidis, and Petr Sosík. On properties of bond-free DNA languages. Theoretical Computer Science, 334(1–3):131–159, April 15, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mutyam:2005:RPS

[120] Madhu Mutyam. Rewriting P systems: improved hierarchies. Theoretical Computer Science, 334(1–3):161– 175, April 15, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sauerhoff:2005:QBP

[121] Martin Sauerhoff and Detlef Sieling. Quantum branching programs and space-bounded nonuniform quantum complexity. Theoretical Computer Science, 334(1–3):177–225, April 15, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Woods:2005:OMC

[122] Damien Woods and Thomas J. Naughton. An optical model of computation. Theoretical Computer Science, 334(1–3):227–258, April 15, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2005:MBR

[123] Guowu Yang, William N. N. Hung, Xiaoyu Song, and Marek Perkowski. Majority-based reversible logic gates. Theoretical Computer Science, 334(1– 3):259–274, April 15, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yamasaki:2005:QVD

[124] Tomohiro Yamasaki, Hirotada Kobayashi, and Hiroshi Imai. Quantum versus deterministic counter automata. Theoretical Computer Science, 334(1–3): 275–297, April 15, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:AIe

[125] Anonymous. Author index. Theoretical Computer Science, 334(1–3): 299, April 15, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBe

[126] Anonymous. Editorial board. Theoretical Computer Science, 334(1–3): v–ix, April 15, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Apostolico:2005:F

[127] Alberto Apostolico and Raffaele Giancarlo. Foreword. Theoretical Computer Science, 335(1):1–2, May 20, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chattaraj:2005:IST

[128] Abhijit Chattaraj and Laxmi Parida. An inexact-suffix-tree-based algorithm for detecting extensible patterns. Theoretical Computer Science, 335(1):3– 14, May 20, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2005:CFE

[129] Lusheng Wang, Hao Zhao, Guozhu Dong, and Jianping Li. On the complexity of finding emerging patterns. Theoretical Computer Science, 335(1): 15–27, May 20, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mauri:2005:APM

[130] Giancarlo Mauri and Giulio Pavesi. Algorithms for pattern matching and discovery in RNA secondary structure. Theoretical Computer Science, 335(1): 29–51, May 20, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Xu:2005:EMR

[131] Ying Xu, Lusheng Wang, Hao Zhao, and Jianping Li. Exact matching of RNA secondary structure patterns. Theoretical Computer Science, 335(1): 53–66, May 20, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Waldispuhl:2005:MPA

[132] Jérôme Waldispühl and Jean-Marc Steyaert. Modeling and predicting allα transmembrane proteins including helix–helix pairing. Theoretical Computer Science, 335(1):67–92, May 20, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choy:2005:CMA

[133] Charles Choy, Jesper Jansson, Kunihiko Sadakane, and Wing-Kin Sung. Computing the maximum agreement of phylogenetic networks. Theoretical Computer Science, 335(1):93–107, May 20, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bafna:2005:PAH

[134] Vineet Bafna, Sorin Istrail, Giuseppe Lancia, and Romeo Rizzi. Polynomial and APX-hard cases of the individual haplotyping problem. Theoretical Computer Science, 335(1):109–125, May 20, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBf

[135] Anonymous. Editorial board. Theoretical Computer Science, 335(1):v– ix, May 20, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aceto:2005:GEF

[136] Luca Aceto, Wan Fokkink, Anna Ingolfsdottir, and Zoltán Ésik. Guest editors’ foreword. Theoretical Computer Science, 335(2–3):127–129, May 23, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeten:2005:BHP

[137] J. C. M. Baeten. A brief history of process algebra. Theoretical Computer Science, 335(2–3):131–146, May 23, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Luttik:2005:DOA

[138] Bas Luttik and Vincent van Oostrom. Decomposition orders — another generalisation of the fundamental theorem of arithmetic. Theoretical Computer Science, 335(2–3):147–186, May 23, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Corradini:2005:MPA

[139] F. Corradini and W. Vogler. Measuring the performance of asynchronous systems with PAFAS. Theoretical Computer Science, 335(2–3):187–213, May 23, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bergstra:2005:PAH

[140] J. A. Bergstra and C. A. Middelburg. Process algebra for hybrid systems. Theoretical Computer Science, 335(2– 3):215–280, May 23, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aldini:2005:UPA

[141] Alessandro Aldini and Marco Bernardo. On the usability of process algebra: an architectural view. Theoretical Computer Science, 335(2–3):281–329, May 23, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bruni:2005:OCD

[142] Roberto Bruni, Ugo Montanari, and Vladimiro Sassone. Observational congruences for dynamically reconfigurable tile systems. Theoretical Computer Science, 335(2–3):331–372, May 23, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Palamidessi:2005:REC

[143] Catuscia Palamidessi and Oltea Mihaela Herescu. A randomized encoding of the π-calculus with mixed choice. Theoretical Computer Science, 335(2– 3):373–404, May 23, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:AIf

[144] Anonymous. Author index. Theoretical Computer Science, 335(2–3):405– 406, May 23, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lenzerini:2005:P

[145] Maurizio Lenzerini. Preface. Theoretical Computer Science, 336(1):1, May 25, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Benedikt:2005:SPX

[146] Michael Benedikt, Wenfei Fan, and Gabriel Kuper. Structural properties of XPath fragments. Theoretical Computer Science, 336(1):3–31, May 25, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calvanese:2005:DCR

[147] Diego Calvanese, Giuseppe De Giacomo, and Moshe Y. Vardi. Decidable containment of recursive queries. Theoretical Computer Science, 336(1): 33–56, May 25, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Deutsch:2005:XQC

[148] Alin Deutsch and Val Tannen. XML queries and constraints, containment and reformulation. Theoretical Computer Science, 336(1):57–87, May 25, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fagin:2005:DES

[149] Ronald Fagin, Phokion G. Kolaitis, Renée J. Miller, and Lucian Popa. Data exchange: semantics and query answering. Theoretical Computer Science, 336 (1):89–124, May 25, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Geerts:2005:DTQ

[150] Floris Geerts and Bart Kuijpers. On the decidability of termination of query evaluation in transitive-closure logics for polynomial constraint databases. Theoretical Computer Science, 336(1): 125–151, May 25, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martens:2005:CTT

[151] Wim Martens and Frank Neven. On the complexity of typechecking topdown XML transformations. Theoretical Computer Science, 336(1):153– 180, May 25, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Toman:2005:RAS

[152] David Toman and Grant Weddell. On reasoning about structural equality in XML: a description logic approach. Theoretical Computer Science, 336(1): 181–203, May 25, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBg

[153] Anonymous. Editorial board. Theoretical Computer Science, 336(1):v– ix, May 25, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Borger:2005:ASM

[154] Egon Börger. Abstract state machines and high-level system design and analysis. Theoretical Computer Science, 336 (2–3):205–207, May 26, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Artho:2005:CTC

[155] Cyrille Artho, Howard Barringer, Allen Goldberg, Klaus Havelund, Sarfraz Khurshid, Mike Lowry, Corina Pasareanu, Grigore Roşu, Koushik Sen, Willem Visser, and Rich Washington. Combining test case generation and runtime verification. Theoretical Computer Science, 336(2–3):209– 234, May 26, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Borger:2005:HLM

[156] Egon Börger, Nicu G. Fruja, Vincenzo Gervasi, and Robert F. Stärk. A highlevel modular definition of the semantics of C#. Theoretical Computer Science, 336(2–3):235–284, May 26, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.only4gurus.net/
miscellaneous/cs03.pdf.

Glasser:2005:FDA

[157] Uwe Glässer and Qian-Ping Gu. Formal description and analysis of a distributed location service for mobile ad hoc networks. Theoretical Computer Science, 336(2–3):285–309, May 26, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gurevich:2005:PU

[158] Yuri Gurevich and Nikolai Tillmann. Partial updates. Theoretical Computer Science, 336(2–3):311–342, May 26, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nanchen:2005:LSM

[159] Stanislas Nanchen and Robert F. Stärk. A logic for secure memory access of abstract state machines. Theoretical Computer Science, 336(2–3): 343–365, May 26, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Roman:2005:PEC

[160] Gruia-Catalin Roman and Jamie Payton. A principled exploration of coordination models. Theoretical Computer Science, 336(2–3):367–401, May 26, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schellhorn:2005:ARG

[161] Gerhard Schellhorn. ASM refinement and generalizations of forward simulation in data refinement: a comparison. Theoretical Computer Science, 336(2– 3):403–435, May 26, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:AIg

[162] Anonymous. Author index. Theoretical Computer Science, 336(2–3):437– 438, May 26, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Giacobazzi:2005:TSA

[163] Roberto Giacobazzi and Isabella Mastroeni. Transforming semantics by abstract interpretation. Theoretical Computer Science, 337(1–3):1–50, June 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Costagliola:2005:ASI

[164] Gennaro Costagliola, Filomena Ferrucci, and Carmine Gravino. Adding symbolic information to picture models: definitions and properties. Theoretical Computer Science, 337(1–3): 51–104, June 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karhumaki:2005:CDC

[165] Juhani Karhumäki, Wojciech Plandowski, and Wojciech Rytter. On the complexity of decidable cases of the commutation problem of languages. Theoretical Computer Science, 337(1–3): 105–118, June 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

McQuillan:2005:GCB

[166] I. McQuillan. The generative capacity of block-synchronized context-free grammars. Theoretical Computer Science, 337(1–3):119–133, June 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheng:2005:SLM

[167] T. C. Edwin Cheng, Hans Kellerer, and Vladimir Kotov. Semi-on-line multiprocessor scheduling with given total processing time. Theoretical Computer Science, 337(1–3):134–146, June 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flocchini:2005:GAR

[168] Paola Flocchini, Giuseppe Prencipe, Nicola Santoro, and Peter Widmayer. Gathering of asynchronous robots with limited visibility. Theoretical Computer Science, 337(1–3):147–168, June 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chuan:2005:FCW

[169] Wai-Fong Chuan. Factors of characteristic words of irrational numbers. Theoretical Computer Science, 337(1–3): 169–182, June 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moriya:2005:SAC

[170] Etsuro Moriya, Dieter Hofbauer, Maria Huber, and Friedrich Otto. On statealternating context-free grammars. Theoretical Computer Science, 337(1– 3):183–216, June 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bille:2005:STE

[171] Philip Bille. A survey on tree edit distance and related problems. Theoretical Computer Science, 337(1–3): 217–239, June 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dias:2005:GBG

[172] Vânia M. F. Dias, Celina M. H. de Figueiredo, and Jayme L. Szwarcfiter. Generating bicliques of a graph in lexicographic order. Theoretical Computer Science, 337(1–3):240–248, June 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elkin:2005:ASP

[173] Michael Elkin and David Peleg. Approximating k-spanner problems for k > 2. Theoretical Computer Science, 337(1–3):249–277, June 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guelev:2005:CDD

[174] Dimitar P. Guelev and Dang Van Hung. On the completeness and decidability of duration calculus with iteration. Theoretical Computer Science, 337(1–3):278–304, June 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2005:AMV

[175] Jianer Chen and Iyad A. Kanj. On approximating minimum vertex cover for graphs with perfect matching. Theoretical Computer Science, 337(1–3): 305–318, June 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aigner:2005:PPT

[176] Martin Aigner, Gianluca De Marco, and Manuela Montangero. The plurality problem with three colors and more. Theoretical Computer Science, 337(1– 3):319–330, June 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manthey:2005:ICH

[177] Bodo Manthey and Rüdiger Reischuk. The intractability of computing the Hamming distance. Theoretical Computer Science, 337(1–3):331–346, June 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoory:2005:CUS

[178] Shlomo Hoory and Stefan Szeider. Computing unsatisfiable k-SAT instances with few occurrences per variable. Theoretical Computer Science, 337(1–3):347–359, June 9, 2005. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chou:2005:CCD

[179] Arthur W. Chou and Ker-I Ko. The computational complexity of distance functions of two-dimensional domains. Theoretical Computer Science, 337(1– 3):360–369, June 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hsieh:2005:ELF

[180] Sun-Yuan Hsieh. Embedding longest fault-free paths onto star graphs with more vertex faults. Theoretical Computer Science, 337(1–3):370–378, June 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bruin:2005:CFL

[181] H. Bruin and O. Volkova. The complexity of Fibonacci-like kneading sequences. Theoretical Computer Science, 337(1–3):379–389, June 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huang:2005:AWE

[182] Li-Sha Huang, Minming Li, and Bo Zhang. Approximation of Walrasian equilibrium in single-minded auctions. Theoretical Computer Science, 337(1– 3):390–398, June 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:AIh

[183] Anonymous. Author index. Theoretical Computer Science, 337(1–3): 399–400, June 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBh

[184] Anonymous. Editorial board. Theoretical Computer Science, 337(1–3): v–ix, June 9, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Borzyszkowski:2005:TST

[185] Andrzej M. Borzyszkowski and Philippe Darondeau. Transition systems without transitions. Theoretical Computer Science, 338(1–3):1–16, June 10, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jeffrey:2005:FAM

[186] Alan Jeffrey and Julian Rathke. A fully abstract may testing semantics for concurrent objects. Theoretical Computer Science, 338(1–3):17–63, June 10, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DAgostino:2005:ABQ

[187] Giovanna D’Agostino and Giacomo Lenzi. An axiomatization of bisimulation quantifiers via the μ-calculus. Theoretical Computer Science, 338(1– 3):64–95, June 10, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fu:2005:QOB

[188] Yuxi Fu. On quasi-open bisimulation. Theoretical Computer Science, 338(1– 3):96–126, June 10, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sakama:2005:ODT

[189] Chiaki Sakama. Ordering default theories and nonmonotonic logic programs. Theoretical Computer Science, 338(1– 3):127–152, June 10, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flanagan:2005:MVM

[190] Cormac Flanagan, Stephen N. Freund, Shaz Qadeer, and Sanjit A. Seshia. Modular verification of multithreaded programs. Theoretical Computer Science, 338(1–3):153–183, June 10, 2005.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Worrell:2005:FSF

[191] James Worrell. On the final sequence of a finitary set functor. Theoretical Computer Science, 338(1–3):184– 199, June 10, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hernest:2005:CAF

[192] Mircea-Dan Hernest and Ulrich Kohlenbach. A complexity analysis of functional interpretations. Theoretical Computer Science, 338(1–3):200–246, June 10, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chevalier:2005:NDP

[193] Yannick Chevalier, Ralf Küsters, Michaël Rusinowitch, and Mathieu Turuani. An NP decision procedure for protocol insecurity with XOR. Theoretical Computer Science, 338(1–3): 247–274, June 10, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fritz:2005:SRA

[194] Carsten Fritz and Thomas Wilke. Simulation relations for alternating Büchi automata. Theoretical Computer Science, 338(1–3):275–314, June 10, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Murawski:2005:FLS

[195] Andrzej S. Murawski. Functions with local state: Regularity and undecidability. Theoretical Computer Science, 338(1–3):315–349, June 10, 2005.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bueno:2005:TTC

[196] Antonio Bueno, Valentín Valero, and Fernando Cuartero. A translation of TPAL_{p} into a class of timedprobabilistic Petri nets. Theoretical Computer Science, 338(1–3):350– 392, June 10, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boreale:2005:MSA

[197] Michele Boreale and Maria Grazia Buscemi. A method for symbolic analysis of security protocols. Theoretical Computer Science, 338(1–3):393– 425, June 10, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:AIi

[198] Anonymous. Author index. Theoretical Computer Science, 338(1–3): 426, June 10, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBi

[199] Anonymous. Editorial board. Theoretical Computer Science, 338(1–3): v–ix, June 10, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harju:2005:P

[200] T. Harju, J. Karhumäki, and A. Restivo. Preface. Theoretical Computer Science, 339(1):1–2, June 11, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adian:2005:DPO

[201] S. I. Adian. Divisibility problem for one relator monoids. Theoretical Computer Science, 339(1):3–6, June 11, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Currie:2005:PAT

[202] James D. Currie. Pattern avoidance: themes and variations. Theoretical Computer Science, 339(1):7– 18, June 11, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rampersad:2005:ALS

[203] Narad Rampersad, Jeffrey Shallit, and Ming wei Wang. Avoiding large squares in infinite binary words. Theoretical Computer Science, 339(1):19– 34, June 11, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gasieniec:2005:SES

[204] Leszek Gaşieniec, Roman Kolpakov, and Igor Potapov. Space efficient search for maximal repetitions. Theoretical Computer Science, 339(1):35– 48, June 11, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Constantinescu:2005:GFW

[205] Sorin Constantinescu and Lucian Ilie. Generalised fine and Wilf’s theorem for arbitrary number of periods. Theoretical Computer Science, 339(1):49– 60, June 11, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Holub:2005:PED

[206] Štěpán Holub. A proof of the extended Duval’s conjecture. Theoretical Computer Science, 339(1):61–67, June 11, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frid:2005:SLA

[207] A. E. Frid. Sequences of linear arithmetical complexity. Theoretical Computer Science, 339(1):68–87, June 11, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Araujo:2005:SWC

[208] Isabel M. Araújo and Véronique Bruyère. Sturmian words and a criterium by Michaux–Villemaire. Theoretical Computer Science, 339(1):88– 102, June 11, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leve:2005:CAF

[209] F. Levé and G. Richomme. On a conjecture about finite fixed points of morphisms. Theoretical Computer Science, 339(1):103–128, June 11, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hivert:2005:ABS

[210] F. Hivert, J.-C. Novelli, and J.-Y. Thibon. The algebra of binary search trees. Theoretical Computer Science, 339(1): 129–165, June 11, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBj

[211] Anonymous. Editorial board. Theoretical Computer Science, 339(1):v– ix, June 11, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2005:MBM

[212] Yijia Chen, Jörg Flum, and Martin Grohe. Machine-based methods in parameterized complexity theory. Theoretical Computer Science, 339(2–3): 167–199, June 12, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maletti:2005:HDC

[213] Andreas Maletti. HASSE diagrams for classes of deterministic bottom-up treeto-tree-series transformations. Theoretical Computer Science, 339(2–3):200– 240, June 12, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2005:PAL

[214] Shengyu Zhang. On the power of Ambainis lower bounds. Theoretical Computer Science, 339(2–3):241– 256, June 12, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lin:2005:SCP

[215] Cheng-Kuan Lin, Hua-Min Huang, and Lih-Hsing Hsu. The super connectivity of the pancake graphs and the super laceability of the star graphs. Theoretical Computer Science, 339(2–3): 257–271, June 12, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bazgan:2005:CSD

[216] Cristina Bazgan, Bruno Escoffier, and Vangelis Th. Paschos. Completeness in standard and differential approximation classes: Poly-(D)APX-and (D)PTAS-completeness. Theoretical Computer Science, 339(2–3):272– 292, June 12, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

He:2005:OSO

[217] Yong He and Yiwei Jiang. Optimal semi-online preemptive algorithms for machine covering on two uniform machines. Theoretical Computer Science, 339(2–3):293–314, June 12, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bagchi:2005:SLM

[218] Amitabha Bagchi, Amitabh Chaudhary, and Petr Kolman. Short length Menger’s theorem and reliable optical routing. Theoretical Computer Science, 339(2–3):315–332, June 12, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Damaschke:2005:QLL

[219] Peter Damaschke and Zhen Zhou. On queuing lengths in on-line switching. Theoretical Computer Science, 339(2– 3):333–343, June 12, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elder:2005:CFC

[220] Murray Elder. A context-free and a 1-counter geodesic language for a Baumslag–Solitar group. Theoretical Computer Science, 339(2–3):344– 371, June 12, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:AIj

[221] Anonymous. Author index. Theoretical Computer Science, 339(2–3):372– 373, June 12, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gorrieri:2005:TFS

[222] Roberto Gorrieri and Fabio Martinelli. Theoretical foundations of security analysis and design II. Theoretical Computer Science, 340(1):1– 2, June 13, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pierro:2005:MCP

[223] Alessandra Di Pierro, Chris Hankin, and Herbert Wiklicky. Measuring the confinement of probabilistic systems. Theoretical Computer Science, 340(1): 3–56, June 13, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Herzog:2005:CID

[224] Jonathan Herzog. A computational interpretation of Dolev–Yao adversaries. Theoretical Computer Science, 340(1): 57–81, June 13, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rohrmair:2005:UDI

[225] Gordon Thomas Rohrmair and Gavin Lowe. Using data-independence in the analysis of intrusion detection systems. Theoretical Computer Science, 340(1): 82–101, June 13, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bhargavan:2005:SWS

[226] Karthikeyan Bhargavan, Cédric Fournet, and Andrew D. Gordon. A semantics for web services authentication. Theoretical Computer Science, 340(1): 102–153, June 13, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zunino:2005:WPE

[227] Roberto Zunino and Pierpaolo Degano. Weakening the perfect encryption assumption in Dolev–Yao adversaries. Theoretical Computer Science, 340(1): 154–178, June 13, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBk

[228] Anonymous. Editorial board. Theoretical Computer Science, 340(1):v– ix, June 13, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deLuca:2005:P

[229] A. de Luca, F. Mignosi, D. Perrin, and G. Rozenberg. Preface. Theoretical Computer Science, 340(2):179– 185, June 27, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EAF

[230] Anonymous. Editorial: From Antonio’s former students. Theoretical Computer Science, 340(2):186–187, June 27, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salomaa:2005:CBS

[231] Arto Salomaa. Connections between subwords and certain matrix mappings. Theoretical Computer Science, 340(2): 188–203, June 27, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Araujo:2005:WDS

[232] Isabel M. Araújo and Véronique Bruyère. Words derivated from Sturmian words. Theoretical Computer Science, 340(2):204–219, June 27, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carpi:2005:CCS

[233] Arturo Carpi and Aldo de Luca. Codes of central Sturmian words. Theoretical Computer Science, 340(2):220– 239, June 27, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Felice:2005:EPF

[234] Clelia De Felice. An enhanced property of factorizing codes. Theoretical Computer Science, 340(2):240–256, June 27, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reghizzi:2005:TRG

[235] Stefano Crespi Reghizzi and Matteo Pradella. Tile rewriting grammars and picture languages. Theoretical Computer Science, 340(2):257–272, June 27, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harju:2005:CBP

[236] Tero Harju and Dirk Nowotka. Counting bordered and primitive words with a fixed weight. Theoretical Computer Science, 340(2):273–279, June 27, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berstel:2005:GRF

[237] Jean Berstel. Growth of repetition-free words — a review. Theoretical Computer Science, 340(2):280–290, June 27, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esik:2005:ARR

[238] Zoltán Ésik and Pascal Weil. Algebraic recognizability of regular tree languages. Theoretical Computer Science, 340(2):291–321, June 27, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karhumaki:2005:CC

[239] Juhani Karhumäki, Michel Latteux, and Ion Petre. Commutation with codes. Theoretical Computer Science, 340(2):322–333, June 27, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Borel:2005:PFB

[240] J.-P. Borel and C. Reutenauer. Palindromic factors of billiard words. Theoretical Computer Science, 340(2):334– 348, June 27, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonizzoni:2005:RSL

[241] Paola Bonizzoni and Giancarlo Mauri. Regular splicing languages and subclasses. Theoretical Computer Science, 340(2):349–363, June 27, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choffrut:2005:CTD

[242] Christian Choffrut and Berke Durak. Collage of two-dimensional words. Theoretical Computer Science, 340(2):364– 380, June 27, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beal:2005:CSC

[243] Marie-Pierre Béal and Dominique Perrin. Codes and sofic constraints. Theoretical Computer Science, 340(2):381– 393, June 27, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bertoni:2005:SSQ

[244] Alberto Bertoni, Carlo Mereghetti, and Beatrice Palano. Small size quantum automata recognizing some regular languages. Theoretical Computer Science, 340(2):394–407, June 27, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anselmo:2005:NOR

[245] Marcella Anselmo, Dora Giammarresi, and Maria Madonia. New operations and regular expressions for twodimensional languages over one-letter alphabet. Theoretical Computer Science, 340(2):408–431, June 27, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Clement:2005:PFD

[246] Julien Clément, Jean-Pierre Duval, Giovanna Guaiana, Dominique Perrin, and Giuseppina Rindone. Parsing with a finite dictionary. Theoretical Computer Science, 340(2):432–442, June 27, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pin:2005:TAT

[247] Jean-Éric Pin and Pedro V. Silva. A topological approach to transductions. Theoretical Computer Science, 340(2): 443–456, June 27, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Auinger:2005:CDP

[248] K. Auinger and B. Steinberg. Constructing divisions into power groups. Theoretical Computer Science, 341(1– 3):1–21, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chlamtac:2005:IAM

[249] Eden Chlamtac and Uriel Feige. Improved approximation of the minimum cover time. Theoretical Computer Science, 341(1–3):22–38, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chang:2005:OPE

[250] Yan-Cheng Chang and Chi-Jen Lu. Oblivious polynomial evaluation and oblivious neural learning. Theoretical Computer Science, 341(1–3):39–54, September 5, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heuberger:2005:AGE

[251] Clemens Heuberger, Rajendra Katti, Helmut Prodinger, and Xiaoyu Ruan. The alternating greedy expansion and applications to computing digit expansions from left-to-right in cryptography. Theoretical Computer Science, 341(1– 3):55–72, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sawada:2005:OVE

[252] J. Sawada. Oracles for vertex elimination orderings. Theoretical Computer Science, 341(1–3):73–90, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gacs:2005:UTA

[253] Peter Gács. Uniform test of algorithmic randomness over a general space. Theoretical Computer Science, 341(1–3):91– 137, September 5, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Inoue:2005:SPO

[254] Atsuyuki Inoue, Akira Ito, Katsushi Inoue, and Tokio Okazaki. Some properties of one-pebble Turing machines with sublogarithmic space. Theoretical Computer Science, 341(1–3):138– 149, September 5, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheng:2005:IMB

[255] Xiaochun Cheng, Dantong Ouyang, Jiang Yunfei, and Chengqi Zhang. An improved model-based method to test circuit faults. Theoretical Computer Science, 341(1–3):150–161, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Collins:2005:CCR

[256] Pieter Collins. Continuity and computability of reachable sets. Theoretical Computer Science, 341(1–3):162– 195, September 5, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lai:2005:SRN

[257] Cheng-Nan Lai and Gen-Huey Chen. Strong Rabin numbers of folded hypercubes. Theoretical Computer Science, 341(1–3):196–215, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schafer:2005:TMS

[258] Guido Schäfer and Naveen Sivadasan. Topology matters: Smoothed competitiveness of metrical task systems. Theoretical Computer Science, 341(1–3): 216–246, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cases:2005:TPD

[259] Blanca Cases and Manuel Alfonseca. Towards a proof of the decidability of the momentary stagnation of the growth function of D. Theoretical Computer Science, 341(1–3):247–262, September 5, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

terBeek:2005:SS

[260] Maurice H. ter Beek, Carlos Martín-Vide, and Victor Mitrana. Synchronized shuffles. Theoretical Computer Science, 341(1–3):263–275, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pirillo:2005:ICS

[261] Giuseppe Pirillo. Inequalities characterizing standard Sturmian and episturmian words. Theoretical Computer Science, 341(1–3):276–292, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berthe:2005:SWA

[262] Valérie Berthé, Srečko Brlek, and Philippe Choquette. Smooth words over arbitrary alphabets. Theoretical Computer Science, 341(1–3):293– 310, September 5, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maitra:2005:RMP

[263] Subhamoy Maitra, Kishan Chand Gupta, and Ayineedi Venkateswarlu. Results on multiples of primitive polynomials and their products over GF(2). Theoretical Computer Science, 341 (1–3):311–343, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dang:2005:CLD

[264] Zhe Dang, Oscar H. Ibarra, and Jianwen Su. On composition and lookahead delegation of e-services modeled by automata. Theoretical Computer Science, 341(1–3):344–363, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nagamochi:2005:RAB

[265] Hiroshi Nagamochi, Kengo Iwata, and Toshimasa Ishii. A robust algorithm for bisecting a triconnected graph with two resource sets. Theoretical Computer Science, 341(1–3):364–378, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tronicek:2005:SSA

[266] Zdeněk Troníček and Ayumi Shinohara. The size of subsequence automaton. Theoretical Computer Science, 341(1–3):379–384, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wu:2005:AWH

[267] Quincy Wu, Chin Lung Lu, and Richard Chia-Tung Lee. The approximability of the weighted Hamiltonian path completion problem on a tree. Theoretical Computer Science, 341(1–3):385–397, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2005:PAA

[268] Wei-Mei Chen. Probabilistic analysis of algorithms for the Dutch national flag problem. Theoretical Computer Science, 341(1–3):398–410, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hung:2005:LTA

[269] Ruo-Wei Hung and Maw-Shang Chang. Linear-time algorithms for the Hamiltonian problems on distance-hereditary graphs. Theoretical Computer Science, 341(1–3):411–440, September 5, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBl

[270] Anonymous. Editorial board. Theoretical Computer Science, 341(1–3):v– ix, September 5, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moggi:2005:ASS

[271] Eugenio Moggi. Applied semantics: Selected topics. Theoretical Computer Science, 342(1):1–2, September 6, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abbott:2005:CCS

[272] Michael Abbott, Thorsten Altenkirch, and Neil Ghani. Containers: Constructing strictly positive types. Theoretical Computer Science, 342(1):3– 27, September 6, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shinwell:2005:MSF

[273] Mark R. Shinwell and Andrew M. Pitts. On a monadic semantics for freshness. Theoretical Computer Science, 342(1): 28–55, September 6, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cachera:2005:EDF

[274] David Cachera, Thomas Jensen, David Pichardie, and Vlad Rusu. Extracting a data flow analyser in constructive logic. Theoretical Computer Science, 342(1):56–78, September 6, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alves:2005:WLL

[275] Sandra Alves and Mário Florido. Weak linearization of the lambda calculus. Theoretical Computer Science, 342(1): 79–103, September 6, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gardner:2005:MDW

[276] Philippa Gardner and Sergio Maffeis. Modelling dynamic web data. Theoretical Computer Science, 342(1):104– 131, September 6, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laird:2005:LBD

[277] J. Laird. Locally Boolean domains. Theoretical Computer Science, 342(1): 132–148, September 6, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ager:2005:FCB

[278] Mads Sig Ager, Olivier Danvy, and Jan Midtgaard. A functional correspondence between monadic evaluators and abstract machines for languages with computational effects. Theoretical Computer Science, 342(1):149–172, September 6, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBm

[279] Anonymous. Editorial board. Theoretical Computer Science, 342(1):v– ix, September 6, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Courcelle:2005:RSG

[280] Bruno Courcelle and Pascal Weil. The recognizability of sets of graphs is a robust property. Theoretical Computer Science, 342(2–3):173–228, September 7, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Haghverdi:2005:BRD

[281] Esfandiar Haghverdi, Paulo Tabuada, and George J. Pappas. Bisimulation relations for dynamical, control, and hybrid systems. Theoretical Computer Science, 342(2–3):229–261, September 7, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Oles:2005:PIP

[282] Frank J. Oles. Precedence–inclusion patterns and relational learning. Theoretical Computer Science, 342(2–3): 262–315, September 7, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cleaveland:2005:PTL

[283] Rance Cleaveland, S. Purushothaman Iyer, and Murali Narasimha. Probabilistic temporal logics via the modal mu-calculus. Theoretical Computer Science, 342(2–3):316–350, September 7, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Loyer:2005:WAL

[284] Yann Loyer and Umberto Straccia. Any-world assumptions in logic programming. Theoretical Computer Science, 342(2–3):351–381, September 7, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nomikos:2005:TST

[285] Christos Nomikos, Panos Rondogiannis, and Manolis Gergatsoulis. Temporal stratification tests for linear and branching-time deductive databases. Theoretical Computer Science, 342(2– 3):382–415, September 7, 2005. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abramsky:2005:GTM

[286] Samson Abramsky and Marios Mavronicolas. Game theory meets theoretical computer science. Theoretical Computer Science, 343(1–2):1–3, October 10, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lopez:2005:MMT

[287] Luis López, Gemma del Rey Almansa, Stéphane Paquelet, and Antonio Fernández. A mathematical model for the TCP Tragedy of the Commons. Theoretical Computer Science, 343(1– 2):4–26, October 10, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ambuhl:2005:ARA

[288] Christoph Ambühl, Andrea E. F. Clementi, Paolo Penna, Gianluca Rossi, and Riccardo Silvestri. On the approximability of the range assignment problem on radio networks in presence of selfish agents. Theoretical Computer Science, 343(1–2):27– 41, October 10, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aichholzer:2005:GT

[289] Oswin Aichholzer, David Bremner, Erik D. Demaine, Ferran Hurtado, Evangelos Kranakis, Hannes Krasser, Suneeta Ramaswami, Saurabh Sethia, and Jorge Urrutia. Games on triangulations. Theoretical Computer Science, 343(1–2):42–71, October 10, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hearn:2005:PCS

[290] Robert A. Hearn and Erik D. Demaine. PSPACE-completeness of sliding-block puzzles and other problems through the nondeterministic constraint logic model of computation. Theoretical Computer Science, 343(1–2):72– 96, October 10, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shoham:2005:NCC

[291] Yoav Shoham and Moshe Tennenholtz. Non-cooperative computation: Boolean functions with correctness and exclusivity. Theoretical Computer Science, 343(1–2):97–113, October 10, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Feigenbaum:2005:CDI

[292] Joan Feigenbaum, Lance Fortnow, David M. Pennock, and Rahul Sami. Computation in a distributed information market. Theoretical Computer Science, 343(1–2):114–132, October 10, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gairing:2005:SCE

[293] M. Gairing, T. Lücking, M. Mavronicolas, B. Monien, and P. Spirakis. Structure and complexity of extreme Nash equilibria. Theoretical Computer Science, 343(1–2):133–157, October 10, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Verbitsky:2005:FOD

[294] Oleg Verbitsky. The first order definability of graphs with separators via the Ehrenfeucht game. Theoretical Computer Science, 343(1–2):158– 176, October 10, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laurent:2005:SVS

[295] Olivier Laurent. Syntax vs. semantics: a polarized approach. Theoretical Computer Science, 343(1–2):177– 206, October 10, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Murawski:2005:GCS

[296] Andrzej S. Murawski. Games for complexity of second-order call-by-name programs. Theoretical Computer Science, 343(1–2):207–236, October 10, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mellies:2005:SAS

[297] Paul-André Melliès. Sequential algorithms and strongly stable functions. Theoretical Computer Science, 343(1– 2):237–281, October 10, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBn

[298] Anonymous. Editorial board. Theoretical Computer Science, 343(1–2):v– ix, October 10, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deBoer:2005:Pb

[299] Frank de Boer and Marcello Bonsangue. Preface. Theoretical Computer Science, 343(3):283–284, October 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bergstra:2005:PPA

[300] J. A. Bergstra and I. Bethke. Polarized process algebra with reactive composition. Theoretical Computer Science, 343(3):285–304, October 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaconescu:2005:BSH

[301] Răzvan Diaconescu. Behavioural specification for hierarchical object composition. Theoretical Computer Science, 343(3):305–331, October 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Groote:2005:PBE

[302] Jan Friso Groote and Tim A. C. Willemse. Parameterised Boolean equation systems. Theoretical Computer Science, 343(3):332–369, October 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gurevich:2005:SEA

[303] Yuri Gurevich, Benjamin Rossman, and Wolfram Schulte. Semantic essence of AsmL. Theoretical Computer Science, 343(3):370–412, October 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pierik:2005:POL

[304] Cees Pierik and Frank S. de Boer. A proof outline logic for object-oriented programming. Theoretical Computer Science, 343(3):413–442, October 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rutten:2005:TCS

[305] J. J. M. M. Rutten. A tutorial on coinductive stream calculus and signal flow graphs. Theoretical Computer Science, 343(3):443–481, October 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stark:2005:FSV

[306] Robert F. Stärk. Formal specification and verification of the C# thread model. Theoretical Computer Science, 343(3):482–508, October 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wehrheim:2005:STV

[307] Heike Wehrheim. Slicing techniques for verification re-use. Theoretical Computer Science, 343(3):509–528, October 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nikoletseas:2005:P

[308] Sotiris Nikoletseas and Jose Rolim. Preface. Theoretical Computer Science, 344(1):1–2, November 11, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Papadimitriou:2005:CRG

[309] Christos H. Papadimitriou and David Ratajczak. On a conjecture related to geometric routing. Theoretical Computer Science, 344(1):3–14, November 11, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Enachescu:2005:SFA

[310] Mihaela Enachescu, Ashish Goel, Ramesh Govindan, and Rajeev Motwani. Scale-free aggregation in sensor networks. Theoretical Computer Science, 344(1):15–29, November 11, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Floreen:2005:EAB

[311] Patrik Floréen, Petteri Kaski, Jukka Kohonen, and Pekka Orponen. Exact and approximate balanced data gathering in energy-constrained sensor networks. Theoretical Computer Science, 344(1):30–46, November 11, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

ODell:2005:TAC

[312] R. O’Dell and R. Wattenhofer. Theoretical aspects of connectivity-based multi-hop positioning. Theoretical Computer Science, 344(1):47–68, November 11, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leone:2005:TDM

[313] Pierre Leone and José Rolim. Towards a dynamical model for wireless sensor networks. Theoretical Computer Science, 344(1):69–85, November 11, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stupp:2005:EUR

[314] Gideon Stupp and Moshe Sidi. The expected uncertainty of range-free localization protocols in sensor networks. Theoretical Computer Science, 344(1): 86–99, November 11, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:Pa

[315] Anonymous. Preface. Theoretical Computer Science, 344(1):v–ix, November 11, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2005:Pb

[316] G. Rozenberg. Preface. Theoretical Computer Science, 344(2–3):101– 102, November 17, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lu:2005:AES

[317] Hongtao Lu, Ruiming Shen, and Fu-Lai Chung. Absolute exponential stability of a class of recurrent neural networks with multiple and variable delays. Theoretical Computer Science, 344(2–3):103–119, November 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2005:DVN

[318] Oscar H. Ibarra. On determinism versus nondeterminism in P systems. Theoretical Computer Science, 344(2–3): 120–133, November 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ying:2005:TCB

[319] Mingsheng Ying. A theory of computation based on quantum logic (I). Theoretical Computer Science, 344(2–3): 134–207, November 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fischer:2005:ODI

[320] Simon Fischer and Ingo Wegener. The one-dimensional Ising model: Mutation versus recombination. Theoretical Computer Science, 344(2–3):208– 225, November 17, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Verlan:2005:BRE

[321] Sergey Verlan. A boundary result on enhanced time-varying distributed H systems with parallel computations. Theoretical Computer Science, 344(2–3):226–242, November 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dorigo:2005:ACO

[322] Marco Dorigo and Christian Blum. Ant colony optimization theory: a survey. Theoretical Computer Science, 344(2–3):243–278, November 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Soloveichik:2005:CPB

[323] David Soloveichik and Erik Winfree. The computational power of Benenson automata. Theoretical Computer Science, 344(2–3):279–297, November 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [2396].

Goulon-Sigwalt-Abram:2005:HNR

[324] Aurélie Goulon-Sigwalt-Abram, Arthur Duprat, and Gérard Dreyfus. From Hopfield nets to recursive networks to graph machines: Numerical machine learning for structured data. Theoretical Computer Science, 344(2–3):298– 334, November 17, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shi:2005:QCT

[325] Yaoyun Shi. Quantum and classical tradeoffs. Theoretical Computer Science, 344(2–3):335–345, November 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jensen:2005:SI

[326] Kurt Jensen and Andreas Podelski. Special issue. Theoretical Computer Science, 345(1):1, November 21, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baier:2005:ECT

[327] Christel Baier, Holger Hermanns, Joost-Pieter Katoen, and Boudewijn R. Haverkort. Efficient computation of time-bounded reachability probabilities in uniform continuous-time Markov decision processes. Theoretical Computer Science, 345(1):2–26, November 21, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lugiez:2005:POS

[328] D. Lugiez, P. Niebert, and S. Zennou. A partial order semantics approach to the clock explosion problem of timed automata. Theoretical Computer Science, 345(1):27–59, November 21, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Geldenhuys:2005:MEF

[329] Jaco Geldenhuys and Antti Valmari. More efficient on-the-fly LTL verification with Tarjan’s algorithm. Theoretical Computer Science, 345(1):60– 82, November 21, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kupferman:2005:CC

[330] Orna Kupferman and Moshe Y. Vardi. From complementation to certification. Theoretical Computer Science, 345(1): 83–100, November 21, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

McMillan:2005:ITP

[331] K. L. McMillan. An interpolating theorem prover. Theoretical Computer Science, 345(1):101–121, November 21, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Su:2005:CPS

[332] Zhendong Su and David Wagner. A class of polynomially solvable range constraints for interval analysis without widenings. Theoretical Computer Science, 345(1):122–138, November 21, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deAlfaro:2005:MCD

[333] Luca de Alfaro, Marco Faella, Thomas A. Henzinger, Rupak Majumdar, and Mariëlle Stoelinga. Model checking discounted temporal properties. Theoretical Computer Science, 345(1):139–170, November 21, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:SI

[334] Anonymous. Special issue. Theoretical Computer Science, 345(1):v–ix, November 21, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koubek:2005:P

[335] V. Koubek and J. Kratochvíl. Preface. Theoretical Computer Science, 345(2–3):171–172, November 22, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bab:2005:OQR

[336] Sebastian Bab and Arfst Nickelsen. One query reducibilities between partial information classes. Theoretical Computer Science, 345(2–3):173–189, November 22, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beal:2005:HSE

[337] Marie-Pierre Béal, Francesca Fiorenzi, and Dominique Perrin. A hierarchy of shift equivalent sofic shifts. Theoretical Computer Science, 345(2–3):190– 205, November 22, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beaudry:2005:CAD

[338] Martin Beaudry, José M. Fernandez, and Markus Holzer. A common algebraic description for probabilistic and quantum computations. Theoretical Computer Science, 345(2–3):206–234, November 22, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bernardi:2005:NDS

[339] Vincent Bernardi, Bruno Durand, Enrico Formenti, and Jarkko Kari. A new dimension sensitive property for cellular automata. Theoretical Computer Science, 345(2–3):235–247, November 22, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beygelzimer:2005:EPC

[340] Alina Beygelzimer and Mitsunori Ogihara. The enumerability of P collapses P to NC. Theoretical Computer Science, 345(2–3):248–259, November 22, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bohler:2005:GP

[341] E. Böhler, C. Glaßer, B. Schwarz, and K. W. Wagner. Generation problems. Theoretical Computer Science, 345(2–3):260–295, November 22, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czeizler:2005:NPW

[342] Elena Czeizler. The non-parametrizability of the word equation xyz. Theoretical Computer Science, 345(2–3):296– 303, November 22, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Domaratzki:2005:DTB

[343] Michael Domaratzki and Kai Salomaa. Decidability of trajectory-based equations. Theoretical Computer Science, 345(2–3):304–330, November 22, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fraigniaud:2005:GEF

[344] Pierre Fraigniaud, David Ilcinkas, Guy Peer, Andrzej Pelc, and David Peleg. Graph exploration by a finite automaton. Theoretical Computer Science, 345(2–3):331–344, November 22, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hemaspaandra:2005:ASI

[345] Edith Hemaspaandra, Lane A. Hemaspaandra, and Harald Hempel. All superlinear inverse schemes are coNPhard. Theoretical Computer Science, 345(2–3):345–358, November 22, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ilie:2005:GRT

[346] Lucian Ilie, Pascal Ochem, and Jeffrey Shallit. A generalization of repetition threshold. Theoretical Computer Science, 345(2–3):359–369, November 22, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kawachi:2005:UTQ

[347] Akinori Kawachi, Hirotada Kobayashi, Takeshi Koshiba, and Raymond H. Putra. Universal test for quantum one-way permutations. Theoretical Computer Science, 345(2–3):370–385, November 22, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lee:2005:RBS

[348] Troy Lee and Andrei Romashchenko. Resource bounded symmetry of information revisited. Theoretical Computer Science, 345(2–3):386–405, November 22, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nordh:2005:CEI

[349] Gustav Nordh. The complexity of equivalence and isomorphism of systems of equations over finite groups. Theoretical Computer Science, 345(2– 3):406–424, November 22, 2005. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okhotin:2005:DC

[350] Alexander Okhotin. The dual of concatenation. Theoretical Computer Science, 345(2–3):425–447, November 22, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Selivanov:2005:RDD

[351] Victor L. Selivanov and Klaus W. Wagner. A reducibility for the dot-depth hierarchy. Theoretical Computer Science, 345(2–3):448–472, November 22, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wich:2005:SA

[352] Klaus Wich. Sublogarithmic ambiguity. Theoretical Computer Science, 345(2–3):473–504, November 22, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cerone:2005:P

[353] Antonio Cerone and Alessandra Di Pierro. Preface. Theoretical Computer Science, 346(1):1–2, November 23, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lermer:2005:TET

[354] Karl Lermer, Colin J. Fidge, and Ian J. Hayes. A theory for execution-time derivation in real-time programs. Theoretical Computer Science, 346(1):3– 27, November 23, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bravetti:2005:QIT

[355] Mario Bravetti, Roberto Gorrieri, Roberto Lucchi, and Gianluigi Zavattaro. Quantitative information in the tuple space coordination model. Theoretical Computer Science, 346(1):28– 57, November 23, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alpuente:2005:SFA

[356] María Alpuente, María del Mar Gallardo, Ernesto Pimentel, and Alicia Villanueva. A semantic framework for the abstract model checking of tccp. Theoretical Computer Science, 346(1): 58–95, November 23, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hurd:2005:PGC

[357] Joe Hurd, Annabelle McIver, and Carroll Morgan. Probabilistic guarded commands mechanized in HOL. Theoretical Computer Science, 346(1):96– 112, November 23, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huth:2005:FSA

[358] Michael Huth. On finite-state approximants for probabilistic computation tree logic. Theoretical Computer Science, 346(1):113–134, November 23, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lluch-Lafuente:2005:QCC

[359] Alberto Lluch-Lafuente and Ugo Montanari. Quantitative μ-calculus and CTL defined over constraint semirings. Theoretical Computer Science, 346(1): 135–160, November 23, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harrison:2005:SES

[360] P. G. Harrison and T. T. Lee. Separable equilibrium state probabilities via time reversal in Markovian process algebra. Theoretical Computer Science, 346(1):161–182, November 23, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:Pb

[361] Anonymous. Preface. Theoretical Computer Science, 346(1):v–ix, November 23, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pergola:2005:P

[362] Elisa Pergola and Simone Rinaldi. Preface. Theoretical Computer Science, 346(2–3):183, November 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pergola:2005:MAD

[363] E. Pergola and S. Rinaldi. In memoriam: Alberto Del Lungo (1965– 2003). Theoretical Computer Science, 346(2–3):184–188, November 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baril:2005:MCL

[364] Jean-Luc Baril and Vincent Vajnovszki. Minimal change list for Lucas strings and some graph theoretic consequences. Theoretical Computer Science, 346(2–3):189–199, November 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brlek:2005:DGT

[365] S. Brlek, G. Labelle, and A. Lacasse. The discrete Green Theorem and some applications in discrete geometry. Theoretical Computer Science, 346(2–3): 200–225, November 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chavanon:2005:GED

[366] Frédéric Chavanon, Matthieu Latapy, Michel Morvan, Eric Rémila, and Laurent Vuillon. Graph encoding of 2 D. Theoretical Computer Science, 346(2–3):226–253, November 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deMier:2005:STB

[367] Anna de Mier and Marc Noy. A solution to the tennis ball problem. Theoretical Computer Science, 346(2–3): 254–264, November 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epifanio:2005:MCF

[368] Chiara Epifanio and Filippo Mignosi. A multidimensional critical factorization theorem. Theoretical Computer Science, 346(2–3):265–280, November 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garduno:2005:ISV

[369] Edgar Garduño and Gabor T. Herman. Implicit surface visualization of reconstructed biological molecules. Theoretical Computer Science, 346(2–3):281– 299, November 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gerard:2005:RTD

[370] Y. Gerard. Reduction from threedimensional discrete tomography to multicommodity flow problem. Theoretical Computer Science, 346(2–3): 300–306, November 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gouyou-Beauchamps:2005:ESC

[371] Dominique Gouyou-Beauchamps and Pierre Leroux. Enumeration of symmetry classes of convex polyominoes on the honeycomb lattice. Theoretical Computer Science, 346(2–3):307– 334, November 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuba:2005:SCN

[372] Attila Kuba and Murice Nivat. A sufficient condition for non-uniqueness in binary tomography with absorption. Theoretical Computer Science, 346(2–3):335–357, November 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Louchard:2005:MRU

[373] Guy Louchard. Monotone runs of uniformly distributed integer random variables: a probabilistic analysis. Theoretical Computer Science, 346(2–3):358– 387, November 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martinez:2005:EIA

[374] Conrado Martínez and Xavier Molinero. Efficient iteration in admissible combinatorial classes. Theoretical Computer Science, 346(2–3):388– 417, November 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Massazza:2005:CCB

[375] P. Massazza and R. Radicioni. On computing the coefficients of bivariate holonomic formal series. Theoretical Computer Science, 346(2–3):418– 438, November 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Picouleau:2005:RCP

[376] C. Picouleau. Reconstruction of convex polyominoes from orthogonal projections of their contours. Theoretical Computer Science, 346(2–3):439– 454, November 28, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sulanke:2005:TDN

[377] Robert A. Sulanke. Three dimensional Narayana and Schröder numbers. Theoretical Computer Science, 346(2–3): 455–468, November 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tijdeman:2005:RSM

[378] Robert Tijdeman. Rauzy substitutions and multi-dimensional Sturmian words. Theoretical Computer Science, 346(2–3):469–489, November 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Petkovic:2005:PVT

[379] Tatjana Petković and Saeed Salehi. Positive varieties of tree languages. Theoretical Computer Science, 347(1– 2):1–35, November 30, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonizzoni:2005:RGT

[380] Paola Bonizzoni, Gianluca Della Vedova, and Riccardo Dondi. Reconciling a gene tree to a species tree under the duplication cost model. Theoretical Computer Science, 347(1–2):36– 53, November 30, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chepoi:2005:ASS

[381] Victor D. Chepoi, Feodor F. Dragan, and Chenyu Yan. Additive sparse spanners for graphs with bounded length of largest induced cycle. Theoretical Computer Science, 347(1–2):54–75, November 30, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gurski:2005:RBN

[382] Frank Gurski and Egon Wanke. On the relationship between NLC-width and linear NLC-width. Theoretical Computer Science, 347(1–2):76–89, November 30, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yamakami:2005:RBI

[383] Tomoyuki Yamakami and Toshio Suzuki. Resource bounded immunity and simplicity. Theoretical Computer Science, 347(1–2):90–129, November 30, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Georgiou:2005:EGR

[384] Chryssis Georgiou, Dariusz R. Kowalski, and Alexander A. Shvartsman. Efficient gossip and robust distributed computation. Theoretical Computer Science, 347(1–2):130–166, November 30, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asveld:2005:FCFa

[385] Peter R. J. Asveld. Fuzzy contextfree languages—Part 1: Generalized fuzzy context-free grammars. Theoretical Computer Science, 347(1–2):167– 190, November 30, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asveld:2005:FCFb

[386] Peter R. J. Asveld. Fuzzy contextfree languages—Part 2: Recognition and parsing algorithms. Theoretical Computer Science, 347(1–2):191–213, November 30, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kiwiel:2005:FRS

[387] Krzysztof C. Kiwiel. On Floyd and Rivest’s SELECT algorithm. Theoretical Computer Science, 347(1–2):214– 238, November 30, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fredriksson:2005:SIT

[388] Kimmo Fredriksson, Gonzalo Navarro, and Esko Ukkonen. Sequential and indexed two-dimensional combinatorial template matching allowing rotations. Theoretical Computer Science, 347(1–2):239–275, November 30, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fulop:2005:LDM

[389] Zoltán Fülöp, Armin Kühnemann, and Heiko Vogler. Linear deterministic multi bottom-up tree transducers. Theoretical Computer Science, 347(1–2): 276–287, November 30, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2005:ASM

[390] Zhaohui Liu and T. C. Edwin Cheng. Approximation schemes for minimizing total (weighted) completion time with release dates on a batch machine. Theoretical Computer Science, 347(1–2):288–298, November 30, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schost:2005:TNE

[391] Éric Schost. There is no efficient reverse derivation mode for discrete derivatives. Theoretical Computer Science, 347(1–2):299–305, November 30, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hsiao:2005:DRS

[392] H. K. Hsiao, Y. T. Yeh, and S. S. Yu. Dependences related to strict binary relations. Theoretical Computer Science, 347(1–2):306–324, November 30, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Miltersen:2005:CCD

[393] Peter Bro Miltersen, Jaikumar Radhakrishnan, and Ingo Wegener. On converting CNF to DNF. Theoretical Computer Science, 347(1–2):325– 335, November 30, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Castiglione:2005:ECP

[394] G. Castiglione, A. Frosini, A. Restivo, and S. Rinaldi. Enumeration of L-convex polyominoes by rows and columns. Theoretical Computer Science, 347(1–2):336–352, November 30, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Munarini:2005:CC

[395] Emanuele Munarini and Damiano Torri. Cayley continuants. Theoretical Computer Science, 347(1–2):353– 369, November 30, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frosini:2005:IPD

[396] Andrea Frosini, Maurice Nivat, and Laurent Vuillon. An introduction to periodical discrete sets from a tomographical perspective. Theoretical Computer Science, 347(1–2):370–392, November 30, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brunetti:2005:RGQ

[397] Sara Brunetti and Alain Daurat. Random generation of Q-convex sets. Theoretical Computer Science, 347(1–2): 393–414, November 30, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vinodchandran:2005:NCC

[398] N. V. Vinodchandran. A note on the circuit complexity of PP. Theoretical Computer Science, 347(1–2):415– 418, November 30, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harju:2005:CPB

[399] Tero Harju, Arto Lepistö, and Dirk Nowotka. A characterization of periodicity of bi-infinite words. Theoretical Computer Science, 347(1–2):419– 422, November 30, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Feng:2005:SS

[400] Feng Feng, Xianzhong Zhao, and Young Bae Jun. *-μ-semirings and *- λ-semirings. Theoretical Computer Science, 347(1–2):423–431, November 30, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daurat:2005:SNC

[401] A. Daurat, Y. Gérard, and M. Nivat. Some necessary clarifications about the chords’ problem and the Partial Digest Problem. Theoretical Computer Science, 347(1–2):432–436, November 30, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Champarnaud:2005:ENR

[402] J.-M. Champarnaud and F. Coulon. Erratum to “NFA reduction algorithms by means of regular inequalities” [Theoret. Comput. Sci. 327 (2004) 241– 253]. Theoretical Computer Science, 347(1–2):437–440, November 30, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [3].

Anonymous:2005:EBo

[403] Anonymous. Editorial board. Theoretical Computer Science, 347(1–2):v– ix, November 30, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abramsky:2005:SAR

[404] Samson Abramsky. A structural approach to reversible computation. Theoretical Computer Science, 347(3):441– 464, December 1, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khasidashvili:2005:CFR

[405] Zurab Khasidashvili and John Glauert. The conflict-free Reduction Geometry. Theoretical Computer Science, 347(3): 465–497, December 1, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duntsch:2005:RTB

[406] Ivo Düntsch and Michael Winter. A representation theorem for Boolean contact algebras. Theoretical Computer Science, 347(3):498–512, December 1, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arimura:2005:P

[407] H. Arimura and S. Jain. Preface. Theoretical Computer Science, 348(1):1– 2, December 2, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Watanabe:2005:SST

[408] Osamu Watanabe. Sequential sampling techniques for algorithmic learning theory. Theoretical Computer Science, 348(1):3–14, December 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lange:2005:IIA

[409] Steffen Lange, Gunter Grieser, and Thomas Zeugmann. Inductive inference of approximations for recursive concepts. Theoretical Computer Science, 348(1):15–40, December 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nessel:2005:LEP

[410] Jochen Nessel and Steffen Lange. Learning erasing pattern languages with queries. Theoretical Computer Science, 348(1):41–57, December 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Satoh:2005:LTR

[411] Ken Satoh. Learning taxonomic relation by case-based reasoning. Theoretical Computer Science, 348(1):58– 69, December 2, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Denis:2005:LPU

[412] François Denis, Rémi Gilleron, and Fabien Letouzey. Learning from positive and unlabeled examples. Theoretical Computer Science, 348(1):70– 83, December 2, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hirata:2005:PHA

[413] Kouichi Hirata. Prediction-hardness of acyclic conjunctive queries. Theoretical Computer Science, 348(1):84– 94, December 2, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DasGupta:2005:ALM

[414] Bhaskar DasGupta and Barbara Hammer. On approximate learning by multi-layered feedforward circuits. Theoretical Computer Science, 348(1): 95–127, December 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:Pc

[415] Anonymous. Preface. Theoretical Computer Science, 348(1):v–ix, December 2, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaz:2005:P

[416] Josep Díaz and Juhani Karhumaki. Preface. Theoretical Computer Science, 348(2–3):129, December 8, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bournez:2005:ECF

[417] Olivier Bournez and Emmanuel Hainry. Elementarily computable functions over the real numbers and R-sub-recursive functions. Theoretical Computer Science, 348(2–3):130–147, December 8, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bulatov:2005:CPF

[418] Andrei Bulatov and Martin Grohe. The complexity of partition functions. Theoretical Computer Science, 348(2–3): 148–186, December 8, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dvorak:2005:LCC

[419] Zdeněk Dvořák, Daniel Král’, and Ondřej Pangrác. Locally consistent constraint satisfaction problems. Theoretical Computer Science, 348(2–3): 187–206, December 8, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Feigenbaum:2005:GPS

[420] Joan Feigenbaum, Sampath Kannan, Andrew McGregor, Siddharth Suri, and Jian Zhang. On graph problems in a semi-streaming model. Theoretical Computer Science, 348(2–3):207– 216, December 8, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fleischer:2005:LTS

[421] Lisa Fleischer. Linear tolls suffice: New bounds and algorithms for tolls in single source networks. Theoretical Computer Science, 348(2–3):217– 225, December 8, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fotakis:2005:SUF

[422] Dimitris Fotakis, Spyros Kontogiannis, and Paul Spirakis. Selfish unsplittable flows. Theoretical Computer Science, 348(2–3):226–239, December 8, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Halperin:2005:MES

[423] Eran Halperin and Richard M. Karp. The minimum-entropy set cover problem. Theoretical Computer Science, 348(2–3):240–250, December 8, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoory:2005:SPM

[424] Shlomo Hoory, Avner Magen, Steven Myers, and Charles Rackoff. Simple permutations mix well. Theoretical Computer Science, 348(2–3):251– 261, December 8, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krauthgamer:2005:BBC

[425] Robert Krauthgamer and James R. Lee. The black-box complexity of nearest-neighbor search. Theoretical Computer Science, 348(2–3):262–276, December 8, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kunc:2005:RSL

[426] Michal Kunc. Regular solutions of language inequalities and well quasiorders. Theoretical Computer Science, 348(2–3):277–293, December 8, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lebhar:2005:COD

[427] Emmanuelle Lebhar and Nicolas Schabanel. Close to optimal decentralized routing in long-range contact networks. Theoretical Computer Science, 348(2–3):294–310, December 8, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanMelkebeek:2005:TLB

[428] Dieter van Melkebeek and Ran Raz. A time lower bound for satisfiability. Theoretical Computer Science, 348(2–3): 311–320, December 8, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Soltys:2005:PHC

[429] Michael Soltys. LA, permutations, and the Hajós Calculus. Theoretical Computer Science, 348(2–3):321– 333, December 8, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vassilvitskii:2005:ECS

[430] Sergei Vassilvitskii and Mihalis Yannakakis. Efficiently computing succinct trade-off curves. Theoretical Computer Science, 348(2–3):334–356, December 8, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Williams:2005:NAO

[431] Ryan Williams. A new algorithm for optimal 2-constraint satisfaction and its implications. Theoretical Computer Science, 348(2–3):357–365, December 8, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nesetril:2005:GC

[432] Jaroslav Nešetřil and Gerhard Woeginger. Graph colorings. Theoretical Computer Science, 349(1):1, December 12, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bodlaender:2005:AGF

[433] Hans L. Bodlaender, Andreas Brandstädt, Dieter Kratsch, Michaël Rao, and Jeremy Spinrad. On algorithms for (P,gem)-free graphs. Theoretical Computer Science, 349(1):2–21, December 12, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bodlaender:2005:ECB

[434] Hans L. Bodlaender and Fedor V. Fomin. Equitable colorings of bounded treewidth graphs. Theoretical Computer Science, 349(1):22–30, December 12, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bulatov:2005:HCD

[435] Andrei A. Bulatov. H-coloring dichotomy revisited. Theoretical Computer Science, 349(1):31–39, December 12, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaz:2005:CCN

[436] Josep Díaz, Vishal Sanwalani, Maria Serna, and Paul G. Spirakis. The chromatic and clique numbers of random scaled sector graphs. Theoretical Computer Science, 349(1):40–51, December 12, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Feder:2005:LMP

[437] Tomás Feder, Pavol Hell, Sulamita Klein, Loana Tito Nogueira, and Fábio Protti. List matrix partitions of chordal graphs. Theoretical Computer Science, 349(1):52–66, December 12, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fiala:2005:CCC

[438] Jiří Fiala and Daniël Paulusma. A complete complexity classification of the role assignment problem. Theoretical Computer Science, 349(1):67– 81, December 12, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kostochka:2005:ECG

[439] A. V. Kostochka and K. Nakprasit. On equitable Δ-coloring of graphs with low average degree. Theoretical Computer Science, 349(1):82–91, December 12, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kohl:2005:LVL

[440] Anja Kohl, Jens Schreyer, Zsolt Tuza, and Margit Voigt. List version of L(d, s)-labelings. Theoretical Computer Science, 349(1):92–98, December 12, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kral:2005:GCC

[441] Daniel Král’. Group coloring is Π_{2}^{p}- complete. Theoretical Computer Science, 349(1):99–111, December 12, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Macajova:2005:FCC

[442] Edita Máčajová and Martin Škoviera. Fano colourings of cubic graphs and the Fulkerson Conjecture. Theoretical Computer Science, 349(1):112–120, December 12, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2005:EBp

[443] Anonymous. Editorial board. Theoretical Computer Science, 349(1):v– ix, December 12, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vandenHerik:2005:F

[444] H. Jaap van den Herik and Hiroyuki Iida. Foreword. Theoretical Computer Science, 349(2):121–122, December 14, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hayward:2005:SHD

[445] Ryan Hayward, Yngvi Björnsson, Michael Johanson, Morgan Kan, Nathan Po, and Jack van Rijswijck.

Solving 7×7 Hex with domination, fillin, and virtual connections. Theoretical Computer Science, 349(2):123– 139, December 14, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bourzutschky:2005:CEM

[446] M. S. Bourzutschky, J. A. Tamplin, and G. McC. Haworth. Chess endgames: 6-man data and strategy. Theoretical Computer Science, 349(2):140–157, December 14, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Andrist:2005:DME

[447] R. B. Andrist and G. Mc. Haworth. Deeper model endgame analysis. Theoretical Computer Science, 349(2):158– 167, December 14, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanderWerf:2005:LSF

[448] Erik C. D. van der Werf, H. Jaap van den Herik, and Jos W. H. M. Uiterwijk. Learning to score final positions in the game of Go. Theoretical Computer Science, 349(2):168– 183, December 14, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nakamura:2005:SAB

[449] Katsuhiko Nakamura. Static analysis based on formal models and incremental computation in Go programming. Theoretical Computer Science, 349(2):184–201, December 14, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gomboc:2005:TEF

[450] D. Gomboc, M. Buro, and T. A. Marsland. Tuning evaluation functions by maximizing concordance. Theoretical Computer Science, 349(2):202–229, December 14, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lieberum:2005:EFG

[451] Jens Lieberum. An evaluation function for the game of amazons. Theoretical Computer Science, 349(2):230– 244, December 14, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Donkers:2005:SEF

[452] H. H. L. M. Donkers, H. J. van den Herik, and J. W. H. M. Uiterwijk. Selecting evaluation functions in Opponent-Model search. Theoretical Computer Science, 349(2):245–267, December 14, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sadikov:2005:BPM

[453] A. Sadikov, I. Bratko, and I. Kononenko. Bias and pathology in minimax search. Theoretical Computer Science, 349(2): 268–281, December 14, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okhotin:2005:USL

[454] Alexander Okhotin. Unresolved systems of language equations: Expressive power and decision problems. Theoretical Computer Science, 349(3):283– 308, December 16, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aldred:2005:PMP

[455] R. E. L. Aldred, M. D. Atkinson, H. P. van Ditmarsch, C. C. Handley, D. A. Holton, and D. J. McCaughan. Permuting machines and priority queues. Theoretical Computer Science, 349(3): 309–317, December 16, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanchard:2005:SRA

[456] F. Blanchard, J. Cervelle, and E. Formenti. Some results about the chaotic behavior of cellular automata. Theoretical Computer Science, 349(3):318– 336, December 16, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epstein:2005:CVS

[457] Leah Epstein and Asaf Levin. The chord version for SONET ADMs minimization. Theoretical Computer Science, 349(3):337–346, December 16, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bjorklund:2005:CSR

[458] Henrik Björklund and Sergei Vorobyov. Combinatorial structure and randomized subexponential algorithms for infinite games. Theoretical Computer Science, 349(3):347–360, December 16, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cieliebak:2005:PDH

[459] Mark Cieliebak, Stephan Eidenbenz, and Paolo Penna. Partial Digest is hard to solve for erroneous input data. Theoretical Computer Science, 349(3):361– 381, December 16, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hemaspaandra:2005:CKE

[460] Edith Hemaspaandra, Holger Spakowski, and Jörg Vogel. The complexity of Kemeny elections. Theoretical Computer Science, 349(3):382–391, December 16, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bourke:2005:ERF

[461] Chris Bourke, John M. Hitchcock, and N. V. Vinodchandran. Entropy rates and finite-state dimension. Theoretical Computer Science, 349(3):392–406, December 16, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wu:2005:ALA

[462] Bang Ye Wu. An analysis of the LPT algorithm for the max–min and the min–ratio partition problems. Theoretical Computer Science, 349(3):407– 419, December 16, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kitaev:2005:SPO

[463] Sergey Kitaev. Segmental partially ordered generalized patterns. Theoretical Computer Science, 349(3):420– 428, December 16, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chuan:2005:LFI

[464] Wai-Fong Chuan and Hui-Ling Ho. Locating factors of the infinite Fibonacci word. Theoretical Computer Science, 349(3):429–442, December 16, 2005.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kutz:2005:CAT

[465] Martin Kutz. Conway’s Angel in three dimensions. Theoretical Computer Science, 349(3):443–451, December 16, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yan:2005:GCP

[466] Weigen Yan, Yeong-Nan Yeh, and Fuji Zhang. Graphical condensation of plane graphs: a combinatorial approach. Theoretical Computer Science, 349(3):452–461, December 16, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demange:2005:CPL

[467] Marc Demange, Tınaz Ekim, and Dominique de Werra. (p, k)-coloring problems in line graphs. Theoretical Computer Science, 349(3):462–474, December 16, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nakata:2005:ETH

[468] Toshio Nakata. On the expected time for Herman’s probabilistic selfstabilizing algorithm. Theoretical Computer Science, 349(3):475–483, December 16, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cazals:2005:ARM

[469] F. Cazals and C. Karande. An algorithm for reporting maximal c-cliques. Theoretical Computer Science, 349(3): 484–490, December 16, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cesa-Bianchi:2006:F

[470] Nicolò Cesa-Bianchi, Rüdiger Reischuk, and Thomas Zeugmann. Foreword. Theoretical Computer Science, 350(1):1–2, January 18, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amano:2006:LMB

[471] Kazuyuki Amano and Akira Maruoka. On learning monotone Boolean functions under the uniform distribution. Theoretical Computer Science, 350(1): 3–12, January 18, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Servedio:2006:LEM

[472] Rocco A. Servedio. On learning embedded midbit functions. Theoretical Computer Science, 350(1):13–23, January 18, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bshouty:2006:MAC

[473] Nader H. Bshouty and Lynn Burroughs. Maximizing agreements and coagnostic learning. Theoretical Computer Science, 350(1):24–39, January 18, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Forster:2006:SPD

[474] Jürgen Forster and Hans Ulrich Simon. On the smallest possible dimension and the largest possible margin of linear arrangements representing given concept classes. Theoretical Computer Science, 350(1):40–48, January 18, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kobler:2006:CLC

[475] Johannes Köbler and Wolfgang Lindner. The complexity of learning concept classes with polynomial general dimension. Theoretical Computer Science, 350(1):49–62, January 18, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Suzuki:2006:OTT

[476] Yusuke Suzuki, Takayoshi Shoudai, Tomoyuki Uchida, and Tetsuhiro Miyahara. Ordered term tree languages which are polynomial time inductively inferable from positive data. Theoretical Computer Science, 350(1):63– 90, January 18, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reidenbach:2006:NLC

[477] Daniel Reidenbach. A non-learnable class of E-pattern languages. Theoretical Computer Science, 350(1):91– 102, January 18, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martin:2006:ULT

[478] Éric Martin, Arun Sharma, and Frank Stephan. Unifying logic, topology and learning in Parametric logic. Theoretical Computer Science, 350(1):103– 124, January 18, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hayashi:2006:MBI

[479] Susumu Hayashi. Mathematics based on incremental learning—Excluded middle and inductive inference. Theoretical Computer Science, 350(1):125– 139, January 18, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fronhofer:2006:HFP

[480] Bertram Fronhöfer and Akihiro Yamamoto. Hypothesis finding with proof theoretical appropriateness criteria. Theoretical Computer Science, 350(1):140–162, January 18, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:F

[481] Anonymous. Foreword. Theoretical Computer Science, 350(1):v–ix, January 18, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sannella:2006:P

[482] Donald Sannella. Preface. Theoretical Computer Science, 350(2–3): 163, February 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bojanczyk:2006:TWA

[483] Mikolaj Bojańczyk and Thomas Colcombet. Tree-walking automata cannot be determinized. Theoretical Computer Science, 350(2–3):164–173, February 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dawar:2006:BGI

[484] Anuj Dawar, Erich Grädel, and Stephan Kreutzer. Backtracking games and inflationary fixed points. Theoretical Computer Science, 350(2–3):174– 187, February 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Deng:2006:TAT

[485] Yuxin Deng and Davide Sangiorgi. Towards an algebraic theory of typed mobile processes. Theoretical Computer Science, 350(2–3):188–212, February 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Faggian:2006:IOL

[486] Claudia Faggian. Interactive observability in Ludics: The geometry of tests. Theoretical Computer Science, 350(2–3):213–233, February 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ghica:2006:SCC

[487] D. R. Ghica, A. S. Murawski, and C.- H. L. Ong. Syntactic control of concurrency. Theoretical Computer Science, 350(2–3):234–251, February 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Haghverdi:2006:CMG

[488] Esfandiar Haghverdi and Philip Scott. A categorical model for the geometry of interaction. Theoretical Computer Science, 350(2–3):252–274, February 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laird:2006:CC

[489] J. Laird. A calculus of coroutines. Theoretical Computer Science, 350(2–3): 275–291, February 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martin:2006:EFP

[490] Keye Martin. Entropy as a fixed point. Theoretical Computer Science, 350(2– 3):292–324, February 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schweikardt:2006:EPM

[491] Nicole Schweikardt. On the expressive power of monadic least fixed point logic. Theoretical Computer Science, 350(2–3):325–344, February 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Serre:2006:GWC

[492] Olivier Serre. Games with winning conditions of high Borel complexity. Theoretical Computer Science, 350(2–3): 345–372, February 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:AIa

[493] Anonymous. Author index. Theoretical Computer Science, 350(2–3): 373, February 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:MIa

[494] Anonymous. Master index. Theoretical Computer Science, 350(2–3):375– 384, February 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daumas:2006:P

[495] Marc Daumas and Nathalie Revol. Preface. Theoretical Computer Science, 351(1):1, February 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Parker:2006:TCD

[496] Matthew W. Parker. Three concepts of decidability for general subsets of uncountable spaces. Theoretical Computer Science, 351(1):2–13, February 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ziegler:2006:SVS

[497] Martin Ziegler. Stability versus speed in a computable algebraic model. Theoretical Computer Science, 351(1):14– 26, February 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zheng:2006:DBC

[498] Xizhong Zheng, Dianchen Lu, and Kejin Bao. Divergence bounded computable real numbers. Theoretical Computer Science, 351(1):27–38, February 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ciaffaglione:2006:CCI

[499] Alberto Ciaffaglione and Pietro Di Gianantonio. A certified, corecursive implementation of exact real numbers. Theoretical Computer Science, 351(1): 39–51, February 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanderHoeven:2006:CER

[500] Joris van der Hoeven. Computations with effective real numbers. Theoretical Computer Science, 351(1):52– 60, February 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hughes:2006:ADS

[501] Jesse Hughes and Milad Niqui. Admissible digit sets. Theoretical Computer Science, 351(1):61–73, February 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Briggs:2006:IER

[502] Keith Briggs. Implementing exact real arithmetic in python, C++ and C. Theoretical Computer Science, 351(1): 74–81, February 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hill:2006:QUB

[503] M. Hill and I. Robinson. Quadrature using 64-bit IEEE arithmetic for integrands over [0, 1] with a singularity at 1. Theoretical Computer Science, 351(1):82–100, February 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kornerup:2006:CSV

[504] Peter Kornerup and Jean-Michel Muller. Choosing starting values for certain Newton–Raphson iterations. Theoretical Computer Science, 351(1): 101–110, February 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bronnimann:2006:DBI

[505] Hervé Brönnimann, Guillaume Melquiond, and Sylvain Pion. The design of the Boost interval arithmetic library. Theoretical Computer Science, 351(1):111– 118, February 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delanoue:2006:UIA

[506] Nicolas Delanoue, Luc Jaulin, and Bertrand Cottenceau. Using interval arithmetic to prove that a set is path-connected. Theoretical Computer Science, 351(1):119–128, February 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:Pa

[507] Anonymous. Preface. Theoretical Computer Science, 351(1):v–ix, February 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maharaj:2006:P

[508] Savi Maharaj, Carron Shankland, and Charles Rattray. Preface. Theoretical Computer Science, 351(2):129–130, February 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garavel:2006:SSR

[509] Hubert Garavel and Wendelin Serwe. State space reduction for process algebra specifications. Theoretical Computer Science, 351(2):131–145, February 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hill:2006:LCM

[510] Gillian Hill and Steven Vickers. A language for configuring multi-level specifications. Theoretical Computer Science, 351(2):146–166, February 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Imine:2006:FDV

[511] Abdessamad Imine, Michaël Rusinowitch, Gérald Oster, and Pascal Molli. Formal design and verification of operational transformation algorithms for copies convergence. Theoretical Computer Science, 351(2):167–183, February 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Knapp:2006:SRM

[512] Alexander Knapp, Stephan Merz, Martin Wirsing, and Júlia Zappe. Specification and refinement of mobile systems in MTLA and mobile UML. Theoretical Computer Science, 351(2):184– 202, February 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuster-Filipe:2006:MCI

[513] Juliana Küster-Filipe. Modelling concurrent interactions. Theoretical Computer Science, 351(2):203–220, February 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moller:2006:AMO

[514] Bernhard Möller and Georg Struth. Algebras of modal operators and partial correctness. Theoretical Computer Science, 351(2):221–239, February 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanRiemsdijk:2006:SPR

[515] M. Birna van Riemsdijk, John-Jules Ch. Meyer, and Frank S. de Boer. Semantics of plan revision in intelligent agents. Theoretical Computer Science, 351(2):240–257, February 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sims:2006:ESL

[516] Élodie-Jane Sims. Extending separation logic with fixpoints and postponed substitution. Theoretical Computer Science, 351(2):258–275, February 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meng:2006:CCR

[517] Sun Meng and Luís S. Barbosa. Components as coalgebras: The refinement dimension. Theoretical Computer Science, 351(2):276–294, February 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Downey:2006:E

[518] Rod Downey, Mike Langston, and Rolf Niedermeier. Editorial. Theoretical Computer Science, 351(3):295, February 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bryant:2006:CUP

[519] David Bryant and Jens Lagergren. Compatibility of unrooted phylogenetic trees is FPT. Theoretical Computer Science, 351(3):296–302, February 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Buss:2006:SWH

[520] Jonathan F. Buss and Tarique Islam. Simplifying the weft hierarchy. Theoretical Computer Science, 351(3):303– 313, February 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2006:MPP

[521] Yijia Chen and Jörg Flum. On miniaturized problems in parameterized complexity theory. Theoretical Computer Science, 351(3):314–336, February 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Damaschke:2006:PET

[522] Peter Damaschke. Parameterized enumeration, transversals, and imperfect phylogeny reconstruction. Theoretical Computer Science, 351(3):337–350, February 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fellows:2006:FSR

[523] Michael R. Fellows, Stefan Szeider, and Graham Wrightson. On finding short resolution refutations and small unsatisfiable subsets. Theoretical Computer Science, 351(3):351–359, February 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Haas:2006:CPT

[524] Robert Haas and Michael Hoffmann. Chordless paths through three vertices. Theoretical Computer Science, 351(3): 360–371, February 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hlineny:2006:TGM

[525] Petr Hliněný and Detlef Seese. Trees, grids, and MSO decidability: From graphs to matroids. Theoretical Computer Science, 351(3):372–393, February 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marx:2006:PGS

[526] Dániel Marx. Parameterized graph separation problems. Theoretical Computer Science, 351(3):394–406, February 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marx:2006:PCP

[527] Dániel Marx. Parameterized coloring problems on chordal graphs. Theoretical Computer Science, 351(3):407– 424, February 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muller-Hannemann:2006:MPC

[528] Matthias Müller-Hannemann and Karsten Weihe. Moving policies in cyclic assembly line scheduling. Theoretical Computer Science, 351(3):425– 436, February 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Prieto:2006:LS

[529] Elena Prieto and Christian Sloper. Looking at the stars. Theoretical Computer Science, 351(3):437–445, February 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Raman:2006:PAF

[530] Venkatesh Raman and Saket Saurabh. Parameterized algorithms for feedback set problems and their duals in tournaments. Theoretical Computer Science, 351(3):446–458, February 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wu:2006:MCD

[531] Weili Wu, Hongwei Du, Xiaohua Jia, Yingshu Li, and Scott C.-H. Huang. Minimum connected dominating sets and maximal independent sets in unit disk graphs. Theoretical Computer Science, 352(1–3):1–7, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hitczenko:2006:DCD

[532] Paweł Hitczenko, Jeremy R. Johnson, and Hung-Jen Huang. Distribution of a class of divide and conquer recurrences arising from the computation of the Walsh–Hadamard transform. Theoretical Computer Science, 352(1–3): 8–30, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Glen:2006:OPC

[533] Amy Glen. Occurrences of palindromes in characteristic Sturmian words. Theoretical Computer Science, 352(1–3): 31–46, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zverovich:2006:SID

[534] Igor Edm. Zverovich. Satgraphs and independent domination. Part 1. Theoretical Computer Science, 352(1–3):

47–56, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tanaka:2006:SMB

[535] Genjiro Tanaka. On syntactic monoids of biunitary submonoids determined by homomorphisms from free semigroups onto completely simple semigroups. Theoretical Computer Science, 352(1–3):57–70, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Xavier:2006:ASK

[536] E. C. Xavier and F. K. Miyazawa. Approximation schemes for knapsack problems with shelf divisions. Theoretical Computer Science, 352(1–3): 71–84, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mubayi:2006:DEP

[537] Dhruv Mubayi, György Turán, and Yi Zhao. The DNF exception problem. Theoretical Computer Science, 352(1– 3):85–96, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Terrier:2006:CPC

[538] Véronique Terrier. Closure properties of cellular automata. Theoretical Computer Science, 352(1–3):97–107, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Palu:2006:SPA

[539] A. Dal Palú, E. Pontelli, and D. Ranjan. Sequential and parallel algorithms for the NCA problem on pure pointer machines. Theoretical Computer Science, 352(1–3):108–135, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Poblete:2006:BTA

[540] Patricio V. Poblete, J. Ian Munro, and Thomas Papadakis. The binomial transform and the analysis of skip lists. Theoretical Computer Science, 352(1– 3):136–158, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferbus-Zanda:2006:KCC

[541] Marie Ferbus-Zanda and Serge Grigorieff. Kolmogorov complexities K_{max}, K_{min} on computable partially ordered sets. Theoretical Computer Science, 352(1–3):159–180, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Papakonstantinou:2006:HCP

[542] Periklis A. Papakonstantinou. Hierarchies for classes of priority algorithms for Job Scheduling. Theoretical Computer Science, 352(1–3):181– 189, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Colcombet:2006:PDE

[543] Thomas Colcombet and Damian Niwiński. On the positional determinacy of edge-labeled games. Theoretical Computer Science, 352(1–3):190– 196, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bao:2006:RBP

[544] Xuewen Bao, Frank K. Hwang, and Qiao Li. Rearrangeability of bit permutation networks. Theoretical Computer Science, 352(1–3):197–214, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zverovich:2006:IDH

[545] Igor É. Zverovich and Olga I. Zverovich. Independent domination in hereditary classes. Theoretical Computer Science, 352(1–3):215–225, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moothathu:2006:SPA

[546] T. K. Subrahmonian Moothathu. Set of periods of additive cellular automata. Theoretical Computer Science, 352(1– 3):226–231, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chang:2006:WSN

[547] F. H. Chang, J. Y. Guo, and F. K. Hwang. Wide-sense nonblocking for multi-log_{d} N networks under various routing strategies. Theoretical Computer Science, 352(1–3):232–239, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huynh:2006:ASM

[548] Trinh N. D. Huynh, Wing-Kai Hon, Tak-Wah Lam, and Wing-Kin Sung. Approximate string matching using compressed suffix arrays. Theoretical Computer Science, 352(1–3):240– 249, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frederickson:2006:EAR

[549] Greg N. Frederickson and Roberto Solis-Oba. Efficient algorithms for robustness in resource allocation and scheduling problems. Theoretical Computer Science, 352(1–3):250–265, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Navarro:2006:MIA

[550] Gonzalo Navarro and Edgar Chávez. A metric index for approximate string matching. Theoretical Computer Science, 352(1–3):266–279, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhao:2006:SAP

[551] Zhu Zhao, Zhongqi Dong, and Yongge Wang. Security analysis of a passwordbased authentication protocol proposed to IEEE 1363. Theoretical Computer Science, 352(1–3):280–287, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Youssef:2006:LCS

[552] A. M. Youssef and G. Gong. On linear complexity of sequences over GF. Theoretical Computer Science, 352(1–3): 288–292, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ballet:2006:ICD

[553] S. Ballet. An improvement of the construction of the D. V. and G. V. Chudnovsky algorithm for multiplication in finite fields. Theoretical Computer Science, 352(1–3):293–305, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brlek:2006:CPS

[554] S. Brlek, S. Dulucq, A. Ladouceur, and L. Vuillon. Combinatorial properties of smooth infinite words. Theoretical Computer Science, 352(1–3):306– 317, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hajiaghayi:2006:MFM

[555] M. T. Hajiaghayi and Tom Leighton. On the max-flow min-cut ratio for directed multicommodity flows. Theoretical Computer Science, 352(1–3):318– 321, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhu:2006:CUT

[556] Daming Zhu and Lusheng Wang. On the complexity of unsigned translocation distance. Theoretical Computer Science, 352(1–3):322–328, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tomescu:2006:CWO

[557] Ioan Tomescu. A characterization of the words occurring as factors in a minimum number of words. Theoretical Computer Science, 352(1–3):329– 331, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Das:2006:RAE

[558] Gautam K. Das, Sandip Das, and Subhas C. Nandy. Range assignment for energy efficient broadcasting in linear radio networks. Theoretical Computer Science, 352(1–3):332–341, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2006:SCD

[559] Oscar H. Ibarra and Zhe Dang. On the solvability of a class of diophantine equations and applications. Theoretical Computer Science, 352(1–3):342– 346, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krumke:2006:ENO

[560] Sven O. Krumke, Willem E. de Paepe, Diana Poensgen, and Leen Stougie. Erratum to “News from the online traveling repairman” [Theoret. Comput. Sci. 295 (1–3) (2003) 279–294]. Theoretical Computer Science, 352(1–3):347– 348, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [1].

Anonymous:2006:EBa

[561] Anonymous. Editorial board. Theoretical Computer Science, 352(1–3): v–ix, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schroder:2006:HPC

[562] Lutz Schröder. The HasCasl prologue: Categorical syntax and semantics of the partial λ-calculus. Theoretical Computer Science, 353(1–3):1– 25, March 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhu:2006:SBP

[563] Zhaohui Zhu. Similarity between preferential models. Theoretical Computer Science, 353(1–3):26–52, March 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hartmann:2006:PFC

[564] Sven Hartmann and Sebastian Link. On a problem of Fagin concerning multivalued dependencies in relational databases. Theoretical Computer Science, 353(1–3):53–62, March 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2006:SPC

[565] Wenhui Zhang. Structure of proofs and the complexity of cut elimination. Theoretical Computer Science, 353(1–3): 63–70, March 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lu:2006:AWV

[566] Shiyong Lu, Arthur Bernstein, and Philip Lewis. Automatic workflow verification and generation. Theoretical Computer Science, 353(1–3):71– 92, March 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goranko:2006:CAD

[567] Valentin Goranko and Govert van Drimmelen. Complete axiomatization and decidability of Alternating-time temporal logic. Theoretical Computer Science, 353(1–3):93–117, March 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mitchell:2006:PPT

[568] John C. Mitchell, Ajith Ramanathan, Andre Scedrov, and Vanessa Teague. A probabilistic polynomial-time process calculus for the analysis of cryptographic protocols. Theoretical Computer Science, 353(1–3):118– 164, March 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kamide:2006:LAL

[569] Norihiro Kamide. Linear and affine logics with temporal, spatial and epistemic operators. Theoretical Computer Science, 353(1–3):165–207, March 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boucheneb:2006:C

[570] Hanifa Boucheneb and Rachid Hadjidj. CTL. Theoretical Computer Science, 353(1–3):208–227, March 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lopez:2006:STI

[571] Natalia López, Manuel Núñez, and Ismael Rodríguez. Specification, testing and implementation relations for symbolic-probabilistic systems. Theoretical Computer Science, 353(1–3): 228–248, March 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laroussinie:2006:ETM

[572] F. Laroussinie, N. Markey, and Ph. Schnoebelen. Efficient timed model checking for discrete-time systems. Theoretical Computer Science, 353(1– 3):249–271, March 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kambites:2006:ASC

[573] Mark Kambites. Automatic semigroups and categories. Theoretical Computer Science, 353(1–3):272– 290, March 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ipate:2006:TAN

[574] Florentin Ipate. Testing against a noncontrollable stream X-machine using state counting. Theoretical Computer Science, 353(1–3):291–316, March 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:EBb

[575] Anonymous. Editorial board. Theoretical Computer Science, 353(1–3): v–ix, March 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nijholt:2006:E

[576] A. Nijholt, G. Scollo, and U. Mönnich. Editorial. Theoretical Computer Science, 354(1):1–3, March 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berline:2006:EGM

[577] C. Berline and A. Salibra. Easiness in graph models. Theoretical Computer Science, 354(1):4–23, March 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harrington:2006:UL

[578] Dana Harrington. Uniqueness logic. Theoretical Computer Science, 354(1):

24–41, March 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Roggenbach:2006:CCN

[579] Markus Roggenbach. CSP-CASL — a new integration of process algebra and algebraic specification. Theoretical Computer Science, 354(1):42– 71, March 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Penn:2006:ETC

[580] Gerald Penn. Efficient transitive closure of sparse matrices over closed semirings. Theoretical Computer Science, 354(1):72–81, March 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kepser:2006:CPL

[581] Stephan Kepser and Uwe Mönnich. Closure properties of linear context-free tree languages with an application to optimality theory. Theoretical Computer Science, 354(1):82–97, March 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gust:2006:MHD

[582] Helmar Gust, Kai-Uwe Kühnberger, and Ute Schmid. Metaphors and heuristic-driven theory projection (HDTP). Theoretical Computer Science, 354(1):98–117, March 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asveld:2006:GAP

[583] Peter R. J. Asveld. Generating all permutations by context-free grammars in Chomsky normal form. Theoretical Computer Science, 354(1):118– 130, March 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kracht:2006:PAM

[584] Marcus Kracht. Partial algebras, meaning categories and algebraization. Theoretical Computer Science, 354(1):131– 141, March 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gramatovici:2006:SDI

[585] Radu Gramatovici and Carlos Martín-Vide. Sorted dependency insertion grammars. Theoretical Computer Science, 354(1):142–152, March 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kundaji:2006:RCB

[586] Rohit N. Kundaji and R. K. Shyamasundar. Refinement calculus: a basis for translation validation, debugging and certification. Theoretical Computer Science, 354(1):153–168, March 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:E

[587] Anonymous. Editorial. Theoretical Computer Science, 354(1):v–ix, March 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garavel:2006:TSI

[588] Hubert Garavel and John Hatcliff. TACAS 2003 Special Issue—Preface. Theoretical Computer Science, 354(2):

169–172, March 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Henzinger:2006:UEF

[589] Thomas A. Henzinger, Orna Kupferman, and Rupak Majumdar. On the universal and existential fragments of the μ-calculus. Theoretical Computer Science, 354(2):173–186, March 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Conchon:2006:SCD

[590] Sylvain Conchon and Sava Krstić. Strategies for combining decision procedures. Theoretical Computer Science, 354(2):187–210, March 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Basu:2006:CAV

[591] Samik Basu and C. R. Ramakrishnan. Compositional analysis for verification of parameterized systems. Theoretical Computer Science, 354(2):211– 229, March 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alur:2006:MSR

[592] Rajeev Alur, Salvatore La Torre, and P. Madhusudan. Modular strategies for recursive game graphs. Theoretical Computer Science, 354(2):230– 249, March 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alur:2006:CGP

[593] Rajeev Alur, Thao Dang, and Franjo Ivančić. Counterexample-guided predicate abstraction of hybrid systems.

Theoretical Computer Science, 354(2): 250–271, March 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abdeddaim:2006:STA

[594] Yasmina Abdeddaı¨ m, Eugene Asarin, and Oded Maler. Scheduling with timed automata. Theoretical Computer Science, 354(2):272–300, March 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fersman:2006:SAF

[595] Elena Fersman, Leonid Mokrushin, Paul Pettersson, and Wang Yi. Schedulability analysis of fixed-priority systems using timed automata. Theoretical Computer Science, 354(2):301– 317, March 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lingas:2006:P

[596] Andrzej Lingas and Leszek Gaşieniec. Preface. Theoretical Computer Science, 354(3):319, April 4, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chlebik:2006:CAB

[597] Miroslav Chlebík and Janka Chlebíková. Complexity of approximating bounded variants of optimization problems. Theoretical Computer Science, 354(3): 320–338, April 4, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asano:2006:IAG

[598] Takao Asano. An improved analysis of Goemans and Williamson’s LPrelaxation for MAX SAT. Theoretical Computer Science, 354(3):339– 353, April 4, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stachowiak:2006:FPC

[599] Grzegorz Stachowiak. Fast periodic correction networks. Theoretical Computer Science, 354(3):354–366, April 4, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hammar:2006:CER

[600] Mikael Hammar, Bengt J. Nilsson, and Mia Persson. Competitive exploration of rectilinear polygons. Theoretical Computer Science, 354(3):367– 378, April 4, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reif:2006:BHV

[601] John H. Reif and Zheng Sun. On boundaries of highly visible spaces and applications. Theoretical Computer Science, 354(3):379–390, April 4, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Antunes:2006:CDC

[602] Luis Antunes, Lance Fortnow, Dieter van Melkebeek, and N. V. Vinodchandran. Computational depth: Concept and applications. Theoretical Computer Science, 354(3):391–404, April 4, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berstel:2006:OPR

[603] Jean Berstel, Luc Boasson, Olivier Carton, Bruno Petazzoni, and Jean-Eric Pin. Operations preserving regular languages. Theoretical Computer Science, 354(3):405–420, April 4, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fokkink:2006:CHM

[604] Wan Fokkink, Rob van Glabbeek, and Paulien de Wind. Compositionality of Hennessy–Milner logic by structural operational semantics. Theoretical Computer Science, 354(3):421– 440, April 4, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kumar:2006:P

[605] Ravi Kumar and Matthieu Latapy. Preface. Theoretical Computer Science, 355(1):1–5, April 6, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DallAsta:2006:ENT

[606] Luca Dall’Asta, Ignacio Alvarez-Hamelin, Alain Barrat, Alexei Vázquez, and Alessandro Vespignani. Exploring networks with traceroute-like probes: Theory and simulations. Theoretical Computer Science, 355(1):6–24, April 6, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Annexstein:2006:BUP

[607] Fred S. Annexstein, Kenneth A. Berman, and Mijhalo A. Jovanović. Broadcasting in unstructured peer-to-peer overlay networks. Theoretical Computer Science, 355(1):25–36, April 6, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Behrisch:2006:ECC

[608] Michael Behrisch and Anusch Taraz. Efficiently covering complex networks with cliques of similar vertices. Theoretical Computer Science, 355(1):37– 47, April 6, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sarshar:2006:SPS

[609] Nima Sarshar, Oscar Boykin, and Vwani Roychowdhury. Scalable percolation search on complex networks. Theoretical Computer Science, 355(1): 48–64, April 6, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fraigniaud:2006:DBB

[610] Pierre Fraigniaud and Philippe Gauron. D2B: a de Bruijn based contentaddressable network. Theoretical Computer Science, 355(1):65–79, April 6, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stauffer:2006:LHE

[611] Alexandre O. Stauffer and Valmir C. Barbosa. Local heuristics and the emergence of spanning subgraphs in complex networks. Theoretical Computer Science, 355(1):80–95, April 6, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duchon:2006:CGT

[612] Philippe Duchon, Nicolas Hanusse, Emmanuelle Lebhar, and Nicolas Schabanel. Could any graph be turned into a small-world? Theoretical Computer Science, 355(1):96–103, April 6, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:Pb

[613] Anonymous. Preface. Theoretical Computer Science, 355(1):v–ix, April 6, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deQueiroz:2006:LLI

[614] Ruy de Queiroz and Dexter Kozen. Logic, Language, Information and Computation. Theoretical Computer Science, 355(2):105–107, April 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alessi:2006:ITL

[615] Fabio Alessi, Franco Barbanera, and Mariangiola Dezani-Ciancaglini. Intersection types and lambda models. Theoretical Computer Science, 355(2):108– 126, April 11, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Allender:2006:NPS

[616] Eric Allender. NL-printable sets and nondeterministic Kolmogorov complexity. Theoretical Computer Science, 355 (2):127–138, April 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bechet:2006:K

[617] Denis Béchet and Annie Foret. k. Theoretical Computer Science, 355(2):139– 152, April 11, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Finger:2006:UPA

[618] Marcelo Finger and Renata Wassermann. The universe of propositional approximations. Theoretical Computer Science, 355(2):153–166, April 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hartmann:2006:AFD

[619] Sven Hartmann, Sebastian Link, and Klaus-Dieter Schewe. Axiomatisations of functional dependencies in the presence of records, lists, sets and multisets. Theoretical Computer Science, 355(2): 167–196, April 11, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [758].

Hella:2006:CQH

[620] Lauri Hella and José María Turull-Torres. Computing queries with higherorder logics. Theoretical Computer Science, 355(2):197–214, April 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lamo:2006:QFL

[621] Yngve Lamo and Michał Walicki. Quantifier-free logic for nondeterministic theories. Theoretical Computer Science, 355(2):215–227, April 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rott:2006:RCU

[622] Hans Rott. Revision by comparison as a unifying framework: Severe withdrawal, irrevocable revision and irrefutable revision. Theoretical Computer Science, 355(2):228–242, April 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zaionc:2006:PDS

[623] Marek Zaionc. Probability distribution for simple tautologies. Theoretical Computer Science, 355(2):243– 260, April 11, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kolliopoulos:2006:AAM

[624] Stavros G. Kolliopoulos and George Steiner. Approximation algorithms for minimizing the total weighted tardiness on a single machine. Theoretical Computer Science, 355(3):261– 273, April 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Malheiro:2006:FDT

[625] António Malheiro. Finite derivation type for Rees matrix semigroups. Theoretical Computer Science, 355(3):274– 290, April 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Xu:2006:MHE

[626] Ke Xu and Wei Li. Many hard examples in exact phase transitions. Theoretical Computer Science, 355(3):291– 302, April 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wu:2006:LTA

[627] Ro-Yu Wu, Jou-Ming Chang, and Yue-Li Wang. A linear time algorithm for binary tree sequences transformation using left-arm and right-arm rotations. Theoretical Computer Science, 355(3): 303–314, April 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeMarco:2006:ADR

[628] Gianluca De Marco, Luisa Gargano, Evangelos Kranakis, Danny Krizanc, Andrzej Pelc, and Ugo Vaccaro. Asynchronous deterministic rendezvous in graphs. Theoretical Computer Science, 355(3):315–326, April 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sanchez-Couso:2006:RBS

[629] José-Ramón Sánchez-Couso and María-Inés Fernández-Camacho. Reductions in binary search trees. Theoretical Computer Science, 355(3):327– 353, April 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jin:2006:PCP

[630] Zemin Jin and Xueliang Li. On the k-path cover problem for cacti. Theoretical Computer Science, 355(3):354– 363, April 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeSimone:2006:ECJ

[631] Caterina De Simone and C. P. de Mello. Edge-colouring of join graphs. Theoretical Computer Science, 355(3):364– 370, April 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flajolet:2006:SWR

[632] Philippe Flajolet, Markus Nebel, and Helmut Prodinger. The scientific works of Rainer Kemp (1949–2004). Theoretical Computer Science, 355(3):371– 381, April 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hitchcock:2006:HDO

[633] John M. Hitchcock. Hausdorff dimension and oracle constructions. Theoretical Computer Science, 355(3):382– 388, April 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bazgan:2006:DCD

[634] Cristina Bazgan, Zsolt Tuza, and Daniel Vanderpooten. Degreeconstrained decompositions of graphs: Bounded treewidth and planarity. Theoretical Computer Science, 355(3):389– 395, April 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bertoni:2006:P

[635] Alberto Bertoni, Zoltan Ésik, and Juhani Karhumäki. Preface. Theoretical Computer Science, 356(1–2): 1–5, May 5, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beal:2006:CUA

[636] Marie-Pierre Béal and Dominique Perrin. Codes, unambiguous automata and sofic systems. Theoretical Computer Science, 356(1–2):6–13, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bertoni:2006:SFT

[637] Alberto Bertoni, Carlo Mereghetti, and Beatrice Palano. Some formal tools for analyzing quantum automata. Theoretical Computer Science, 356(1–2): 14–25, May 5, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brijder:2006:RGP

[638] Robert Brijder, Hendrik Jan Hoogeboom, and Grzegorz Rozenberg. Reducibility of gene patterns in ciliates using the breakpoint graph. Theoretical Computer Science, 356(1–2): 26–45, May 5, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brzozowski:2006:RCN

[639] Janusz. A. Brzozowski. Representation of a class of nondeterministic semiautomata by canonical words. Theoretical Computer Science, 356(1–2): 46–57, May 5, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Castiglione:2006:RAC

[640] Giusi Castiglione, Antonio Restivo, and Roberto Vaglica. A reconstruction algorithm for L-convex polyominoes. Theoretical Computer Science, 356(1– 2):58–72, May 5, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chaubard:2006:AWP

[641] Laura Chaubard, Jean-Éric Pin, and Howard Straubing. Actions, wreath products of C-varieties and concatenation product. Theoretical Computer Science, 356(1–2):73–89, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cherubini:2006:PLT

[642] Alessandra Cherubini, Stefano Crespi Reghizzi, Matteo Pradella, and Pierluigi San Pietro. Picture languages: Tiling systems versus tile rewriting grammars. Theoretical Computer Science, 356(1–2):90–103, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DAlessandro:2006:SCF

[643] Flavio D’Alessandro, Benedetto Intrigila, and Stefano Varricchio. On the structure of the counting function of sparse context-free languages. Theoretical Computer Science, 356(1–2): 104–117, May 5, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deLuca:2006:SCF

[644] Aldo de Luca and Alessandro De Luca. Some characterizations of finite Sturmian words. Theoretical Computer Science, 356(1–2):118–125, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diekert:2006:LGT

[645] Volker Diekert and Paul Gastin. From local to global temporal logics over Mazurkiewicz traces. Theoretical Computer Science, 356(1–2):126–135, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esik:2006:CCL

[646] Zoltán Ésik. Characterizing CTL-like logics on finite trees. Theoretical Computer Science, 356(1–2):136–152, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goldwurm:2006:PSV

[647] Massimiliano Goldwurm and Violetta Lonati. Pattern statistics and Vandermonde matrices. Theoretical Computer Science, 356(1–2):153–169, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grigorieff:2006:SBD

[648] Serge Grigorieff. Synchronization of a bounded degree graph of cellular automata with nonuniform delays in time D⌊log_{m} D⌋. Theoretical Computer Science, 356(1–2):170–185, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harju:2006:UFP

[649] Tero Harju and Dirk Nowotka. On unique factorizations of primitive words. Theoretical Computer Science, 356(1–2):186–189, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2006:PBM

[650] Oscar H. Ibarra and Bala Ravikumar. On partially blind multihead finite automata. Theoretical Computer Science, 356(1–2):190–199, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karhumaki:2006:CC

[651] Juhani Karhumäki, Michal Kunc, and Alexander Okhotin. Computing by commuting. Theoretical Computer Science, 356(1–2):200–211, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Latteux:2006:IBL

[652] Michel Latteux, Aurélien Lemay, Yves Roos, and Alain Terlutte. Identification of biRFSA languages. Theoretical Computer Science, 356(1–2): 212–223, May 5, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lombardy:2006:S

[653] Sylvain Lombardy and Jacques Sakarovitch. Sequential? Theoretical Computer Science, 356(1–2):224–244, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Neraud:2006:CPC

[654] Jean Néraud. Completing prefix codes in submonoids. Theoretical Computer Science, 356(1–2):245–254, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zielonka:2006:TSM

[655] Wiesław Zielonka. Time-stamps for Mazurkiewicz traces. Theoretical Computer Science, 356(1–2):255–262, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:PC

[656] Anonymous. Professor Choffrut. Theoretical Computer Science, 356(1–2): v–ix, May 5, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:Pd

[657] Anonymous. Preface. Theoretical Computer Science, 356(1–2):xiv, May 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeten:2006:P

[658] Jos Baeten and Flavio Corradini. Preface. Theoretical Computer Science, 356 (3):263–264, May 30, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanGlabbeek:2006:EHD

[659] R. J. van Glabbeek. On the expressiveness of higher dimensional automata. Theoretical Computer Science, 356(3): 265–290, May 30, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [922].

Corradini:2006:FCS

[660] F. Corradini, M. R. Di Berardini, and W. Vogler. Fairness of components in system computations. Theoretical Computer Science, 356(3):291– 324, May 30, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Andova:2006:BBP

[661] Suzana Andova and Tim A. C. Willemse. Branching bisimulation for probabilistic systems: Characteristics and decidability. Theoretical Computer Science, 356(3):325–355, May 30, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beffara:2006:CNS

[662] Emmanuel Beffara and François Maurel. Concurrent nets: a study of prefixing in process calculi. Theoretical Computer Science, 356(3):356–373, May 30, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Finkel:2006:LEP

[663] A. Finkel, G. Geeraerts, J.-F. Raskin, and L. Van Begin. On the ωlanguage expressive power of extended Petri nets. Theoretical Computer Science, 356(3):374–386, May 30, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeNicola:2006:EPK

[664] Rocco De Nicola, Daniele Gorla, and Rosario Pugliese. On the expressive power of KLAIM-based calculi. Theoretical Computer Science, 356(3):387– 421, May 30, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bundgaard:2006:CEN

[665] Mikkel Bundgaard, Thomas Hildebrandt, and Jens Chr. Godskesen. A CPS encoding of name-passing in Higher-order mobile embedded resources. Theoretical Computer Science, 356(3):422–439, May 30, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guan:2006:NPA

[666] Xudong Guan. Name-passing in an ambient-like calculus and its proof using spatial logic. Theoretical Computer Science, 356(3):440–467, May 30, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Phillips:2006:LER

[667] Iain Phillips and Maria Grazia Vigliotti. Leader election in rings of ambient processes. Theoretical Computer Science, 356(3):468–494, May 30, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Artemov:2006:P

[668] Sergei Artemov and Michael Mislove. Preface. Theoretical Computer Science, 357(1–3):1–3, July 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Artemov:2006:JCK

[669] Sergei Artemov. Justified common knowledge. Theoretical Computer Science, 357(1–3):4–22, July 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brezhnev:2006:MKE

[670] Vladimir Brezhnev and Roman Kuznets. Making knowledge explicit: How hard it is. Theoretical Computer Science, 357(1–3):23–34, July 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Buss:2006:PSF

[671] Samuel R. Buss. Polynomial-size Frege and resolution proofs of st. Theoretical Computer Science, 357(1–3): 35–52, July 25, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dershowitz:2006:ACP

[672] Nachum Dershowitz and Claude Kirchner. Abstract canonical presentations. Theoretical Computer Science, 357(1– 3):53–69, July 25, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hyland:2006:CES

[673] Martin Hyland, Gordon Plotkin, and John Power. Combining effects: Sum and tensor. Theoretical Computer Science, 357(1–3):70–99, July 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Japaridze:2006:TC

[674] Giorgi Japaridze. From truth to computability I. Theoretical Computer Science, 357(1–3):100–135, July 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krupski:2006:CRL

[675] Nikolai V. Krupski. On the complexity of the reflected logic of proofs. Theoretical Computer Science, 357(1–3): 136–142, July 25, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krupski:2006:RLP

[676] Vladimir N. Krupski. Referential logic of proofs. Theoretical Computer Science, 357(1–3):143–166, July 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Naumov:2006:LS

[677] Pavel Naumov. Logic of subtyping. Theoretical Computer Science, 357(1– 3):167–185, July 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pentus:2006:LCN

[678] Mati Pentus. Lambek calculus is NPcomplete. Theoretical Computer Science, 357(1–3):186–201, July 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schwichtenberg:2006:APT

[679] Helmut Schwichtenberg. An arithmetic for polynomial-time computation. Theoretical Computer Science, 357(1–3): 202–214, July 25, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Slavnov:2006:GSL

[680] Sergey Slavnov. Geometrical semantics for linear logic (multiplicative fragment). Theoretical Computer Science, 357(1–3):215–229, July 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeJaeger:2006:CCL

[681] Frédéric De Jaeger, Martín Escardó, and Gabriele Santini. On the computational content of the Lawson topology. Theoretical Computer Science, 357 (1–3):230–240, July 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fajstrup:2006:ATC

[682] Lisbeth Fajstrup, Martin Raußen, and Eric Goubault. Algebraic topology and concurrency. Theoretical Computer Science, 357(1–3):241–278, July 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:Pe

[683] Anonymous. Preface. Theoretical Computer Science, 357(1–3):v–ix, July 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2006:Pa

[684] G. Rozenberg. Preface. Theoretical Computer Science, 358(1):1–2, July 31, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Broadbent:2006:PNL

[685] Anne Broadbent and André Allan Méthot. On the power of non-local boxes. Theoretical Computer Science, 358(1):3–14, July 31, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Miakisz:2006:QGT

[686] Katarzyna Miakisz, Edward W. Piotrowski, and Jan Sładkowski. Quantization of games: Towards quantum artificial intelligence. Theoretical Computer Science, 358(1):15–22, July 31, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Potgieter:2006:ZMH

[687] Petrus H. Potgieter. Zeno machines and hypercomputation. Theoretical Computer Science, 358(1):23–33, July 31, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garcez:2006:CCI

[688] Artur S. d’Avila Garcez, Luís C. Lamb, and Dov M. Gabbay. Connectionist computations of intuitionistic reasoning. Theoretical Computer Science, 358 (1):34–55, July 31, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liao:2006:GRD

[689] Bo Liao and Kequan Ding. A 3D graphical representation of DNA sequences and its application. Theoretical Computer Science, 358(1):56–64, July 31, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2006:DEB

[690] Lvzhou Li and Daowen Qiu. Determination of equivalence between quantum sequential machines. Theoretical Computer Science, 358(1):65–74, July 31, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Loos:2006:ADS

[691] Remco Loos. An alternative definition of splicing. Theoretical Computer Science, 358(1):75–87, July 31, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2006:CCS

[692] Oscar H. Ibarra and Gheorghe Păun. Characterizations of context-sensitive languages and other language classes in terms of symport/antiport P systems. Theoretical Computer Science, 358(1): 88–103, July 31, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kumar:2006:AME

[693] Rajeev Kumar and Nilanjan Banerjee. Analysis of a Multiobjective Evolutionary Algorithm on the 0–1 knapsack problem. Theoretical Computer Science, 358(1):104–120, July 31, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rogers:2006:PTS

[694] Alex Rogers, Adam Prügel-Bennett, and Nicholas R. Jennings. Phase transitions and symmetry breaking in genetic algorithms with crossover. Theoretical Computer Science, 358(1):121– 141, July 31, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czeizler:2006:PPC

[695] Elena Czeizler and Eugen Czeizler. On the power of parallel communicating Watson–Crick automata systems. Theoretical Computer Science, 358(1):142– 147, July 31, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:Pf

[696] Anonymous. Preface. Theoretical Computer Science, 358(1):v–ix, July 31, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gardner:2006:E

[697] Philippa Gardner and Nobuko Yoshida. Editorial. Theoretical Computer Science, 358(2–3):149, August 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bollig:2006:MPA

[698] Benedikt Bollig and Martin Leucker. Message-passing automata are expressively equivalent to EMSO logic. Theoretical Computer Science, 358(2–3): 150–172, August 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Varacca:2006:PES

[699] Daniele Varacca, Hagen Völzer, and Glynn Winskel. Probabilistic event structures and domains. Theoretical Computer Science, 358(2–3):173– 199, August 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mellies:2006:AGT

[700] Paul-André Melliès. Asynchronous games 2: The true concurrency of innocence. Theoretical Computer Science, 358(2–3):200–228, August 7, 2006.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amadio:2006:RCS

[701] Roberto M. Amadio and Silvano Dal Zilio. Resource control for synchronous cooperative threads. Theoretical Computer Science, 358(2–3):229–254, August 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bojanczyk:2006:CEE

[702] Mikolaj Bojańczyk and Igor Walukiewicz. Characterizing EF and EX tree logics. Theoretical Computer Science, 358(2– 3):255–272, August 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Markey:2006:MCR

[703] Nicolas Markey and Jean-François Raskin. Model checking restricted sets of timed paths. Theoretical Computer Science, 358(2–3):273–292, August 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caires:2006:EQU

[704] Luís Caires and Etienne Lozes. Elimination of quantifiers and undecidability in spatial logics for concurrency. Theoretical Computer Science, 358(2–3): 293–314, August 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kucera:2006:GAC

[705] Antonín Kučera and Philippe Schnoebelen. A general approach to comparing infinite-state systems with their finite-state specifications. Theoretical Computer Science, 358(2–3):315– 333, August 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Homeister:2006:LBR

[706] Matthias Homeister. Lower bounds for restricted read-once parity branching programs. Theoretical Computer Science, 359(1–3):1–14, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kamae:2006:MPC

[707] Teturo Kamae, Hui Rao, and Yu-Mei Xue. Maximal pattern complexity of two-dimensional words. Theoretical Computer Science, 359(1–3):15– 27, August 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wu:2006:CVD

[708] Yongcheng Wu and Klaus Weihrauch. A computable version of the Daniell– Stone theorem on integration and linear functionals. Theoretical Computer Science, 359(1–3):28–42, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balogh:2006:EBG

[709] József Balogh, Dhruv Mubayi, and András Pluhár. On the edgebandwidth of graph products. Theoretical Computer Science, 359(1–3): 43–57, August 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Verhoeven:2006:LBC

[710] Yves F. Verhoeven. A lower bound on the competitivity of memoryless algorithms for a generalization of the CNN problem. Theoretical Computer Science, 359(1–3):58–68, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2006:EAF

[711] Chiuyuan Chen, James K. Lan, and Wen-Shiang Tang. An efficient algorithm to find a double-loop network that realizes a given L-shape. Theoretical Computer Science, 359(1–3):69– 76, August 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tiplea:2006:PNR

[712] Ferucio Laurenţiu Ţiplea and Aurora Ţiplea. Petri net reactive modules. Theoretical Computer Science, 359(1– 3):77–100, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arnold:2006:SGA

[713] Fredrick Arnold and Benjamin Steinberg. Synchronizing groups and automata. Theoretical Computer Science, 359(1–3):101–110, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leykin:2006:NMD

[714] Anton Leykin, Jan Verschelde, and Ailing Zhao. Newton’s method with deflation for isolated singularities of polynomial systems. Theoretical Computer Science, 359(1–3):111–122, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Doerr:2006:IBS

[715] Benjamin Doerr, Nils Hebbinghaus, and Sören Werth. Improved bounds and schemes for the declustering problem. Theoretical Computer Science, 359(1–3):123–132, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sadahiro:2006:MPT

[716] Taizo Sadahiro. Multiple points of tilings associated with Pisot numeration systems. Theoretical Computer Science, 359(1–3):133–147, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DiGiacomo:2006:K

[717] Emilio Di Giacomo, Walter Didimo, Giuseppe Liotta, and Matthew Suderman. k. Theoretical Computer Science, 359(1–3):148–175, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gu:2006:NDP

[718] Xiaoyang Gu. A note on dimensions of polynomial size circuits. Theoretical Computer Science, 359(1–3):176– 187, August 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gurski:2006:VDP

[719] Frank Gurski and Egon Wanke. Vertex disjoint paths on clique-width bounded graphs. Theoretical Computer Science, 359(1–3):188–199, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garg:2006:ACB

[720] Vijay K. Garg. Algorithmic combinatorics based on slicing posets. Theoretical Computer Science, 359(1–3):200– 213, August 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fici:2006:WAT

[721] Gabriele Fici, Filippo Mignosi, Antonio Restivo, and Marinella Sciortino. Word assembly through minimal forbidden words. Theoretical Computer Science, 359(1–3):214–230, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Higgins:2006:US

[722] Peter M. Higgins and Christopher J. Saker. Unavoidable sets. Theoretical Computer Science, 359(1–3):231– 238, August 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lavault:2006:AAL

[723] Christian Lavault and Guy Louchard. Asymptotic analysis of a leader election algorithm. Theoretical Computer Science, 359(1–3):239–254, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barbe:2006:AAS

[724] A. Barbé and F. von Haeseler. Averages of automatic sequences. Theoretical Computer Science, 359(1–3):255– 281, August 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bjorner:2006:RMF

[725] Anders Björner and Bruce E. Sagan. Rationality of the Möbius function of a composition poset. Theoretical Computer Science, 359(1–3):282– 298, August 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Xu:2006:PDB

[726] Guangjun Xu, Liying Kang, Erfang Shan, and Min Zhao. Power domination in block graphs. Theoretical Computer Science, 359(1–3):299– 305, August 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

McTague:2006:APD

[727] Carl S. McTague and James P. Crutchfield. Automated pattern detection— An algorithm for constructing optimally synchronizing multi-regular language filters. Theoretical Computer Science, 359(1–3):306–328, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nebel:2006:FSM

[728] Markus E. Nebel. Fast string matching by using probabilities: On an optimal mismatch variant of Horspool’s algorithm. Theoretical Computer Science, 359(1–3):329–343, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Perez-Diaz:2006:DBP

[729] Sonia Pérez-Díaz, Juana Sendra, and J. Rafael Sendra. Distance bounds of ϵ-points on hypersurfaces. Theoretical Computer Science, 359(1–3):344– 368, August 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Escoffier:2006:CAC

[730] Bruno Escoffier and Vangelis Th. Paschos. Completeness in approximation classes beyond APX. Theoretical Computer Science, 359(1–3):369– 377, August 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gorecki:2006:DTM

[731] Paweł Górecki and Jerzy Tiuryn. DLStrees: a model of evolutionary scenarios. Theoretical Computer Science, 359(1–3):378–399, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Efraimidis:2006:ASS

[732] Pavlos S. Efraimidis and Paul G. Spirakis. Approximation schemes for scheduling and covering on unrelated machines. Theoretical Computer Science, 359(1–3):400–417, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epstein:2006:CCS

[733] Leah Epstein and Asaf Levin. The conference call search problem in wireless networks. Theoretical Computer Science, 359(1–3):418–429, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chan:2006:NRA

[734] Wun-Tat Chan, Tak-Wah Lam, Kin-Shing Liu, and Prudence W. H. Wong. New resource augmentation analysis of the total stretch of SRPT and SJF in multiprocessor scheduling. Theoretical Computer Science, 359(1–3):430– 439, August 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ryabko:2006:AKC

[735] Boris Ryabko, Jaakko Astola, and Alex Gammerman. Application of Kolmogorov complexity and universal codes to identity testing and nonparametric testing of serial independence for time series. Theoretical Computer Science, 359(1–3):440–448, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Szabo:2006:CCR

[736] Sándor Szabó. Completing codes and the Rédei property of groups. Theoretical Computer Science, 359(1–3):449– 454, August 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dinu:2006:EAR

[737] Liviu P. Dinu and Florin Manea. An efficient approach for the rank aggregation problem. Theoretical Computer Science, 359(1–3):455–461, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:EBc

[738] Anonymous. Editorial board. Theoretical Computer Science, 359(1–3):v– ix, August 14, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Friggens:2006:MPT

[739] David Friggens and Robert Goldblatt. A modal proof theory for final polynomial coalgebras. Theoretical Computer Science, 360(1–3):1–22, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Binkley:2006:TFD

[740] Dave Binkley, Sebastian Danicic, Tibor Gyimóthy, Mark Harman, Ákos Kiss, and Bogdan Korel. Theoretical foundations of dynamic program slicing. Theoretical Computer Science, 360(1–3): 23–41, August 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bruscoli:2006:SPS

[741] Paola Bruscoli and Alessio Guglielmi. On structuring proof search for first order linear logic. Theoretical Computer Science, 360(1–3):42–76, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manes:2006:BRC

[742] Ernie Manes. Boolean restriction categories and taut monads. Theoretical Computer Science, 360(1–3):77– 95, August 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demri:2006:LIP

[743] Stéphane Demri. LTL over integer periodicity constraints. Theoretical Computer Science, 360(1–3):96– 123, August 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guiraud:2006:TPP

[744] Yves Guiraud. Two polygraphic presentations of Petri nets. Theoretical Computer Science, 360(1–3):124– 146, August 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhu:2006:CTI

[745] Zhaohui Zhu, Rong Zhang, and Shan Lu. A characterization theorem for injective model classes axiomatized by general rules. Theoretical Computer Science, 360(1–3):147–171, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lasota:2006:DPE

[746] Slawomir Lasota. Decidability of performance equivalence for basic parallel processes. Theoretical Computer Science, 360(1–3):172–192, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dowek:2006:EBD

[747] Gilles Dowek and Ying Jiang. Eigenvariables, bracketing and the decidability of positive minimal predicate logic. Theoretical Computer Science, 360(1– 3):193–208, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Luttgen:2006:BSU

[748] Gerald Lüttgen and Walter Vogler. Bisimulation on speed: a unified approach. Theoretical Computer Science, 360(1–3):209–227, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dougherty:2006:NFB

[749] Daniel J. Dougherty and Claudio Gutiérrez. Normal forms for binary relations. Theoretical Computer Science, 360(1–3):228–246, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nguyen:2006:MLP

[750] Linh Anh Nguyen. Multimodal logic programming. Theoretical Computer Science, 360(1–3):247–288, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tzevelekos:2006:IDC

[751] Nikos Tzevelekos. Investigations on the Dual Calculus. Theoretical Computer Science, 360(1–3):289–326, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hosoya:2006:BOI

[752] Haruo Hosoya and Makoto Murata. Boolean operations and inclusion test for attribute-element constraints. Theoretical Computer Science, 360(1–3): 327–351, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bozzelli:2006:MCP

[753] Laura Bozzelli. Model checking for process rewrite systems and a class of action-based regular properties. Theoretical Computer Science, 360(1–3): 352–372, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanBreugel:2006:ACB

[754] Franck van Breugel and James Worrell. Approximating and computing behavioural distances in probabilistic transition systems. Theoretical Computer Science, 360(1–3):373– 385, August 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bruni:2006:SFG

[755] Roberto Bruni and José Meseguer. Semantic foundations for generalized rewrite theories. Theoretical Computer Science, 360(1–3):386–414, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Michelbrink:2006:IFP

[756] Markus Michelbrink. Interfaces as functors, programs as coalgebras — a final coalgebra theorem in intensional type theory. Theoretical Computer Science, 360(1–3):415–439, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boreale:2006:PFP

[757] Michele Boreale and Fabio Gadducci. Processes as formal power series: a coinductive approach to denotational semantics. Theoretical Computer Science, 360(1–3):440–458, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hartmann:2006:EAF

[758] Sven Hartmann, Sebastian Link, and Klaus-Dieter Schewe. Erratum to “Axiomatisations of functional dependencies in the presence of records, lists, sets and multisets”. Theoretical Computer Science, 360(1–3):459, August 21, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [619].

Anonymous:2006:AIb

[759] Anonymous. Author index. Theoretical Computer Science, 360(1–3): 461, August 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:MI

[760] Anonymous. Master index. Theoretical Computer Science, 360(1–3):463– 472, August 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:EBd

[761] Anonymous. Editorial board. Theoretical Computer Science, 360(1–3):v– ix, August 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vose:2006:P

[762] Michael D. Vose. Preface. Theoretical Computer Science, 361(1):1, August 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Whitley:2006:SSL

[763] Darrell Whitley and Jonathan Rowe. Subthreshold-seeking local search. Theoretical Computer Science, 361(1): 2–17, August 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arnold:2006:WME

[764] Dirk V. Arnold. Weighted multirecombination evolution strategies. Theoretical Computer Science, 361(1):18– 37, August 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jagerskupper:2006:HUI

[765] Jens Jägersküpper. How the (1+1) ES using isotropic mutations minimizes positive definite quadratic forms. Theoretical Computer Science, 361(1):38– 56, August 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Toussaint:2006:CRS

[766] Marc Toussaint. Compact representations as a search strategy: Compression EDAs. Theoretical Computer Science, 361(1):57–71, August 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mitavskiy:2006:SRA

[767] Boris Mitavskiy and Jonathan Rowe. Some results about the Markov chains associated to GPs and general EAs. Theoretical Computer Science, 361(1): 72–110, August 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rowe:2006:DCG

[768] Jonathan E. Rowe, Michael D. Vose, and Alden H. Wright. Differentiable coarse graining. Theoretical Computer Science, 361(1):111–129, August 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:EBe

[769] Anonymous. Editorial board. Theoretical Computer Science, 361(1):v– ix, August 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:Pg

[770] Anonymous. Preface. Theoretical Computer Science, 361(1):v–ix, August 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jansen:2006:P

[771] Klaus Jansen and Roberto Solis-Oba. Preface. Theoretical Computer Science, 361(2–3):131–132, September 1, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marx:2006:MSM

[772] Dániel Marx. Minimum sum multicoloring on the edges of trees. Theoretical Computer Science, 361(2–3):133– 149, September 1, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krysta:2006:EAA

[773] Piotr Krysta and Krzysztof Loryś. Efficient approximation algorithms for the achromatic number. Theoretical Computer Science, 361(2–3):150–171, September 1, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demaine:2006:CCG

[774] Erik D. Demaine, Dotan Emanuel, Amos Fiat, and Nicole Immorlica. Correlation clustering in general weighted graphs. Theoretical Computer Science, 361(2–3):172–187, September 1, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gortz:2006:ACV

[775] Inge Li Gørtz and Anthony Wirth. Asymmetry in k-center variants. Theoretical Computer Science, 361(2–3): 188–199, September 1, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Awerbuch:2006:TWC

[776] Baruch Awerbuch, Yossi Azar, Yossi Richter, and Dekel Tsur. Tradeoffs in worst-case equilibria. Theoretical Computer Science, 361(2–3):200–209, September 1, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deKlerk:2006:PMP

[777] Etienne de Klerk, Monique Laurent, and Pablo A. Parrilo. A PTAS for the minimization of polynomials of fixed degree over the simplex. Theoretical Computer Science, 361(2–3):210– 225, September 1, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Decker:2006:AAS

[778] T. Decker, T. Lücking, and B. Monien. A 5/4-approximation algorithm for scheduling identical malleable tasks. Theoretical Computer Science, 361(2– 3):226–240, September 1, 2006. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Uma:2006:RBC

[779] R. N. Uma, Joel Wein, and David P. Williamson. On the relationship between combinatorial and LP-based lower bounds for NP-hard scheduling problems. Theoretical Computer Science, 361(2–3):241–256, September 1, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ageev:2006:OBS

[780] A. A. Ageev, A. V. Fishkin, A. V. Kononov, and S. V. Sevastyanov. Open block scheduling in optical communication networks. Theoretical Computer Science, 361(2–3):257–274, September 1, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fotakis:2006:IAF

[781] Dimitris Fotakis. Incremental algorithms for Facility Location and kmedian. Theoretical Computer Science, 361(2–3):275–313, September 1, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Azar:2006:LBT

[782] Yossi Azar, Amir Epstein, and Leah Epstein. Load balancing of temporary tasks in the ℓ_{p} norm. Theoretical Computer Science, 361(2–3):314– 328, September 1, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gutierrez:2006:HWM

[783] Sandra Gutiérrez, Sven O. Krumke, Nicole Megow, and Tjark Vredeveld. How to whack moles. Theoretical Computer Science, 361(2–3):329–341, September 1, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demaine:2006:OST

[784] Erik D. Demaine, Sándor P. Fekete, and Shmuel Gal. Online searching with turn cost. Theoretical Computer Science, 361(2–3):342–355, September 1, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:AIc

[785] Anonymous. Author index. Theoretical Computer Science, 361(2–3):356– 357, August 21, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fates:2006:FAB

[786] Nazim Fatès, Éric Thierry, Michel Morvan, and Nicolas Schabanel. Fully asynchronous behavior of double-quiescent elementary cellular automata. Theoretical Computer Science, 362(1–3):1– 16, October 11, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

El-Fakih:2006:PSP

[787] Khaled El-Fakih, Nina Yevtushenko, Sergey Buffalov, and Gregor v. Bochmann. Progressive solutions to a parallel automata equation. Theoretical Computer Science, 362(1–3):17–32, October 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rossello:2006:AVR

[788] Francesc Rosselló and Gabriel Valiente. An algebraic view of the relation between largest common subtrees and smallest common supertrees. Theoretical Computer Science, 362(1–3):33– 53, October 11, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hemaspaandra:2006:IPN

[789] Lane A. Hemaspaandra, Kari Pasanen, and Jörg Rothe. If P ≠ NP then some strongly noninvertible functions are invertible. Theoretical Computer Science, 362(1–3):54–62, October 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2006:CAM

[790] Guizhen Yang. Computational aspects of mining maximal frequent patterns. Theoretical Computer Science, 362(1– 3):63–85, October 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bollig:2006:PGD

[791] Beate Bollig, Stephan Waack, and Philipp Woelfel. Parity graph-driven read-once branching programs and an exponential lower bound for integer multiplication. Theoretical Computer Science, 362(1–3):86–99, October 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garcia:2006:V

[792] Pedro García and José Ruiz. V. Theoretical Computer Science, 362(1–3): 100–114, October 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ji:2006:SLD

[793] Min Ji, Yong He, and T. C. E. Cheng. Scheduling linear deteriorating jobs with an availability constraint on a single machine. Theoretical Computer Science, 362(1–3):115–126, October 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boyar:2006:MRP

[794] Joan Boyar, Leah Epstein, Lene M. Favrholdt, Jens S. Kohrt, Kim S. Larsen, Morten M. Pedersen, and Sanne Wøhlk. The maximum resource bin packing problem. Theoretical Computer Science, 362(1–3):127– 139, October 11, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Malvestuto:2006:MIS

[795] Francesco M. Malvestuto, Mauro Mezzini, and Marina Moscarini. Minimal invariant sets in a vertex-weighted graph. Theoretical Computer Science, 362(1–3):140–161, October 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheng:2006:IAM

[796] Chih-Huai Cheng, Kuan-Yu Chen, Wen-Chin Tien, and Kun-Mao Chao. Improved algorithms for the kmaximum-sums problems. Theoretical Computer Science, 362(1–3):162– 170, October 11, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Neary:2006:SFU

[797] Turlough Neary and Damien Woods. Small fast universal Turing machines. Theoretical Computer Science, 362(1– 3):171–195, October 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gasieniec:2006:DMM

[798] Leszek Ga̧sieniec, Evangelos Kranakis, Andrzej Pelc, and Qin Xin. Deterministic M2M multicast in radio networks. Theoretical Computer Science, 362(1–3):196–206, October 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Picantin:2006:FTD

[799] Matthieu Picantin. Finite transducers for divisibility monoids. Theoretical Computer Science, 362(1–3):207– 221, October 11, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salomaa:2006:ICQ

[800] Arto Salomaa. Independence of certain quantities indicating subword occurrences. Theoretical Computer Science, 362(1–3):222–231, October 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kambites:2006:WPR

[801] Mark Kambites. Word problems recognisable by deterministic blind monoid automata. Theoretical Computer Science, 362(1–3):232–237, October 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hussein:2006:UNO

[802] Mohamed Eid Hussein and Uwe Schwiegelshohn. Utilization of nonclairvoyant online schedules. Theoretical Computer Science, 362(1–3):238– 247, October 11, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crochemore:2006:LRB

[803] Maxime Crochemore, Costas S. Iliopoulos, Manal Mohamed, and Marie-France Sagot. Longest repeats with a block of k don’t cares. Theoretical Computer Science, 362(1–3):248– 254, October 11, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bretto:2006:CTG

[804] Alain Bretto, Alain Faisant, and Thierry Vallée. Compatible topologies on graphs: an application to graph isomorphism problem complexity. Theoretical Computer Science, 362(1–3): 255–272, October 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheng:2006:MAS

[805] T. C. E. Cheng, C. T. Ng, and J. J. Yuan. Multi-agent scheduling on a single machine to minimize total weighted number of tardy jobs. Theoretical Computer Science, 362(1–3):273–281, October 11, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deLuca:2006:PCO

[806] Aldo de Luca and Alessandro De Luca. Pseudopalindrome closure operators in free monoids. Theoretical Computer Science, 362(1–3):282–300, October 11, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:EBf

[807] Anonymous. Editorial board. Theoretical Computer Science, 362(1–3):v– ix, October 11, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chwa:2006:P

[808] Kyung-Yong Chwa and J. Ian Munro. Preface. Theoretical Computer Science, 363(1):1, October 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ratsaby:2006:CH

[809] Joel Ratsaby. Complexity of hyperconcepts. Theoretical Computer Science, 363(1):2–10, October 25, 2006.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adamy:2006:CO

[810] Udo Adamy, Michael Hoffmann, József Solymosi, and Miloš Stojaković. Coloring octrees. Theoretical Computer Science, 363(1):11–17, October 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Azar:2006:IAO

[811] Yossi Azar, Amos Fiat, Meital Levy, and N. S. Narayanaswamy. An improved algorithm for online coloring of intervals with bandwidth. Theoretical Computer Science, 363(1):18– 27, October 25, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tomita:2006:WCT

[812] Etsuji Tomita, Akira Tanaka, and Haruhisa Takahashi. The worst-case time complexity for generating all maximal cliques and computational experiments. Theoretical Computer Science, 363(1):28–42, October 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tsukiji:2006:CPR

[813] Tatsuie Tsukiji and Zhi-Zhong Chen. Computing phylogenetic roots with bounded degrees and errors is NPcomplete. Theoretical Computer Science, 363(1):43–59, October 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jansson:2006:ILP

[814] Jesper Jansson and Wing-Kin Sung. Inferring a level-1 phylogenetic network from a dense set of rooted triplets. Theoretical Computer Science, 363(1):60– 68, October 25, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leung:2006:EAO

[815] H. F. Leung, Z. S. Peng, and H. F. Ting. An efficient algorithm for online square detection. Theoretical Computer Science, 363(1):69–75, October 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Woelfel:2006:CMO

[816] Philipp Woelfel. A construction method for optimally universal hash families and its consequences for the existence of RBIBDs. Theoretical Computer Science, 363(1):76–84, October 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lefmann:2006:LTD

[817] Hanno Lefmann. Large triangles in the d-dimensional unit cube. Theoretical Computer Science, 363(1):85– 98, October 25, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2006:AMW

[818] Yong Zhang, Qi Ge, Rudolf Fleischer, Tao Jiang, and Hong Zhu. Approximating the minimum weight weak vertex cover. Theoretical Computer Science, 363(1):99–105, October 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:Ph

[819] Anonymous. Preface. Theoretical Computer Science, 363(1):v–ix, October 25, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Farre:2006:E

[820] Jacques Farré and Igor Litovsky. Editorial. Theoretical Computer Science, 363(2):107, October 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bailly:2006:CCP

[821] Arnaud Bailly, Mireille Clerbout, and Isabelle Simplot-Ryl. Component composition preserving behavioral contracts based on communication traces. Theoretical Computer Science, 363(2): 108–123, October 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bastien:2006:PNF

[822] Cédric Bastien, Jurek Czyzowicz, Wojciech Fraczak, and Wojciech Rytter. Prime normal form and equivalence of simple grammars. Theoretical Computer Science, 363(2):124– 134, October 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Campeanu:2006:ICM

[823] Cezar Câmpeanu, Andrei Păun, and Jason R. Smith. Incremental construction of minimal deterministic finite cover automata. Theoretical Computer Science, 363(2):135–148, October 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2006:DCS

[824] Oscar H. Ibarra and Hsu-Chun Yen. Deterministic catalytic systems are not universal. Theoretical Computer Science, 363(2):149–161, October 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jurdzinski:2006:RAR

[825] Tomasz Jurdziński and Friedrich Otto. Restarting automata with restricted utilization of auxiliary symbols. Theoretical Computer Science, 363(2):162– 181, October 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klein:2006:EDA

[826] Joachim Klein and Christel Baier. Experiments with deterministic ωautomata for formulas of linear temporal logic. Theoretical Computer Science, 363(2):182–195, October 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lohrey:2006:CTA

[827] Markus Lohrey and Sebastian Maneth. The complexity of tree automata and XPath on grammar-compressed trees. Theoretical Computer Science, 363(2): 196–210, October 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rytter:2006:SSG

[828] Wojciech Rytter. The structure of subword graphs and suffix trees of Fibonacci words. Theoretical Computer Science, 363(2):211–223, October 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Althoff:2006:ODB

[829] Christoph Schulte Althoff, Wolfgang Thomas, and Nico Wallmeier. Observations on determinization of Büchi automata. Theoretical Computer Science, 363(2):224–233, October 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tamm:2006:SRT

[830] Hellis Tamm, Matti Nykänen, and Esko Ukkonen. On size reduction techniques for multitape automata. Theoretical Computer Science, 363(2):234– 246, October 28, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Megiddo:2006:P

[831] Nimrod Megiddo, Yinfeng Xu, and Binhai Zhu. Preface. Theoretical Computer Science, 363(3):247, October 31, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cai:2006:CAF

[832] Mao cheng Cai, Xiaotie Deng, and Zhongfei Li. Computation of arbitrage in frictional bond markets. Theoretical Computer Science, 363(3):248– 256, October 31, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

He:2006:IAT

[833] Yong He, Weiya Zhong, and Huikun Gu. Improved algorithms for two single machine scheduling problems. Theoretical Computer Science, 363(3):257– 265, October 31, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lin:2006:GBW

[834] Mingen Lin, Zhiyong Lin, and Jinhui Xu. Graph bandwidth of weighted caterpillars. Theoretical Computer Science, 363(3):266–277, October 31, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2006:PCT

[835] Jianer Chen and Fenghui Zhang. On product covering in 3-tier supply chain models: Natural complete problems for W. Theoretical Computer Science, 363(3):278–288, October 31, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gabor:2006:AAF

[836] Adriana F. Gabor and Jan-Kees C. W. van Ommeren. Approximation algorithms for facility location problems with a special class of subadditive cost functions. Theoretical Computer Science, 363(3):289–300, October 31, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gavalda:2006:F

[837] Richard Gavaldà and Eiji Takimoto. Foreword. Theoretical Computer Science, 364(1):1–2, November 2, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nouretdinov:2006:CCT

[838] Ilia Nouretdinov and Vladimir Vovk. Criterion of calibration for transductive confidence machine with limited feedback. Theoretical Computer Science, 364(1):3–9, November 2, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vovk:2006:WCP

[839] Vladimir Vovk. Well-calibrated predictions from on-line compression models. Theoretical Computer Science, 364(1): 10–26, November 2, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hutter:2006:GCU

[840] Marcus Hutter. On generalized computable universal priors and their convergence. Theoretical Computer Science, 364(1):27–41, November 2, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zilles:2006:AIC

[841] Sandra Zilles. An approach to intrinsic complexity of uniform learning. Theoretical Computer Science, 364(1):42– 61, November 2, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martin:2006:OVD

[842] Eric Martin, Arun Sharma, and Frank Stephan. On ordinal VC-dimension and some notions of complexity. Theoretical Computer Science, 364(1):62– 76, November 2, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zeugmann:2006:LLS

[843] Thomas Zeugmann. From learning in the limit to stochastic finite learning. Theoretical Computer Science, 364(1): 77–97, November 2, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Uemura:2006:LEP

[844] Jin Uemura and Masako Sato. Learning of erasing primitive formal systems from positive examples. Theoretical Computer Science, 364(1):98– 114, November 2, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Case:2006:LSR

[845] John Case, Sanjay Jain, Rüdiger Reischuk, Frank Stephan, and Thomas Zeugmann. Learning a subclass of regular patterns in polynomial time. Theoretical Computer Science, 364(1):115– 131, November 2, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kitagawa:2006:SEK

[846] Genshiro Kitagawa. Signal extraction and knowledge discovery based on statistical modeling. Theoretical Computer Science, 364(1):132–142, November 2, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:EBg

[847] Anonymous. Editorial board. Theoretical Computer Science, 364(1):v– ix, November 2, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deQueiroz:2006:P

[848] Ruy de Queiroz and Patrick Cégielski. Preface. Theoretical Computer Science, 364(2):143–145, November 6, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bellin:2006:CPT

[849] Gianluigi Bellin, Martin Hyland, Edmund Robinson, and Christian Urban. Categorical proof theory of classical propositional calculus. Theoretical Computer Science, 364(2):146– 165, November 6, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrhard:2006:DIN

[850] T. Ehrhard and L. Regnier. Differential interaction nets. Theoretical Computer Science, 364(2):166–195, November 6, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Finkel:2006:DPL

[851] Olivier Finkel. On decidability properties of local sentences. Theoretical Computer Science, 364(2):196–211, November 6, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hartmann:2006:DIF

[852] Sven Hartmann and Sebastian Link. Deciding implication for functional dependencies in complex-value databases. Theoretical Computer Science, 364(2): 212–240, November 6, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kaminski:2006:DTM

[853] Michael Kaminski and Julia Rubin-Mosin. Default theories over monadic languages. Theoretical Computer Science, 364(2):241–253, November 6, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Power:2006:GMC

[854] John Power. Generic models for computational effects. Theoretical Computer Science, 364(2):254–269, November 6, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hofmann:2006:P

[855] Martin Hofmann and Hans-Wolfgang Loidl. Preface. Theoretical Computer Science, 364(3):271–272, November 8, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Besson:2006:PCC

[856] Frédéric Besson, Thomas Jensen, and David Pichardie. Proof-carrying code from certified abstract interpretation and fixpoint compression. Theoretical Computer Science, 364(3):273–291, November 8, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laud:2006:TSE

[857] Peeter Laud, Tarmo Uustalu, and Varmo Vene. Type systems equivalent to data-flow analyses for imperative languages. Theoretical Computer Science, 364(3):292–310, November 8, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kennedy:2006:SNP

[858] Andrew Kennedy. Securing the .NET programming model. Theoretical Computer Science, 364(3):311–317, November 8, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cansell:2006:FIC

[859] Dominique Cansell and Dominique Méry. Formal and incremental construction of distributed algorithms: On the distributed reference counting algorithm. Theoretical Computer Science, 364(3):318–337, November 8, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Niehren:2006:CLC

[860] J. Niehren, J. Schwinghammer, and G. Smolka. A concurrent lambda calculus with futures. Theoretical Computer Science, 364(3):338–356, November 8, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deBoer:2006:P

[861] Frank de Boer and Marcello Bonsangue. Preface. Theoretical Computer Science, 365(1–2):1, November 10, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barbosa:2006:TPC

[862] Luís S. Barbosa and José N. Oliveira. Transposing partial components—An exercise on coalgebraic refinement. Theoretical Computer Science, 365(1– 2):2–22, November 10, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Johnsen:2006:CTS

[863] Einar Broch Johnsen, Olaf Owe, and Ingrid Chieh Yu. Creol: a typesafe object-oriented model for distributed concurrent systems. Theoretical Computer Science, 365(1–2):23– 66, November 10, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chatterjee:2006:GSE

[864] Krishnendu Chatterjee, Thomas A. Henzinger, and Marcin Jurdziński. Games with secure equilibria. Theoretical Computer Science, 365(1–2):67– 82, November 10, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheung:2006:SPP

[865] Ling Cheung, Nancy Lynch, Roberto Segala, and Frits Vaandrager. Switched PIOA: Parallel composition via distributed scheduling. Theoretical Computer Science, 365(1–2):83–108, November 10, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jifeng:2006:RRC

[866] He Jifeng, Xiaoshan Li, and Zhiming Liu. rCOS: a refinement calculus of object systems. Theoretical Computer Science, 365(1–2):109–142, November 10, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Naumann:2006:TIM

[867] David A. Naumann and Mike Barnett. Towards imperative modules: Reasoning about invariants and sharing of mutable state. Theoretical Computer Science, 365(1–2):143–168, November 10, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:EBh

[868] Anonymous. Editorial board. Theoretical Computer Science, 365(1–2):v– ix, November 10, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kopperman:2006:F

[869] Ralph Kopperman, Prakash Panangaden, Michael B. Smyth, Dieter Spreen, and Julian Webster. Foreword. Theoretical Computer Science, 365(3):169–170, November 12, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abbes:2006:PTB

[870] Samy Abbes and Klaus Keimel. Projective topology on bifinite domains and applications. Theoretical Computer Science, 365(3):171–183, November 12, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abe:2006:CTI

[871] K. Abe, J. Bisceglio, D. R. Ferguson, T. J. Peters, A. C. Russell, and T. Sakkalis. Computational topology for isotopic surface reconstruction. Theoretical Computer Science, 365(3): 184–198, November 12, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fajstrup:2006:CLP

[872] Lisbeth Fajstrup. Cubical local partial orders on cubically subdivided spaces—Existence and construction. Theoretical Computer Science, 365(3): 199–205, November 12, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gratus:2006:SVI

[873] Jonathan Gratus and Timothy Porter. A spatial view of information. Theoretical Computer Science, 365(3):206– 215, November 12, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hamrin:2006:TCE

[874] G. Hamrin and V. Stoltenberg-Hansen. Two categories of effective continuous cpos. Theoretical Computer Science, 365(3):216–236, November 12, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kunzi:2006:PQM

[875] H.-P. A. Künzi, H. Pajoohesh, and M. P. Schellekens. Partial quasimetrics. Theoretical Computer Science, 365(3):237–246, November 12, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Raussen:2006:DDM

[876] Martin Raußen. Deadlocks and dihomotopy in mutual exclusion models. Theoretical Computer Science, 365(3): 247–257, November 12, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Selivanov:2006:TDS

[877] Victor L. Selivanov. Towards a descriptive set theory for domain-like structures. Theoretical Computer Science, 365(3):258–282, November 12, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fiadeiro:2006:P

[878] José Fiadeiro and Jan Rutten. Preface. Theoretical Computer Science, 366(1– 2):1–2, November 16, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Milius:2006:CTS

[879] Stefan Milius and Lawrence S. Moss. The category-theoretic solution of recursive program schemes. Theoretical Computer Science, 366(1–2):3–59, November 16, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See corrigendum [1707].

Aceto:2006:BFB

[880] Luca Aceto, Wan Fokkink, Anna Ingolfsdottir, and Sumit Nain. Bisimilarity is not finitely based over BPA with interrupt. Theoretical Computer Science, 366(1–2):60–81, November 16, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adamek:2006:IAC

[881] J. Adámek. The intersection of algebra and coalgebra. Theoretical Computer Science, 366(1–2):82–97, November 16, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bruni:2006:BAS

[882] Roberto Bruni, Ivan Lanese, and Ugo Montanari. A basic algebra of stateless connectors. Theoretical Computer Science, 366(1–2):98–120, November 16, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hausmann:2006:CAS

[883] Daniel Hausmann, Till Mossakowski, and Lutz Schröder. A coalgebraic approach to the semantics of the ambient calculus. Theoretical Computer Science, 366(1–2):121–143, November 16, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hyland:2006:DLT

[884] Martin Hyland and John Power. Discrete Lawvere theories and computational effects. Theoretical Computer Science, 366(1–2):144–162, November 16, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Thati:2006:CSR

[885] Prasanna Thati and José Meseguer. Complete symbolic reachability analysis using back-and-forth narrowing. Theoretical Computer Science, 366(1– 2):163–179, November 16, 2006. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:EBi

[886] Anonymous. Editorial board. Theoretical Computer Science, 366(1–2):v– ix, November 16, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esik:2006:P

[887] Zoltán Ésik. Preface. Theoretical Computer Science, 366(3):181, November 20, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Domaratzki:2006:CDM

[888] Michael Domaratzki and Kai Salomaa. Codes defined by multiple sets of trajectories. Theoretical Computer Science, 366(3):182–193, November 20, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Domosi:2006:APL

[889] Pál Dömösi and Géza Horváth. Alternative proof of the Lyndon– Schützenberger Theorem. Theoretical Computer Science, 366(3):194–198, November 20, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Droste:2006:SIF

[890] Manfred Droste and Dietrich Kuske. Skew and infinitary formal power series. Theoretical Computer Science, 366(3):199–227, November 20, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Droste:2006:WTA

[891] Manfred Droste and Heiko Vogler. Weighted tree automata and weighted logics. Theoretical Computer Science, 366(3):228–247, November 20, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maletti:2006:CTS

[892] Andreas Maletti. Compositions of tree series transformations. Theoretical Computer Science, 366(3):248–271, November 20, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mraz:2006:MCG

[893] F. Mráz, F. Otto, M. Plátek, and T. Jurdziński. Marcus t-contextual grammars and cut hierarchies and monotonicity for restarting automata. Theoretical Computer Science, 366(3):272– 296, November 20, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Truthe:2006:FST

[894] Bianca Truthe. The finiteness of synchronous, tabled picture languages is decidable. Theoretical Computer Science, 366(3):297–315, November 20, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Degano:2006:P

[895] Pierpaolo Degano and Luca Viganò. Preface. Theoretical Computer Science, 367(1–2):1, November 24, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abadi:2006:DKS

[896] Martín Abadi and Véronique Cortier. Deciding knowledge in security protocols under equational theories. Theoretical Computer Science, 367(1–2):2– 32, November 24, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Backes:2006:CAC

[897] Michael Backes, Anupam Datta, Ante Derek, John C. Mitchell, and Mathieu Turuani. Compositional analysis of contract-signing protocols. Theoretical Computer Science, 367(1–2):33– 56, November 24, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Butler:2006:FAK

[898] Frederick Butler, Iliano Cervesato, Aaron D. Jaggard, Andre Scedrov, and Christopher Walstad. Formal analysis of Kerberos 5. Theoretical Computer Science, 367(1–2):57–87, November 24, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caleiro:2006:SAS

[899] Carlos Caleiro, Luca Viganò, and David Basin. On the semantics of Alice&Bob specifications of security protocols. Theoretical Computer Science, 367(1–2):88–122, November 24, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chatzikokolakis:2006:PIR

[900] Konstantinos Chatzikokolakis and Catuscia Palamidessi. Probable innocence revisited. Theoretical Computer Science, 367(1–2):123–138, November 24, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cremers:2006:ISE

[901] C. J. F. Cremers, S. Mauw, and E. P. de Vink. Injective synchronisation: an extension of the authentication hierarchy. Theoretical Computer Science, 367(1–2):139–161, November 24, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Escobar:2006:RBI

[902] Santiago Escobar, Catherine Meadows, and José Meseguer. A rewriting-based inference system for the NRL Protocol Analyzer and its meta-logical properties. Theoretical Computer Science, 367(1–2):162–202, November 24, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nanz:2006:FSA

[903] Sebastian Nanz and Chris Hankin. A framework for security analysis of mobile wireless networks. Theoretical Computer Science, 367(1–2):203–227, November 24, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ramanujam:2006:RQE

[904] R. Ramanujam and S. P. Suresh. A (restricted) quantifier elimination for security protocols. Theoretical Computer Science, 367(1–2):228–256, November 24, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Steel:2006:FAP

[905] Graham Steel. Formal analysis of PIN block attacks. Theoretical Computer Science, 367(1–2):257–270, November 24, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:EBj

[906] Anonymous. Editorial board. Theoretical Computer Science, 367(1–2):v– ix, November 24, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2006:Pb

[907] G. Rozenberg. Preface. Theoretical Computer Science, 367(3):271–272, December 1, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2006:CFN

[908] Yiguang Liu, Zhisheng You, and Liping Cao. A concise functional neural network computing the largest modulus eigenvalues and their corresponding eigenvectors of a real skew matrix. Theoretical Computer Science, 367(3):273– 285, December 1, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lv:2006:GCO

[909] Jian Cheng Lv, Zhang Yi, and K. K. Tan. Global convergence of Oja’s PCA learning algorithm with a non-zero-approaching adaptive learning rate. Theoretical Computer Science, 367(3): 286–307, December 1, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Izmerly:2006:CCA

[910] Oleg Izmerly and Tal Mor. Chosen ciphertext attacks on lattice-based public key encryption and modern (nonquantum) cryptography in a quantum environment. Theoretical Computer Science, 367(3):308–323, December 1, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carpentieri:2006:GSB

[911] Marco Carpentieri. A genetic system based on simulated crossover of sequences of two-bit genes. Theoretical Computer Science, 367(3):324– 335, December 1, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khan:2006:SMQ

[912] Faisal Shah Khan and Marek Perkowski. Synthesis of multi-qudit hybrid and dvalued quantum logic circuits by decomposition. Theoretical Computer Science, 367(3):336–346, December 1, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [1690].

Anonymous:2006:EBk

[913] Anonymous. Editorial board. Theoretical Computer Science, 367(3):iii– vii, December 1, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Drossopoulou:2006:FMD

[914] Sophia Drossopoulou, Giovanni Lagorio, and Susan Eisenbach. A flexible model for dynamic linking in Java and C#. Theoretical Computer Science, 368(1–2):1–29, December 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krotzsch:2006:GUS

[915] Markus Krötzsch. Generalized ultrametric spaces in quantitative domain theory. Theoretical Computer Science, 368(1–2):30–49, December 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kaminski:2006:IUS

[916] Michael Kaminski. Invariance under stuttering in a temporal logic of actions. Theoretical Computer Science, 368(1–2):50–63, December 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vasconcelos:2006:TCM

[917] Vasco T. Vasconcelos, Simon J. Gay, and António Ravara. Type checking a multithreaded functional language with session types. Theoretical Computer Science, 368(1–2):64– 87, December 5, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Afrati:2006:RQU

[918] Foto Afrati, Chen Li, and Prasenjit Mitra. Rewriting queries using views in the presence of arithmetic comparisons. Theoretical Computer Science, 368(1–2):88–123, December 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2006:LAS

[919] Yi-Xiang Chen and Achim Jung. A logical approach to stable domains. Theoretical Computer Science, 368(1–2): 124–148, December 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2006:FFR

[920] Zaiyue Zhang, Yuefei Sui, Cungen Cao, and Guohua Wu. A formal fuzzy reasoning system and reasoning mechanism based on propositional modal logic. Theoretical Computer Science, 368(1–2):149–160, December 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delaune:2006:URA

[921] Stéphanie Delaune. An undecidability result for AGh. Theoretical Computer Science, 368(1–2):161–167, December 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanGlabbeek:2006:EEH

[922] R. J. van Glabbeek. Erratum to “On the expressiveness of higher dimensional automata”: [Theoret. Comput. Sci. 356 (2006) 265–290]. Theoretical Computer Science, 368(1–2):168– 194, December 5, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [659].

Anonymous:2006:EBl

[923] Anonymous. Editorial board. Theoretical Computer Science, 368(1–2):v– ix, December 5, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sahinalp:2006:P

[924] Suleyman Cenk Sahinalp, Ugur Dogrusoz, and S. Muthukrishnan. Preface. Theoretical Computer Science, 368(3): 195, December 10, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amir:2006:FTD

[925] Amihood Amir, Oren Kapah, and Dekel Tsur. Faster two-dimensional pattern matching with rotations. Theoretical Computer Science, 368(3):196– 204, December 10, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Davydov:2006:CMR

[926] Eugene Davydov and Serafim Batzoglou. A computational model for RNA multiple structural alignment. Theoretical Computer Science, 368(3): 205–216, December 10, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lonardi:2006:FBR

[927] Stefano Lonardi, Wojciech Szpankowski, and Qiaofeng Yang. Finding biclusters by random projections. Theoretical Computer Science, 368(3):217– 230, December 10, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Geary:2006:SOR

[928] Richard F. Geary, Naila Rahman, Rajeev Raman, and Venkatesh Raman. A simple optimal representation for balanced parentheses. Theoretical Computer Science, 368(3):231– 246, December 10, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jurdzinski:2006:DNM

[929] T. Jurdziński, F. Mráz, F. Otto, and M. Plátek. Degrees of nonmonotonicity for restarting automata. Theoretical Computer Science, 369(1– 3):1–34, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pikhurko:2006:EBG

[930] Oleg Pikhurko and Jerzy Wojciechowski. Edge-bandwidth of grids and tori. Theoretical Computer Science, 369(1–3):35–43, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2006:ESS

[931] Jing-Chao Chen. Efficient sample sort and the average case analysis of PEsort. Theoretical Computer Science, 369(1–3):44–66, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bordihn:2006:IST

[932] Henning Bordihn, Henning Fernau, Markus Holzer, Vincenzo Manca, and Carlos Martín-Vide. Iterated sequential transducers as language generating devices. Theoretical Computer Science, 369(1–3):67–81, December 15, 2006.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Eiter:2006:RUM

[933] Thomas Eiter and Georg Gottlob. Reasoning under minimal upper bounds in propositional logic. Theoretical Computer Science, 369(1–3):82–115, December 15, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gairing:2006:PAP

[934] Martin Gairing, Thomas Lücking, Marios Mavronicolas, and Burkhard Monien. The price of anarchy for polynomial social cost. Theoretical Computer Science, 369(1–3):116–135, December 15, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wu:2006:IDT

[935] Bang Ye Wu. On the intercluster distance of a tree metric. Theoretical Computer Science, 369(1–3):136– 141, December 15, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Terrier:2006:LCC

[936] Véronique Terrier. Low complexity classes of multidimensional cellular automata. Theoretical Computer Science, 369(1–3):142–156, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Angel:2006:TAS

[937] Eric Angel, Evripidis Bampis, and Fanny Pascual. Truthful algorithms for scheduling selfish tasks on parallel machines. Theoretical Computer Science, 369(1–3):157–168, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blundo:2006:VCS

[938] Carlo Blundo, Stelvio Cimato, and Alfredo De Santis. Visual cryptography schemes with optimal pixel expansion. Theoretical Computer Science, 369(1–3):169–182, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Domaratzki:2006:APS

[939] Michael Domaratzki, Petr Sosík, and Alfonso Rodríguez-Patón. Algebraic properties of substitution on trajectories. Theoretical Computer Science, 369(1–3):183–196, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rostamabadi:2006:LUA

[940] Farshad Rostamabadi and Mohammad Ghodsi. Label updating to avoid pointshaped obstacles in fixed model. Theoretical Computer Science, 369(1–3): 197–210, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Travers:2006:CMP

[941] Stephen Travers. The complexity of membership problems for circuits over sets of integers. Theoretical Computer Science, 369(1–3):211–229, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kellerer:2006:FPA

[942] Hans Kellerer and Vitaly A. Strusevich. A fully polynomial approximation scheme for the single machine weighted total tardiness problem with a common due date. Theoretical Computer Science, 369(1–3):230–238, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bebek:2006:DDG

[943] G. Bebek, P. Berenbrink, C. Cooper, T. Friedetzky, J. Nadeau, and S. C. Sahinalp. The degree distribution of the generalized duplication model. Theoretical Computer Science, 369(1– 3):239–249, December 15, 2006. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hou:2006:LEL

[944] Jianfeng Hou, Guizhen Liu, and Jiansheng Cai. List edge and list total colorings of planar graphs without 4cycles. Theoretical Computer Science, 369(1–3):250–255, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yan:2006:EST

[945] Weigen Yan and Yeong-Nan Yeh. Enumeration of subtrees of trees. Theoretical Computer Science, 369(1–3):256– 268, December 15, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bilo:2006:SCM

[946] V. Bilò, M. Flammini, G. Melideo, L. Moscardelli, and A. Navarra. Sharing the cost of multicast transmissions in wireless networks. Theoretical Computer Science, 369(1–3):269– 284, December 15, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epstein:2006:RSP

[947] Leah Epstein and Alex Kesselman. On the remote server problem or more about TCP acknowledgments. Theoretical Computer Science, 369(1–3):285– 299, December 15, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoffmann:2006:GCA

[948] Michael Hoffmann and Richard M. Thomas. A geometric characterization of automatic semigroups. Theoretical Computer Science, 369(1–3):300– 313, December 15, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Forys:2006:AAP

[949] Wit Foryś and Tomasz Krawczyk. An algorithmic approach to the problem of a semiretract base. Theoretical Computer Science, 369(1–3):314–322, December 15, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arora:2006:SRA

[950] Amrinder Arora, Fanchun Jin, and Hyeong-Ah Choi. Scheduling resource allocation with timeslot penalty for changeover. Theoretical Computer Science, 369(1–3):323–337, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koch:2006:CAS

[951] Ronald Koch and Ines Spenke. Complexity and approximability of ksplittable flows. Theoretical Computer Science, 369(1–3):338–347, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dolev:2006:DLB

[952] Shlomi Dolev, Roberto Segala, and Alexander Shvartsman. Dynamic load balancing with group communication. Theoretical Computer Science, 369(1– 3):348–360, December 15, 2006. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pion:2006:CRB

[953] Sylvain Pion and Chee K. Yap. Constructive root bound for k-ary rational input numbers. Theoretical Computer Science, 369(1–3):361–376, December 15, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Azriel:2006:QLR

[954] D. Azriel and D. Berend. On a question of Leiss regarding the Hanoi Tower problem. Theoretical Computer Science, 369(1–3):377–383, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Levin:2006:PMT

[955] Asaf Levin and Danny Segev. Partial multicuts in trees. Theoretical Computer Science, 369(1–3):384–395, December 15, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Acuna:2006:CVB

[956] Vicente Acuña, Gilles Didier, and Alejandro Maass. Coding with variable block maps. Theoretical Computer Science, 369(1–3):396–405, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Branco:2006:TAA

[957] Mário J. J. Branco. Two algebraic approaches to variants of the concatenation product. Theoretical Computer Science, 369(1–3):406–426, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gonzalez:2006:PED

[958] Teofilo F. Gonzalez and David Serena. Pairwise edge disjoint shortest paths in the n-cube. Theoretical Computer Science, 369(1–3):427–435, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuppusamy:2006:NAI

[959] Lakshmanan Kuppusamy. A note on ambiguity of internal contextual grammars. Theoretical Computer Science, 369(1–3):436–441, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2006:CDC

[960] Wei-Mei Chen. Cost distribution of the Chang–Roberts leader election algorithm and related problems. Theoretical Computer Science, 369(1–3):442– 447, December 15, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dai:2006:ERS

[961] Wenqiang Dai, Yinfeng Xu, and Binhai Zhu. On the edge l_{∞} radius of Saitou and Nei’s method for phylogenetic reconstruction. Theoretical Computer Science, 369(1–3):448–455, December 15, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhao:2006:NFA

[962] Wenbo Zhao, Peng Zhang, and Tao Jiang. A network flow approach to the Minimum Common Integer Partition Problem. Theoretical Computer Science, 369(1–3):456–462, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gajardo:2006:CIT

[963] A. Gajardo and E. Goles. Crossing information in two-dimensional Sandpiles. Theoretical Computer Science, 369(1–3):463–469, December 15, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2006:EBm

[964] Anonymous. Editorial board. Theoretical Computer Science, 369(1–3):v– ix, December 15, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nicolas:2007:LCS

[965] François Nicolas and Eric Rivals. Longest common subsequence problem for unoriented and cyclic strings. Theoretical Computer Science, 370(1–3):1– 18, February 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:2007:EAA

[966] Jin Wook Kim and Kunsoo Park. An efficient alignment algorithm for masked sequences. Theoretical Computer Science, 370(1–3):19–33, February 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gajardo:2007:OHM

[967] A. Gajardo and J. Mazoyer. One Head Machines from a symbolic approach. Theoretical Computer Science, 370(1– 3):34–47, February 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Avigad:2007:QER

[968] Jeremy Avigad and Yimu Yin. Quantifier elimination for the reals with a predicate for the powers of two. Theoretical Computer Science, 370(1–3):48– 59, February 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Glasser:2007:CDN

[969] Christian Glaßer, Alan L. Selman, and Liyu Zhang. Canonical disjoint NPpairs of propositional proof systems. Theoretical Computer Science, 370(1– 3):60–73, February 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Giakoumakis:2007:AMP

[970] Vassilis Giakoumakis and Stephan Olariu. All minimal prime extensions of hereditary classes of graphs. Theoretical Computer Science, 370(1–3):74– 93, February 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Saunders:2007:SSP

[971] Shane Saunders and Tadao Takaoka. Solving shortest paths efficiently on nearly acyclic directed graphs. Theoretical Computer Science, 370(1–3):94– 109, February 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Han:2007:OSR

[972] Yo-Sub Han and Derick Wood. Obtaining shorter regular expressions from finite-state automata. Theoretical Computer Science, 370(1–3):110–120, February 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pan:2007:IPP

[973] Jun-Jie Pan and Gerard J. Chang. Induced-path partition on graphs with special blocks. Theoretical Computer Science, 370(1–3):121–130, February 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Almeida:2007:ATA

[974] Jorge Almeida and Marc Zeitoun. An automata-theoretic approach to the word problem for ω-terms over R. Theoretical Computer Science, 370(1–3): 131–169, February 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leupold:2007:LGI

[975] Peter Leupold. Languages generated by iterated idempotency. Theoretical Computer Science, 370(1–3):170– 185, February 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Catalano:2007:CDS

[976] Dario Catalano and Rosario Gennaro. Cramer-damgård signatures revisited: Efficient flat-tree signatures based on factoring. Theoretical Computer Science, 370(1–3):186–200, February 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frosini:2007:BMU

[977] Andrea Frosini and Maurice Nivat. Binary matrices under the microscope: a tomographical problem. Theoretical Computer Science, 370(1–3):201– 217, February 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chandran:2007:RBA

[978] L. Sunil Chandran and L. Shankar Ram. On the relationship between ATSP and the cycle cover problem. Theoretical Computer Science, 370(1–3):218–228, February 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brandstadt:2007:NAC

[979] Andreas Brandstädt, Van Bang Le, and Suhail Mahfud. New applications of clique separator decomposition for the Maximum Weight Stable Set problem. Theoretical Computer Science, 370(1–3):229–239, February 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheng:2007:STC

[980] Yongxi Cheng, Xi Chen, and Yiqun Lisa Yin. On searching a table consistent with division poset. Theoretical Computer Science, 370(1–3):240– 253, February 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Patt-Shamir:2007:NEA

[981] Boaz Patt-Shamir. A note on efficient aggregate queries in sensor networks. Theoretical Computer Science, 370(1–3):254–264, February 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuba:2007:LRI

[982] Markus Kuba and Alois Panholzer. The left-right-imbalance of binary search trees. Theoretical Computer Science, 370(1–3):265–278, February 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pelc:2007:FCB

[983] Andrzej Pelc and David Peleg. Feasibility and complexity of broadcasting with random transmission failures. Theoretical Computer Science, 370(1–3):279–292, February 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lu:2007:IHA

[984] Chi-Jen Lu, Shi-Chun Tsai, and Hsin-Lung Wu. Improved hardness amplification in NP. Theoretical Computer Science, 370(1–3):293–298, February 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hsieh:2007:FML

[985] Sun-Yuan Hsieh. Finding maximal leafagreement isomorphic descendent subtrees from phylogenetic trees with different species. Theoretical Computer Science, 370(1–3):299–308, February 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBa

[986] Anonymous. Editorial board. Theoretical Computer Science, 370(1–3):v– ix, February 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Back:2007:CNS

[987] Thomas Bäck and Benedikt Löwe. Computing and the natural sciences at CiE 2005. Theoretical Computer Science, 371(1–2):1–3, February 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beggs:2007:CNS

[988] Edwin J. Beggs and John V. Tucker. Can Newtonian systems, bounded in space, time, mass and energy compute all functions? Theoretical Computer Science, 371(1–2):4–19, February 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bernardini:2007:QSP

[989] Francesco Bernardini, Marian Gheorghe, and Natalio Krasnogor. Quorum sensing P systems. Theoretical Computer Science, 371(1–2):20–33, February 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garcez:2007:CML

[990] Artur S. d’Avila Garcez, Luís C. Lamb, and Dov M. Gabbay. Connectionist modal logic: Representing modalities in neural networks. Theoretical Computer Science, 371(1–2):34– 53, February 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gutierrez-Naranjo:2007:USS

[991] Miguel A. Gutiérrez-Naranjo, Mario J. Pérez-Jiménez, and Francisco J. Romero-Campero. A uniform solution to SAT using membrane creation. Theoretical Computer Science, 371(1–2): 54–61, February 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Michalak:2007:IDD

[992] Krzysztof Michalak and Halina Kwaśnicka. Influence of data dimensionality on the quality of forecasts given by a multilayer perceptron. Theoretical Computer Science, 371(1–2):62–71, February 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manea:2007:ANS

[993] Florin Manea, Carlos Martín-Vide, and Victor Mitrana. Accepting networks of splicing processors: Complexity results. Theoretical Computer Science, 371(1–2):72–82, February 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [1135].

Krishna:2007:URP

[994] Shankara Narayanan Krishna. Universality results for P systems based on brane calculi operations. Theoretical Computer Science, 371(1–2):83– 105, February 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Trautteur:2007:NDV

[995] Giuseppe Trautteur and Guglielmo Tamburrini. A note on discreteness and virtuality in analog computing. Theoretical Computer Science, 371(1–2): 106–114, February 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tucker:2007:CAN

[996] John V. Tucker and Jeffery I. Zucker. Computability of analog networks. Theoretical Computer Science, 371(1– 2):115–146, February 22, 2007. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBb

[997] Anonymous. Editorial board. Theoretical Computer Science, 371(1–2):v– ix, February 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Eiter:2007:P

[998] Thomas Eiter and Leonid Libkin. Preface. Theoretical Computer Science, 371 (3):147, March 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koltun:2007:ADR

[999] Vladlen Koltun and Christos H. Papadimitriou. Approximately dominating representatives. Theoretical Computer Science, 371(3):148–154, March 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Atserias:2007:CQE

[1000] Albert Atserias. Conjunctive query evaluation by search-tree revisited.

Theoretical Computer Science, 371(3): 155–168, March 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calvanese:2007:VBQ

[1001] Diego Calvanese, Giuseppe De Giacomo, Maurizio Lenzerini, and Moshe Y. Vardi. View-based query processing: On the relationship between rewriting, answering and losslessness. Theoretical Computer Science, 371(3): 169–182, March 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

VandenBussche:2007:WDS

[1002] Jan Van den Bussche, Dirk Van Gucht, and Stijn Vansummeren. Welldefinedness and semantic type-checking for the nested relational calculus. Theoretical Computer Science, 371(3):183– 199, March 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Deutsch:2007:RQU

[1003] Alin Deutsch, Bertram Ludäscher, and Alan Nash. Rewriting queries using views with access patterns under integrity constraints. Theoretical Computer Science, 371(3):200–226, March 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Matias:2007:OWB

[1004] Yossi Matias and Daniel Urieli. Optimal workload-based weighted wavelet synopses. Theoretical Computer Science, 371(3):227–246, March 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balcazar:2007:HAS

[1005] José L. Balcázar and Gemma C. Garriga. Horn axiomatizations for sequential data. Theoretical Computer Science, 371(3):247–264, March 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dolinka:2007:AIB

[1006] Igor Dolinka. Axiomatizing the identities of binoid languages. Theoretical Computer Science, 372(1):1–14, March 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leve:2007:QSW

[1007] F. Levé and G. Richomme. Quasiperiodic Sturmian words and morphisms. Theoretical Computer Science, 372(1): 15–25, March 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berman:2007:OTM

[1008] Piotr Berman, Marek Karpinski, and Yakov Nekrich. Optimal trade-off for Merkle tree traversal. Theoretical Computer Science, 372(1):26–36, March 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bell:2007:MID

[1009] Paul Bell and Igor Potapov. On the membership of invertible diagonal and scalar matrices. Theoretical Computer Science, 372(1):37–45, March 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pivato:2007:RCL

[1010] Marcus Pivato. RealLife: The continuum limit of Larger than Life cellular automata. Theoretical Computer Science, 372(1):46–68, March 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tan:2007:OSO

[1011] Zhiyi Tan and Yong Wu. Optimal semionline algorithms for machine covering. Theoretical Computer Science, 372(1): 69–80, March 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ishii:2007:MCS

[1012] Toshimasa Ishii, Hitoshi Fujita, and Hiroshi Nagamochi. Minimum cost source location problem with local 3vertex-connectivity requirements. Theoretical Computer Science, 372(1):81– 93, March 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ye:2007:LSM

[1013] Deshi Ye and Guochuan Zhang. Online scheduling mesh jobs with dependencies. Theoretical Computer Science, 372(1):94–102, March 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dou:2007:RMT

[1014] Wenqing Dou and Enyu Yao. On rearrangeable multirate three-stage Clos networks. Theoretical Computer Science, 372(1):103–107, March 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gurski:2007:CRG

[1015] Frank Gurski. Characterizations for restricted graphs of NLC-width 2. Theoretical Computer Science, 372(1):108– 114, March 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferragina:2007:SSS

[1016] Paolo Ferragina and Rossano Venturini. A simple storage scheme for strings achieving entropy bounds. Theoretical Computer Science, 372(1):115– 121, March 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBc

[1017] Anonymous. Editorial board. Theoretical Computer Science, 372(1):v– ix, March 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paun:2007:FBW

[1018] Gheorghe Păun and Mario J. Pérez-Jiménez. Fourth Brainstorming Week on Membrane Computing. Theoretical Computer Science, 372(2–3):123– 124, March 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Busi:2007:UWS

[1019] Nadia Busi. Using well-structured transition systems to decide divergence for catalytic P systems. Theoretical Computer Science, 372(2–3):125– 135, March 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cavaliere:2007:MRC

[1020] Matteo Cavaliere, Rudolf Freund, Marion Oswald, and Dragoş Sburlan. Multiset random context grammars, checkers, and transducers. Theoretical Computer Science, 372(2–3):136– 151, March 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Csuhaj-Varju:2007:SUA

[1021] Erzsébet Csuhaj-Varjú, Maurice Margenstern, György Vaszil, and Sergey Verlan. On small universal antiport P systems. Theoretical Computer Science, 372(2–3):152–164, March 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fontana:2007:DSD

[1022] Federico Fontana and Vincenzo Manca.

Discrete solutions to differential equations by metabolic P systems. Theoretical Computer Science, 372(2–3):165– 182, March 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gutierrez-Naranjo:2007:DPM

[1023] Miguel A. Gutiérrez-Naranjo, Mario J.

Pérez-Jiménez, and Agustín Riscos-Núñez. On the degree of parallelism in membrane systems. Theoretical Computer Science, 372(2–3):183– 195, March 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2007:NFS

[1024] Oscar H. Ibarra, Andrei Păun, Gheorghe Păun, Alfonso Rodríguez-Patón, Petr Sosík, and Sara Woodworth. Normal forms for spiking neural P systems. Theoretical Computer Science, 372(2– 3):196–217, March 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leporati:2007:TQA

[1025] Alberto Leporati and Sara Felloni. Three “quantum” algorithms to solve 3-SAT. Theoretical Computer Science, 372(2–3):218–241, March 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muskulus:2007:CCC

[1026] Michael Muskulus, Daniela Besozzi, Robert Brijder, Paolo Cazzaniga, Sanne Houweling, Dario Pescini, and Grzegorz Rozenberg. Cycles and communicating classes in membrane systems and molecular dynamics. Theoretical Computer Science, 372(2–3): 242–266, March 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBd

[1027] Anonymous. Editorial board. Theoretical Computer Science, 372(2–3): i–v, March 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Danicic:2007:ELF

[1028] Sebastian Danicic, Mark Harman, Rob Hierons, John Howroyd, and Michael R. Laurence. Equivalence of linear, free, liberal, structured program schemas is decidable in polynomial time. Theoretical Computer Science, 373(1–2):1–18, March 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Luttgen:2007:CPF

[1029] Gerald Lüttgen and Walter Vogler. Conjunction on processes: Full abstraction via ready-tree semantics. Theoretical Computer Science, 373(1–2):19– 40, March 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barbanera:2007:SAA

[1030] Franco Barbanera, Michele Bugliesi, Mariangiola Dezani-Ciancaglini, and Vladimiro Sassone. Space-aware ambients and processes. Theoretical Computer Science, 373(1–2):41–69, March 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Clavel:2007:RME

[1031] Manuel Clavel, José Meseguer, and Miguel Palomino. Reflection in membership equational logic, many-sorted equational logic, Horn logic with equality, and rewriting logic. Theoretical Computer Science, 373(1–2):70– 91, March 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Deng:2007:APF

[1032] Yuxin Deng and Catuscia Palamidessi.

Axiomatizations for probabilistic finitestate behaviors. Theoretical Computer Science, 373(1–2):92–114, March 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lago:2007:CTA

[1033] Ugo Dal Lago, Angelo Montanari, and Gabriele Puppis. Compact and tractable automaton-based representations of time granularities. Theoretical Computer Science, 373(1–2):115– 141, March 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stubbe:2007:TDD

[1034] Isar Stubbe. Towards “dynamic domains”: Totally continuous cocomplete Q-categories. Theoretical Computer Science, 373(1–2):142–160, March 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBe

[1035] Anonymous. Editorial board. Theoretical Computer Science, 373(1–2): v–ix, March 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mosses:2007:P

[1036] Peter D. Mosses and Irek Ulidowski. Preface. Theoretical Computer Science, 373(3):161–162, April 5, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Andrei:2007:RLF

[1037] Oana Andrei, Gabriel Ciobanu, and Dorel Lucanu. A rewriting logic framework for operational semantics of membrane systems. Theoretical Computer Science, 373(3):163–181, April 5, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hym:2007:ARD

[1038] Samuel Hym and Matthew Hennessy. Adding recursion to Dpi. Theoretical Computer Science, 373(3):182– 212, April 5, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meseguer:2007:RLS

[1039] José Meseguer and Grigore Roşu. The rewriting logic semantics project. Theoretical Computer Science, 373(3):213– 237, April 5, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mousavi:2007:SFM

[1040] MohammadReza Mousavi, Michel A.

Reniers, and Jan Friso Groote. SOS formats and meta-theory: 20 years after. Theoretical Computer Science, 373 (3):238–272, April 5, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Saabas:2007:CNS

[1041] Ando Saabas and Tarmo Uustalu. A compositional natural semantics and Hoare logic for low-level languages. Theoretical Computer Science, 373(3): 273–302, April 5, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Murakami:2007:FAC

[1042] M. Murakami, M. Hara, M. Yamamoto, and S. Tani. Fast algorithms for computing Jones polynomials of certain links. Theoretical Computer Science, 374(1–3):1–24, April 20, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nicolay:2007:AFL

[1043] S. Nicolay and M. Rigo. About frequencies of letters in generalized automatic sequences. Theoretical Computer Science, 374(1–3):25–40, April 20, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wu:2007:KPP

[1044] Biao Wu and Enyue Yao. kpartitioning problems with partition matroid constraint. Theoretical Computer Science, 374(1–3):41–48, April 20, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tian:2007:LSD

[1045] Ji Tian, Ruyan Fu, and Jinjiang Yuan. On-line scheduling with delivery time on a single batch machine. Theoretical Computer Science, 374(1–3):49– 57, April 20, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Simpson:2007:IPW

[1046] Jamie Simpson. Intersecting periodic words. Theoretical Computer Science, 374(1–3):58–65, April 20, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheng:2007:LGP

[1047] Yongxi Cheng. Lattice grids and prisms are antimagic. Theoretical Computer Science, 374(1–3):66–73, April 20, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Csuhaj-Varju:2007:SAS

[1048] Erzsébet Csuhaj-Varjú, Ion Petre, and György Vaszil. Self-assembly of strings and languages. Theoretical Computer Science, 374(1–3):74–81, April 20, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chang:2007:AGC

[1049] Gerard J. Chang, Sen-Peng Eu, and Chung-Heng Yeh. On the (n, t)- antipodal Gray codes. Theoretical Computer Science, 374(1–3):82– 90, April 20, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rodrigues:2007:MAF

[1050] Estela M. Rodrigues, Marie-France Sagot, and Yoshiko Wakabayashi. The maximum agreement forest problem: Approximation algorithms and computational experiments. Theoretical Computer Science, 374(1–3):91–110, April 20, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Eriksen:2007:RTM

[1051] Niklas Eriksen. Reversal and transposition medians. Theoretical Computer Science, 374(1–3):111–126, April 20, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuske:2007:WAC

[1052] Dietrich Kuske. Weighted asynchronous cellular automata. Theoretical Computer Science, 374(1–3):127– 148, April 20, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alon:2007:AMC

[1053] Noga Alon, Andrzej Lingas, and Martin Wahlen. Approximating the maximum clique minor and some subgraph homeomorphism problems. Theoretical Computer Science, 374(1–3):149– 158, April 20, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2007:SJA

[1054] L. L. Liu, C. T. Ng, and T. C. E. Cheng. Scheduling jobs with agreeable processing times and due dates on a single batch processing machine. Theoretical Computer Science, 374(1–3): 159–169, April 20, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Her:2007:EMD

[1055] Jun-Ho Her and R. S. Ramakrishna. An external-memory depth-first search algorithm for general grid graphs. Theoretical Computer Science, 374(1–3): 170–180, April 20, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yan:2007:CRL

[1056] Qiqi Yan. Classifying regular languages by a split game. Theoretical Computer Science, 374(1–3):181–190, April 20, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2007:HAM

[1057] Wenbin Chen and Jiangtao Meng. Hardness of approximating the Minimum Solutions of Linear Diophantine Equations. Theoretical Computer Science, 374(1–3):191–195, April 20, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fu:2007:OSP

[1058] Ruyan Fu, Tian Ji, Jinjiang Yuan, and Yixun Lin. Online scheduling in a parallel batch processing system to minimize makespan using restarts. Theoretical Computer Science, 374(1–3):

196–202, April 20, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kralovic:2007:RGS

[1059] R. Královič and P. Ružička. Ranks of graphs: The size of acyclic orientation cover for deadlock-free packet routing. Theoretical Computer Science, 374(1–3):203–213, April 20, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maurer:2007:CRP

[1060] Ina Mäurer. Characterizations of recognizable picture series. Theoretical Computer Science, 374(1–3):214– 228, April 20, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Catalano:2007:HCT

[1061] Dario Catalano and Ivan Visconti. Hybrid commitments and their applications to zero-knowledge proof systems. Theoretical Computer Science, 374(1– 3):229–260, April 20, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cimato:2007:CVC

[1062] S. Cimato, R. De Prisco, and A. De Santis. Colored visual cryptography without color darkening. Theoretical Computer Science, 374(1–3):261– 276, April 20, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mycka:2007:NCF

[1063] Jerzy Mycka and José Félix Costa. A new conceptual framework for analog computation. Theoretical Computer Science, 374(1–3):277–290, April 20, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBf

[1064] Anonymous. Editorial board. Theoretical Computer Science, 374(1–3): v–ix, April 20, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Danvy:2007:P

[1065] Olivier Danvy, Peter O’Hearn, and Philip Wadler. Preface. Theoretical Computer Science, 375(1–3):1–2, May 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hermida:2007:FFP

[1066] C. Hermida and R. D. Tennent. A fibrational framework for possible-world semantics of Algol. Theoretical Computer Science, 375(1–3):3–19, May 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hyland:2007:CAE

[1067] Martin Hyland, Paul Blain Levy, Gordon Plotkin, and John Power. Combining algebraic effects with continuations. Theoretical Computer Science, 375(1– 3):20–40, May 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Filinski:2007:RBM

[1068] Andrzej Filinski. On the relations between monadic semantics. Theoretical Computer Science, 375(1–3): 41–75, May 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biernacka:2007:SCB

[1069] Małgorzata Biernacka and Olivier Danvy. A syntactic correspondence between context-sensitive calculi and abstract machines. Theoretical Computer Science, 375(1–3):76–108, May 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jones:2007:SAS

[1070] C. B. Jones. Splitting atoms safely.

Theoretical Computer Science, 375(1– 3):109–119, May 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jones:2007:FAL

[1071] Neil D. Jones and Nils Andersen. Flow analysis of lazy higher-order functional programs. Theoretical Computer Science, 375(1–3):120–136, May 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Might:2007:AES

[1072] Matthew Might and Olin Shivers. Analyzing the environment structure of higher-order languages using frame strings. Theoretical Computer Science, 375(1–3):137–168, May 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sumii:2007:BDS

[1073] Eijiro Sumii and Benjamin C. Pierce. A bisimulation for dynamic sealing. Theoretical Computer Science, 375(1–3): 169–192, May 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Freyd:2007:CAR

[1074] Peter Freyd. Core algebra revisited. Theoretical Computer Science, 375(1– 3):193–200, May 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wadler:2007:GRI

[1075] Philip Wadler. The Girard–Reynolds isomorphism (second edition). Theoretical Computer Science, 375(1–3): 201–226, May 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brookes:2007:SCS

[1076] Stephen Brookes. A semantics for concurrent separation logic. Theoretical Computer Science, 375(1–3):227– 270, May 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

OHearn:2007:RCL

[1077] Peter W. O’Hearn. Resources, concurrency, and local reasoning. Theoretical Computer Science, 375(1–3): 271–307, May 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2007:RSL

[1078] Hongseok Yang. Relational separation logic. Theoretical Computer Science, 375(1–3):308–334, May 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Morris:2007:FET

[1079] F. Lockwood Morris. A few exercises in theorem processing. Theoretical Computer Science, 375(1–3):335–345, May 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Oles:2007:BSJ

[1080] Frank J. Oles. On being a student of John Reynolds. Theoretical Computer Science, 375(1–3):346–350, May 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBg

[1081] Anonymous. Editorial board. Theoretical Computer Science, 375(1–3): v–ix, May 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2007:DLT

[1082] Oscar H. Ibarra. Developments in language theory. Theoretical Computer Science, 376(1–2):1–2, May 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:2007:EMR

[1083] A. Ehrenfeucht and G. Rozenberg.

Events and modules in reaction systems. Theoretical Computer Science, 376(1–2):3–16, May 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gurevich:2007:CAS

[1084] Yuri Gurevich, Margus Veanes, and Charles Wallace. Can abstract state machines be useful in language theory? Theoretical Computer Science, 376(1– 2):17–29, May 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ananichev:2007:SAL

[1085] D. S. Ananichev, M. V. Volkov, and Yu. I. Zaks. Synchronizing automata with a letter of deficiency 2. Theoretical Computer Science, 376(1–2): 30–41, May 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bastien:2007:ESF

[1086] Cédric Bastien, Jurek Czyzowicz, Wojciech Fraczak, and Wojciech Rytter. Equivalence of simple functions. Theoretical Computer Science, 376(1–2): 42–51, May 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carton:2007:GRS

[1087] Olivier Carton. The growth ratio of synchronous rational relations is unique. Theoretical Computer Science, 376(1–2):52–59, May 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Han:2007:EPD

[1088] Yo-Sub Han, Arto Salomaa, Kai Salomaa, Derick Wood, and Sheng Yu. On the existence of prime decompositions. Theoretical Computer Science, 376(1– 2):60–69, May 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krieger:2007:CEF

[1089] Dalia Krieger. On critical exponents in fixed points of non-erasing morphisms. Theoretical Computer Science, 376(1– 2):70–88, May 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kufleitner:2007:PFT

[1090] Manfred Kufleitner. Polynomials, fragments of temporal logic and the variety DA over traces. Theoretical Computer Science, 376(1–2):89–100, May 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kutrib:2007:CDN

[1091] Martin Kutrib and Andreas Malcher.

Context-dependent nondeterminism for pushdown automata. Theoretical Computer Science, 376(1–2):101– 111, May 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okhotin:2007:LEC

[1092] Alexander Okhotin and Oksana Yakimova. Language equations with complementation: Decision problems. Theoretical Computer Science, 376(1–2): 112–126, May 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ravikumar:2007:SVT

[1093] Bala Ravikumar. On some variations of two-way probabilistic finite automata models. Theoretical Computer Science, 376(1–2):127–136, May 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBh

[1094] Anonymous. Editorial board. Theoretical Computer Science, 376(1–2): v–ix, May 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cerioli:2007:F

[1095] Maura Cerioli, Tiziana Margaria, and Michel Wermelinger. Foreword. Theoretical Computer Science, 376(3):137– 138, May 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deLara:2007:AGT

[1096] Juan de Lara, Roswitha Bardohl, Hartmut Ehrig, Karsten Ehrig, Ulrike Prange, and Gabriele Taentzer. Attributed graph transformation with node type inheritance. Theoretical Computer Science, 376(3):139–163, May 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Furia:2007:ACP

[1097] Carlo A. Furia, Matteo Rossi, Dino Mandrioli, and Angelo Morzenti. Automated compositional proofs for realtime systems. Theoretical Computer Science, 376(3):164–184, May 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Roman:2007:MAB

[1098] Gruia-Catalin Roman, Christine Julien, and Jamie Payton. Modeling adaptive behaviors in Context UNITY. Theoretical Computer Science, 376(3):185– 204, May 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Naumann:2007:OPE

[1099] David A. Naumann. Observational purity and encapsulation. Theoretical Computer Science, 376(3):205– 224, May 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salehi:2007:TAV

[1100] Saeed Salehi and Magnus Steinby. Tree algebras and varieties of tree languages. Theoretical Computer Science, 377(1– 3):1–24, May 31, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2007:ALP

[1101] Zhao Zhang and Hao Li. Algorithms for long paths in graphs. Theoretical Computer Science, 377(1–3):25–34, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Avidor:2007:MMC

[1102] Adi Avidor and Michael Langberg. The multi-multiway cut problem. Theoretical Computer Science, 377(1–3): 35–42, May 31, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Uchida:2007:ABG

[1103] Jiro Uchida, Wei Chen, and Koichi Wada. Acknowledged broadcasting and gossiping in ad hoc radio networks. Theoretical Computer Science, 377(1– 3):43–54, May 31, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epstein:2007:PCF

[1104] Leah Epstein, Yanir Kleiman, Jiří Sgall, and Rob van Stee. Paging with connections: FIFO strikes again. Theoretical Computer Science, 377(1–3): 55–64, May 31, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marco:2007:BBM

[1105] Ana Marco and José-Javier Martínez. Bernstein–Bezoutian matrices and curve implicitization. Theoretical Computer Science, 377(1–3):65–72, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DAlessandro:2007:WQO

[1106] Flavio D’Alessandro, Gwénaël Richomme, and Stefano Varricchio. Well quasi-orders generated by a wordshuffle rewriting. Theoretical Computer Science, 377(1–3):73–92, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beyersdorff:2007:CRD

[1107] Olaf Beyersdorff. Classes of representable disjoint NP. Theoretical Computer Science, 377(1–3):93–109, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tan:2007:SOS

[1108] Zhiyi Tan and Yong He. Semi-online scheduling problems on two identical machines with inexact partial information. Theoretical Computer Science, 377(1–3):110–125, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Becher:2007:TUA

[1109] Verónica Becher, Santiago Figueira, and Rafael Picchi. Turing’s unpublished algorithm for normal numbers. Theoretical Computer Science, 377(1– 3):126–138, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nock:2007:SIG

[1110] Richard Nock and Frank Nielsen. Selfimproved gaps almost everywhere for the agnostic approximation of monomials. Theoretical Computer Science, 377 (1–3):139–150, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lin:2007:RAS

[1111] Tien-Ching Lin and D. T. Lee. Randomized algorithm for the sum selection problem. Theoretical Computer Science, 377(1–3):151–156, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Uscka-Wehlou:2007:DLI

[1112] Hanna Uscka-Wehlou. Digital lines with irrational slopes. Theoretical Computer Science, 377(1–3):157–169, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Park:2007:PPH

[1113] Jung-Heum Park, Hyeong-Seok Lim, and Hee-Chul Kim. Panconnectivity and pancyclicity of hypercube-like interconnection networks with faulty elements. Theoretical Computer Science, 377(1–3):170–180, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diedrich:2007:FSA

[1114] Florian Diedrich and Klaus Jansen.

Faster and simpler approximation algorithms for mixed packing and covering problems. Theoretical Computer Science, 377(1–3):181–204, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pivato:2007:DPK

[1115] Marcus Pivato. Defect particle kinematics in one-dimensional cellular automata. Theoretical Computer Science, 377(1–3):205–228, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bengtsson:2007:RMS

[1116] Fredrik Bengtsson and Jingsen Chen. Ranking k maximum sums. Theoretical Computer Science, 377(1–3):229– 237, May 31, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lai:2007:EGG

[1117] Yung-Ling Lai and Chang-Sin Tian. An extremal graph with given bandwidth. Theoretical Computer Science, 377(1– 3):238–242, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rogers:2007:CTP

[1118] A. Rogers, R. K. Dash, S. D. Ramchurn, P. Vytelingum, and N. R. Jennings. Coordinating team players within a noisy Iterated Prisoner’s Dilemma tournament. Theoretical Computer Science, 377(1–3):243– 259, May 31, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rao:2007:MPP

[1119] Michaël Rao. MSOL partitioning problems on graphs of bounded treewidth and clique-width. Theoretical Computer Science, 377(1–3):260–267, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rybalov:2007:SGU

[1120] Alexander Rybalov. On the strongly generic undecidability of the Halting Problem. Theoretical Computer Science, 377(1–3):268–270, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaz:2007:MCM

[1121] Josep Díaz and Marcin Kamiński.

MAX-CUT and MAX-BISECTION are NP-hard on unit disk graphs. Theoretical Computer Science, 377(1–3): 271–276, May 31, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cleary:2007:RUB

[1122] Sean Cleary, Fabrizio Luccio, and Linda Pagli. Refined upper bounds for right-arm rotation distances. Theoretical Computer Science, 377(1–3): 277–281, May 31, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yokomori:2007:EPT

[1123] Takashi Yokomori. Erratum to “Polynomial-time identification of very simple grammars from positive data” [Theoret. Comput. Sci. 298 (2003) 179–206]. Theoretical Computer Science, 377(1–3):282–283, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [2].

Anonymous:2007:EBi

[1124] Anonymous. Editorial board. Theoretical Computer Science, 377(1–3): v–ix, May 31, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2007:Pa

[1125] G. Rozenberg. Preface. Theoretical Computer Science, 378(1):1–2, June 3, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daley:2007:CDB

[1126] Mark Daley and Michael Domaratzki. On codes defined by bio-operations. Theoretical Computer Science, 378(1): 3–16, June 3, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brun:2007:ACT

[1127] Yuriy Brun. Arithmetic computation in the tile assembly model: Addition and multiplication. Theoretical Computer Science, 378(1):17–31, June 3, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Neumann:2007:RLS

[1128] Frank Neumann and Ingo Wegener. Randomized local search, evolutionary algorithms, and the minimum spanning tree problem. Theoretical Computer Science, 378(1):32–40, June 3, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ambainis:2007:IAQ

[1129] Andris Ambainis, Kazuo Iwama, Akinori Kawachi, Rudy Raymond, and Shigeru Yamashita. Improved algorithms for quantum identification of Boolean oracles. Theoretical Computer Science, 378(1):41–53, June 3, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Verel:2007:FLC

[1130] S. Verel, P. Collard, M. Tomassini, and L. Vanneschi. Fitness landscape of the cellular automata majority problem: View from the “Olympus”. Theoretical Computer Science, 378(1):54–77, June 3, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Talbi:2007:BSS

[1131] El-Ghazali Talbi and Benjamin Weinberg. Breaking the search space symmetry in partitioning problems: an application to the graph coloring problem. Theoretical Computer Science, 378(1):78–86, June 3, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Peng:2007:CAO

[1132] Dezhong Peng and Zhang Yi. Convergence analysis of the OJAn MCA learning algorithm by the deterministic discrete time method. Theoretical Computer Science, 378(1):87–100, June 3, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kashefi:2007:SZK

[1133] Elham Kashefi and Iordanis Kerenidis. Statistical Zero Knowledge and quantum one-way functions. Theoretical Computer Science, 378(1):101– 116, June 3, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ciobanu:2007:PSM

[1134] Gabriel Ciobanu, Linqiang Pan, Gheorghe Păun, and Mario J. Pérez-Jiménez. P systems with minimal parallelism. Theoretical Computer Science, 378(1):117–130, June 3, 2007.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manea:2007:EAN

[1135] Florin Manea, Carlos Martín-Vide, and Victor Mitrana. Erratum to: “Accepting networks of splicing processors: Complexity results” [Theoret. Comput. Sci. 371 (2007) 72–82]. Theoretical Computer Science, 378(1):131, June 3, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [993].

Anonymous:2007:EBj

[1136] Anonymous. Editorial board. Theoretical Computer Science, 378(1):v– ix, June 3, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Deng:2007:P

[1137] Xiaotie Deng. Preface. Theoretical Computer Science, 378(2):133, June 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ye:2007:EME

[1138] Yinyu Ye. Exchange market equilibria with Leontief’s utility: Freedom of pricing leads to rationality. Theoretical Computer Science, 378(2):134– 142, June 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garg:2007:PDA

[1139] Dinesh Garg, Kamal Jain, Kunal Talwar, and Vijay V. Vazirani. A primal-dual algorithm for computing Fisher equilibrium in the absence of gross substitutability property. Theoretical Computer Science, 378(2):143– 152, June 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kapoor:2007:ABM

[1140] Sanjiv Kapoor, Aranyak Mehta, and Vijay Vazirani. An auction-based market equilibrium algorithm for a production model. Theoretical Computer Science, 378(2):153–164, June 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fischer:2007:SCW

[1141] Simon Fischer and Berthold Vöcking.

On the structure and complexity of worst-case equilibria. Theoretical Computer Science, 378(2):165–174, June 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Feigenbaum:2007:SCP

[1142] Joan Feigenbaum, David R. Karger, Vahab S. Mirrokni, and Rahul Sami. Subjective-cost policy routing. Theoretical Computer Science, 378(2):175– 189, June 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheng:2007:NAB

[1143] Gang Cheng, Ping Li, and Peng Shi.

A new algorithm based on copulas for VaR valuation with empirical calculations. Theoretical Computer Science, 378(2):190–197, June 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhu:2007:EIB

[1144] Robert W. Zhu, Guomin Yang, and Duncan S. Wong. An efficient identitybased key exchange protocol with KGS forward secrecy for low-power devices. Theoretical Computer Science, 378(2): 198–207, June 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2007:P

[1145] Zhi-Zhong Chen, Xiaotie Deng, and Ding zhu Du. Preface. Theoretical Computer Science, 378(3):209– 210, June 9, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ganguly:2007:CDI

[1146] Sumit Ganguly. Counting distinct items over update streams. Theoretical Computer Science, 378(3):211– 222, June 9, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2007:LIS

[1147] Erdong Chen, Linji Yang, and Hao Yuan. Longest increasing subsequences in windows based on canonical antichain partition. Theoretical Computer Science, 378(3):223–236, June 9, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ben-Moshe:2007:EAC

[1148] Boaz Ben-Moshe, Binay Bhattacharya, Qiaosheng Shi, and Arie Tamir. Efficient algorithms for center problems in cactus networks. Theoretical Computer Science, 378(3):237–252, June 9, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ackermann:2007:DMB

[1149] Heiner Ackermann, Alantha Newman, Heiko Röglin, and Berthold Vöcking.

Decision-making based on approximate and smoothed Pareto curves. Theoretical Computer Science, 378(3):253– 270, June 9, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kao:2007:ACA

[1150] Ming-Yang Kao, Xiang-Yang Li, and Weizhao Wang. Average case analysis for tree labelling schemes. Theoretical Computer Science, 378(3):271– 291, June 9, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manthey:2007:SAB

[1151] Bodo Manthey and Rüdiger Reischuk.

Smoothed analysis of binary search trees. Theoretical Computer Science, 378(3):292–315, June 9, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2007:CAM

[1152] Lan Liu, Chen Xi, Jing Xiao, and Tao Jiang. Complexity and approximation of the minimum recombinant haplotype configuration problem. Theoretical Computer Science, 378(3):316– 330, June 9, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Groote:2007:OSP

[1153] Jan Friso Groote and Marc Voorhoeve. Operational semantics for Petri net components. Theoretical Computer Science, 379(1–2):1–19, June 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Japaridze:2007:TCI

[1154] Giorgi Japaridze. From truth to computability II. Theoretical Computer Science, 379(1–2):20–52, June 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martins:2007:CPC

[1155] Manuel A. Martins. Closure properties for the class of behavioral models. Theoretical Computer Science, 379(1– 2):53–83, June 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kanovich:2007:SPU

[1156] Max Kanovich and Jacqueline Vauzeilles. Strong planning under uncertainty in domains with numerous but identical elements (a generic approach). Theoretical Computer Science, 379(1– 2):84–119, June 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marcial-Romero:2007:SSL

[1157] J. Raymundo Marcial-Romero and Martín H. Escardó. Semantics of a sequential language for exact realnumber computation. Theoretical Computer Science, 379(1–2):120–141, June 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chadha:2007:RAP

[1158] R. Chadha, L. Cruz-Filipe, P. Mateus, and A. Sernadas. Reasoning about probabilistic sequential programs. Theoretical Computer Science, 379(1–2): 142–165, June 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vaux:2007:DC

[1159] Lionel Vaux. The differential λμ calculus. Theoretical Computer Science, 379 (1–2):166–209, June 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaconescu:2007:UPW

[1160] Răzvan Diaconescu and Petros Stefaneas. Ultraproducts and possible worlds semantics in institutions. Theoretical Computer Science, 379(1–2): 210–230, June 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asarin:2007:AAP

[1161] Eugene Asarin, Gerardo Schneider, and Sergio Yovine. Algorithmic analysis of polygonal hybrid systems, part I: Reachability. Theoretical Computer Science, 379(1–2):231–265, June 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dawar:2007:GAM

[1162] Anuj Dawar and Stephan Kreutzer.

Generalising automaticity to modal properties of finite structures. Theoretical Computer Science, 379(1–2): 266–285, June 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bozzelli:2007:CRB

[1163] Laura Bozzelli. Complexity results on branching-time pushdown model checking. Theoretical Computer Science, 379(1–2):286–297, June 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vagvolgyi:2007:LR

[1164] Sándor Vágvölgyi. Losing recognizability. Theoretical Computer Science, 379 (1–2):298–304, June 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBk

[1165] Anonymous. Editorial board. Theoretical Computer Science, 379(1–2): v–ix, June 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeten:2007:Pa

[1166] Jos C. M. Baeten, Jan Karel Lenstra, and Gerhard J. Woeginger. Preface. Theoretical Computer Science, 379(3): 305, June 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berger:2007:DDF

[1167] Noam Berger, Béla Bollobás, Christian Borgs, Jennifer Chayes, and Oliver Riordan. Degree distribution of the FKP network model. Theoretical Computer Science, 379(3):306–316, June 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bansal:2007:PTA

[1168] Vipul Bansal, Aseem Agrawal, and Varun S. Malhotra. Polynomial time algorithm for an optimal stable assignment with multiple partners. Theoretical Computer Science, 379(3):317– 328, June 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jagerskupper:2007:AAB

[1169] Jens Jägersküpper. Algorithmic analysis of a basic evolutionary algorithm for continuous optimization. Theoretical Computer Science, 379(3):329– 347, June 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bleichenbacher:2007:DIR

[1170] Daniel Bleichenbacher, Aggelos Kiayias, and Moti Yung. Decoding interleaved Reed–Solomon codes over noisy channels. Theoretical Computer Science, 379(3):348–360, June 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khachiyan:2007:DBG

[1171] Leonid Khachiyan, Endre Boros, Khaled Elbassioni, Vladimir Gurvich, and Kazuhisa Makino. Dual-bounded generating problems: Efficient and inefficient points for discrete probability distributions and sparse boxes for multidimensional data. Theoretical Computer Science, 379(3):361–376, June 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bodirsky:2007:GLP

[1172] Manuel Bodirsky, Clemens Gröpl, and Mihyun Kang. Generating labeled planar graphs uniformly at random. Theoretical Computer Science, 379(3):377– 386, June 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hall:2007:MFT

[1173] Alex Hall, Steffen Hippler, and Martin Skutella. Multicommodity flows over time: Efficient algorithms and complexity. Theoretical Computer Science, 379(3):387–404, June 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gal:2007:CPC

[1174] Anna Gál and Peter Bro Miltersen. The cell probe complexity of succinct data structures. Theoretical Computer Science, 379(3):405–417, June 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Matias:2007:EPL

[1175] Yossi Matias and Ely Porat. Efficient pebbling for list traversal synopses with application to program rollback. Theoretical Computer Science, 379(3):418– 436, June 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Italiano:2007:P

[1176] Giuseppe F. Italiano and Catuscia Palamidessi. Preface. Theoretical Computer Science, 380(1–2):1, June 21, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Avin:2007:CTM

[1177] Chen Avin and Gunes Ercal. On the cover time and mixing time of random geometric graphs. Theoretical Computer Science, 380(1–2):2–22, June 21, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czeizler:2007:TLB

[1178] Eugen Czeizler and Jarkko Kari. A tight linear bound on the synchronization delay of bijective automata. Theoretical Computer Science, 380(1–2):

23–36, June 21, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czumaj:2007:FAF

[1179] Artur Czumaj, Mirosław Kowaluk, and Andrzej Lingas. Faster algorithms for finding lowest common ancestors in directed acyclic graphs. Theoretical Computer Science, 380(1–2):37–46, June 21, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dietzfelbinger:2007:BAD

[1180] Martin Dietzfelbinger and Christoph Weidling. Balanced allocation and dictionaries with tightly packed constant size bins. Theoretical Computer Science, 380(1–2):47–68, June 21, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Droste:2007:WAW

[1181] Manfred Droste and Paul Gastin.

Weighted automata and weighted logics. Theoretical Computer Science, 380 (1–2):69–86, June 21, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gairing:2007:FCA

[1182] Martin Gairing, Burkhard Monien, and Andreas Woclaw. A faster combinatorial approximation algorithm for scheduling unrelated parallel machines. Theoretical Computer Science, 380(1– 2):87–99, June 21, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hromkovic:2007:CSN

[1183] Juraj Hromkovič and Georg Schnitger. Comparing the size of NFAs with and without ϵ-transitions. Theoretical Computer Science, 380(1–2):100– 114, June 21, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koiran:2007:QQC

[1184] Pascal Koiran, Vincent Nesme, and Natacha Portier. The quantum query complexity of the abelian hidden subgroup problem. Theoretical Computer Science, 380(1–2):115–126, June 21, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Patrascu:2007:DBP

[1185] Mihai Pat̆raşcu and Corina E. Tarniţă. On dynamic bit-probe complexity. Theoretical Computer Science, 380(1– 2):127–142, June 21, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanBreugel:2007:RDM

[1186] Franck van Breugel, Claudio Hermida, Michael Makkai, and James Worrell. Recursively defined metric spaces without contraction. Theoretical Computer Science, 380(1–2):143–163, June 21, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See addendum [2399].

Pous:2007:NTW

[1187] Damien Pous. New up-to techniques for weak bisimulation. Theoretical Computer Science, 380(1–2):164–180, June 21, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mislove:2007:DRV

[1188] Michael Mislove. Discrete random variables over domains. Theoretical Computer Science, 380(1–2):181– 198, June 21, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grohe:2007:TLB

[1189] Martin Grohe, Christoph Koch, and Nicole Schweikardt. Tight lower bounds for query processing on streaming and external memory data. Theoretical Computer Science, 380(1–2): 199–217, June 21, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBl

[1190] Anonymous. Editorial board. Theoretical Computer Science, 380(1–2): v–ix, June 21, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brlek:2007:P

[1191] Srecko Brlek and Christophe Reutenauer.

Preface. Theoretical Computer Science, 380(3):219, June 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adamczewski:2007:RPC

[1192] Boris Adamczewski and Jean-Paul Allouche. Reversals and palindromes in continued fractions. Theoretical Computer Science, 380(3):220–237, June 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ambroz:2007:AAE

[1193] Petr Ambroz and Christiane Frougny. On alpha-adic expansions in Pisot bases. Theoretical Computer Science, 380(3):238–250, June 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arnoux:2007:FSS

[1194] Pierre Arnoux, Valérie Berthé, Thomas Fernique, and Damien Jamet. Functional stepped surfaces, flips, and generalized substitutions. Theoretical Computer Science, 380(3):251– 265, June 28, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balazi:2007:FVP

[1195] Peter Balázi, Zuzana Masáková, and Edita Pelantová. Factor versus palindromic complexity of uniformly recurrent infinite words. Theoretical Computer Science, 380(3):266–275, June 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berthe:2007:DRS

[1196] Valérie Berthé and Bertrand Nouvel. Discrete rotations and symbolic dynamics. Theoretical Computer Science, 380(3):276–285, June 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Borel:2007:GCF

[1197] Jean-Pierre Borel. A geometrical characterization of factors of multidimensional Billiard words and some applications. Theoretical Computer Science, 380(3):286–303, June 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cassaigne:2007:ACS

[1198] J. Cassaigne and A. E. Frid. On the arithmetical complexity of Sturmian words. Theoretical Computer Science, 380(3):304–316, June 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fernique:2007:LRS

[1199] Thomas Fernique. Local rule substitutions and stepped surfaces. Theoretical Computer Science, 380(3):317– 329, June 28, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Glen:2007:PCS

[1200] Amy Glen. Powers in a class of A-strict standard episturmian words. Theoretical Computer Science, 380(3):330– 354, June 28, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Halava:2007:EDM

[1201] Vesa Halava, Tero Harju, Juhani Karhumäki, and Michel Latteux. Extension of the decidability of the marked PCP to instances with unique blocks. Theoretical Computer Science, 380(3):355–362, June 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Holub:2007:SWE

[1202] Stepán Holub and Juha Kortelainen.

On systems of word equations with simple loop sets. Theoretical Computer Science, 380(3):363–372, June 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ilie:2007:NNS

[1203] Lucian Ilie. A note on the number of squares in a word. Theoretical Computer Science, 380(3):373–376, June 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lepisto:2007:TIW

[1204] Arto Lepistö, Francesco Pappalardi, and Kalle Saari. Transposition invariant words. Theoretical Computer Science, 380(3):377–387, June 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ochem:2007:LFI

[1205] Pascal Ochem. Letter frequency in infinite repetition-free words. Theoretical Computer Science, 380(3):388– 392, June 28, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Richomme:2007:CML

[1206] G. Richomme. Conjugacy of morphisms and Lyndon decomposition of standard Sturmian words. Theoretical Computer Science, 380(3):393– 400, June 28, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

terBeek:2007:IUS

[1207] Maurice H. ter Beek and Jetty Kleijn. Infinite unfair shuffles and associativity. Theoretical Computer Science, 380 (3):401–410, June 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cervelle:2007:SSA

[1208] Julien Cervelle, Enrico Formenti, and Benoît Masson. From sandpiles to sand automata. Theoretical Computer Science, 381(1–3):1–28, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2007:NUP

[1209] Feng Wang, Hongwei David Du, Xiaohua Jia, Ping Deng, Weili Wu, and David MacCallum. Non-unique probe selection and group testing. Theoretical Computer Science, 381(1–3):29– 32, August 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jurgensen:2007:FAE

[1210] Helmut Jürgensen, Ludwig Staiger, and Hideki Yamasaki. Finite automata encoding geometric figures. Theoretical Computer Science, 381(1–3):33– 43, August 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koukopoulos:2007:IIN

[1211] Dimitrios Koukopoulos, Marios Mavronicolas, and Paul Spirakis. The increase of the instability of networks due to Quasi-Static link capacities. Theoretical Computer Science, 381(1–3):44– 56, August 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deFigueiredo:2007:CSP

[1212] C. M. H. de Figueiredo, L. Faria, S. Klein, and R. Sritharan. On the complexity of the sandwich problems for strongly chordal graphs and chordal bipartite graphs. Theoretical Computer Science, 381(1–3):57–67, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gomez:2007:ORD

[1213] Domingo Gómez, Jaime Gutierrez, and Álvar Ibeas. Optimal routing in double loop networks. Theoretical Computer Science, 381(1–3):68–85, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bassino:2007:ERG

[1214] Frédérique Bassino and Cyril Nicaud. Enumeration and random generation of accessible automata. Theoretical Computer Science, 381(1–3):86– 104, August 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Doty:2007:PD

[1215] David Doty and Jared Nichols. Pushdown dimension. Theoretical Computer Science, 381(1–3):105–123, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2007:SAA

[1216] Mingxia Chen, Jianbo Li, Jianping Li, Weidong Li, and Lusheng Wang. Some approximation algorithms for the clique partition problem in weighted interval graphs. Theoretical Computer Science, 381(1–3):124–133, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amiraslani:2007:BFG

[1217] A. Amiraslani, D. A. Aruliah, and R. M. Corless. Block LU factors of generalized companion matrix pencils.

Theoretical Computer Science, 381(1– 3):134–147, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ko:2007:JCP

[1218] Ker-I Ko and Fuxiang Yu. Jordan curves with polynomial inverse moduli of continuity. Theoretical Computer Science, 381(1–3):148–161, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fleiner:2007:EAG

[1219] Tamás Fleiner, Robert W. Irving, and David F. Manlove. Efficient algorithms for generalized Stable Marriage and Roommates problems. Theoretical Computer Science, 381(1–3):162– 176, August 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krieger:2007:ERN

[1220] Dalia Krieger and Jeffrey Shallit. Every real number greater than 1 is a critical exponent. Theoretical Computer Science, 381(1–3):177–182, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Parnas:2007:AMV

[1221] Michal Parnas and Dana Ron. Approximating the minimum vertex cover in sublinear time and a connection to distributed algorithms. Theoretical Computer Science, 381(1–3):183– 196, August 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alvarez:2007:CTP

[1222] Carme Àlvarez, Rafel Cases, Josep Díaz, Jordi Petit, and Maria Serna. Communication tree problems. Theoretical Computer Science, 381(1–3): 197–217, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lin:2007:SCS

[1223] Cheng-Kuan Lin, Jimmy J. M. Tan, D. Frank Hsu, and Lih-Hsing Hsu. On the spanning connectivity and spanning laceability of hypercube-like networks. Theoretical Computer Science, 381(1–3):218–229, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wu:2007:LAP

[1224] Jian-Liang Wu, Jian-Feng Hou, and Gui-Zhen Liu. The linear arboricity of planar graphs with no short cycles. Theoretical Computer Science, 381(1– 3):230–233, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

He:2007:BSB

[1225] Cheng He, Yixun Lin, and Jinjiang Yuan. Bicriteria scheduling on a batching machine to minimize maximum lateness and makespan. Theoretical Computer Science, 381(1–3):234– 240, August 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shang:2007:AMC

[1226] Weiping Shang, Pengjun Wan, Frances Yao, and Xiaodong Hu. Algorithms for minimum m-connected k-tuple dominating set problem. Theoretical Computer Science, 381(1–3):241–247, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asdre:2007:NCR

[1227] Katerina Asdre and Stavros D.

Nikolopoulos. NP-completeness results for some problems on subclasses of bipartite and chordal graphs. Theoretical Computer Science, 381(1–3):248– 259, August 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brandstadt:2007:IMC

[1228] Andreas Brandstädt, Elaine M. Eschen, and R. Sritharan. The induced matching and chain subgraph cover problems for convex bipartite graphs. Theoretical Computer Science, 381(1– 3):260–265, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lavallee:2007:ZFA

[1229] Sylvain Lavallée and Christophe Reutenauer. On a zeta function associated with automata and codes. Theoretical Computer Science, 381(1–3): 266–273, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choffrut:2007:SSC

[1230] Christian Choffrut, Flavio D’Alessandro, and Stefano Varricchio. On the separability of sparse context-free languages and of bounded rational relations. Theoretical Computer Science, 381(1–3): 274–279, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonifaci:2007:AQM

[1231] Vincenzo Bonifaci. An adversarial queueing model for online server routing. Theoretical Computer Science, 381(1–3):280–287, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hsieh:2007:ASI

[1232] Sun-Yuan Hsieh and Shih-Cheng Yang. Approximating the selected-internal Steiner tree. Theoretical Computer Science, 381(1–3):288–291, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2007:MLN

[1233] Xinmao Wang and Yaokun Wu. Minimum light number of lit-only σ-game on a tree. Theoretical Computer Science, 381(1–3):292–300, August 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBm

[1234] Anonymous. Editorial board. Theoretical Computer Science, 381(1–3):v– ix, August 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DiPierro:2007:PQA

[1235] Alessandra Di Pierro and Herbert Wiklicky. Preface: Quantitative aspects of programming languages. Theoretical Computer Science, 382(1):1– 2, August 28, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aldini:2007:MLR

[1236] Alessandro Aldini and Marco Bernardo.

Mixing logics and rewards for the component-oriented specification of performance measures. Theoretical Computer Science, 382(1):3–23, August 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Remke:2007:CMC

[1237] Anne Remke, Boudewijn R. Haverkort, and Lucia Cloth. CSL model checking algorithms for QBDs. Theoretical Computer Science, 382(1):24– 41, August 28, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeNicola:2007:MCM

[1238] Rocco De Nicola, Joost-Pieter Katoen, Diego Latella, Michele Loreti, and Mieke Massink. Model checking mobile stochastic logic. Theoretical Computer Science, 382(1):42– 70, August 28, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hasan:2007:FSU

[1239] Osman Hasan and Sofiène Tahar. Formalization of the Standard Uniform random variable. Theoretical Computer Science, 382(1):71–83, August 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:PQA

[1240] Anonymous. Preface: Quantitative aspects of programming languages. Theoretical Computer Science, 382(1):v– ix, August 28, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cryan:2007:P

[1241] Mary Cryan and Martin Farach-Colton. Preface. Theoretical Computer Science, 382(2):85, August 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Badoiu:2007:UAB

[1242] Mihai Bădoiu, Richard Cole, Erik D. Demaine, and John Iacono. A unified access bound on comparison-based dynamic dictionaries. Theoretical Computer Science, 382(2):86–96, August 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caminiti:2007:CLT

[1243] Saverio Caminiti, Irene Finocchi, and Rossella Petreschi. On coding labeled trees. Theoretical Computer Science, 382(2):97–108, August 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carton:2007:CRS

[1244] Olivier Carton and Chloé Rispal. Complementation of rational sets on scattered linear orderings of finite rank. Theoretical Computer Science, 382(2): 109–119, August 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaz:2007:BBW

[1245] J. Díaz, M. J. Serna, and N. C. Wormald. Bounds on the bisection width for random d-regular graphs. Theoretical Computer Science, 382(2):

120–130, August 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gutierrez:2007:CBM

[1246] Claudio Gutierrez, Flavio Gutierrez, and Maria-Cecilia Rivara. Complexity of the bisection method. Theoretical Computer Science, 382(2):131– 138, August 31, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khachiyan:2007:DHB

[1247] Leonid Khachiyan, Endre Boros, Khaled Elbassioni, and Vladimir Gurvich. On the dualization of hypergraphs with bounded edge-intersections and other related classes of hypergraphs. Theoretical Computer Science, 382(2): 139–150, August 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pruhs:2007:ASC

[1248] Kirk Pruhs and Gerhard J. Woeginger. Approximation schemes for a class of subset selection problems. Theoretical Computer Science, 382(2):151– 156, August 31, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2007:IPN

[1249] Ke Yang. On the (im)possibility of non-interactive correlation distillation. Theoretical Computer Science, 382(2): 157–166, August 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ben-David:2007:F

[1250] Shai Ben-David, JohnCase, and Thomas Zeugmann. Foreword. Theoretical Computer Science, 382(3):167– 169, September 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martin:2007:DCB

[1251] Eric Martin, Arun Sharma, and Frank Stephan. On the data consumption benefits of accepting increased uncertainty. Theoretical Computer Science, 382(3):170–182, September 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Besombes:2007:LTL

[1252] Jérôme Besombes and Jean-Yves Marion. Learning tree languages from positive examples and membership queries. Theoretical Computer Science, 382(3): 183–197, September 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sloan:2007:RTF

[1253] Robert H. Sloan, Balázs Szörényi, and György Turán. Revising threshold functions. Theoretical Computer Science, 382(3):198–208, September 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bulatov:2007:LIC

[1254] Andrei Bulatov, Hubie Chen, and Víctor Dalmau. Learning intersectionclosed classes with signatures. Theoretical Computer Science, 382(3):209– 220, September 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cesa-Bianchi:2007:ARL

[1255] Nicolò Cesa-Bianchi. Applications of regularized least squares to pattern classification. Theoretical Computer Science, 382(3):221–231, September 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ambroladze:2007:CPC

[1256] Amiran Ambroladze, Emilio Parrado-Hernández, and John Shawe-Taylor. Complexity of pattern classes and the Lipschitz property. Theoretical Computer Science, 382(3):232–246, September 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hutter:2007:SPM

[1257] Marcus Hutter and Andrej Muchnik.

On semimeasures predicting Martin-Löf random sequences. Theoretical Computer Science, 382(3):247–261, September 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Simon:2007:CWS

[1258] Hans Ulrich Simon. On the complexity of working set selection. Theoretical Computer Science, 382(3):262– 279, September 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kralovic:2007:P

[1259] Rastislav Královič. Preface. Theoretical Computer Science, 383(1):1, September 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:D

[1260] Anonymous. Dedication. Theoretical Computer Science, 383(1):3, September 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chlebus:2007:CAB

[1261] Bogdan S. Chlebus and Mariusz A. Rokicki. Centralized asynchronous broadcast in radio networks. Theoretical Computer Science, 383(1):5–22, September 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DiSalvo:2007:SFE

[1262] Aleksej Di Salvo and Guido Proietti. Swapping a failing edge of a shortest paths tree by minimizing the average stretch factor. Theoretical Computer Science, 383(1):23–33, September 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dourisboure:2007:SBT

[1263] Yon Dourisboure, Feodor F. Dragan, Cyril Gavoille, and Chenyu Yan. Spanners for bounded tree-length graphs. Theoretical Computer Science, 383(1): 34–44, September 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gasieniec:2007:TEC

[1264] Leszek Ga̧sieniec, Igor Potapov, and Qin Xin. Time efficient centralized gossiping in radio networks. Theoretical Computer Science, 383(1):45–58, September 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Georgiou:2007:LLR

[1265] Chryssis Georgiou, Peter M. Musial, and Alexander A. Shvartsman. Longlived Rambo: Trading knowledge for communication. Theoretical Computer Science, 383(1):59–85, September 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Luccio:2007:FVS

[1266] Flaminia L. Luccio and Jop F. Sibeyn.

Feedback vertex sets in mesh-based networks. Theoretical Computer Science, 383(1):86–101, September 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Matichin:2007:AAH

[1267] Rachel Matichin and David Peleg. Approximation algorithm for hotlink assignment in the greedy model. Theoretical Computer Science, 383(1):102– 110, September 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:Pa

[1268] Anonymous. Preface. Theoretical Computer Science, 383(1):v–ix, September 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Burgin:2007:P

[1269] Mark Burgin and Cristian S. Calude.

Preface. Theoretical Computer Science, 383(2–3):111–114, September 18, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pollett:2007:CPW

[1270] Chris Pollett and Norman Danner. Circuit principles and weak pigeonhole variants. Theoretical Computer Science, 383(2–3):115–131, September 18, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calvin:2007:LBC

[1271] James M. Calvin. A lower bound on complexity of optimization on the Wiener space. Theoretical Computer Science, 383(2–3):132–139, September 18, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salomaa:2007:SCC

[1272] Arto Salomaa, Kai Salomaa, and Sheng Yu. State complexity of combined operations. Theoretical Computer Science, 383(2–3):140–152, September 18, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hemaspaandra:2007:QMT

[1273] Lane A. Hemaspaandra and Mayur Thakur. Query-monotonic Turing reductions. Theoretical Computer Science, 383(2–3):153–186, September 18, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Staiger:2007:KCI

[1274] Ludwig Staiger. The Kolmogorov complexity of infinite words. Theoretical Computer Science, 383(2–3):187–199, September 18, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Eberbach:2007:CPA

[1275] Eugene Eberbach. The $-calculus process algebra for problem solving: a paradigmatic shift in handling hard computational problems. Theoretical Computer Science, 383(2–3):200–243, September 18, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Burgin:2007:ACC

[1276] Mark Burgin. Algorithmic complexity as a criterion of unsolvability. Theoretical Computer Science, 383(2–3):244– 259, September 18, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wiedermann:2007:AAC

[1277] Jiří Wiedermann. Autopoietic automata: Complexity issues in offspringproducing evolving processes. Theoretical Computer Science, 383(2–3):260– 269, September 18, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kelly:2007:ORE

[1278] Kevin T. Kelly. Ockham’s razor, empirical complexity, and truth-finding efficiency. Theoretical Computer Science, 383(2–3):270–289, September 18, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cooper:2007:PTA

[1279] Barry Cooper and Angsheng Li. Preface: Theory and applications of models of computation. Theoretical Computer Science, 384(1):1, September 24, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arpe:2007:LJP

[1280] Jan Arpe and Rüdiger Reischuk. Learning juntas in the presence of noise. Theoretical Computer Science, 384(1):2–21, September 24, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cai:2007:VHT

[1281] Jin-Yi Cai and Vinay Choudhary. Valiant’s Holant Theorem and matchgate tensors. Theoretical Computer Science, 384(1):22–32, September 24, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hutter:2007:UPB

[1282] Marcus Hutter. On universal prediction and Bayesian confirmation. Theoretical Computer Science, 384(1):33– 48, September 24, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jain:2007:IC

[1283] Sanjay Jain, Jochen Nessel, and Frank Stephan. Invertible classes. Theoretical Computer Science, 384(1):49–65, September 24, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hellerstein:2007:PLA

[1284] Lisa Hellerstein and Rocco A. Servedio. On PAC learning algorithms for rich Boolean function classes. Theoretical Computer Science, 384(1):66– 76, September 24, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muchnik:2007:NRD

[1285] Andrej Muchnik, Alexander Shen, Mikhail Ustinov, Nikolai Vereshchagin, and Michael Vyugin. Non-reducible descriptions for conditional Kolmogorov complexity. Theoretical Computer Science, 384(1):77–86, September 24, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sun:2007:BSW

[1286] Xiaoming Sun. Block sensitivity of weakly symmetric functions. Theoretical Computer Science, 384(1):87–91, September 24, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tan:2007:LTA

[1287] Xuehou Tan. A linear-time 2approximation algorithm for the watchman route problem for simple polygons. Theoretical Computer Science, 384(1): 92–103, September 24, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Urquhart:2007:WVS

[1288] Alasdair Urquhart. Width versus size in resolution proofs. Theoretical Computer Science, 384(1):104–110, September 24, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Xia:2007:CCC

[1289] Mingji Xia, Peng Zhang, and Wenbo Zhao. Computational complexity of counting problems on 3-regular planar graphs. Theoretical Computer Science, 384(1):111–125, September 24, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2007:NAA

[1290] Peng Zhang. A new approximation algorithm for the k-facility location problem. Theoretical Computer Science, 384(1):126–135, September 24, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:PTA

[1291] Anonymous. Preface: Theory and applications of models of computation. Theoretical Computer Science, 384(1): v–ix, September 24, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pelc:2007:P

[1292] Andrzej Pelc, David Peleg, and Michel Raynal. Preface. Theoretical Computer Science, 384(2–3):137–138, October 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bermond:2007:TGP

[1293] Jean-Claude Bermond, Laurent Braud, and David Coudert. Traffic grooming on the path. Theoretical Computer Science, 384(2–3):139–151, October 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caragiannis:2007:TBO

[1294] Ioannis Caragiannis, Aleksei V. Fishkin, Christos Kaklamanis, and Evi Papaioannou. A tight bound for online colouring of disk graphs. Theoretical Computer Science, 384(2–3):152– 160, October 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Clementi:2007:BHM

[1295] Andrea E. F. Clementi, Miriam Di Ianni, Massimo Lauria, Angelo Monti, Gianluca Rossi, and Riccardo Silvestri. On the bounded-hop MST problem on random Euclidean instances. Theoretical Computer Science, 384(2–3):161– 167, October 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dinitz:2007:TAB

[1296] Yefim Dinitz and Noam Solomon. Two absolute bounds for distributed bit complexity. Theoretical Computer Science, 384(2–3):168–183, October 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hinkelmann:2007:CUN

[1297] Markus Hinkelmann and Andreas Jakoby. Communications in unknown networks: Preserving the secret of topology. Theoretical Computer Science, 384(2–3):184–200, October 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klasing:2007:HAR

[1298] Ralf Klasing, Euripides Markou, Tomasz Radzik, and Fabiano Sarracco. Hardness and approximation results for Black Hole Search in arbitrary networks. Theoretical Computer Science, 384(2–3):201–221, October 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Prencipe:2007:IGS

[1299] Giuseppe Prencipe. Impossibility of gathering by a set of autonomous mobile robots. Theoretical Computer Science, 384(2–3):222–231, October 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Santoro:2007:ASN

[1300] Nicola Santoro and Peter Widmayer. Agreement in synchronous networks with ubiquitous faults. Theoretical Computer Science, 384(2–3):232– 249, October 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shalom:2007:MNA

[1301] Mordechai Shalom and Shmuel Zaks. Minimization of the number of ADMs in SONET rings with maximum throughput with implications to the traffic grooming problem. Theoretical Computer Science, 384(2–3):250– 262, October 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2007:OGS

[1302] Rui Wang and Francis C. M. Lau. Optimal gossiping in square 2D meshes. Theoretical Computer Science, 384(2– 3):263–286, October 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2007:MMM

[1303] Xueliang Li and Xiaoyan Zhang. On the minimum monochromatic or multicolored subgraph partition problems. Theoretical Computer Science, 385(1– 3):1–10, October 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Andre:2007:PCS

[1304] Y. André, A. C. Caron, D. Debarbieux, Y. Roos, and S. Tison. Path constraints in semistructured data. Theoretical Computer Science, 385(1–3):11– 33, October 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Das:2007:MCU

[1305] Shantanu Das, Paola Flocchini, Shay Kutten, Amiya Nayak, and Nicola Santoro. Map construction of unknown graphs by multiple agents. Theoretical Computer Science, 385(1–3):34– 48, October 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Thai:2007:AAC

[1306] My T. Thai, Ning Zhang, Ravi Tiwari, and Xiaochun Xu. On approximation algorithms of k-connected mdominating sets in disk graphs. Theoretical Computer Science, 385(1–3):49– 59, October 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Broersma:2007:CDS

[1307] Hajo Broersma and Xueliang Li. On the complexity of dominating set problems related to the minimum allones problem. Theoretical Computer Science, 385(1–3):60–70, October 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2007:SCP

[1308] Weifan Wang and Yongzhu Chen. A sufficient condition for a planar graph to be class 1. Theoretical Computer Science, 385(1–3):71–77, October 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galbiati:2007:AAM

[1309] Giulia Galbiati and Francesco Maffioli. Approximation algorithms for maximum cut with limited unbalance. Theoretical Computer Science, 385(1–3): 78–87, October 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hsieh:2007:FCG

[1310] Ming Yu Hsieh and Shi-Chun Tsai. On the fairness and complexity of generalized k-in-a-row games. Theoretical Computer Science, 385(1–3):88– 100, October 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khan:2007:SRS

[1311] Maleq Khan, Gopal Pandurangan, and V. S. Anil Kumar. A simple randomized scheme for constructing low-weight k-connected spanning subgraphs with applications to distributed algorithms. Theoretical Computer Science, 385(1– 3):101–114, October 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Popov:2007:MGR

[1312] V. Yu. Popov. Multiple genome rearrangement by swaps and by element duplications. Theoretical Computer Science, 385(1–3):115–126, October 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Na:2007:AIL

[1313] Joong Chae Na and Kunsoo Park.

Alphabet-independent linear-time construction of compressed suffix arrays using o(n log n)-bit working space. Theoretical Computer Science, 385(1– 3):127–136, October 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carpi:2007:DCL

[1314] Arturo Carpi. On Dejean’s conjecture over large alphabets. Theoretical Computer Science, 385(1–3):137– 151, October 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epifanio:2007:LM

[1315] C. Epifanio, A. Gabriele, F. Mignosi, A. Restivo, and M. Sciortino. Languages with mismatches. Theoretical Computer Science, 385(1–3):152– 166, October 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pavan:2007:PRI

[1316] A. Pavan, Alan L. Selman, Samik Sengupta, and N. V. Vinodchandran. Polylogarithmic-round interactive proofs for coNP collapse the exponential hierarchy. Theoretical Computer Science, 385(1–3):167–178, October 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanchet-Sadri:2007:PWC

[1317] F. Blanchet-Sadri and Nathan D. Wetzler. Partial words and the critical factorization theorem revisited. Theoretical Computer Science, 385(1–3):179– 192, October 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Becher:2007:RRC

[1318] Verónica Becher and Serge Grigorieff. Random reals à la Chaitin with or without prefix-freeness. Theoretical Computer Science, 385(1–3):193– 201, October 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berger:2007:ACE

[1319] André Berger, Takuro Fukunaga, Hiroshi Nagamochi, and Ojas Parekh. Approximability of the capacitated bedge dominating set problem. Theoretical Computer Science, 385(1–3):202– 213, October 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bedaride:2007:CRT

[1320] Nicolas Bedaride. Classification of rotations on the torus T^{2}. Theoretical Computer Science, 385(1–3):214– 225, October 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kratsch:2007:EAM

[1321] Dieter Kratsch and Mathieu Liedloff. An exact algorithm for the minimum dominating clique problem. Theoretical Computer Science, 385(1–3):226– 240, October 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blin:2007:ECI

[1322] Guillaume Blin, Guillaume Fertin, and Stéphane Vialette. Extracting constrained 2-interval subsets in 2-interval sets. Theoretical Computer Science, 385(1–3):241–263, October 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Malvestuto:2007:AAI

[1323] Francesco M. Malvestuto, Mauro Mezzini, and Marina Moscarini. An analytical approach to the inference of summary data of additive type. Theoretical Computer Science, 385(1–3): 264–285, October 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fang:2007:BPF

[1324] Jywe-Fei Fang. The bipanconnectivity and m-panconnectivity of the folded hypercube. Theoretical Computer Science, 385(1–3):286–300, October 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBn

[1325] Anonymous. Editorial board. Theoretical Computer Science, 385(1–3):v– ix, October 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2007:Pb

[1326] G. Rozenberg. Preface. Theoretical Computer Science, 386(1–2):1–2, October 28, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barrett:2007:PEP

[1327] Chris Barrett, Harry B. Hunt III, Madhav V. Marathe, S. S. Ravi, Daniel J. Rosenkrantz, Richard E. Stearns, and Mayur Thakur. Predecessor existence problems for finite discrete dynamical systems. Theoretical Computer Science, 386(1–2):3–37, October 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [1529].

Qiu:2007:ATB

[1328] Daowen Qiu. Automata theory based on quantum logic: Reversibilities and pushdown automata. Theoretical Computer Science, 386(1–2):38–56, October 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carnell:2007:PCS

[1329] Andrew Carnell and Daniel Richardson. Parallel computation in spiking neural nets. Theoretical Computer Science, 386(1–2):57–72, October 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jansen:2007:CSA

[1330] Thomas Jansen and Ingo Wegener. A comparison of simulated annealing with a simple evolutionary algorithm on pseudo-boolean functions of unitation. Theoretical Computer Science, 386(1–2):73–93, October 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Benfold:2007:OPS

[1331] W. Benfold, J. Hallam, and A. Prügel-Bennett. Optimal parameters for search using a barrier tree Markov model. Theoretical Computer Science, 386(1–2):94–113, October 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Storch:2007:FLC

[1332] Tobias Storch. Finding large cliques in sparse semi-random graphs by simple randomized search heuristics. Theoretical Computer Science, 386(1–2):114– 131, October 28, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Loos:2007:CTS

[1333] Remco Loos and Mitsunori Ogihara.

Complexity theory for splicing systems. Theoretical Computer Science, 386(1– 2):132–150, October 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Feng:2007:PRC

[1334] Yuan Feng, Runyao Duan, Zhengfeng Ji, and Mingsheng Ying. Proof rules for the correctness of quantum programs. Theoretical Computer Science, 386(1–2):151–166, October 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:Pb

[1335] Anonymous. Preface. Theoretical Computer Science, 386(1–2):v–ix, October 28, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeten:2007:Pb

[1336] Jos C. M. Baeten and Iain C. C.

Phillips. Preface. Theoretical Computer Science, 386(3):167–168, November 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Froschle:2007:CVT

[1337] Sibylle Fröschle and Sławomir Lasota. Causality versus true-concurrency. Theoretical Computer Science, 386(3): 169–187, November 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deAlfaro:2007:CRG

[1338] Luca de Alfaro, Thomas A. Henzinger, and Orna Kupferman. Concurrent reachability games. Theoretical Computer Science, 386(3):188–217, November 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cacciagrano:2007:SSA

[1339] D. Cacciagrano, F. Corradini, and C. Palamidessi. Separation of synchronous and asynchronous communication via testing. Theoretical Computer Science, 386(3):218–235, November 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Briais:2007:OBR

[1340] Sébastien Briais and Uwe Nestmann. Open bisimulation, revisited. Theoretical Computer Science, 386(3):236– 271, November 1, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Simon:2007:GEF

[1341] Hans Ulrich Simon and Etsuji Tomita. Guest editors’ foreword. Theoretical Computer Science, 387(1):1–3, November 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Watanabe:2007:SCM

[1342] Kazuho Watanabe and Sumio Watanabe. Stochastic complexity for mixture of exponential families in generalized variational Bayes. Theoretical Computer Science, 387(1):4–17, November 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Palmer:2007:PLP

[1343] Nick Palmer and Paul W. Goldberg.

PAC-learnability of probabilistic deterministic finite state automata in terms of variation distance. Theoretical Computer Science, 387(1):18–31, November 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bennet:2007:LAE

[1344] Rotem Bennet and Nader H. Bshouty.

Learning attribute-efficiently with corrupt oracles. Theoretical Computer Science, 387(1):32–50, November 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jain:2007:GCL

[1345] Sanjay Jain, Steffen Lange, and Sandra Zilles. A general comparison of language learning from examples and from queries. Theoretical Computer Science, 387(1):51–66, November 6, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jain:2007:LML

[1346] Sanjay Jain and Efim Kinber. Learning multiple languages in groups. Theoretical Computer Science, 387(1):67– 76, November 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vovk:2007:NAC

[1347] Vladimir Vovk. Non-asymptotic calibration and resolution. Theoretical Computer Science, 387(1):77–89, November 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBo

[1348] Anonymous. Editorial board. Theoretical Computer Science, 387(1):v– ix, November 6, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leung:2007:P

[1349] Hing Leung and Giovanni Pighizzini. Preface. Theoretical Computer Science, 387(2):91–92, November 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Almeida:2007:EGS

[1350] Marco Almeida, Nelma Moreira, and Rogério Reis. Enumeration and generation with a string automata representation. Theoretical Computer Science, 387(2):93–102, November 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biegler:2007:RRR

[1351] Franziska Biegler, Michael J. Burrell, and Mark Daley. Regulated RNA rewriting: Modelling RNA editing with guided insertion. Theoretical Computer Science, 387(2):103–112, November 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biegler:2007:IHI

[1352] Franziska Biegler, Ian McQuillan, and Kai Salomaa. An infinite hierarchy induced by depth synchronization. Theoretical Computer Science, 387(2):113– 124, November 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cordy:2007:ERA

[1353] Brendan Cordy and Kai Salomaa. On the existence of regular approximations. Theoretical Computer Science, 387(2):125–135, November 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dassow:2007:NCS

[1354] Jürgen Dassow and Bianca Truthe. On the number of components for some parallel communicating grammar systems. Theoretical Computer Science, 387(2):136–146, November 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Domaratzki:2007:TCL

[1355] Michael Domaratzki and Kai Salomaa.

Transition complexity of language operations. Theoretical Computer Science, 387(2):147–154, November 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gruber:2007:AST

[1356] Hermann Gruber and Markus Holzer.

On the average state and transition complexity of finite languages. Theoretical Computer Science, 387(2):155– 166, November 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gruber:2007:SHH

[1357] Hermann Gruber, Markus Holzer, and Martin Kutrib. The size of Higman– Haines sets. Theoretical Computer Science, 387(2):167–176, November 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mereghetti:2007:QAS

[1358] Carlo Mereghetti and Beatrice Palano. Quantum automata for some multiperiodic languages. Theoretical Computer Science, 387(2):177–186, November 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Niessner:2007:CCD

[1359] Frank Nießner and Ulrich Ultes-Nitsche. A complete characterization of deterministic regular liveness properties. Theoretical Computer Science, 387(2):187–195, November 12, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferragina:2007:F

[1360] Paolo Ferragina, Giovanni Manzini, and S. Muthukrishnan. Foreword. Theoretical Computer Science, 387(3):197– 199, November 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fenwick:2007:BWC

[1361] Peter Fenwick. Burrows–Wheeler compression: Principles and reflections. Theoretical Computer Science, 387(3): 200–219, November 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kaplan:2007:SAB

[1362] Haim Kaplan, Shir Landau, and Elad Verbin. A simpler analysis of Burrows– Wheeler-based compression. Theoretical Computer Science, 387(3):220–235, November 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Giancarlo:2007:FPB

[1363] R. Giancarlo, A. Restivo, and M. Sciortino. From first principles to the Burrows and Wheeler transform and beyond, via combinatorial optimization. Theoretical Computer Science, 387(3):236–248, November 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karkkainen:2007:FBS

[1364] Juha Kärkkäinen. Fast BWT in small space by blockwise suffix sorting. Theoretical Computer Science, 387(3):249– 257, November 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Larsson:2007:FSS

[1365] N. Jesper Larsson and Kunihiko Sadakane. Faster suffix sorting. Theoretical Computer Science, 387(3):258– 272, November 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vo:2007:CTD

[1366] Binh Dao Vo and Kiem-Phong Vo.

Compressing table data with column dependency. Theoretical Computer Science, 387(3):273–283, November 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barbay:2007:ASS

[1367] Jérémy Barbay, Alexander Golynski, J. Ian Munro, and S. Srinivasa Rao. Adaptive searching in succinctly encoded binary relations and tree-structured documents. Theoretical Computer Science, 387(3):284–297, November 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mantaci:2007:EBW

[1368] S. Mantaci, A. Restivo, G. Rosone, and M. Sciortino. An extension of the Burrows–Wheeler Transform. Theoretical Computer Science, 387(3):298– 312, November 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gupta:2007:CDS

[1369] Ankur Gupta, Wing-Kai Hon, Rahul Shah, and Jeffrey Scott Vitter. Compressed data structures: Dictionaries and data-aware measures. Theoretical Computer Science, 387(3):313–331, November 22, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Makinen:2007:RSR

[1370] Veli Mäkinen and Gonzalo Navarro. Rank and select revisited and extended. Theoretical Computer Science, 387(3):332–347, November 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Golynski:2007:OLB

[1371] Alexander Golynski. Optimal lower bounds for rank and select indexes. Theoretical Computer Science, 387(3):

348–359, November 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lai:2007:CDC

[1372] Hongliang Lai and Dexue Zhang.

Complete and directed complete Ωcategories. Theoretical Computer Science, 388(1–3):1–25, December 5, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Geniet:2007:WFT

[1373] Dominique Geniet and Gaëlle Largeteau.

WCET free time analysis of hard realtime systems on multiprocessors: a regular language-based model. Theoretical Computer Science, 388(1–3):26– 52, December 5, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Spoto:2007:OCI

[1374] Fausto Spoto. Optimality and condensing of information flow through linear refinement. Theoretical Computer Science, 388(1–3):53–82, December 5, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cirstea:2007:MCC

[1375] Corina Cĭrstea and Dirk Pattinson.

Modular construction of complete coalgebraic logics. Theoretical Computer Science, 388(1–3):83–108, December 5, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Johnson:2007:FUV

[1376] Michael Johnson and Robert Rosebrugh. Fibrations and universal view updatability. Theoretical Computer Science, 388(1–3):109–129, December 5, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adamek:2007:API

[1377] Jiří Adámek, Stefan Milius, and Jiří Velebil. Algebras with parametrized iterativity. Theoretical Computer Science, 388(1–3):130–151, December 5, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Birkedal:2007:DTM

[1378] L. Birkedal, R. E. Møgelberg, and R. L. Petersen. Domain-theoretical models of parametric polymorphism. Theoretical Computer Science, 388(1–3):152– 172, December 5, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Plummer:2007:SEM

[1379] Andrew R. Plummer. S4 enriched multimodal categorial grammars are context-free. Theoretical Computer Science, 388(1–3):173–180, December 5, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chatterjee:2007:CGT

[1380] Krishnendu Chatterjee. Concurrent games with tail objectives. Theoretical Computer Science, 388(1–3):181– 198, December 5, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Komenda:2007:CDE

[1381] Jan Komenda and Jan H. van Schuppen. Control of discrete-event systems with modular or distributed structure. Theoretical Computer Science, 388(1–3):199–226, December 5, 2007.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cook:2007:VBP

[1382] Byron Cook, Daniel Kroening, and Natasha Sharygina. Verification of Boolean programs with unbounded thread creation. Theoretical Computer Science, 388(1–3):227–242, December 5, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schaefer:2007:CRC

[1383] Mark Schaefer and Walter Vogler.

Component refinement and CSCsolving for STG decomposition. Theoretical Computer Science, 388(1–3): 243–266, December 5, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vigliotti:2007:TSR

[1384] Maria Grazia Vigliotti, Iain Phillips, and Catuscia Palamidessi. Tutorial on separation results in process calculi via leader election problems. Theoretical Computer Science, 388(1–3):267– 289, December 5, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Voigtlander:2007:SSP

[1385] Janis Voigtländer and Patricia Johann.

Selective strictness and parametricity in structural operational semantics, inequationally. Theoretical Computer Science, 388(1–3):290–318, December 5, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ponse:2007:BLC

[1386] Alban Ponse and Mark B. van der Zwaag. Belnap’s logic and conditional composition. Theoretical Computer Science, 388(1–3):319–336, December 5, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBp

[1387] Anonymous. Editorial board. Theoretical Computer Science, 388(1–3):v– ix, December 5, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khosravi:2007:QPV

[1388] Ramtin Khosravi and Mohammad Ghodsi. Query-point visibility constrained shortest paths in simple polygons. Theoretical Computer Science, 389(1–2):1– 11, December 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bernat:2007:CIW

[1389] Julien Bernat, Zuzana Masáková, and Edita Pelantová. On a class of infinite words with affine factor complexity. Theoretical Computer Science, 389(1–2):12–25, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aspnes:2007:TTD

[1390] James Aspnes, Joan Feigenbaum, Aleksandr Yampolskiy, and Sheng Zhong. Towards a theory of data entanglement. Theoretical Computer Science, 389(1–2):26–43, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gronau:2007:HIP

[1391] Ilan Gronau and Shlomo Moran. On the hardness of inferring phylogenies from triplet-dissimilarities. Theoretical Computer Science, 389(1–2):44–55, December 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jurdzinski:2007:CGR

[1392] Tomasz Jurdziński. On complexity of grammars related to the safety problem. Theoretical Computer Science, 389(1–2):56–72, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meduna:2007:CSC

[1393] Alexander Meduna and Jiří Techet.

Canonical scattered context generators of sentences with their parses. Theoretical Computer Science, 389(1–2):73– 81, December 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Le:2007:SPD

[1394] Minh Ha Le and Thi Ha Duong Phan.

Strict partitions and discrete dynamical systems. Theoretical Computer Science, 389(1–2):82–90, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeSimone:2007:ECR

[1395] Caterina De Simone and Anna Galluccio. Edge-colouring of regular graphs of large degree. Theoretical Computer Science, 389(1–2):91–99, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Herranz:2007:IBR

[1396] Javier Herranz. Identity-based ring signatures from RSA. Theoretical Computer Science, 389(1–2):100–117, December 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tan:2007:MSP

[1397] Bo Tan. Mirror substitutions and palindromic sequences. Theoretical Computer Science, 389(1–2):118–124, December 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Michail:2007:RRM

[1398] Dimitrios Michail. Reducing rankmaximal to maximum weight matching. Theoretical Computer Science, 389(1–2):125–132, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martinez:2007:ATS

[1399] Fábio Viduani Martinez, José Coelho de Pina, and José Soares. Algorithms for terminal Steiner trees. Theoretical Computer Science, 389(1–2):133– 142, December 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Xu:2007:TCP

[1400] Yinfeng Xu, Wenqiang Dai, Naoki Katoh, and Makoto Ohsaki. Triangulating a convex polygon with fewer number of non-standard bars. Theoretical Computer Science, 389(1–2):143– 151, December 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kloster:2007:SAP

[1401] Oddvar Kloster. A solution to the Angel Problem. Theoretical Computer Science, 389(1–2):152–161, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harkins:2007:USW

[1402] Ryan C. Harkins and John M. Hitchcock. Upward separations and weaker hypotheses in resource-bounded measure. Theoretical Computer Science, 389(1–2):162–171, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koiran:2007:CTP

[1403] Pascal Koiran and Sylvain Perifel. The complexity of two problems on arithmetic circuits. Theoretical Computer Science, 389(1–2):172–181, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2007:TAK

[1404] Hsiao-Fei Liu and Kun-Mao Chao.

A tight analysis of the Katriel– Bodlaender algorithm for online topological ordering. Theoretical Computer Science, 389(1–2):182–189, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jain:2007:LLP

[1405] Sanjay Jain and Efim Kinber. Learning languages from positive data and a limited number of short counterexamples. Theoretical Computer Science, 389(1–2):190–218, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alekseev:2007:NHG

[1406] V. E. Alekseev, R. Boliac, D. V. Korobitsyn, and V. V. Lozin. NP-hard graph problems and boundary classes of graphs. Theoretical Computer Science, 389(1–2):219–236, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Halava:2007:RCW

[1407] Vesa Halava, Tero Harju, and Tomi Kärki. Relational codes of words. Theoretical Computer Science, 389(1–2): 237–249, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2007:HMS

[1408] Rui Wang, Francis C. M. Lau, and Yan Yan Liu. On the hardness of minimizing space for all-shortest-path interval routing schemes. Theoretical Computer Science, 389(1–2):250– 264, December 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manea:2007:FPW

[1409] Florin Manea and Robert Mercaş. Freeness of partial words. Theoretical Computer Science, 389(1–2):265–277, December 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Strothmann:2007:AAD

[1410] Dirk Strothmann. The affix array data structure and its applications to RNA secondary structure analysis. Theoretical Computer Science, 389(1–2):278– 294, December 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brandstadt:2007:CSN

[1411] Andreas Brandstädt and Chính T.

Hoàng. On clique separators, nearly chordal graphs, and the Maximum Weight Stable Set Problem. Theoretical Computer Science, 389(1–2):295– 306, December 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Han:2007:PFR

[1412] Yo-Sub Han, Yajun Wang, and Derick Wood. Prefix-free regular languages and pattern matching. Theoretical Computer Science, 389(1–2):307– 317, December 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moshe:2007:SCT

[1413] Yossi Moshe. On the subword complexity of Thue–Morse polynomial extractions. Theoretical Computer Science, 389(1–2):318–329, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Le:2007:CCG

[1414] Van Bang Le, Bert Randerath, and Ingo Schiermeyer. On the complexity of 4-coloring graphs without long induced paths. Theoretical Computer Science, 389(1–2):330–335, December 10, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2007:EBq

[1415] Anonymous. Editorial board. Theoretical Computer Science, 389(1–2):v– ix, December 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sannella:2007:SLF

[1416] Donald Sannella and Vladimiro Sassone. Semantic and logical foundations of global computing: Papers from the EU-FET global computing initiative (2001–2005). Theoretical Computer Science, 389(3):337–340, December 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ahern:2007:FJR

[1417] Alexander Ahern and Nobuko Yoshida. Formalising Java RMI with explicit code mobility. Theoretical Computer Science, 389(3):341–410, December 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aspinall:2007:PLR

[1418] David Aspinall, Lennart Beringer, Martin Hofmann, Hans-Wolfgang Loidl, and Alberto Momigliano. A program logic for resources. Theoretical Computer Science, 389(3):411– 445, December 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baldan:2007:SFO

[1419] P. Baldan, A. Bracciali, and R. Bruni. A semantic framework for open processes. Theoretical Computer Science, 389(3):446–483, December 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Briais:2007:FSP

[1420] Sébastien Briais and Uwe Nestmann. A formal semantics for protocol narrations. Theoretical Computer Science, 389(3):484–511, December 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chatzikokolakis:2007:FAP

[1421] Konstantinos Chatzikokolakis and Catuscia Palamidessi. A framework for analyzing probabilistic protocols and its application to the Partial Secrets Exchange. Theoretical Computer Science, 389(3):512–527, December 15, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krukow:2007:CFP

[1422] Karl Krukow and Andrew Twigg. The complexity of fixed point models of trust in distributed networks. Theoretical Computer Science, 389(3):528– 549, December 15, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asarinp:2008:AAP

[1423] Eugene Asarinp, Gordon Pace, Gerardo Schneider, and Sergio Yovine. Algorithmic analysis of polygonal hybrid systems, Part II: Phase portrait and tools. Theoretical Computer Science, 390(1):1–26, January 22, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klasing:2008:GAO

[1424] Ralf Klasing, Euripides Markou, and Andrzej Pelc. Gathering asynchronous oblivious mobile robots in a ring. Theoretical Computer Science, 390(1):27– 39, January 22, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Apostolico:2008:MP

[1425] Alberto Apostolico, Laxmi Parida, and Simona E. Rombo. Motif patterns in 2D. Theoretical Computer Science, 390 (1):40–55, January 22, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chepoi:2008:RAA

[1426] Victor Chepoi, Karim Nouioua, and Yann Vaxès. A rounding algorithm for approximating minimum Manhattan networks. Theoretical Computer Science, 390(1):56–69, January 22, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Park:2008:PEP

[1427] Jung-Heum Park. Panconnectivity and edge-pancyclicity of faulty recursive circulant G. Theoretical Computer Science, 390(1):70–80, January 22, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Morsy:2008:IAA

[1428] Ehab Morsy and Hiroshi Nagamochi. An improved approximation algorithm for capacitated multicast routings in networks. Theoretical Computer Science, 390(1):81–91, January 22, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeMol:2008:TSC

[1429] Liesbeth De Mol. Tag systems and Collatz-like functions. Theoretical Computer Science, 390(1):92–101, January 22, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Atanasiu:2008:PMA

[1430] Adrian Atanasiu, Radu Atanasiu, and Ion Petre. Parikh matrices and amiable words. Theoretical Computer Science, 390(1):102–109, January 22, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shi:2008:OPA

[1431] Yongqiang Shi and Deshi Ye. Online bin packing with arbitrary release times. Theoretical Computer Science, 390(1):110–119, January 22, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheng:2008:CNU

[1432] Yongxi Cheng, Ker-I Ko, and Weili Wu. On the complexity of non-unique probe selection. Theoretical Computer Science, 390(1):120–125, January 22, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:EBa

[1433] Anonymous. Editorial board. Theoretical Computer Science, 390(1):v– ix, January 22, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sassone:2008:FSS

[1434] Vladimiro Sassone. Foundations of Software Science and Computational Structures: Selected papers from FOS-SACS 2005. Theoretical Computer Science, 390(2–3):127–128, January 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abbes:2008:TCP

[1435] Samy Abbes and Albert Benveniste.

True-concurrency probabilistic models: Markov nets and a law of large numbers. Theoretical Computer Science, 390(2–3):129–170, January 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jeffrey:2008:FAP

[1436] Alan Jeffrey and Julian Rathke. Full abstraction for polymorphic π-calculus. Theoretical Computer Science, 390(2– 3):171–196, January 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Larsen:2008:ORM

[1437] Kim Guldstrand Larsen and Jacob Illum Rasmussen. Optimal reachability for multi-priced timed automata. Theoretical Computer Science, 390(2– 3):197–213, January 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Murawski:2008:TOI

[1438] Andrzej S. Murawski and Igor Walukiewicz. Third-order Idealized Algol with iteration is decidable. Theoretical Computer Science, 390(2–3):214– 229, January 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schroder:2008:ECM

[1439] Lutz Schröder. Expressivity of coalgebraic modal logic: The limits and beyond. Theoretical Computer Science, 390(2–3):230–247, January 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stark:2008:FAM

[1440] Ian Stark. Free-algebra models for the π-calculus. Theoretical Computer Science, 390(2–3):248–270, January 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berthe:2008:PSI

[1441] Valérie Berthé, Pierre Lecomte, and Michel Rigo. Preface to the special issue dedicated to combinatorics, automata and number theory. Theoretical Computer Science, 391(1–2):1– 2, February 4, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bell:2008:UBM

[1442] Paul Bell and Igor Potapov. On undecidability bounds for matrix decision problems. Theoretical Computer Science, 391(1–2):3–13, February 4, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czeizler:2008:MCT

[1443] Elena Czeizler. Multiple constraints on three and four words. Theoretical Computer Science, 391(1–2):14– 19, February 4, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Durand:2008:CST

[1444] F. Durand. Cobham–Semenov theorem and N^{d}-subshifts. Theoretical Computer Science, 391(1–2):20– 38, February 4, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frettloh:2008:SDT

[1445] D. Frettlöh. Self-dual tilings with respect to star-duality. Theoretical Computer Science, 391(1–2):39–50, February 4, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Glen:2008:CFW

[1446] Amy Glen. A characterization of fine words over a finite alphabet. Theoretical Computer Science, 391(1–2):51– 60, February 4, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Honkala:2008:CPP

[1447] Juha Honkala. Cancellation and periodicity properties of iterated morphisms. Theoretical Computer Science, 391(1–2):61–64, February 4, 2008. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koutschan:2008:RLT

[1448] Christoph Koutschan. Regular languages and their generating functions: The inverse problem. Theoretical Computer Science, 391(1–2):65–74, February 4, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brlek:2008:SOD

[1449] S. Brlek, G. Labelle, and A. Lacasse. Shuffle operations on discrete paths. Theoretical Computer Science, 391(1– 2):75–89, February 4, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Neraud:2008:CCC

[1450] Jean Néraud. Completing circular codes in regular submonoids. Theoretical Computer Science, 391(1–2):90– 98, February 4, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Petrov:2008:ARC

[1451] I. V. Petrov. An algorithm for recognition of n-collapsing words. Theoretical Computer Science, 391(1–2):99– 108, February 4, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Louchard:2008:GAC

[1452] Guy Louchard and Helmut Prodinger.

Generalized approximate counting revisited. Theoretical Computer Science, 391(1–2):109–125, February 4, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kao:2008:WAR

[1453] Jui-Yi Kao, Narad Rampersad, Jeffrey Shallit, and Manuel Silva. Words avoiding repetitions in arithmetic progressions. Theoretical Computer Science, 391(1–2):126–137, February 4, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schratzberger:2008:CAB

[1454] B. Schratzberger. A conversion algorithm based on the technique of singularization. Theoretical Computer Science, 391(1–2):138–149, February 4, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brignall:2008:SPD

[1455] Robert Brignall, Nik Ruškuc, and Vincent Vatter. Simple permutations: Decidability and unavoidable substructures. Theoretical Computer Science, 391(1–2):150–163, February 4, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bernat:2008:SI

[1456] Julien Bernat. Symmetrized β-integers. Theoretical Computer Science, 391(1– 2):164–177, February 4, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Puzynina:2008:PTD

[1457] S. A. Puzynina and S. V. Avgustinovich. On periodicity of two-dimensional words. Theoretical Computer Science, 391(1–2):178–187, February 4, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:EBb

[1458] Anonymous. Editorial board. Theoretical Computer Science, 391(1–2):v– ix, February 4, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Priami:2008:P

[1459] Corrado Priami. Preface. Theoretical Computer Science, 391(3):189, February 14, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cardelli:2008:PRS

[1460] Luca Cardelli. On process rate semantics. Theoretical Computer Science, 391(3):190–215, February 14, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cavaliere:2008:MBM

[1461] Matteo Cavaliere, Radu Mardare, and Sean Sedwards. A multiset-based model of synchronizing agents: Computability and robustness. Theoretical Computer Science, 391(3):216–238, February 14, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heath:2008:PMC

[1462] John Heath, Marta Kwiatkowska, Gethin Norman, David Parker, and Oksana Tymchyshyn. Probabilistic model checking of complex biological pathways. Theoretical Computer Science, 391(3):239–257, February 14, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Siebert:2008:TCL

[1463] Heike Siebert and Alexander Bockmayr. Temporal constraints in the logical analysis of regulatory networks. Theoretical Computer Science, 391(3): 258–275, February 14, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amir-Kroll:2008:GGP

[1464] Hila Amir-Kroll, Avital Sadot, Irun R.

Cohen, and David Harel. GemCell: a generic platform for modeling multicellular biological systems. Theoretical Computer Science, 391(3):276–290, February 14, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:Pa

[1465] Anonymous. Preface. Theoretical Computer Science, 391(3):i–v, February 14, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Buse:2008:E

[1466] Laurent Busé, Mohamed Elkadi, and Bernard Mourrain. Editorial. Theoretical Computer Science, 392(1–3):1– 2, February 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galligo:2008:P

[1467] André Galligo. Preface. Theoretical Computer Science, 392(1–3):3–4, February 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biasotti:2008:RGS

[1468] S. Biasotti, D. Giorgi, M. Spagnuolo, and B. Falcidieno. Reeb graphs for shape analysis and applications. Theoretical Computer Science, 392(1–3):5– 22, February 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cox:2008:MCI

[1469] David A. Cox. The moving curve ideal and the Rees algebra. Theoretical Computer Science, 392(1–3):23– 36, February 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dahan:2008:COR

[1470] Xavier Dahan, Xin Jin, Marc Moreno Maza, and Éric Schost. Change of order for regular chains in positive dimension. Theoretical Computer Science, 392(1– 3):37–65, February 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Damon:2008:SRS

[1471] James Damon. Swept regions and surfaces: Modeling and volumetric properties. Theoretical Computer Science, 392(1–3):66–91, February 28, 2008.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Haviv:2008:UAN

[1472] D. Haviv and Y. Yomdin. Uniform approximation of near-singular surfaces. Theoretical Computer Science, 392(1– 3):92–100, February 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kostov:2008:MS

[1473] Vladimir Petrov Kostov. On multiplier sequences. Theoretical Computer Science, 392(1–3):101–112, February 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lombardi:2008:SRF

[1474] Henri Lombardi and Claude Quitté.

Seminormal rings (following Thierry Coquand). Theoretical Computer Science, 392(1–3):113–127, February 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Simpson:2008:ACC

[1475] Carlos Simpson. Algebraic cycles from a computational point of view. Theoretical Computer Science, 392(1–3):128– 140, February 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sir:2008:CSR

[1476] Zbyněk Šír, Jens Gravesen, and Bert Jüttler. Curves and surfaces represented by polynomial support functions. Theoretical Computer Science, 392(1–3):141–157, February 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tsigaridas:2008:CRR

[1477] Elias P. Tsigaridas and Ioannis Z. Emiris. On the complexity of real root isolation using continued fractions. Theoretical Computer Science, 392(1–3):158–173, February 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yengui:2008:MUM

[1478] Ihsen Yengui. Making the use of maximal ideals constructive. Theoretical Computer Science, 392(1–3):174– 178, February 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:E

[1479] Anonymous. Editorial. Theoretical Computer Science, 392(1–3):v–ix, February 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bu:2008:UCA

[1480] Tian-Ming Bu, Qi Qi, and Aries Wei Sun. Unconditional competitive auctions with copy and budget constraints. Theoretical Computer Science, 393(1– 3):1–13, March 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bein:2008:FAA

[1481] Wolfgang Bein, José R. Correa, and Xin Han. A fast asymptotic approximation scheme for bin packing with rejection. Theoretical Computer Science, 393(1–3):14–22, March 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bucci:2008:DGE

[1482] Michelangelo Bucci, Aldo de Luca, Alessandro De Luca, and Luca Q. Zamboni. On different generalizations of episturmian words. Theoretical Computer Science, 393(1–3):23– 36, March 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferrante:2008:MNE

[1483] Alessandro Ferrante and Mimmo Parente. Mixed Nash equilibria in selfish routing problems with dynamic constraints. Theoretical Computer Science, 393(1–3):37–53, March 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berman:2008:AOS

[1484] Piotr Berman and Bhaskar DasGupta.

Approximating the online set multicover problems via randomized winnowing. Theoretical Computer Science, 393(1–3):54–71, March 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bui-Xuan:2008:CGS

[1485] Binh-Minh Bui-Xuan, Michel Habib, and Christophe Paul. Competitive graph searches. Theoretical Computer Science, 393(1–3):72–80, March 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2008:ECW

[1486] Hao Li, Huifang Miao, Li Liu, Lian Li, and Heping Zhang. Energy conservation in wireless sensor networks and connectivity of graphs. Theoretical Computer Science, 393(1–3):81– 89, March 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Suzuki:2008:MOG

[1487] Tomoko Suzuki, Taisuke Izumi, Fukuhito Ooshita, Hirotsugu Kakugawa, and Toshimitsu Masuzawa. Move-optimal gossiping among mobile agents. Theoretical Computer Science, 393(1–3):90–101, March 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lu:2008:SMS

[1488] Lingfa Lu, Jinjiang Yuan, and Liqi Zhang. Single machine scheduling with release dates and job delivery to minimize the makespan. Theoretical Computer Science, 393(1–3):102– 108, March 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Buchholz:2008:BRW

[1489] Peter Buchholz. Bisimulation relations for weighted automata. Theoretical Computer Science, 393(1–3):109– 123, March 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nedev:2008:MDG

[1490] Z. Nedev and S. Muthukrishnan. The Magnus–Derek game. Theoretical Computer Science, 393(1–3):124– 132, March 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nagamochi:2008:AVS

[1491] Hiroshi Nagamochi and Takaharu Ohnishi. Approximating a vehicle scheduling problem with time windows and handling times. Theoretical Computer Science, 393(1–3):133– 146, March 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bozapalidis:2008:GA

[1492] Symeon Bozapalidis and Antonios Kalampakas. Graph automata. Theoretical Computer Science, 393(1–3): 147–165, March 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brlek:2008:SWL

[1493] S. Brlek, D. Jamet, and G. Paquin.

Smooth words on 2-letter alphabets having same parity. Theoretical Computer Science, 393(1–3):166–181, March 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Broersma:2008:CCP

[1494] Hajo Broersma, Matthew Johnson, Daniël Paulusma, and Iain A. Stewart. The computational complexity of the parallel knock-out problem. Theoretical Computer Science, 393(1–3):182– 195, March 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berstel:2008:APS

[1495] Jean Berstel and Christophe Reutenauer.

Another proof of Soittola’s theorem. Theoretical Computer Science, 393(1– 3):196–203, March 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuo:2008:PMS

[1496] Wen-Hung Kuo and Dar-Li Yang. Parallel-machine scheduling with time dependent processing times. Theoretical Computer Science, 393(1–3):204– 210, March 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Angelelli:2008:SOS

[1497] Enrico Angelelli, Maria Grazia Speranza, and Zsolt Tuza. Semi-online scheduling on two uniform processors. Theoretical Computer Science, 393(1– 3):211–219, March 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferrante:2008:HOP

[1498] Alessandro Ferrante, Gopal Pandurangan, and Kihong Park. On the hardness of optimization in power-law graphs. Theoretical Computer Science, 393(1– 3):220–230, March 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kujala:2008:COB

[1499] Jussi Kujala and Tapio Elomaa. The cost of offline binary search tree algorithms and the complexity of the request sequence. Theoretical Computer Science, 393(1–3):231–239, March 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Xavier:2008:CCP

[1500] E. C. Xavier and F. K. Miyazawa. The class constrained bin packing problem with applications to video-on-demand. Theoretical Computer Science, 393(1– 3):240–259, March 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daude:2008:PSA

[1501] Hervé Daudé, Marc Mézard, Thierry Mora, and Riccardo Zecchina. Pairs of SAT-assignments in random Boolean formulæ. Theoretical Computer Science, 393(1–3):260–279, March 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huang:2008:ANW

[1502] Y. B. Huang. About the number of C^{∞}-words of form w̃ × w. Theoretical Computer Science, 393(1–3):280– 286, March 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tian:2008:BLA

[1503] Ji Tian, Ruyan Fu, and Jinjiang Yuan.

A best on-line algorithm for single machine scheduling with small delivery times. Theoretical Computer Science, 393(1–3):287–293, March 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Segal:2008:CCF

[1504] Michael Segal and Eli Zeitlin. Computing closest and farthest points for a query segment. Theoretical Computer Science, 393(1–3):294–300, March 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:EBc

[1505] Anonymous. Editorial board. Theoretical Computer Science, 393(1–3): v–ix, March 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Courcelle:2008:MDC

[1506] Bruno Courcelle and Christian Delhommé. The modular decomposition of countable graphs. definition and construction in monadic secondorder logic. Theoretical Computer Science, 394(1–2):1–38, March 31, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2008:CLR

[1507] Yu Zhang. Cryptographic logical relations. Theoretical Computer Science, 394(1–2):39–63, March 31, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laird:2008:DSC

[1508] J. Laird. Decidability and syntactic control of interference. Theoretical Computer Science, 394(1–2):64– 83, March 31, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calders:2008:IFS

[1509] Toon Calders. Itemset frequency satisfiability: Complexity and axiomatization. Theoretical Computer Science, 394(1–2):84–111, March 31, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garcia:2008:SCC

[1510] Flavio D. Garcia and Peter van Rossum. Sound and complete computational interpretation of symbolic hashes in the standard model. Theoretical Computer Science, 394(1–2):112– 133, March 31, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kretinsky:2008:PNL

[1511] Mojmír Křetínský, Vojtěch Řehák, and Jan Strejček. Petri nets are less expressive than state-extended PA. Theoretical Computer Science, 394(1–2):134– 140, March 31, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:EBd

[1512] Anonymous. Editorial board. Theoretical Computer Science, 394(1–2): v–ix, March 31, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beckmann:2008:GEC

[1513] Arnold Beckmann, Edwin Beggs, and Benedikt Löwe. From Gödel to Einstein: Computability between logic and physics at CiE 2006. Theoretical Computer Science, 394(3):141–143, April 8, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Avron:2008:CDV

[1514] Arnon Avron. Constructibility and decidability versus domain independence and absoluteness. Theoretical Computer Science, 394(3):144–158, April 8, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cockshott:2008:PCH

[1515] Paul Cockshott, Lewis Mackenzie, and Greg Michaelson. Physical constraints on hypercomputation. Theoretical Computer Science, 394(3):159– 174, April 8, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fouche:2008:DGB

[1516] Willem L. Fouché. Dynamics of a generic Brownian motion: Recursive aspects. Theoretical Computer Science, 394(3):175–186, April 8, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kendon:2008:OCN

[1517] Viv Kendon and Olivier Maloyer. Optimal computation with non-unitary quantum walks. Theoretical Computer Science, 394(3):187–196, April 8, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koepke:2008:MCO

[1518] Peter Koepke and Ryan Siders. Minimality considerations for ordinal computers modeling constructibility. Theoretical Computer Science, 394(3):197– 207, April 8, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nagy:2008:IVC

[1519] Benedek Nagy and Sándor Vályi. Interval-valued computations and their connection with PSPACE. Theoretical Computer Science, 394(3):208– 222, April 8, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Welch:2008:BLN

[1520] Philip D. Welch. Bounding lemmata for non-deterministic halting times of transfinite Turing machines. Theoretical Computer Science, 394(3):223– 228, April 8, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2008:Pa

[1521] G. Rozenberg. Preface. Theoretical Computer Science, 395(1):1–2, April 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brun:2008:NPT

[1522] Yuriy Brun. Nondeterministic polynomial time factoring in the tile assembly model. Theoretical Computer Science, 395(1):3–23, April 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Langille:2008:SVP

[1523] Miika Langille and Ion Petre. Sequential vs. parallel complexity in simple gene assembly. Theoretical Computer Science, 395(1):24–30, April 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brun:2008:SNC

[1524] Yuriy Brun. Solving NP-complete problems in the tile assembly model. Theoretical Computer Science, 395(1): 31–46, April 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gao:2008:GCP

[1525] Shile Gao and Kequan Ding. A graphical criterion of planarity for RNA secondary structures with pseudoknots in Rivas–Eddy class. Theoretical Computer Science, 395(1):47–56, April 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zaharakis:2008:MSN

[1526] Ioannis D. Zaharakis and Achilles D. Kameas. Modeling spiking neural networks. Theoretical Computer Science, 395(1):57–76, April 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barbuti:2008:CSB

[1527] Roberto Barbuti, Andrea Maggiolo-Schettini, Paolo Milazzo, and Simone Tini. Compositional semantics and behavioral equivalences for P Systems. Theoretical Computer Science, 395(1): 77–100, April 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Morita:2008:RCC

[1528] Kenichi Morita. Reversible computing and cellular automata — a survey. Theoretical Computer Science, 395(1): 101–131, April 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barrett:2008:EPP

[1529] Chris Barrett, Harry B. Hunt III, Madhav V. Marathe, S. S. Ravi, Daniel J. Rosenkrantz, Richard E. Stearns, and Mayur Thakur. Errata for the paper “Predecessor existence problems for finite discrete dynamical systems” [Theoret. Comput. Sci. 386 (1–2) (2007) 3– 37]. Theoretical Computer Science, 395 (1):132–133, April 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [1327].

Anonymous:2008:Pb

[1530] Anonymous. Preface. Theoretical Computer Science, 395(1):v–ix, April 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Giancarlo:2008:FSI

[1531] Raffaele Giancarlo and Stefano Lonardi.

Foreword: Special issue in honor of the 60th birthday of Professor Alberto Apostolico: Work is for people who do not know how to: SAIL — String Algorithms, Information and Learning. Theoretical Computer Science, 395(2– 3):135–136, May 1, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pizzi:2008:FPM

[1532] Cinzia Pizzi and Esko Ukkonen. Fast profile matching algorithms — a survey. Theoretical Computer Science, 395 (2–3):137–157, May 1, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Comin:2008:DSV

[1533] Matteo Comin and Laxmi Parida. Detection of subtle variations as consensus motifs. Theoretical Computer Science, 395(2–3):158–170, May 1, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lemaitre:2008:STU

[1534] Claire Lemaitre and Marie-France Sagot. A small trip in the untranquil world of genomes: a survey on the detection and analysis of genome rearrangement breakpoints. Theoretical Computer Science, 395(2–3):171– 192, May 1, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alekseyev:2008:MBR

[1535] Max A. Alekseyev and Pavel A. Pevzner. Multi-break rearrangements and chromosomal evolution. Theoretical Computer Science, 395(2–3):193– 202, May 1, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jacquet:2008:EHM

[1536] Philippe Jacquet, Gadiel Seroussi, and Wojciech Szpankowski. On the entropy of a hidden Markov process. Theoretical Computer Science, 395(2–3): 203–219, May 1, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schurmann:2008:CSA

[1537] Klaus-Bernd Schürmann and Jens Stoye. Counting suffix arrays and strings. Theoretical Computer Science, 395(2–3):220–234, May 1, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nederhof:2008:CDR

[1538] Mark-Jan Nederhof and Giorgio Satta. Computation of distances for regular and context-free probabilistic languages. Theoretical Computer Science, 395(2–3):235–254, May 1, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Iliopoulos:2008:ACV

[1539] Costas S. Iliopoulos and M. Sohel Rahman. Algorithms for computing variants of the longest common subsequence problem. Theoretical Computer Science, 395(2–3):255–267, May 1, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:2008:CSR

[1540] Jin Wook Kim, Amihood Amir, Gad M. Landau, and Kunsoo Park. Computing similarity of run-length encoded strings with affine gap penalty. Theoretical Computer Science, 395(2– 3):268–282, May 1, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crochemore:2008:AIP

[1541] Maxime Crochemore, Danny Hermelin, Gad M. Landau, Dror Rawitz, and Stéphane Vialette. Approximating the 2-interval pattern problem. Theoretical Computer Science, 395(2–3):283– 297, May 1, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amir:2008:PMW

[1542] Amihood Amir, Eran Chencinski, Costas Iliopoulos, Tsvi Kopelowitz, and Hui Zhang. Property matching and weighted matching. Theoretical Computer Science, 395(2–3):298–310, May 1, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2008:ILB

[1543] WenBin Chen, Jiangtao Meng, and Dengpan Yin. An improved lower bound for approximating minimum GCD multiplier in ℓ_{∞} norm (GCDM_{∞}). Theoretical Computer Science, 396(1– 3):1–9, May 10, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salinas:2008:CS

[1544] L. Salinas and E. Goles. Covering by squares. Theoretical Computer Science, 396(1–3):10–27, May 10, 2008.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2008:IMM

[1545] Longcheng Liu and Enyu Yao. Inverse min–max spanning tree problem under the Weighted sum-type Hamming distance. Theoretical Computer Science, 396(1–3):28–34, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dukes:2008:CGC

[1546] W. M. B. Dukes, M. F. Flanagan, T. Mansour, and V. Vajnovszki. Combinatorial Gray codes for classes of pattern avoiding permutations. Theoretical Computer Science, 396(1–3): 35–49, May 10, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shieh:2008:JMA

[1547] Min-Zheng Shieh and Shi-Chun Tsai. Jug measuring: Algorithms and complexity. Theoretical Computer Science, 396(1–3):50–62, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Escoffier:2008:BDA

[1548] Bruno Escoffier and Jérôme Monnot. A better differential approximation ratio for symmetric TSP. Theoretical Computer Science, 396(1–3):63–70, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Couceiro:2008:QOB

[1549] Miguel Couceiro and Maurice Pouzet. On a quasi-ordering on Boolean functions. Theoretical Computer Science, 396(1–3):71–87, May 10, 2008.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fenner:2008:MNP

[1550] Trevor Fenner, Mark Levene, and George Loizou. Modelling the navigation potential of a web page. Theoretical Computer Science, 396(1–3): 88–96, May 10, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Defago:2008:NUC

[1551] Xavier Défago and Samia Souissi. Nonuniform circle formation algorithm for oblivious mobile robots with convergence toward uniformity. Theoretical Computer Science, 396(1–3):97– 112, May 10, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dunne:2008:CDR

[1552] Paul E. Dunne and Yann Chevaleyre.

The complexity of deciding reachability properties of distributed negotiation schemes. Theoretical Computer Science, 396(1–3):113–144, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yuster:2008:APD

[1553] Raphael Yuster. All-pairs disjoint paths from a common ancestor in Õ(n^{ω}) time. Theoretical Computer Science, 396(1–3):145–150, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2008:OCC

[1554] Xueliang Li and Wenli Zhou. The 2ndorder conditional 3-coloring of clawfree graphs. Theoretical Computer Science, 396(1–3):151–157, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Irani:2008:PAS

[1555] Sandy Irani and Vitus Leung. Probabilistic analysis for scheduling with conflicts. Theoretical Computer Science, 396(1–3):158–179, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2008:AOD

[1556] Chen Wang, Myung-Ah Park, James Willson, Yongxi Cheng, Andras Farago, and Weili Wu. On approximate optimal dual power assignment for biconnectivity and edge-biconnectivity. Theoretical Computer Science, 396(1– 3):180–190, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2008:WVP

[1557] Ying-You Chen, Dyi-Rong Duh, Tai-Ling Ye, and Jung-Sheng Fu. Weak-vertex-pancyclicity of (n, k)-star graphs. Theoretical Computer Science, 396(1–3):191–199, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hung:2008:CAM

[1558] Regant Y. S. Hung and H. F. Ting. Competitive analysis of most-request-first for scheduling broadcasts with start-up delay. Theoretical Computer Science, 396(1–3):200–211, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chandler:2008:PPC

[1559] David B. Chandler, Maw-Shang Chang, Ton Kloks, Jiping Liu, and Sheng-Lung Peng. Partitioned probe comparability graphs. Theoretical Computer Science, 396(1–3):212–222, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boyar:2008:TBM

[1560] Joan Boyar and René Peralta. Tight bounds for the multiplicative complexity of symmetric functions. Theoretical Computer Science, 396(1–3):223– 246, May 10, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goles:2008:CBP

[1561] E. Goles and L. Salinas. Comparison between parallel and serial dynamics of Boolean networks. Theoretical Computer Science, 396(1–3):247– 253, May 10, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Currie:2008:PPT

[1562] James D. Currie. Palindrome positions in ternary square-free words. Theoretical Computer Science, 396(1–3): 254–257, May 10, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2008:PCG

[1563] Guizhen Liu and Ping Li. Paths in circuit graphs of matroids. Theoretical Computer Science, 396(1–3):258– 263, May 10, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kari:2008:WUI

[1564] Lila Kari and Petr Sosík. On the weight of universal insertion grammars. Theoretical Computer Science, 396(1–3): 264–270, May 10, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Buss:2008:NHF

[1565] Samuel R. Buss and Jan Hoffmann. The NP-hardness of finding a directed acyclic graph for regular resolution. Theoretical Computer Science, 396(1– 3):271–276, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Neven:2008:STS

[1566] Gregory Neven. A simple transitive signature scheme for directed trees. Theoretical Computer Science, 396(1–3): 277–282, May 10, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lu:2008:UPB

[1567] Lingfa Lu, Liqi Zhang, and Jinjiang Yuan. The unbounded parallel batch machine scheduling with release dates and rejection to minimize makespan. Theoretical Computer Science, 396(1– 3):283–289, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Myung:2008:CPP

[1568] Young-Soo Myung. On the clique partitioning problem in weighted interval graphs. Theoretical Computer Science, 396(1–3):290–293, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Verbitsky:2008:OCP

[1569] Oleg Verbitsky. On the obfuscation complexity of planar graphs. Theoretical Computer Science, 396(1–3):294– 300, May 10, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:EBe

[1570] Anonymous. Editorial board. Theoretical Computer Science, 396(1–3): v–ix, May 10, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Case:2008:F

[1571] John Case, Takeshi Shinohara, Thomas Zeugmann, and Sandra Zilles. Foreword. Theoretical Computer Science, 397(1–3):1–3, May 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zeugmann:2008:LRF

[1572] Thomas Zeugmann and Sandra Zilles.

Learning recursive functions: a survey. Theoretical Computer Science, 397(1– 3):4–56, May 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grieser:2008:RII

[1573] Gunter Grieser. Reflective inductive inference of recursive functions. Theoretical Computer Science, 397(1–3): 57–69, May 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Freivalds:2008:QII

[1574] R. Freivalds and R. F. Bonner. Quantum inductive inference by finite automata. Theoretical Computer Science, 397(1–3):70–76, May 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Poland:2008:NBC

[1575] Jan Poland. Nonstochastic bandits: Countable decision set, unbounded costs and reactive environments. Theoretical Computer Science, 397(1–3): 77–93, May 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balbach:2008:MTU

[1576] Frank J. Balbach. Measuring teachability using variants of the teaching dimension. Theoretical Computer Science, 397(1–3):94–113, May 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jain:2008:AVP

[1577] Sanjay Jain, Eric Martin, and Frank Stephan. Absolute versus probabilistic classification in a logical setting. Theoretical Computer Science, 397(1–3): 114–128, May 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rao:2008:SCT

[1578] M. R. K. Krishna Rao. Some classes of term rewriting systems inferable from positive data. Theoretical Computer Science, 397(1–3):129–149, May 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ng:2008:DEL

[1579] Yen Kaow Ng and Takeshi Shinohara. Developments from enquiries into the learnability of the pattern languages from positive data. Theoretical Computer Science, 397(1–3):150– 165, May 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reidenbach:2008:DPI

[1580] Daniel Reidenbach. Discontinuities in pattern inference. Theoretical Computer Science, 397(1–3):166–193, May 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lange:2008:LIF

[1581] Steffen Lange, Thomas Zeugmann, and Sandra Zilles. Learning indexed families of recursive languages from positive data: a survey. Theoretical Computer Science, 397(1–3):194–232, May 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jain:2008:LES

[1582] Sanjay Jain and Efim Kinber. Learning and extending sublanguages. Theoretical Computer Science, 397(1–3): 233–246, May 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:Fb

[1583] Anonymous. Foreword. Theoretical Computer Science, 397(1–3):v–ix, May 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berardi:2008:CTA

[1584] Stefano Berardi and Ugo de’Liguoro.

Calculi, types and applications: Essays in honour of M. Coppo, M. Dezani-Ciancaglini and S. Ronchi Della Rocca. Theoretical Computer Science, 398(1– 3):1–11, May 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barendregt:2008:TRP

[1585] Henk Barendregt. Towards the range property for the lambda theory H. Theoretical Computer Science, 398(1–3): 12–15, May 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klop:2008:LCP

[1586] Jan Willem Klop, Vincent van Oostrom, and Roel de Vrijer. Lambda calculus with patterns. Theoretical Computer Science, 398(1–3):16–31, May 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lago:2008:WLC

[1587] Ugo Dal Lago and Simone Martini. The weak lambda calculus as a reasonable machine. Theoretical Computer Science, 398(1–3):32–50, May 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paolini:2008:PE

[1588] Luca Paolini. Parametric λ-theories. Theoretical Computer Science, 398(1– 3):51–62, May 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boudol:2008:SNT

[1589] Gérard Boudol. On strong normalization and type inference in the intersection type discipline. Theoretical Computer Science, 398(1–3):63–81, May 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanBakel:2008:HIT

[1590] Steffen van Bakel. The heart of intersection type assignment: Normalisation proofs revisited. Theoretical Computer Science, 398(1–3):82–94, May 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bono:2008:TLC

[1591] Viviana Bono, Betti Venneri, and Lorenzo Bettini. A typed lambda calculus with intersection types. Theoretical Computer Science, 398(1–3): 95–113, May 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dougherty:2008:CSN

[1592] Daniel J. Dougherty, Silvia Ghilezan, and Pierre Lescanne. Characterizing strong normalization in the Curien– Herbelin symmetric lambda calculus: Extending the Coppo–Dezani heritage. Theoretical Computer Science, 398(1– 3):114–128, May 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alessi:2008:IFM

[1593] Fabio Alessi. An irregular filter model.

Theoretical Computer Science, 398(1– 3):129–149, May 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DiGianantonio:2008:TAS

[1594] Pietro Di Gianantonio, Furio Honsell, and Marina Lenisa. A type assignment system for game semantics. Theoretical Computer Science, 398(1–3):

150–169, May 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoyrup:2008:CMC

[1595] Mathieu Hoyrup, Arda Kolçak, and Giuseppe Longo. Computability and the morphological complexity of some dynamics on continuous domains. Theoretical Computer Science, 398(1–3): 170–182, May 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Margaria:2008:ACM

[1596] Ines Margaria and Maddalena Zacchi. Access control in mobile ambient calculi: a comparative view. Theoretical Computer Science, 398(1–3):183– 202, May 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Compagnoni:2008:RBA

[1597] Adriana Compagnoni, Elsa L. Gunter, and Philippe Bidinger. Role-based access control for boxed ambients. Theoretical Computer Science, 398(1–3): 203–216, May 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Castagna:2008:SSP

[1598] Giuseppe Castagna, Rocco De Nicola, and Daniele Varacca. Semantic subtyping for the pi-calculus. Theoretical Computer Science, 398(1–3):217– 242, May 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liquori:2008:EFJ

[1599] Luigi Liquori and Arnaud Spiwack.

Extending FeatherTrait Java with Interfaces. Theoretical Computer Science, 398(1–3):243–260, May 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:EBf

[1600] Anonymous. Editorial board. Theoretical Computer Science, 398(1–3): v–ix, May 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flocchini:2008:P

[1601] Paola Flocchini and Leszek A.

Ga̧sieniec. Preface. Theoretical Computer Science, 399(1–2):1–2, June 3, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bar-Noy:2008:ODM

[1602] Amotz Bar-Noy, Richard E. Ladner, and Tami Tamir. Optimal delay for media-on-demand with pre-loading and pre-buffering. Theoretical Computer Science, 399(1–2):3–11, June 3, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blin:2008:DCN

[1603] Lélia Blin, Pierre Fraigniaud, Nicolas Nisse, and Sandrine Vial. Distributed chasing of network intruders. Theoretical Computer Science, 399(1–2): 12–37, June 3, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calamoneri:2008:MEB

[1604] Tiziana Calamoneri, Andrea E. F. Clementi, Miriam Di Ianni, Massimo Lauria, Angelo Monti, and Riccardo Silvestri. Minimum-Energy Broadcast and disk cover in grid wireless networks. Theoretical Computer Science, 399(1–2):38–53, June 3, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chalopin:2008:ERI

[1605] Jérémie Chalopin. Election and rendezvous with incomparable labels. Theoretical Computer Science, 399(1–2): 54–70, June 3, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cohen:2008:LSA

[1606] Reuven Cohen and David Peleg. Local spreading algorithms for autonomous robot systems. Theoretical Computer Science, 399(1–2):71–82, June 3, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Derbel:2008:FDD

[1607] Bilel Derbel and Cyril Gavoille. Fast deterministic distributed algorithms for sparse spanners. Theoretical Computer Science, 399(1–2):83–100, June 3, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dobrev:2008:FDF

[1608] Stefan Dobrev, Rastislav Královič, Richard Královič, and Nicola Santoro. On fractional dynamic faults with thresholds. Theoretical Computer Science, 399(1–2):101–117, June 3, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goddard:2008:DKS

[1609] Wayne Goddard, Stephen T. Hedetniemi, David P. Jacobs, and Vilmar Trevisan. Distance-k knowledge in self-stabilizing algorithms. Theoretical Computer Science, 399(1–2):118– 127, June 3, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kosowski:2008:MED

[1610] Adrian Kosowski. The maximum edgedisjoint paths problem in complete graphs. Theoretical Computer Science, 399(1–2):128–140, June 3, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kowalski:2008:HMA

[1611] Dariusz R. Kowalski and Adam Malinowski. How to meet in anonymous network. Theoretical Computer Science, 399(1–2):141–156, June 3, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:Pc

[1612] Anonymous. Preface. Theoretical Computer Science, 399(1–2):v–ix, June 3, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fomin:2008:FSI

[1613] Fedor V. Fomin, Pierre Fraigniaud, and Dimitrios M. Thilikos. Forewords: Special issue on graph searching. Theoretical Computer Science, 399(3):157, June 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alspach:2008:TCG

[1614] Brian Alspach, Danny Dyer, Denis Hanson, and Boting Yang. Time constrained graph searching. Theoretical Computer Science, 399(3):158– 168, June 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mazoit:2008:MND

[1615] Frédéric Mazoit and Nicolas Nisse. Monotonicity of non-deterministic graph searching. Theoretical Computer Science, 399(3):169–178, June 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Isler:2008:RIC

[1616] Volkan Isler and Nikhil Karnad. The role of information in the cop-robber game. Theoretical Computer Science, 399(3):179–190, June 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Messinger:2008:CNB

[1617] M. E. Messinger, R. J. Nowakowski, and P. Prałat. Cleaning a network with brushes. Theoretical Computer Science, 399(3):191–205, June 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hunter:2008:DMK

[1618] Paul Hunter and Stephan Kreutzer. Digraph measures: Kelly decompositions, games, and orderings. Theoretical Computer Science, 399(3):206– 219, June 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dumitrescu:2008:OVL

[1619] Adrian Dumitrescu, Ichiro Suzuki, and Paweł Żyliński. Offline variants of the “lion and man” problem: —Some problems and techniques for measuring crowdedness and for safe path planning—. Theoretical Computer Science, 399(3):220–235, June 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fomin:2008:ABG

[1620] Fedor V. Fomin and Dimitrios M. Thilikos. An annotated bibliography on guaranteed graph searching. Theoretical Computer Science, 399(3):236– 245, June 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sablik:2008:DDC

[1621] Mathieu Sablik. Directional dynamics for cellular automata: a sensitivity to initial condition approach. Theoretical Computer Science, 400(1–3): 1–18, June 9, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blumensath:2008:SGC

[1622] Achim Blumensath. On the structure of graphs in the Caucal hierarchy. Theoretical Computer Science, 400(1–3): 19–45, June 9, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [2398].

Moser:2008:MFD

[1623] Philippe Moser. Martingale families and dimension in P. Theoretical Computer Science, 400(1–3):46–61, June 9, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moshe:2008:SQR

[1624] Yossi Moshe. On some questions regarding k-regular and k-context-free sequences. Theoretical Computer Science, 400(1–3):62–69, June 9, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pina:2008:RWM

[1625] Claribet Piña and Carlos Uzcátegui. Reconstruction of a word from a multiset of its factors. Theoretical Computer Science, 400(1–3):70–83, June 9, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheng:2008:UAF

[1626] Christine Cheng, Eric McDermid, and Ichiro Suzuki. A unified approach to finding good stable matchings in the hospitals/residents setting. Theoretical Computer Science, 400(1–3): 84–99, June 9, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kirillovs:2008:PCC

[1627] Jurijs Kirillovs. Polyomino coloring and complex numbers. Theoretical Computer Science, 400(1–3):100– 112, June 9, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ambroz:2008:MIP

[1628] P. Ambroź, Z. Masáková, and E. Pelantová. Matrices of 3-iet preserving morphisms. Theoretical Computer Science, 400(1–3):113–136, June 9, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jing:2008:HAT

[1629] Caixia Jing, Guochun Tang, and Xingsan Qian. Heuristic algorithms for two machine re-entrant flow shop. Theoretical Computer Science, 400(1– 3):137–143, June 9, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berstel:2008:EBD

[1630] Jean Berstel and Christophe Reutenauer.

Extension of Brzozowski’s derivation calculus of rational expressions to series over the free partially commutative monoids. Theoretical Computer Science, 400(1–3):144–158, June 9, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Parekh:2008:AAP

[1631] Ojas Parekh. Approximation algorithms for partially covering with edges. Theoretical Computer Science, 400(1–3):159–168, June 9, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krieger:2008:SIW

[1632] Dalia Krieger. On stabilizers of infinite words. Theoretical Computer Science, 400(1–3):169–181, June 9, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rhodes:2008:TMB

[1633] John Rhodes and Pedro V. Silva. Turing machines and bimachines. Theoretical Computer Science, 400(1–3): 182–224, June 9, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ceccherini-Silberstein:2008:GCH

[1634] Tullio Ceccherini-Silberstein and Michel Coornaert. A generalization of the Curtis–Hedlund theorem. Theoretical Computer Science, 400(1–3):225– 229, June 9, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shao:2008:IUB

[1635] Zhendong Shao and David Zhang. Improved upper bounds on the L(2, 1)- labeling of the skew and converse skew product graphs. Theoretical Computer Science, 400(1–3):230–233, June 9, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:EBg

[1636] Anonymous. Editorial board. Theoretical Computer Science, 400(1–3): v–ix, June 9, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karagiorgos:2008:CDM

[1637] Gregory Karagiorgos and Nikolaos M. Missirlis. Convergence of the diffusion method for weighted torus graphs using Fourier analysis. Theoretical Computer Science, 401(1–3):1–16, July 23, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Michel:2008:RBT

[1638] Christian J. Michel, Giuseppe Pirillo, and Mario A. Pirillo. A relation between trinucleotide comma-free codes and trinucleotide circular codes. Theoretical Computer Science, 401(1–3): 17–26, July 23, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hemaspaandra:2008:EDA

[1639] Lane A. Hemaspaandra, Jörg Rothe, and Amitabh Saxena. Enforcing and defying associativity, commutativity, totality, and strong noninvertibility for worst-case one-way functions. Theoretical Computer Science, 401(1–3): 27–35, July 23, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cohen:2008:GSH

[1640] David A. Cohen, Martin C. Cooper, and Peter G. Jeavons. Generalising submodularity and horn clauses: Tractable optimization problems defined by tournament pair multimorphisms. Theoretical Computer Science, 401(1–3):36–51, July 23, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2008:ODC

[1641] Zhi-Zhong Chen, Ze Feng, Minming Li, and Frances Yao. Optimizing deletion cost for secure multicast key management. Theoretical Computer Science, 401(1–3):52–61, July 23, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marx:2008:CUL

[1642] Dániel Marx. Complexity of unique list colorability. Theoretical Computer Science, 401(1–3):62–76, July 23, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ting:2008:NOS

[1643] Hing-Fung Ting. A near optimal scheduler for on-demand data broadcasts.

Theoretical Computer Science, 401(1– 3):77–84, July 23, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huo:2008:OSE

[1644] Yumei Huo, Joseph Y.-T. Leung, and Xin Wang. Online scheduling of equalprocessing-time task systems. Theoretical Computer Science, 401(1–3): 85–95, July 23, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brenner:2008:GSC

[1645] Janina Brenner and Guido Schäfer. Group-strategyproof cost sharing mechanisms for makespan and other scheduling problems. Theoretical Computer Science, 401(1–3):96–106, July 23, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Giroudeau:2008:CAP

[1646] Rodolphe Giroudeau, Jean-Claude Konig, Farida Kamila Moulai, and Jérôme Palaysi. Complexity and approximation for precedence constrained scheduling problems with large communication delays. Theoretical Computer Science, 401(1–3):107–119, July 23, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laber:2008:RCA

[1647] Eduardo Sany Laber. A randomized competitive algorithm for evaluating priced AND/OR trees. Theoretical Computer Science, 401(1–3):120– 130, July 23, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chepoi:2008:AAF

[1648] Victor Chepoi, Bertrand Estellon, and Yann Vaxès. Approximation algorithms for forests augmentation ensuring two disjoint paths of bounded length. Theoretical Computer Science, 401(1–3):131–143, July 23, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonifaci:2008:CUN

[1649] Vincenzo Bonifaci, Ugo Di Iorio, and Luigi Laura. The complexity of uniform Nash equilibria and related regular subgraph problems. Theoretical Computer Science, 401(1–3):144– 152, July 23, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kanj:2008:STT

[1650] Iyad A. Kanj, Luay Nakhleh, Cuong Than, and Ge Xia. Seeing the trees and their branches in the network is hard. Theoretical Computer Science, 401(1– 3):153–164, July 23, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Puglisi:2008:HMR

[1651] Simon J. Puglisi, Jamie Simpson, and W. F. Smyth. How many runs can a string contain? Theoretical Computer Science, 401(1–3):165–171, July 23, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bachmat:2008:AST

[1652] Eitan Bachmat, Tao Kai Lam, and Avner Magen. Analysis of set-up time models: a metric perspective. Theoretical Computer Science, 401(1–3):

172–180, July 23, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tsai:2008:EMM

[1653] Chang-Hsiung Tsai. Embedding of meshes in Möbius cubes. Theoretical Computer Science, 401(1–3):181– 190, July 23, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hernich:2008:RCR

[1654] André Hernich and Nicole Schweikardt. Reversal complexity revisited. Theoretical Computer Science, 401(1–3): 191–205, July 23, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Costa:2008:BIV

[1655] José Carlos Costa and Conceição Nogueira. On bases of identities for the ω-variety generated by locally testable semigroups. Theoretical Computer Science, 401(1–3):206–216, July 23, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kauers:2008:SDE

[1656] Manuel Kauers. Solving difference equations whose coefficients are not transcendental. Theoretical Computer Science, 401(1–3):217–227, July 23, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chang:2008:CTF

[1657] Ching-Lueh Chang, Yuh-Dauh Lyuu, and Yen-Wu Ti. The complexity of Tarski’s fixed point theorem. Theoretical Computer Science, 401(1–3):

228–235, July 23, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ilcinkas:2008:SPN

[1658] David Ilcinkas. Setting port numbers for fast graph exploration. Theoretical Computer Science, 401(1–3):236– 242, July 23, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuske:2008:STF

[1659] Dietrich Kuske. Schützenberger’s theorem on formal power series follows from Kleene’s theorem. Theoretical Computer Science, 401(1–3):243– 248, July 23, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:EBh

[1660] Anonymous. Editorial board. Theoretical Computer Science, 401(1–3): v–ix, July 23, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nikoletseas:2008:P

[1661] Sotiris Nikoletseas and Jose Rolim.

Preface. Theoretical Computer Science, 402(1):1, July 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bertossi:2008:ECT

[1662] Alan A. Bertossi, Stephan Olariu, and Cristina M. Pinotti. Efficient corona training protocols for sensor networks. Theoretical Computer Science, 402(1): 2–15, July 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cichon:2008:AIA

[1663] J. Cichoń, M. Kutyłowski, and M. Zawada. Adaptive initialization algorithm for ad hoc radio networks with carrier sensing. Theoretical Computer Science, 402(1):16–28, July 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Halldorsson:2008:MIW

[1664] Magnús M. Halldórsson and Takeshi Tokuyama. Minimizing interference of a wireless ad-hoc network in a plane. Theoretical Computer Science, 402(1): 29–42, July 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bilo:2008:CMI

[1665] Davide Bilò and Guido Proietti. On the complexity of minimizing interference in ad-hoc and sensor networks. Theoretical Computer Science, 402(1): 43–55, July 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nikoletseas:2008:ESN

[1666] Sotiris Nikoletseas and Paul Spirakis. Efficient sensor network design for continuous monitoring of moving objects. Theoretical Computer Science, 402(1): 56–66, July 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flocchini:2008:SDM

[1667] Paola Flocchini, Giuseppe Prencipe, and Nicola Santoro. Self-deployment of mobile sensors on a ring. Theoretical Computer Science, 402(1):67– 80, July 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:Pd

[1668] Anonymous. Preface. Theoretical Computer Science, 402(1):v–ix, July 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Montanari:2008:P

[1669] Ugo Montanari and Donald Sannella.

Preface. Theoretical Computer Science, 402(2–3):81, August 8, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Askarov:2008:CMF

[1670] Aslan Askarov, Daniel Hedin, and Andrei Sabelfeld. Cryptographicallymasked flows. Theoretical Computer Science, 402(2–3):82–101, August 8, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bruni:2008:PSM

[1671] Roberto Bruni and Ivan Lanese. Parametric synchronizations in mobile nominal calculi. Theoretical Computer Science, 402(2–3):102–119, August 8, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caires:2008:SBT

[1672] Luís Caires. Spatial-behavioral types for concurrency and resource control in distributed systems. Theoretical Computer Science, 402(2–3):120– 141, August 8, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caragiannis:2008:SMP

[1673] Ioannis Caragiannis, Christos Kaklamanis, Panagiotis Kanellopoulos, and Evi Papaioannou. Scheduling to maximize participation. Theoretical Computer Science, 402(2–3):142–155, August 8, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dezani-Ciancaglini:2008:STD

[1674] Mariangiola Dezani-Ciancaglini, Silvia Ghilezan, Jovanka Pantović, and Daniele Varacca. Security types for dynamic web data. Theoretical Computer Science, 402(2–3):156–171, August 8, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hahnle:2008:IST

[1675] Reiner Hähnle, Jing Pan, Philipp Rümmer, and Dennis Walter. Integration of a security type system into a program logic. Theoretical Computer Science, 402(2–3):172–189, August 8, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hanusse:2008:MSA

[1676] Nicolas Hanusse, Dimitris Kavvadias, Evangelos Kranakis, and Danny Krizanc. Memoryless search algorithms in a network with faulty advice. Theoretical Computer Science, 402(2–3):190– 198, August 8, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pous:2008:UBP

[1677] Damien Pous. Using bisimulation proof techniques for the analysis of distributed abstract machines. Theoretical Computer Science, 402(2–3):199– 220, August 8, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Silvestri:2008:LCO

[1678] Francesco Silvestri. On the limits of cache-oblivious rational permutations. Theoretical Computer Science, 402(2– 3):221–233, August 8, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wehrman:2008:TSO

[1679] Ian Wehrman, David Kitchin, William R.

Cook, and Jayadev Misra. A timed semantics of Orc. Theoretical Computer Science, 402(2–3):234–248, August 8, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2008:Pb

[1680] G. Rozenberg. Preface. Theoretical Computer Science, 403(1):1–2, August 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2008:GDN

[1681] Zhiguo Yang and Daoyi Xu. Global dynamics for non-autonomous reactiondiffusion neural networks with timevarying delays. Theoretical Computer Science, 403(1):3–10, August 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Timmis:2008:TAA

[1682] J. Timmis, A. Hone, T. Stibor, and E. Clark. Theoretical advances in artificial immune systems. Theoretical Computer Science, 403(1):11– 32, August 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Borisovsky:2008:CEA

[1683] P. A. Borisovsky and A. V. Eremeev. Comparing evolutionary algorithms to the (1 + 1)-EA. Theoretical Computer Science, 403(1):33– 41, August 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2008:DEO

[1684] Lvzhou Li and Daowen Qiu. Determining the equivalence for oneway quantum finite automata. Theoretical Computer Science, 403(1):42– 51, August 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fages:2008:AIT

[1685] François Fages and Sylvain Soliman. Abstract interpretation and types for systems biology. Theoretical Computer Science, 403(1):52–70, August 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Toffoli:2008:WHC

[1686] Tommaso Toffoli, Silvio Capobianco, and Patrizia Mentrasti. When–and how–can a cellular automaton be rewritten as a lattice gas? Theoretical Computer Science, 403(1):71– 88, August 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rao:2008:IA

[1687] M. V. Panduranga Rao. Interference automata. Theoretical Computer Science, 403(1):89–103, August 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Witt:2008:PSV

[1688] Carsten Witt. Population size versus runtime of a simple evolutionary algorithm. Theoretical Computer Science, 403(1):104–120, August 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fouche:2008:UPQ

[1689] Willem Fouché, Johannes Heidema, Glyn Jones, and Petrus H. Potgieter. Universality and programmability of quantum computers. Theoretical Computer Science, 403(1):121– 129, August 20, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khan:2008:ESM

[1690] Faisal Shah Khan and Marek Perkowski.

Erratum to: “Synthesis of multi-qudit hybrid and d-valued quantum logic circuits by decomposition” [Theoret. Comput. Sci. 367 (3) (2006) 336–346]. Theoretical Computer Science, 403(1): 130–131, August 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [912].

Anonymous:2008:Pe

[1691] Anonymous. Preface. Theoretical Computer Science, 403(1):v–ix, August 20, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dufourd:2008:PGT

[1692] Jean-François Dufourd. Polyhedra genus theorem and Euler formula: a hypermap-formalized intuitionistic proof. Theoretical Computer Science, 403(2–3):133–159, August 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baaz:2008:CAF

[1693] Matthias Baaz, Stefan Hetzl, Alexander Leitsch, Clemens Richter, and Hendrik Spohr. CERES: an analysis of Fürstenberg’s proof of the infinity of primes. Theoretical Computer Science, 403(2–3):160–175, August 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ipate:2008:TDP

[1694] Florentin Ipate and Mike Holcombe. Testing data processing-oriented systems from stream X-machine models. Theoretical Computer Science, 403(2– 3):176–191, August 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2008:SCA

[1695] Sanjiang Li and Mingsheng Ying. Soft constraint abstraction based on semiring homomorphism. Theoretical Computer Science, 403(2–3):192–201, August 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berard:2008:WTA

[1696] B. Bérard, F. Cassez, S. Haddad, D. Lime, and O. H. Roux. When are Timed Automata weakly timed bisimilar to Time Petri Nets? Theoretical Computer Science, 403(2–3):202– 220, August 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stuber:2008:WMD

[1697] Torsten Stüber and Heiko Vogler.

Weighted monadic datalog. Theoretical Computer Science, 403(2–3):221– 238, August 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meseguer:2008:EA

[1698] José Meseguer, Miguel Palomino, and Narciso Martí-Oliet. Equational abstractions. Theoretical Computer Science, 403(2–3):239–264, August 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Olteanu:2008:WSD

[1699] Dan Olteanu, Christoph Koch, and Lyublena Antova. World-set decompositions: Expressiveness and efficient algorithms. Theoretical Computer Science, 403(2–3):265–284, August 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calves:2008:PNU

[1700] Christophe Calvès and Maribel Fernández.

A polynomial nominal unification algorithm. Theoretical Computer Science, 403(2–3):285–306, August 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Payet:2008:LDT

[1701] Étienne Payet. Loop detection in term rewriting using the eliminating unfoldings. Theoretical Computer Science, 403(2–3):307–327, August 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ciabattoni:2008:DE

[1702] Agata Ciabattoni and George Metcalfe. Density elimination. Theoretical Computer Science, 403(2–3):328– 346, August 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrhard:2008:UTE

[1703] Thomas Ehrhard and Laurent Regnier. Uniformity and the Taylor expansion of ordinary lambda-terms. Theoretical Computer Science, 403(2–3):347– 372, August 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rabinovich:2008:AHT

[1704] Alexander Rabinovich. Arity hierarchy for temporal logics. Theoretical Computer Science, 403(2–3):373– 381, August 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bozzelli:2008:VWF

[1705] Laura Bozzelli, Salvatore La Torre, and Adriano Peron. Verification of well-formed communicating recursive state machines. Theoretical Computer Science, 403(2–3):382–405, August 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Plummer:2008:SEM

[1706] Andrew R. Plummer. S4 enriched multimodal categorial grammars are context-free: Corrigendum. Theoretical Computer Science, 403(2–3):406– 408, August 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Milius:2008:CCT

[1707] Stefan Milius and Lawrence S. Moss.

Corrigendum to: “The category theoretic solution of recursive program schemes” [Theoret. Comput. Sci. 366 (2006) 3–59]. Theoretical Computer Science, 403(2–3):409–415, August 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [879].

Rozenberg:2008:Pc

[1708] G. Rozenberg. Preface. Theoretical Computer Science, 404(1–2):1–2, September 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Busi:2008:F

[1709] Nadia Busi and Claudio Zandron. Foreword. Theoretical Computer Science, 404(1–2):3–4, September 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cardelli:2008:BMS

[1710] Luca Cardelli. Bitonal membrane systems: Interactions of biological membranes. Theoretical Computer Science, 404(1–2):5–18, September 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paun:2008:MCB

[1711] Gheorghe Păun. Membrane computing and brane calculi. Old, new, and future bridges. Theoretical Computer Science, 404(1–2):19–25, September 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brijder:2008:MSP

[1712] Robert Brijder, Matteo Cavaliere, Agustín Riscos-Núñez, Grzegorz Rozenberg, and Dragoş Sburlan. Membrane systems with proteins embedded in membranes. Theoretical Computer Science, 404(1–2):26–39, September 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cavaliere:2008:DPM

[1713] Matteo Cavaliere and Sean Sedwards. Decision problems in membrane systems with peripheral proteins, transport and evolution. Theoretical Computer Science, 404(1–2):40–51, September 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Csuhaj-Varju:2008:MBA

[1714] Erzsébet Csuhaj-Varjú and György Vaszil. (Mem)brane automata. Theoretical Computer Science, 404(1–2):52– 60, September 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Curien:2008:CSA

[1715] Pierre-Louis Curien, Vincent Danos, Jean Krivine, and Min Zhang. Computational self-assembly. Theoretical Computer Science, 404(1–2):61–75, September 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaz-Pernil:2008:UFT

[1716] Daniel Díaz-Pernil, Miguel A. Gutiérrez-Naranjo, Mario J. Pérez-Jiménez, and Agustín Riscos-Núñez. A uniform family of tissue P systems with cell division solving 3-COL in a linear time. Theoretical Computer Science, 404(1–2):76– 87, September 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Franco:2008:DCI

[1717] Giuditta Franco and Maurice Margenstern. A DNA computing inspired computational model. Theoretical Computer Science, 404(1–2):88– 96, September 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Geisweiller:2008:RCD

[1718] Nil Geisweiller, Jane Hillston, and Marco Stenico. Relating continuous and discrete PEPA models of signalling pathways. Theoretical Computer Science, 404(1–2):97–111, September 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kleijn:2008:PMS

[1719] Jetty Kleijn and Maciej Koutny. Processes of membrane systems with promoters and inhibitors. Theoretical Computer Science, 404(1–2):112–126, September 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laneve:2008:SCP

[1720] Cosimo Laneve and Fabien Tarissan.

A simple calculus for proteins and cells. Theoretical Computer Science, 404(1–2):127–141, September 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manca:2008:MAP

[1721] Vincenzo Manca. The metabolic algorithm for P systems: Principles and applications. Theoretical Computer Science, 404(1–2):142–155, September 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Romanel:2008:DCS

[1722] A. Romanel and C. Priami. On the decidability and complexity of the structural congruence for beta-binders. Theoretical Computer Science, 404(1–2): 156–169, September 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Verlan:2008:GCP

[1723] Sergey Verlan, Francesco Bernardini, Marian Gheorghe, and Maurice Margenstern. Generalized communicating P systems. Theoretical Computer Science, 404(1–2):170–184, September 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:Pf

[1724] Anonymous. Preface. Theoretical Computer Science, 404(1–2):v–ix, September 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Margaria:2008:P

[1725] Tiziana Margaria and Bernhard Steffen. Preface. Theoretical Computer Science, 404(3):185, September 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hussain:2008:MCM

[1726] Altaf Hussain and Michael Huth. On model checking multiple hybrid views. Theoretical Computer Science, 404(3): 186–201, September 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ben-David:2008:EFA

[1727] Shoham Ben-David, Dana Fisman, and Sitvanit Ruah. Embedding finite automata within regular expressions. Theoretical Computer Science, 404(3):202–218, September 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Copty:2008:PAR

[1728] Shady Copty, Shai Fine, Shmuel Ur, Elad Yom-Tov, and Avi Ziv. A probabilistic alternative to regression suites. Theoretical Computer Science, 404(3): 219–234, September 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calder:2008:AAT

[1729] M. Calder and A. Miller. An automatic abstraction technique for verifying featured, parameterised systems. Theoretical Computer Science, 404(3):235– 255, September 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ivancic:2008:ESB

[1730] Franjo Ivančić, Zijiang Yang, Malay K.

Ganai, Aarti Gupta, and Pranav Ashar. Efficient SAT-based bounded model checking for software verification. Theoretical Computer Science, 404(3):256–274, September 28, 2008.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bensalem:2008:AGP

[1731] Saddek Bensalem, Doron Peled, Hongyang Qu, and Stavros Tripakis. Automatic generation of path conditions for concurrent timed systems. Theoretical Computer Science, 404(3): 275–292, September 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schonherr:2008:SFV

[1732] Jens Schönherr, Martin Freibothe, Bernd Straube, and Jörg Bormann. Semi-formal verification of the steady state behavior of mixed-signal circuits by SAT-based property checking. Theoretical Computer Science, 404(3):293– 307, September 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kopperman:2008:F

[1733] Ralph Kopperman, Prakash Panangaden, Michael B. Smyth, and Dieter Spreen. Foreword. Theoretical Computer Science, 405(1–2):1–2, October 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bridges:2008:ACN

[1734] Douglas Bridges, Hajime Ishihara, Peter M. Schuster, and Luminiţa Vîţă. Apartness, compactness and nearness. Theoretical Computer Science, 405(1– 2):3–10, October 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ciraulo:2008:FFT

[1735] Francesco Ciraulo and Giovanni Sambin. Finitary formal topologies and Stone’s representation theorem. Theoretical Computer Science, 405(1–2):11– 23, October 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hardie:2008:FRS

[1736] K. A. Hardie and P. J. Witbooi. Finite relational structure models of topological spaces and maps. Theoretical Computer Science, 405(1–2):24–34, October 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huang:2008:CET

[1737] Fangping Huang and Jihua Liang. On computational environments of topological spaces. Theoretical Computer Science, 405(1–2):35–40, October 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jordan:2008:MTT

[1738] K. E. Jordan, Lance E. Miller, E. L. F.

Moore, T. J. Peters, and Alexander Russell. Modeling time and topology for animation and visualization with examples on parametric geometry. Theoretical Computer Science, 405(1– 2):41–49, October 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kreinovich:2008:CCD

[1739] Vladik Kreinovich and Olga Kosheleva. Computational complexity of determining which statements about causality hold in different space–time models. Theoretical Computer Science, 405(1–2):50–63, October 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kunzi:2008:SST

[1740] Hans-Peter A. Künzi and Vladik Kreinovich. Static space–times naturally lead to quasi-pseudometrics. Theoretical Computer Science, 405(1–2): 64–72, October 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lawson:2008:MSD

[1741] Jimmie D. Lawson. Metric spaces and F S-domains. Theoretical Computer Science, 405(1–2):73–74, October 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martin:2008:TIT

[1742] Keye Martin. Topology in information theory in topology. Theoretical Computer Science, 405(1–2):75–87, October 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Porter:2008:ECM

[1743] Timothy Porter. Enriched categories and models for spaces of evolving states. Theoretical Computer Science, 405(1–2):88–100, October 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schuster:2008:ZSF

[1744] Peter Schuster. The Zariski spectrum as a formal geometry. Theoretical Computer Science, 405(1–2):101– 115, October 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Selivanov:2008:FHR

[1745] Victor L. Selivanov. Fine hierarchies and m-reducibilities in theoretical computer science. Theoretical Computer Science, 405(1–2):116–163, October 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Slapal:2008:QUD

[1746] Josef Šlapal. A quotient-universal digital topology. Theoretical Computer Science, 405(1–2):164–175, October 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Spreen:2008:ISR

[1747] Dieter Spreen, Luoshan Xu, and Xuxin Mao. Information systems revisited — the general continuous case. Theoretical Computer Science, 405(1–2):176– 187, October 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Surya:2008:CST

[1748] Sumati Surya. Causal set topology.

Theoretical Computer Science, 405(1– 2):188–197, October 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tsuiki:2008:LTS

[1749] Hideki Tsuiki and Yasunao Hattori.

Lawson topology of the space of formal balls and the hyperbolic topology. Theoretical Computer Science, 405(1– 2):198–205, October 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:F

[1750] Anonymous. Foreword. Theoretical Computer Science, 405(1–2):v– ix, October 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Long:2008:P

[1751] Philip M. Long and Frank Stephan. Preface. Theoretical Computer Science, 405(3):207–208, October 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Atici:2008:LUD

[1752] Alp Atıcı and Rocco A. Servedio. Learning unions of ω(1)-dimensional rectangles. Theoretical Computer Science, 405(3):209–222, October 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kontorovich:2008:KML

[1753] Leonid (Aryeh) Kontorovich, Corinna Cortes, and Mehryar Mohri. Kernel methods for learning languages. Theoretical Computer Science, 405(3):223– 236, October 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maurer:2008:USS

[1754] Andreas Maurer. Unsupervised slow subspace-learning from stationary processes. Theoretical Computer Science, 405(3):237–255, October 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Poland:2008:CDB

[1755] Jan Poland. Consistency of discrete Bayesian learning. Theoretical Computer Science, 405(3):256–273, October 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ryabko:2008:PLR

[1756] Daniil Ryabko and Marcus Hutter. On the possibility of learning in reactive environments with arbitrary dependence. Theoretical Computer Science, 405(3):274–284, October 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vovk:2008:LSC

[1757] Vladimir Vovk. Leading strategies in competitive on-line prediction. Theoretical Computer Science, 405(3):285– 296, October 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Batenburg:2008:P

[1758] Kees Joost Batenburg, Antal Nagy, and Maurice Nivat. Preface. Theoretical Computer Science, 406(1–2): 1, October 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Batenburg:2008:MAK

[1759] Kees Joost Batenburg, Antal Nagy, and Maurice Nivat. In memoriam: Attila Kuba (1953–2006). Theoretical Computer Science, 406(1–2):2–7, October 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Andres:2008:SFD

[1760] Eric Andres. The supercover of an mflat is a discrete analytical object. Theoretical Computer Science, 406(1–2):8– 14, October 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balazs:2008:FGS

[1761] Péter Balázs. A framework for generating some discrete sets with disjoint components by using uniform distributions. Theoretical Computer Science, 406(1–2):15–23, October 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brimkov:2008:PCI

[1762] Valentin E. Brimkov and Reneta P. Barneva. On the polyhedral complexity of the integer points in a hyperball. Theoretical Computer Science, 406(1– 2):24–30, October 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brlek:2008:DSM

[1763] S. Brlek, G. Labelle, and A. Lacasse. Discrete sets with minimal moment of inertia. Theoretical Computer Science, 406(1–2):31–42, October 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brocchi:2008:RBM

[1764] S. Brocchi, A. Frosini, and C. Picouleau. Reconstruction of binary matrices under fixed size neighborhood constraints. Theoretical Computer Science, 406(1–2):43–54, October 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brunetti:2008:RCL

[1765] Sara Brunetti and Alain Daurat. Reconstruction of convex lattice sets from tomographic projections in quartic time. Theoretical Computer Science, 406(1–2):55–62, October 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brunetti:2008:RBP

[1766] S. Brunetti, A. Del Lungo, P. Gritzmann, and S. de Vries. On the reconstruction of binary and permutation matrices under (binary) tomographic constraints. Theoretical Computer Science, 406(1–2):63–71, October 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Coeurjolly:2008:FMM

[1767] David Coeurjolly, Jérôme Hulin, and Isabelle Sivignon. Finding a minimum medial axis of a discrete shape is NP-hard. Theoretical Computer Science, 406(1–2):72–79, October 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dulio:2008:CDP

[1768] Paolo Dulio. Convex decomposition of U-polygons. Theoretical Computer Science, 406(1–2):80–89, October 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frosini:2008:SIM

[1769] Andrea Frosini, Maurice Nivat, and Simone Rinaldi. Scanning integer matrices by means of two rectangular windows. Theoretical Computer Science, 406(1–2):90–96, October 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kong:2008:MND

[1770] T. Yung Kong. Minimal non-deletable sets and minimal non-codeletable sets in binary images. Theoretical Computer Science, 406(1–2):97–118, October 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Palagyi:2008:FPS

[1771] Kálmán Palágyi. A 3D fully parallel surface-thinning algorithm. Theoretical Computer Science, 406(1–2):119– 135, October 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Santoyo:2008:AOI

[1772] Miguel Santoyo and Ernesto Vallejo. Additivity obstructions for integral matrices and pyramids. Theoretical Computer Science, 406(1–2):136–145, October 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vainsencher:2008:IOP

[1773] Daniel Vainsencher and Alfred M. Bruckstein. On isoperimetrically optimal polyforms. Theoretical Computer Science, 406(1–2):146–159, October 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zdunek:2008:IRA

[1774] Rafal Zdunek. On image reconstruction algorithms for binary electromagnetic geotomography. Theoretical Computer Science, 406(1–2):160–170, October 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:Pg

[1775] Anonymous. Preface. Theoretical Computer Science, 406(1–2):v–ix, October 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kaklamanis:2008:P

[1776] Christos Kaklamanis. Preface. Theoretical Computer Science, 406(3):171– 172, October 31, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ambrosio:2008:DMA

[1777] Pasquale Ambrosio and Vincenzo Auletta. Deterministic monotone algorithms for scheduling on related machines. Theoretical Computer Science, 406(3):173–186, October 31, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lucking:2008:NMS

[1778] Thomas Lücking, Marios Mavronicolas, Burkhard Monien, and Manuel Rode. A new model for selfish routing. Theoretical Computer Science, 406(3): 187–206, October 31, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chlebik:2008:STP

[1779] Miroslav Chlebík and Janka Chlebíková.

The Steiner tree problem on graphs: Inapproximability results. Theoretical Computer Science, 406(3):207– 214, October 31, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nikoletseas:2008:LIS

[1780] S. Nikoletseas, C. Raptopoulos, and P. Spirakis. Large independent sets in general random intersection graphs. Theoretical Computer Science, 406(3): 215–224, October 31, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klasing:2008:CBA

[1781] Ralf Klasing, Nelson Morales, and Stéphane Pérennes. On the complexity of bandwidth allocation in radio networks. Theoretical Computer Science, 406(3):225–239, October 31, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alvarez:2008:HLC

[1782] Carme Àlvarez, Josep Díaz, Jordi Petit, José Rolim, and Maria Serna. High level communication functionalities for wireless sensor networks. Theoretical Computer Science, 406(3):240– 247, October 31, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cooper:2008:RAJ

[1783] Colin Cooper, Ralf Klasing, and Tomasz Radzik. A randomized algorithm for the joining protocol in dynamic distributed networks. Theoretical Computer Science, 406(3):248– 262, October 31, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Eckhardt:2008:CNA

[1784] Stefan Eckhardt, Sven Kosub, Moritz G. Maaß, Hanjo Täubig, and Sebastian Wernicke. Combinatorial network abstraction by trees and distances. Theoretical Computer Science, 407(1–3):1– 20, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhu:2008:ELC

[1785] Junlei Zhu and Yuehua Bu. Equitable list colorings of planar graphs without short cycles. Theoretical Computer Science, 407(1–3):21–28, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Margenstern:2008:DPH

[1786] Maurice Margenstern. The domino problem of the hyperbolic plane is undecidable. Theoretical Computer Science, 407(1–3):29–84, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epstein:2008:OUC

[1787] Leah Epstein, Asaf Levin, and Rob van Stee. Online unit clustering: Variations on a theme. Theoretical Computer Science, 407(1–3):85–96, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fournier:2008:URP

[1788] Hervé Fournier and Guillaume Malod.

Universal relations and #P-completeness. Theoretical Computer Science, 407(1– 3):97–109, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ma:2008:FTE

[1789] Meijie Ma, Guizhen Liu, and Jun-Ming Xu. Fault-tolerant embedding of paths in crossed cubes. Theoretical Computer Science, 407(1–3):110–116, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gonidec:2008:CFI

[1790] Marion Le Gonidec. On complexity functions of infinite words associated with generalized Dyck languages. Theoretical Computer Science, 407(1–3): 117–133, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Efthymiou:2008:RSC

[1791] Charilaos Efthymiou and Paul G. Spirakis. Random sampling of colourings of sparse random graphs with a constant number of colours. Theoretical Computer Science, 407(1–3):134– 154, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bostan:2008:SSL

[1792] Alin Bostan, Claude-Pierre Jeannerod, and Éric Schost. Solving structured linear systems with large displacement rank. Theoretical Computer Science, 407(1–3):155–181, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bu:2008:AOA

[1793] Tian-Ming Bu, Xiaotie Deng, and Qi Qi. Arbitrage opportunities across sponsored search markets. Theoretical Computer Science, 407(1–3):182– 191, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garcia:2008:UAN

[1794] Pedro García, Manuel Vázquez de Parga, Gloria I. Álvarez, and José Ruiz. Universal automata and NFA learning. Theoretical Computer Science, 407(1–3):192–202, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epstein:2008:OIC

[1795] Leah Epstein and Meital Levy. Online interval coloring with packing constraints. Theoretical Computer Science, 407(1–3):203–212, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeSantis:2008:NCP

[1796] Alfredo De Santis, Anna Lisa Ferrara, and Barbara Masucci. New constructions for provably-secure time-bound hierarchical key assignment schemes. Theoretical Computer Science, 407(1– 3):213–230, November 6, 2008. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krumke:2008:B

[1797] Sven O. Krumke, Willem E. de Paepe, Jörg Rambau, and Leen Stougie. Bincoloring. Theoretical Computer Science, 407(1–3):231–241, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beveridge:2008:GCI

[1798] Andrew Beveridge, Tom Bohman, Alan Frieze, and Oleg Pikhurko. Game chromatic index of graphs with given restrictions on degrees. Theoretical Computer Science, 407(1–3):242–249, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salomaa:2008:SHA

[1799] Arto Salomaa. Subword histories and associated matrices. Theoretical Computer Science, 407(1–3):250– 257, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lee:2008:ARA

[1800] Jonathan K. Lee, Jens Palsberg, and Fernando Magno Quintão Pereira. Aliased register allocation for straightline programs is NP-complete. Theoretical Computer Science, 407(1–3):258– 273, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chiarandini:2008:HAH

[1801] M. Chiarandini, I. S. Kotsireas, C. Koukouvinos, and L. Paquete. Heuristic algorithms for Hadamard matrices with two circulant cores. Theoretical Computer Science, 407(1–3):274– 277, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klima:2008:VMA

[1802] Ondřej Klíma and Libor Polák. On varieties of meet automata. Theoretical Computer Science, 407(1–3):278– 289, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Prieur:2008:LCC

[1803] Élise Prieur and Thierry Lecroq. Online construction of compact suffix vectors and maximal repeats. Theoretical Computer Science, 407(1–3):290– 301, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Buchsbaum:2008:NRF

[1804] A. L. Buchsbaum, R. Giancarlo, and B. Racz. New results for finding common neighborhoods in massive graphs in the data stream model. Theoretical Computer Science, 407(1–3):302– 309, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2008:DDS

[1805] Jiang Chen and Ke Yi. A dynamic data structure for top-k queries on uncertain data. Theoretical Computer Science, 407(1–3):310–317, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fu:2008:FFH

[1806] Jung-Sheng Fu. Fault-free Hamiltonian cycles in twisted cubes with conditional link faults. Theoretical Computer Science, 407(1–3):318–329, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kozik:2008:FSF

[1807] Marcin Kozik. A finite set of functions with an EXPTIME-complete composition problem. Theoretical Computer Science, 407(1–3):330–341, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sanches:2008:PTS

[1808] C. A. A. Sanches, N. Y. Soma, and H. H. Yanasse. Parallel time and space upper-bounds for the subset-sum problem. Theoretical Computer Science, 407(1–3):342–348, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2008:AFW

[1809] Hsiao-Fei Liu and Kun-Mao Chao. Algorithms for finding the weightconstrained k-longest paths in a tree and the length-constrained k maximum-sum segments of a sequence. Theoretical Computer Science, 407(1– 3):349–358, November 6, 2008. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maneva:2008:STC

[1810] Elitza Maneva and Alistair Sinclair. On the satisfiability threshold and clustering of solutions of random 3-SAT formulas. Theoretical Computer Science, 407(1–3):359–369, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2008:NFD

[1811] Guomin Yang, Jing Chen, Duncan S. Wong, Xiaotie Deng, and Dongsheng Wang. A new framework for the design and analysis of identity-based identification schemes. Theoretical Computer Science, 407(1–3):370–388, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2008:DOS

[1812] Rongjun Chen, Wanzhen Huang, and Guochun Tang. Dense open-shop schedules with release times. Theoretical Computer Science, 407(1–3):389– 399, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Benedetto:2008:TPA

[1813] Katherine P. Benedetto and Nicholas A.

Loehr. Tiling problems, automata, and tiling graphs. Theoretical Computer Science, 407(1–3):400–411, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flocchini:2008:APF

[1814] Paola Flocchini, Giuseppe Prencipe, Nicola Santoro, and Peter Widmayer. Arbitrary pattern formation by asynchronous, anonymous, oblivious robots. Theoretical Computer Science, 407(1– 3):412–447, November 6, 2008. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Buning:2008:CCQ

[1815] Hans Kleine Büning and Xishun Zhao.

Computational complexity of quantified Boolean formulas with fixed maximal deficiency. Theoretical Computer Science, 407(1–3):448–457, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Latapy:2008:MMT

[1816] Matthieu Latapy. Main-memory triangle computations for very large (sparse (power-law)) graphs. Theoretical Computer Science, 407(1–3):458– 473, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Uchizawa:2008:ELB

[1817] Kei Uchizawa and Eiji Takimoto. Exponential lower bounds on the size of constant-depth threshold circuits with small energy complexity. Theoretical Computer Science, 407(1–3):474– 487, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balogh:2008:BDH

[1818] J. Balogh, S. L. Bezrukov, L. H. Harper, and A. Seress. On the bandwidth of 3-dimensional Hamming graphs. Theoretical Computer Science, 407(1–3):488–495, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhao:2008:TSC

[1819] Qianchuan Zhao, Jianfeng Mao, and Tao Ye. Time separations of cyclic event rule systems with min–max timing constraints. Theoretical Computer Science, 407(1–3):496–510, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huang:2008:LBN

[1820] Scott C.-H. Huang, Frances Yao, Minming Li, and Weili Wu. Lower bounds and new constructions on secure group communication schemes. Theoretical Computer Science, 407(1–3):511–523, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2008:RRP

[1821] Hung-Lung Wang and Kun-Mao Chao. The 2-radius and 2-radiian problems on trees. Theoretical Computer Science, 407(1–3):524–531, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2008:MDS

[1822] Boting Yang and Yi Cao. Monotonicity in digraph search problems. Theoretical Computer Science, 407(1–3):532– 544, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lin:2008:TWD

[1823] Hong-Yiu Lin, Yuh-Dauh Lyuu, Tak-Man Ma, and Yen-Wu Ti. Testing whether a digraph contains Hfree k-induced subgraphs. Theoretical Computer Science, 407(1–3):545– 553, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hong:2008:ELR

[1824] Jin-Ju Hong and Gen-Huey Chen. Efficient on-line repetition detection. Theoretical Computer Science, 407(1–3): 554–563, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cazals:2008:NPR

[1825] F. Cazals and C. Karande. A note on the problem of reporting maximal cliques. Theoretical Computer Science, 407(1–3):564–568, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deLuca:2008:RST

[1826] Aldo de Luca, Amy Glen, and Luca Q.

Zamboni. Rich, Sturmian, and trapezoidal words. Theoretical Computer Science, 407(1–3):569–573, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Borowiecki:2008:MCG

[1827] Mieczysław Borowiecki and Katarzyna Jesse-Józefczyk. Matching cutsets in graphs of diameter 2. Theoretical Computer Science, 407(1–3):574–582, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Du:2008:CDB

[1828] Ye Du. On the complexity of deciding bimatrix games similarity. Theoretical Computer Science, 407(1–3):583– 586, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Porat:2008:PMP

[1829] Benny Porat, Ely Porat, and Asaf Zur. Pattern matching with pair correlation distance. Theoretical Computer Science, 407(1–3):587–590, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Horvath:2008:ACF

[1830] Gábor Horváth, Chrystopher L. Nehaniv, and Csaba Szabó. An assertion concerning functionally complete algebras and NP-completeness. Theoretical Computer Science, 407(1–3):591– 595, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lehtonen:2008:TUV

[1831] Eero Lehtonen. Two undecidable variants of Collatz’s problems. Theoretical Computer Science, 407(1–3):596– 600, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:EBi

[1832] Anonymous. Editorial board. Theoretical Computer Science, 407(1–3):v– ix, November 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2008:Pd

[1833] G. Rozenberg. Preface. Theoretical Computer Science, 408(1):1–2, November 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calder:2008:P

[1834] Muffy Calder and Stephen Gilmore.

Preface. Theoretical Computer Science, 408(1):3, November 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Girolami:2008:BID

[1835] Mark Girolami. Bayesian inference for differential equations. Theoretical Computer Science, 408(1):4–16, November 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Credi:2008:NCM

[1836] A. Credi, M. Garavelli, C. Laneve, S. Pradalier, S. Silvi, and G. Zavattaro. nanoκ: a calculus for the modeling and simulation of nano devices. Theoretical Computer Science, 408(1):17– 30, November 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khanin:2008:CME

[1837] Raya Khanin and Desmond J.

Higham. Chemical Master Equation and Langevin regimes for a gene transcription model. Theoretical Computer Science, 408(1):31–40, November 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ciocchetta:2008:MCT

[1838] Federica Ciocchetta, Jane Hillston, Martin Kos, and David Tollervey. Modelling co-transcriptional cleavage in the synthesis of yeast pre-rRNA. Theoretical Computer Science, 408(1):41– 54, November 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fages:2008:TLC

[1839] François Fages and Aurélien Rizk. On temporal logic constraint solving for analyzing numerical data time series. Theoretical Computer Science, 408(1): 55–65, November 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bracciali:2008:SSC

[1840] Andrea Bracciali, Marcello Brunelli, Enrico Cataldo, and Pierpaolo Degano. Synapses as stochastic concurrent systems. Theoretical Computer Science, 408(1):66–82, November 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dematte:2008:EBP

[1841] L. Dematté, C. Priami, A. Romanel, and O. Soyer. Evolving BlenX programs to simulate the evolution of biological networks. Theoretical Computer Science, 408(1):83–96, November 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:Ph

[1842] Anonymous. Preface. Theoretical Computer Science, 408(1):v–ix, November 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amato:2008:P

[1843] Nancy M. Amato, Der-Tsai Lee, Andrea Pietracaprina, and Roberto Tamassia. Preface. Theoretical Computer Science, 408(2–3):97–98, November 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Preparata:2008:UDM

[1844] Franco P. Preparata. The unpredictable deviousness of models. Theoretical Computer Science, 408(2–3):99– 105, November 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Apostolico:2008:IDI

[1845] Alberto Apostolico and Claudia Tagliacollo. Incremental discovery of the irredundant motif bases for all suffixes of a string in O. Theoretical Computer Science, 408(2–3):106–115, November 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Allulli:2008:PLL

[1846] Luca Allulli, Giorgio Ausiello, Vincenzo Bonifaci, and Luigi Laura. On the power of lookahead in on-line server routing problems. Theoretical Computer Science, 408(2–3):116–128, November 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Badent:2008:DCG

[1847] Melanie Badent, Emilio Di Giacomo, and Giuseppe Liotta. Drawing colored graphs on colored points. Theoretical Computer Science, 408(2–3):129– 142, November 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bhatt:2008:ATT

[1848] Sandeep N. Bhatt, Gianfranco Bilardi, and Geppino Pucci. Area-time tradeoffs for universal VLSI circuits. Theoretical Computer Science, 408(2–3): 143–150, November 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bock:2008:CDM

[1849] Mary Ellen Bock, Claudio Garutti, and Concettina Guerra. Cavity detection and matching for binding site recognition. Theoretical Computer Science, 408(2–3):151–162, November 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boissonnat:2008:ADL

[1850] Jean-Daniel Boissonnat, Camille Wormser, and Mariette Yvinec. Anisotropic diagrams: Labelle Shewchuk approach revisited. Theoretical Computer Science, 408(2–3):163–173, November 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aleardi:2008:SRP

[1851] L. Castelli Aleardi, O. Devillers, and G. Schaeffer. Succinct representations of planar maps. Theoretical Computer Science, 408(2–3):174–187, November 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Codenotti:2008:CEH

[1852] Bruno Codenotti, Amin Saberi, Kasturi Varadarajan, and Yinyu Ye. The complexity of equilibria: Hardness results for economies via a correspondence with games. Theoretical Computer Science, 408(2–3):188–198, November 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goodrich:2008:PAD

[1853] Michael T. Goodrich. Pipelined algorithms to detect cheating in longterm grid computations. Theoretical Computer Science, 408(2–3):199–207, November 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Iwama:2008:RMB

[1854] Kazuo Iwama, Hiroki Morizumi, and Jun Tarui. Reductions for monotone Boolean circuits. Theoretical Computer Science, 408(2–3):208–212, November 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Katriel:2008:CUU

[1855] Irit Katriel, Claire Kenyon-Mathieu, and Eli Upfal. Commitment under uncertainty: Two-stage stochastic matching problems. Theoretical Computer Science, 408(2–3):213–223, November 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Papamanthou:2008:ACP

[1856] Charalampos Papamanthou and Ioannis G. Tollis. Algorithms for computing a parameterized st. Theoretical Computer Science, 408(2–3):224– 240, November 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rachlin:2008:NAR

[1857] Eric Rachlin and John E. Savage. Nanowire addressing with randomized-contact decoders. Theoretical Computer Science, 408(2–3):241– 261, November 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hines:2008:MS

[1858] Peter Hines. Machine semantics. Theoretical Computer Science, 409(1):1– 23, December 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demri:2008:VQC

[1859] Stéphane Demri and Régis Gascon. Verification of qualitative Z constraints. Theoretical Computer Science, 409(1):24–40, December 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Avellone:2008:OTP

[1860] Alessandro Avellone, Guido Fiorino, and Ugo Moscato. Optimization techniques for propositional intuitionistic logic and their implementation. Theoretical Computer Science, 409(1):41– 58, December 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Acciai:2008:RPC

[1861] Lucia Acciai and Michele Boreale. Responsiveness in process calculi. Theoretical Computer Science, 409(1):59– 93, December 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gambino:2008:ITW

[1862] Nicola Gambino and Richard Garner.

The identity type weak factorisation system. Theoretical Computer Science, 409(1):94–109, December 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2008:CST

[1863] Xiaojun Liu and Edward A. Lee.

CPO semantics of timed interactive actor networks. Theoretical Computer Science, 409(1):110–125, December 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reisig:2008:CKA

[1864] W. Reisig. The computable kernel of Abstract State Machines. Theoretical Computer Science, 409(1):126– 136, December 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cohen:2008:EPV

[1865] David A. Cohen, Peter G. Jeavons, and Stanislav Živný. The expressive power of valued constraints: Hierarchies and collapses. Theoretical Computer Science, 409(1):137–153, December 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2008:EBj

[1866] Anonymous. Editorial board. Theoretical Computer Science, 409(1):v– ix, December 6, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bini:2008:P

[1867] Dario Andrea Bini, Victor Y. Pan, and Jan Verschelde. Preface. Theoretical Computer Science, 409(2):155– 157, December 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bella:2008:CQP

[1868] T. Bella, Y. Eidelman, I. Gohberg, and V. Olshevsky. Computations with quasiseparable polynomials and matrices. Theoretical Computer Science, 409(2): 158–179, December 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cuyt:2008:NAS

[1869] Annie Cuyt and Wen shin Lee. A new algorithm for sparse interpolation of multivariate polynomials. Theoretical Computer Science, 409(2):180–185, December 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Emiris:2008:RAN

[1870] Ioannis Z. Emiris and Elias P. Tsigaridas. Real algebraic numbers and polynomial systems of small degree. Theoretical Computer Science, 409(2):186– 199, December 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2008:AGP

[1871] Bin Li, Jiawang Nie, and Lihong Zhi.

Approximate GCDs of polynomials and sparse SOS relaxations. Theoretical Computer Science, 409(2):200–210, December 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leibon:2008:FHT

[1872] Gregory Leibon, Daniel N. Rockmore, Wooram Park, Robert Taintor, and Gregory S. Chirikjian. A fast Hermite transform. Theoretical Computer Science, 409(2):211–228, December 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mourrain:2008:SNF

[1873] Bernard Mourrain and Philippe Trébuchet. Stable normal forms for polynomial system solving. Theoretical Computer Science, 409(2):229–240, December 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Normandin:2008:GWS

[1874] Marc Eric Normandin, Adam Vajda, and Sree Ram Valluri. Gravitational wave signal templates, pattern recognition, and reciprocal Eulerian gamma functions. Theoretical Computer Science, 409(2):241–254, December 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pan:2008:SAL

[1875] V. Y. Pan, D. Grady, B. Murphy, G. Qian, R. E. Rosholt, and A. D. Ruslanov. Schur aggregation for linear systems and determinants. Theoretical Computer Science, 409(2):255– 268, December 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Peyrl:2008:CSS

[1876] Helfried Peyrl and Pablo A. Parrilo. Computing sum of squares decompositions with rational coefficients. Theoretical Computer Science, 409(2):269– 281, December 17, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sekigawa:2008:NPZ

[1877] Hiroshi Sekigawa. The nearest polynomial with a zero in a given domain. Theoretical Computer Science, 409(2): 282–291, December 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sharma:2008:CRR

[1878] Vikram Sharma. Complexity of real root isolation using continued fractions. Theoretical Computer Science, 409(2): 292–310, December 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shirayanagi:2008:NGB

[1879] Kiyoshi Shirayanagi and Hiroshi Sekigawa. A new Gröbner basis conversion method based on stabilization techniques. Theoretical Computer Science, 409(2):311–317, December 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zeng:2008:NEM

[1880] Zhonggang Zeng. A numerical elimination method for polynomial computations. Theoretical Computer Science, 409(2):318–331, December 17, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fisher:2008:RPS

[1881] Edgar Fisher and Nándor Sieben. Rectangular polyomino set weak (1, 2)- achievement games. Theoretical Computer Science, 409(3):333–340, December 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ganchev:2008:GIS

[1882] Antoniy Ganchev, Lata Narayanan, and Sunil Shende. Games to induce specified equilibria. Theoretical Computer Science, 409(3):341–350, December 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cieliebak:2008:SVP

[1883] Mark Cieliebak, Alexander Hall, Riko Jacob, and Marc Nunkesser. Sequential vector packing. Theoretical Computer Science, 409(3):351–363, December 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abbasi:2008:CQA

[1884] Sarmad Abbasi and Numan Sheikh.

Complexity of question/answer games. Theoretical Computer Science, 409(3): 364–381, December 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baril:2008:PGL

[1885] J. L. Baril and J. M. Pallo. The pruning-grafting lattice of binary trees. Theoretical Computer Science, 409(3): 382–393, December 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elzinga:2008:ASC

[1886] Cees Elzinga, Sven Rahmann, and Hui Wang. Algorithms for subsequence combinatorics. Theoretical Computer Science, 409(3):394–404, December 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yan:2008:QIN

[1887] Weigen Yan and Fuji Zhang. A quadratic identity for the number of perfect matchings of plane graphs. Theoretical Computer Science, 409(3):405– 410, December 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Deckelbaum:2008:SOR

[1888] Alan Deckelbaum. Simulating onereversal multicounter machines by partially blind multihead finite automata. Theoretical Computer Science, 409(3): 411–416, December 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gerede:2008:MCD

[1889] Cagdas E. Gerede, Oscar H. Ibarra, Bala Ravikumar, and Jianwen Su. Minimum-cost delegation in service composition. Theoretical Computer Science, 409(3):417–431, December 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Currie:2008:LBP

[1890] James D. Currie and Terry I. Visentin. Long binary patterns are Abelian 2avoidable. Theoretical Computer Science, 409(3):432–437, December 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amir:2008:GL

[1891] Amihood Amir, Tzvika Hartman, Oren Kapah, B. Riva Shalom, and Dekel Tsur. Generalized LCS. Theoretical Computer Science, 409(3):438–449, December 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tsai:2008:EFT

[1892] Ping-Ying Tsai, Jung-Sheng Fu, and Gen-Huey Chen. Edge-fault-tolerant Hamiltonicity of pancake graphs under the conditional fault model. Theoretical Computer Science, 409(3):450– 460, December 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kassel:2008:PMF

[1893] Christian Kassel and Christophe Reutenauer. A palindromization map for the free group. Theoretical Computer Science, 409(3):461–470, December 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jukna:2008:ETR

[1894] Stasys Jukna. Expanders and timerestricted branching programs. Theoretical Computer Science, 409(3):471– 476, December 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shen:2008:DGP

[1895] Xiaojun Shen, Jianyu Lou, Weifa Liang, and Junzhou Luo. Deadline guaranteed packet scheduling for overloaded traffic in input-queued switches. Theoretical Computer Science, 409(3): 477–485, December 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bille:2008:FCR

[1896] Philip Bille and Martin Farach-Colton. Fast and compact regular expression matching. Theoretical Computer Science, 409(3):486–496, December 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abouelaoualim:2008:PTE

[1897] A. Abouelaoualim, K. Ch. Das, L. Faria, Y. Manoussakis, C. Martinhon, and R. Saad. Paths and trails in edge-colored graphs. Theoretical Computer Science, 409(3):497– 510, December 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berenbrink:2008:WBB

[1898] Petra Berenbrink, Tom Friedetzky, Zengjian Hu, and Russell Martin. On weighted balls-into-bins games. Theoretical Computer Science, 409(3):511– 520, December 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chan:2008:DPU

[1899] Joseph Wun-Tat Chan, Tak-Wah Lam, and Prudence W. H. Wong. Dynamic bin packing of unit fractions items. Theoretical Computer Science, 409(3): 521–529, December 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fuster-Sabater:2008:SAS

[1900] A. Fúster-Sabater and P. Caballero-Gil. Strategic attack on the shrinking generator. Theoretical Computer Science, 409(3):530–536, December 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Silva:2008:RSP

[1901] Pedro V. Silva. Rational subsets of partially reversible monoids. Theoretical Computer Science, 409(3):537– 548, December 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leroux:2008:SPD

[1902] Jérôme Leroux. Structural Presburger digit vector automata. Theoretical Computer Science, 409(3):549– 556, December 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jeandel:2008:PCP

[1903] Emmanuel Jeandel and Nicolas Ollinger.

Playing with Conway’s problem. Theoretical Computer Science, 409(3):557– 564, December 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asveld:2008:GAP

[1904] Peter R. J. Asveld. Generating all permutations by context-free grammars in Greibach normal form. Theoretical Computer Science, 409(3):565–577, December 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Glen:2008:QLE

[1905] Amy Glen, Florence Levé, and Gwénaël Richomme. Quasiperiodic and Lyndon episturmian words. Theoretical Computer Science, 409(3):578–600, December 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berry:2008:SPT

[1906] Anne Berry, Elias Dahlhaus, Pinar Heggernes, and Geneviève Simonet. Sequential and parallel triangulating algorithms for Elimination Game and new insights on Minimum Degree. Theoretical Computer Science, 409(3):601– 616, December 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jansen:2008:NLB

[1907] Maurice J. Jansen and Kenneth W. Regan. A nonlinear lower bound for constant depth arithmetical circuits via the discrete uncertainty principle. Theoretical Computer Science, 409(3):617– 622, December 28, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heggernes:2009:SEM

[1908] Pinar Heggernes and Charis Papadopoulos. Single-edge monotonic sequences of graphs and linear-time algorithms for minimal completions and deletions. Theoretical Computer Science, 410(1):1–15, January 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choffrut:2009:FTA

[1909] Christian Choffrut and Serge Grigorieff. Finite n-tape automata over possibly infinite alphabets: Extending a theorem of Eilenberg et al. Theoretical Computer Science, 410(1):16– 34, January 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Suchan:2009:MIC

[1910] Karol Suchan and Ioan Todinca. Minimal interval completion through graph exploration. Theoretical Computer Science, 410(1):35–43, January 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Currie:2009:CBM

[1911] James D. Currie and Ali Aberkane.

A cyclic binary morphism avoiding Abelian fourth powers. Theoretical Computer Science, 410(1):44–52, January 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fellows:2009:PCM

[1912] Michael R. Fellows, Danny Hermelin, Frances Rosamond, and Stéphane Vialette. On the parameterized complexity of multiple-interval graph problems. Theoretical Computer Science, 410(1):53–61, January 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shoaran:2009:FTC

[1913] Maryam Shoaran and Alex Thomo.

Fault-tolerant computation of distributed regular path queries. Theoretical Computer Science, 410(1):62– 77, January 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2009:MLS

[1914] Ruifang Liu, Zhonghua Lu, and Jinlong Shu. The minimal Laplacian spectral radius of trees with a given diameter. Theoretical Computer Science, 410(1): 78–83, January 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baswana:2009:APN

[1915] Surender Baswana, Vishrut Goyal, and Sandeep Sen. All-pairs nearly 2-approximate shortest paths in ∨(n^{2}polylogn) time. Theoretical Computer Science, 410(1):84–93, January 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ikeda:2009:HCT

[1916] Satoshi Ikeda, Izumi Kubo, and Masafumi Yamashita. The hitting and cover times of random walks on finite graphs using local degree information. Theoretical Computer Science, 410(1):94– 100, January 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Faliszewski:2009:CPI

[1917] Piotr Faliszewski and Lane Hemaspaandra. The complexity of powerindex comparison. Theoretical Computer Science, 410(1):101–107, January 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Masopust:2009:DCS

[1918] Tomáš Masopust. On the descriptional complexity of scattered context grammars. Theoretical Computer Science, 410(1):108–112, January 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2009:EBa

[1919] Anonymous. Editorial board. Theoretical Computer Science, 410(1):v– ix, January 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonsangue:2009:P

[1920] Marcello Bonsangue, Einar Broch Johnsen, Amy Murphy, and Jan Vitek. Preface. Theoretical Computer Science, 410(2–3):113, February 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bidinger:2009:PCR

[1921] Philippe Bidinger and Adriana Compagnoni. Pict correctness revisited. Theoretical Computer Science, 410(2– 3):114–127, February 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deBoer:2009:SVC

[1922] F. S. de Boer. A shared-variable concurrency analysis of multi-threaded object-oriented programs. Theoretical Computer Science, 410(2–3):128– 141, February 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Capecchi:2009:ASM

[1923] Sara Capecchi, Mario Coppo, Mariangiola Dezani-Ciancaglini, Sophia Drossopoulou, and Elena Giachino. Amalgamating sessions and methods in object-oriented languages with generics. Theoretical Computer Science, 410(2–3):142–167, February 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Field:2009:RDO

[1924] John Field, Maria-Cristina Marinescu, and Christian Stefansen. Reactors: a data-oriented synchronous/ asynchronous programming model for distributed applications. Theoretical Computer Science, 410(2–3):168– 201, February 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Haller:2009:SAU

[1925] Philipp Haller and Martin Odersky. Scala Actors: Unifying thread-based and event-based programming. Theoretical Computer Science, 410(2–3): 202–220, February 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jacquet:2009:FAM

[1926] Jean-Marie Jacquet and Isabelle Linden. Fully abstract models and refinements as tools to compare agents in timed coordination languages. Theoretical Computer Science, 410(2–3):221– 253, February 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Olveczky:2009:FMP

[1927] Peter Csaba Ölveczky and Stian Thorvaldsen. Formal modeling, performance estimation, and model checking of wireless sensor network algorithms in Real-Time Maude. Theoretical Computer Science, 410(2–3):254–280, February 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2009:Pa

[1928] G. Rozenberg. Preface. Theoretical Computer Science, 410(4–5):281– 282, February 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonizzoni:2009:F

[1929] Paola Bonizzoni, S. Barry Cooper, Benedikt Löwe, and Andrea Sorbi. Foreword. Theoretical Computer Science, 410(4–5):283–284, February 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zandron:2009:NB

[1930] Claudio Zandron. Nadia Busi (1968– 2007). Theoretical Computer Science, 410(4–5):285, February 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Busi:2009:CEG

[1931] Nadia Busi and Claudio Zandron.

Computational expressiveness of Genetic Systems. Theoretical Computer Science, 410(4–5):286–293, February 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Condon:2009:CPN

[1932] Anne Condon and Hosna Jabbari.

Computational prediction of nucleic acid secondary structure: Methods, applications, and challenges. Theoretical Computer Science, 410(4–5):294– 301, February 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DHondt:2009:QAG

[1933] Ellie D’Hondt. Quantum approaches to graph colouring. Theoretical Computer Science, 410(4–5):302–309, February 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:2009:ITR

[1934] A. Ehrenfeucht and G. Rozenberg. Introducing time in reaction systems. Theoretical Computer Science, 410(4– 5):310–322, February 17, 2009. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garzillo:2009:CVB

[1935] Carmine Garzillo and Giuseppe Trautteur. Computational virtuality in biological systems. Theoretical Computer Science, 410(4–5):323–331, February 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jonoska:2009:CCS

[1936] Nataša Jonoska and Gregory L. Mc-Colm. Complexity classes for selfassembling flexible tiles. Theoretical Computer Science, 410(4–5):332– 346, February 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kjos-Hanssen:2009:EDP

[1937] Bjørn Kjos-Hanssen and Anil Nerode. Effective dimension of points visited by Brownian motion. Theoretical Computer Science, 410(4–5):347–354, February 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krishna:2009:MCT

[1938] Shankara Narayanan Krishna. Membrane computing with transport and embedded proteins. Theoretical Computer Science, 410(4–5):355–375, February 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ladyman:2009:WDI

[1939] James Ladyman. What does it mean to say that a physical system implements a computation? Theoretical Computer Science, 410(4–5):376– 383, February 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lathrop:2009:SSA

[1940] James I. Lathrop, Jack H. Lutz, and Scott M. Summers. Strict self-assembly of discrete Sierpinski triangles. Theoretical Computer Science, 410(4–5): 384–405, February 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Loos:2009:SRF

[1941] Remco Loos, Florin Manea, and Victor Mitrana. On small, reduced, and fast universal accepting networks of splicing processors. Theoretical Computer Science, 410(4–5):406–416, February 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manea:2009:TCO

[1942] Florin Manea, Victor Mitrana, and Takashi Yokomori. Two complementary operations inspired by the DNA hairpin formation: Completion and reduction. Theoretical Computer Science, 410(4–5):417–425, February 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Welch:2009:CDT

[1943] P. D. Welch. Characteristics of discrete transfinite time Turing machine models: Halting times, stabilization times, and Normal Form theorems. Theoretical Computer Science, 410(4–5):426– 442, February 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Woods:2009:CSU

[1944] Damien Woods and Turlough Neary. The complexity of small universal Turing machines: a survey. Theoretical Computer Science, 410(4–5):443– 450, February 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2009:Pa

[1945] Anonymous. Preface. Theoretical Computer Science, 410(4–5):i–vi, February 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lavallee:2009:EP

[1946] Ivan Lavallée and Alexander A. Shvartsman. Editors’ preface. Theoretical Computer Science, 410(6–7):451– 452, February 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Awerbuch:2009:RRN

[1947] Baruch Awerbuch and Christian Scheideler. Robust random number generation for peer-to-peer systems. Theoretical Computer Science, 410(6–7):453– 466, February 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bazzi:2009:HCS

[1948] Rida A. Bazzi, Young ri Choi, and Mohamed G. Gouda. Hop chains: Secure routing and the establishment of distinct identities. Theoretical Computer Science, 410(6–7):467–480, February 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czyzowicz:2009:GFF

[1949] Jurek Czyzowicz, Leszek Ga̧sieniec, and Andrzej Pelc. Gathering few fat mobile robots in the plane. Theoretical Computer Science, 410(6–7):481– 499, February 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demirbas:2009:GLQ

[1950] Murat Demirbas, Anish Arora, and Vinodkrishnan Kulathumani. Glance: a lightweight querying service for wireless sensor networks. Theoretical Computer Science, 410(6–7):500–513, February 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dolev:2009:ECS

[1951] Shlomi Dolev and Nir Tzachar. Empire of colonies: Self-stabilizing and selforganizing distributed algorithm. Theoretical Computer Science, 410(6–7): 514–532, February 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Englert:2009:CUP

[1952] Burkhard Englert. On the cost of uniform protocols whose memory consumption is adaptive to interval contention. Theoretical Computer Science, 410(6–7):533–545, February 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gilbert:2009:MMS

[1953] Seth Gilbert, Rachid Guerraoui, and Calvin Newport. Of malicious motes and suspicious sensors: On the efficiency of malicious interference in wireless networks. Theoretical Computer Science, 410(6–7):546–569, February 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guerraoui:2009:TTE

[1954] Rachid Guerraoui, Maurice Herlihy, and Bastian Pochon. A topological treatment of early-deciding setagreement. Theoretical Computer Science, 410(6–7):570–580, February 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Johnen:2009:RSS

[1955] Colette Johnen and Le Huy Nguyen. Robust self-stabilizing weight-based clustering algorithm. Theoretical Computer Science, 410(6–7):581–594, February 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mavronicolas:2009:CPE

[1956] Marios Mavronicolas, Loizos Michael, and Paul Spirakis. Computing on a partially eponymous ring. Theoretical Computer Science, 410(6–7):595– 613, February 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mittal:2009:STD

[1957] Neeraj Mittal, Kuppahalli L. Phaneesh, and Felix C. Freiling. Safe termination detection in an asynchronous distributed system when processes may crash and recover. Theoretical Computer Science, 410(6–7):614– 628, February 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moser:2009:TRT

[1958] Heinrich Moser. Towards a real-time distributed computing model. Theoretical Computer Science, 410(6–7):629– 659, February 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2009:CSC

[1959] Deying Li, Hongwei Du, Peng-Jun Wan, Xiaofeng Gao, Zhao Zhang, and Weili Wu. Construction of strongly connected dominating sets in asymmetric multihop wireless networks. Theoretical Computer Science, 410(8–10): 661–669, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kiwi:2009:AQT

[1960] M. Kiwi, M. Soto, and C. Thraves. Adversarial queuing theory with setups. Theoretical Computer Science, 410(8– 10):670–687, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gao:2009:DDR

[1961] Yong Gao. The degree distribution of random k-trees. Theoretical Computer Science, 410(8–10):688–695, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Djelloul:2009:TLD

[1962] Selma Djelloul. Treewidth and logical definability of graph products. Theoretical Computer Science, 410(8–10): 696–710, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bein:2009:QTA

[1963] Wolfgang W. Bein, Lawrence L. Larmore, Linda Morales, and I. Hal Sudborough. A quadratic time 2approximation algorithm for block sorting. Theoretical Computer Science, 410(8–10):711–717, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guo:2009:MEL

[1964] Jiong Guo. A more effective linear kernelization for cluster editing. Theoretical Computer Science, 410(8–10): 718–726, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2009:ACP

[1965] Yuchen Zhang and Xiaoming Sun. The antimagicness of the Cartesian product of graphs. Theoretical Computer Science, 410(8–10):727–735, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Susilo:2009:SFS

[1966] Willy Susilo. Short fail-stop signature scheme based on factorization and discrete logarithm assumptions. Theoretical Computer Science, 410(8–10):736– 744, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kaporis:2009:POS

[1967] A. C. Kaporis and P. G. Spirakis.

The price of optimum in Stackelberg games on arbitrary single commodity networks and latency functions. Theoretical Computer Science, 410(8–10): 745–755, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dai:2009:AAM

[1968] Decheng Dai and Changyuan Yu. A 5 + ϵ-approximation algorithm for minimum weighted dominating set in unit disk. Theoretical Computer Science, 410(8–10):756–765, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tsai:2009:FFL

[1969] Ping-Ying Tsai, Jung-Sheng Fu, and Gen-Huey Chen. Fault-free longest paths in star networks with conditional link faults. Theoretical Computer Science, 410(8–10):766–775, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ng:2009:TSO

[1970] C. T. Ng, Zhiyi Tan, Yong He, and T. C. E. Cheng. Two semi-online scheduling problems on two uniform machines. Theoretical Computer Science, 410(8–10):776–792, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanchet-Sadri:2009:GTF

[1971] F. Blanchet-Sadri, Robert Mercaş, and Geoffrey Scott. A generalization of Thue freeness for partial words. Theoretical Computer Science, 410(8–10): 793–800, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kung:2009:BBH

[1972] Tzu-Liang Kung, Cheng-Kuan Lin, Tyne Liang, Lih-Hsing Hsu, and Jimmy J. M. Tan. On the bipanpositionable bipanconnectedness of hypercubes. Theoretical Computer Science, 410(8–10): 801–811, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2009:ACS

[1973] Zhao Zhang, Xiaofeng Gao, and Weili Wu. Algorithms for connected set cover problem and fault-tolerant connected set cover problem. Theoretical Computer Science, 410(8–10):812– 817, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See note [2397].

Chen:2009:PAN

[1974] Jianer Chen, Iyad A. Kanj, Jie Meng, Ge Xia, and Fenghui Zhang. On the pseudo-achromatic number problem. Theoretical Computer Science, 410(8– 10):818–829, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Qi:2009:SMP

[1975] Xianglai Qi, Shiguo Zhou, and Jinjiang Yuan. Single machine parallelbatch scheduling with deteriorating jobs. Theoretical Computer Science, 410(8–10):830–836, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ouangraoua:2009:CED

[1976] Aïda Ouangraoua and Pascal Ferraro.

A constrained edit distance algorithm between semi-ordered trees. Theoretical Computer Science, 410(8–10):837– 846, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bouvel:2009:VTD

[1977] Mathilde Bouvel and Dominique Rossin. A variant of the tandem duplication — random loss model of genome rearrangement. Theoretical Computer Science, 410(8–10):847–858, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asodi:2009:CMG

[1978] Vera Asodi and Christopher Umans.

The complexity of the matroid– greedoid partition problem. Theoretical Computer Science, 410(8–10):859– 866, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chan:2009:EAT

[1979] Chi-Yuan Chan, Shan-Chyun Ku, Chi-Jen Lu, and Biing-Feng Wang. Efficient algorithms for two generalized 2median problems and the group median problem on trees. Theoretical Computer Science, 410(8–10):867– 876, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2009:SCM

[1980] Jianping Li, Weidong Li, Tongquan Zhang, and Zhongxu Zhang. The subdivision-constrained minimum spanning tree problem. Theoretical Computer Science, 410(8–10):877–885, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferrante:2009:FPS

[1981] Alessandro Ferrante, Gennaro Parlato, Francesco Sorrentino, and Carmine Ventre. Fast payment schemes for truthful mechanisms with verification. Theoretical Computer Science, 410(8– 10):886–899, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Matsubara:2009:EAC

[1982] Wataru Matsubara, Shunsuke Inenaga, Akira Ishino, Ayumi Shinohara, Tomoyuki Nakamura, and Kazuo Hashimoto. Efficient algorithms to compute compressed longest common substrings and compressed palindromes. Theoretical Computer Science, 410(8–10):900–913, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koga:2009:DTA

[1983] Hisashi Koga. Dynamic TCP acknowledgment with sliding window. Theoretical Computer Science, 410(8–10): 914–925, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hsieh:2009:CED

[1984] Sun-Yuan Hsieh and Chang-Jen Tu.

Constructing edge-disjoint spanning trees in locally twisted cubes. Theoretical Computer Science, 410(8–10): 926–932, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kontogiannis:2009:SSS

[1985] Spyros C. Kontogiannis and Paul G.

Spirakis. On the support size of stable strategies in random games. Theoretical Computer Science, 410(8–10): 933–942, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Halava:2009:OFI

[1986] Vesa Halava, Tero Harju, Tomi Kärki, and Patrice Séébold. Overlap-freeness in infinite partial words. Theoretical Computer Science, 410(8–10):943– 948, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cardinal:2009:IAB

[1987] Jean Cardinal, Stefan Langerman, and Eythan Levy. Improved approximation bounds for edge dominating set in dense graphs. Theoretical Computer Science, 410(8–10):949–957, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gagie:2009:CDS

[1988] Travis Gagie. Compressed depth sequences. Theoretical Computer Science, 410(8–10):958–962, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hong:2009:ABC

[1989] Sung-Pil Hong, Myoung-Ju Park, and Soo Y. Chang. Approximation of the k-batch consolidation problem. Theoretical Computer Science, 410(8–10): 963–967, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanchet-Sadri:2009:TAS

[1990] F. Blanchet-Sadri, Raphaël M. Jungers, and Justin Palumbo. Testing avoidability on sets of partial words is hard. Theoretical Computer Science, 410(8–10): 968–972, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Saez:2009:QAC

[1991] Pablo Sáez. A quadratic algorithm for the 2-cyclic robotic scheduling problem. Theoretical Computer Science, 410(8–10):973–976, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Orlovich:2009:IID

[1992] Yury L. Orlovich, Valery S. Gordon, and Dominique de Werra. On the inapproximability of independent domination in 2 P. Theoretical Computer Science, 410(8–10):977–982, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pizzi:2009:K

[1993] Cinzia Pizzi. k. Theoretical Computer Science, 410(8–10):983–987, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tsai:2009:ECI

[1994] Tsung-Hsi Tsai. Efficient computation of the iteration of functions. Theoretical Computer Science, 410(8–10):988– 993, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wahlen:2009:CAH

[1995] Martin Wahlén. On the complexity of approximating the Hadwiger number. Theoretical Computer Science, 410(8– 10):994–996, March 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Honkala:2009:SIM

[1996] Juha Honkala. On the simplification of infinite morphic words. Theoretical Computer Science, 410(8–10):997– 1000, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cooper:2009:PAC

[1997] S. Barry Cooper and Hong Zhu.

Preface: Algorithms, complexity and models of computation. Theoretical Computer Science, 410(11):1001– 1002, March 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Danos:2009:HLB

[1998] Vincent Danos and Linus J. Schumacher. How liquid is biological signalling? Theoretical Computer Science, 410(11):1003–1012, March 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2009:AOT

[1999] Minming Li, Ze Feng, Nan Zang, Ronald L. Graham, and Frances F. Yao. Approximately optimal trees for group key management with batch updates. Theoretical Computer Science, 410(11):1013–1021, March 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Feng:2009:AAM

[2000] Wangsen Feng, Li’ang Zhang, and Hanpin Wang. Approximation algorithm for maximum edge coloring. Theoretical Computer Science, 410(11):1022– 1029, March 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jakoby:2009:IAD

[2001] Andreas Jakoby, Maciej Liśkiewicz, Rüdiger Reischuk, and Christian Schindelhauer. Improving the average delay of sorting. Theoretical Computer Science, 410(11):1030–1041, March 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2009:EDA

[2002] Angsheng Li. Elementary differences among jump classes. Theoretical Computer Science, 410(11):1042– 1053, March 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Morizumi:2009:LSL

[2003] Hiroki Morizumi and Jun Tarui. Linear-size log-depth negation-limited inverter for k-tonic binary sequences. Theoretical Computer Science, 410(11): 1054–1060, March 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kawaguchi:2009:DSG

[2004] Akifumi Kawaguchi and Hiroshi Nagamochi. Drawing slicing graphs with face areas. Theoretical Computer Science, 410(11):1061–1072, March 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sun:2009:TIR

[2005] He Sun and Chung Keung Poon. Two improved range-efficient algorithms for F_{0}-estimation. Theoretical Computer Science, 410(11):1073–1080, March 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhao:2009:CSA

[2006] Yingchao Zhao and Shang hua Teng.

Combinatorial and spectral aspects of nearest neighbor graphs in doubling dimensional and nearly-Euclidean spaces. Theoretical Computer Science, 410(11):1081–1092, March 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2009:AAS

[2007] Peng Zhang and Mingji Xia. An approximation algorithm to the k-Steiner Forest problem. Theoretical Computer Science, 410(11):1093–1098, March 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yao:2009:NUC

[2008] Andrew C. C. Yao, Frances F. Yao, and Yunlei Zhao. A note on universal composable zero-knowledge in the common reference string model. Theoretical Computer Science, 410(11):1099– 1108, March 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Popescu:2009:SAI

[2009] Andrei Popescu, Traian Florin Şerbănuţă, and Grigore Roşu. A semantic approach to interpolation. Theoretical Computer Science, 410(12–13):1109– 1128, March 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gumm:2009:CF

[2010] H. Peter Gumm. Copower functors. Theoretical Computer Science, 410(12–13):1129–1142, March 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bova:2009:CRL

[2011] Simone Bova and Franco Montagna. The consequence relation in the logic of commutative GBL. Theoretical Computer Science, 410(12–13):1143– 1158, March 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gabbay:2009:SSU

[2012] Murdoch J. Gabbay. A study of substitution, using nominal techniques and Fraenkel–Mostowksi sets. Theoretical Computer Science, 410(12–13):1159– 1189, March 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lorenz:2009:ESP

[2013] Robert Lorenz, Gabriel Juhás, Robin Bergenthum, Jörg Desel, and Sebastian Mauser. Executability of scenarios in Petri nets. Theoretical Computer Science, 410(12–13):1190–1216, March 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schroder:2009:HIH

[2014] Lutz Schröder and Till Mossakowski.

HasCasl: Integrated higher-order specification and program development. Theoretical Computer Science, 410(12–13):1217–1260, March 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bergstra:2009:MES

[2015] J. A. Bergstra, Y. Hirshfeld, and J. V.

Tucker. Meadows and the equational specification of division. Theoretical Computer Science, 410(12–13):1261– 1271, March 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kwiatkowska:2009:PMA

[2016] Marta Kwiatkowska, Gethin Norman, David Parker, and Maria Grazia Vigliotti. Probabilistic Mobile Ambients. Theoretical Computer Science, 410(12–13):1272–1303, March 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Prencipe:2009:P

[2017] Giuseppe Prencipe and Shmuel Zaks.

Preface. Theoretical Computer Science, 410(14):1305–1306, March 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nisse:2009:GSA

[2018] Nicolas Nisse and David Soguet. Graph searching with advice. Theoretical Computer Science, 410(14):1307– 1318, March 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Broersma:2009:UBA

[2019] Hajo Broersma, Matthew Johnson, and Daniël Paulusma. Upper bounds and algorithms for parallel knock-out numbers. Theoretical Computer Science, 410(14):1319–1327, March 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gafni:2009:ARS

[2020] Eli Gafni, Achour Mostéfaoui, Michel Raynal, and Corentin Travers. From adaptive renaming to set agreement. Theoretical Computer Science, 410(14): 1328–1335, March 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manne:2009:NSS

[2021] Fredrik Manne, Morten Mjelde, Laurence Pilard, and Sébastien Tixeuil. A new self-stabilizing maximal matching algorithm. Theoretical Computer Science, 410(14):1336–1345, March 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Korteweg:2009:DAS

[2022] Peter Korteweg, Alberto Marchetti-Spaccamela, Leen Stougie, and Andrea Vitaletti. Data aggregation in sensor networks: Balancing communication and delay costs. Theoretical Computer Science, 410(14):1346–1354, March 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Efrima:2009:DAP

[2023] Asaf Efrima and David Peleg. Distributed algorithms for partitioning a swarm of autonomous mobile robots. Theoretical Computer Science, 410(14): 1355–1368, March 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Even:2009:OCS

[2024] Guy Even, Tamir Levi, and Ami Litman. Optimal conclusive sets for comparator networks. Theoretical Computer Science, 410(14):1369– 1376, March 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kralovic:2009:RAC

[2025] Rastislav Královič and Richard Královič.

Rapid almost-complete broadcasting in faulty networks. Theoretical Computer Science, 410(14):1377–1387, March 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czyzowicz:2009:LEC

[2026] J. Czyzowicz, S. Dobrev, E. Kranakis, J. Opatrny, and J. Urrutia. Local edge colouring of Yao-like subgraphs of Unit Disk Graphs. Theoretical Computer Science, 410(14):1388–1400, March 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Korman:2009:NMG

[2027] Amos Korman and Shay Kutten. A note on models for graph representations. Theoretical Computer Science, 410(14):1401–1412, March 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2009:Pb

[2028] G. Rozenberg. Preface. Theoretical Computer Science, 410(15):1413– 1414, April 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jonoska:2009:Pa

[2029] Nataša Jonoska and Jarkko Kari. Preface. Theoretical Computer Science, 410 (15):1415–1416, April 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Greenberg:2009:CRM

[2030] Sam Greenberg and Dana Randall. Convergence rates of Markov chains for some self-assembly and non-saturated Ising models. Theoretical Computer Science, 410(15):1417–1427, April 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reif:2009:APD

[2031] John H. Reif and Sudheer Sahu. Autonomous programmable DNA nanorobotic devices using DNAzymes. Theoretical Computer Science, 410(15): 1428–1439, April 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grayson:2009:DDC

[2032] N. E. Grayson, A. Taormina, and R. Twarock. DNA duplex cage structures with icosahedral symmetry. Theoretical Computer Science, 410(15): 1440–1447, April 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jonoska:2009:ERS

[2033] Nataša Jonoska, Nadrian C. Seeman, and Gang Wu. On existence of reporter strands in DNA-based graph structures. Theoretical Computer Science, 410(15):1448–1460, April 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brun:2009:PFT

[2034] Yuriy Brun and Dustin Reishus. Path finding in the tile assembly model. Theoretical Computer Science, 410(15): 1461–1472, April 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2009:Pc

[2035] G. Rozenberg. Preface. Theoretical Computer Science, 410(16):1473– 1474, April 2, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jonoska:2009:Pb

[2036] Nataša Jonoska and Jarkko Kari. Preface. Theoretical Computer Science, 410 (16):1475–1476, April 2, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anselmo:2009:DUT

[2037] Marcella Anselmo and Maria Madonia.

Deterministic and unambiguous twodimensional languages over one-letter alphabet. Theoretical Computer Science, 410(16):1477–1485, April 2, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brijder:2009:PQR

[2038] Robert Brijder and Hendrik Jan Hoogeboom. Perfectly quilted rectangular snake tilings. Theoretical Computer Science, 410(16):1486–1494, April 2, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Becker:2009:PWT

[2039] Florent Becker. Pictures worth a thousand tiles, a geometrical programming language for self-assembly. Theoretical Computer Science, 410(16):1495– 1515, April 2, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lukkarila:2009:WDT

[2040] Ville Lukkarila. The 4-way deterministic tiling problem is undecidable. Theoretical Computer Science, 410(16): 1516–1533, April 2, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goodman-Strauss:2009:RPS

[2041] Chaim Goodman-Strauss. Regular production systems and triangle tilings. Theoretical Computer Science, 410(16): 1534–1549, April 2, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2009:Pb

[2042] Anonymous. Preface. Theoretical Computer Science, 410(16):i–vi, April 2, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Spirakis:2009:P

[2043] Paul G. Spirakis, Marios Mavronicolas, and Spyros C. Kontogiannis. Preface. Theoretical Computer Science, 410 (17):1551, April 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ackermann:2009:PNE

[2044] Heiner Ackermann, Heiko Röglin, and Berthold Vöcking. Pure Nash equilibria in player-specific and weighted congestion games. Theoretical Computer Science, 410(17):1552–1563, April 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bilo:2009:DMD

[2045] Davide Bilò, Luciano Gualà, and Guido Proietti. Dynamic mechanism design. Theoretical Computer Science, 410(17): 1564–1572, April 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2009:MEH

[2046] Xi Chen, Li-Sha Huang, and Shang-Hua Teng. Market equilibria with hybrid linear-Leontief utilities. Theoretical Computer Science, 410(17):1573– 1580, April 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daskalakis:2009:NAN

[2047] Constantinos Daskalakis, Aranyak Mehta, and Christos Papadimitriou. A note on approximate Nash equilibria. Theoretical Computer Science, 410(17): 1581–1588, April 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Immorlica:2009:CMS

[2048] Nicole Immorlica, Li (Erran) Li, Vahab S. Mirrokni, and Andreas S. Schulz. Coordination mechanisms for selfish scheduling. Theoretical Computer Science, 410(17):1589–1598, April 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kontogiannis:2009:PAA

[2049] Spyros C. Kontogiannis, Panagiota N. Panagopoulou, and Paul G. Spirakis. Polynomial algorithms for approximating Nash equilibria of bimatrix games. Theoretical Computer Science, 410(17): 1599–1606, April 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Penna:2009:SPT

[2050] Paolo Penna, Guido Proietti, and Peter Widmayer. Strongly polynomialtime truthful mechanisms in one shot. Theoretical Computer Science, 410(17): 1607–1615, April 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arge:2009:P

[2051] Lars Arge, Christian Cachin, and Andrzej Tarlecki. Preface. Theoretical Computer Science, 410(18):1617, April 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cai:2009:HAP

[2052] Jin-Yi Cai and Pinyan Lu. Holographic algorithms: The power of dimensionality resolved. Theoretical Computer Science, 410(18):1618–1628, April 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Larose:2009:UAH

[2053] Benoît Larose and Pascal Tesson. Universal algebra and hardness results for constraint satisfaction problems. Theoretical Computer Science, 410(18): 1629–1647, April 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blomer:2009:SMS

[2054] Johannes Blömer and Stefanie Naewe.

Sampling methods for shortest vectors, closest vectors and successive minima. Theoretical Computer Science, 410(18): 1648–1665, April 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Atserias:2009:ASE

[2055] Albert Atserias, Andrei Bulatov, and Anuj Dawar. Affine systems of equations and counting infinitary logic. Theoretical Computer Science, 410(18): 1666–1683, April 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bodirsky:2009:MIV

[2056] Manuel Bodirsky, Hubie Chen, Jan Kára, and Timo von Oertzen. Maximal infinite-valued constraint languages. Theoretical Computer Science, 410(18): 1684–1693, April 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boigelot:2009:GCT

[2057] Bernard Boigelot and Julien Brusten.

A generalization of Cobham’s theorem to automata over real numbers.

Theoretical Computer Science, 410(18): 1694–1703, April 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fiore:2009:CFA

[2058] Marcelo Fiore and Chung-Kil Hur. On the construction of free algebras for equational systems. Theoretical Computer Science, 410(18):1704– 1729, April 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ishai:2009:PMS

[2059] Yuval Ishai, Tal Malkin, Martin J. Strauss, and Rebecca N. Wright. Private multiparty sampling and approximation of vector combinations. Theoretical Computer Science, 410(18): 1730–1745, April 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hutter:2009:P

[2060] Marcus Hutter and Rocco A. Servedio. Preface. Theoretical Computer Science, 410(19):1747–1748, April 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maier:2009:OCN

[2061] Markus Maier, Matthias Hein, and Ulrike von Luxburg. Optimal construction of k-nearest-neighbor graphs for identifying noisy clusters. Theoretical Computer Science, 410(19):1749– 1764, April 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chang:2009:MPS

[2062] Kevin L. Chang. Multiple pass streaming algorithms for learning mixtures of distributions in R^{d}. Theoretical Computer Science, 410(19):1765– 1780, April 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vyugin:2009:CER

[2063] Vladimir V. V’yugin. On calibration error of randomized forecasting algorithms. Theoretical Computer Science, 410(19):1781–1795, April 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jain:2009:PLR

[2064] Sanjay Jain, Frank Stephan, and Nan Ye. Prescribed learning of r.e. classes. Theoretical Computer Science, 410(19): 1796–1806, April 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yoshinaka:2009:LEV

[2065] Ryo Yoshinaka. Learning efficiency of very simple grammars from positive data. Theoretical Computer Science, 410(19):1807–1825, April 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mahmud:2009:UTL

[2066] M. M. Hassan Mahmud. On universal transfer learning. Theoretical Computer Science, 410(19):1826– 1846, April 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shin:2009:PSP

[2067] Kilho Shin and Tetsuji Kuboyama. Polynomial summaries of positive semidefinite kernels. Theoretical Computer Science, 410(19):1847–1862, April 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Case:2009:PII

[2068] John Case and Samuel E. Moelius III. Parallelism increases iterative learning power. Theoretical Computer Science, 410(19):1863–1875, April 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Audibert:2009:EET

[2069] Jean-Yves Audibert, Rémi Munos, and Csaba Szepesvári. Exploration– exploitation tradeoff using variance estimates in multi-armed bandits. Theoretical Computer Science, 410(19): 1876–1902, April 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Feldman:2009:SML

[2070] Vitaly Feldman and Shrenik Shah. Separating models of learning with faulty teachers. Theoretical Computer Science, 410(19):1903–1912, April 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2009:Pd

[2071] G. Rozenberg. Preface. Theoretical Computer Science, 410(20):1913– 1914, May 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hirvensalo:2009:P

[2072] Mika Hirvensalo. Preface. Theoretical Computer Science, 410(20):1915, May 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ambainis:2009:ICQ

[2073] Andris Ambainis and Nikolajs Nahimovs. Improved constructions of quantum automata. Theoretical Computer Science, 410(20):1916–1922, May 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Freivalds:2009:ICM

[2074] Rūsiņš Freivalds, Māris Ozols, and Laura Mančinska. Improved constructions of mixed state quantum automata. Theoretical Computer Science, 410(20):1923–1931, May 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yakaryilmaz:2009:EPA

[2075] Abuzer Yakaryılmaz and A. C. Cem Say. Efficient probability amplification in two-way quantum finite automata. Theoretical Computer Science, 410(20): 1932–1941, May 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Golovkins:2009:CLR

[2076] Marats Golovkins, Maksim Kravtsev, and Vasilijs Kravcevs. On a class of languages recognizable by probabilistic reversible decide-and-halt automata. Theoretical Computer Science, 410(20):1942–1951, May 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dzelme-Berzina:2009:MLQ

[2077] Ilze Dzelme-Bērziņa. Mathematical logic and quantum finite state automata. Theoretical Computer Science, 410(20):1952–1959, May 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2009:Pc

[2078] Anonymous. Preface. Theoretical Computer Science, 410(20):i–vi, May 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meduna:2009:IHL

[2079] Alexander Meduna and Jiří Techet. An infinite hierarchy of language families generated by scattered context grammars with n-limited derivations. Theoretical Computer Science, 410(21–23): 1961–1969, May 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fraigniaud:2009:UAS

[2080] Pierre Fraigniaud, Cyril Gavoille, Adrian Kosowski, Emmanuelle Lebhar, and Zvi Lotker. Universal augmentation schemes for network navigability. Theoretical Computer Science, 410(21–23):1970–1981, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2009:PCC

[2081] Guanghui Wang and Guizhen Liu. Paths, cycles and circular colorings in digraphs. Theoretical Computer Science, 410(21–23):1982–1985, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Oswald:2009:SCO

[2082] Marcus Oswald and Gerhard Reinelt.

The simultaneous consecutive ones problem. Theoretical Computer Science, 410(21–23):1986–1992, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elias:2009:FNJ

[2083] Isaac Elias and Jens Lagergren. Fast neighbor joining. Theoretical Computer Science, 410(21–23):1993–2000, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2009:IST

[2084] Jinn-Shyong Yang, Jou-Ming Chang, Shyue-Ming Tang, and Yue-Li Wang. On the independent spanning trees of recursive circulant graphs G. Theoretical Computer Science, 410(21–23): 2001–2010, May 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Glasser:2009:NMS

[2085] Christian Glaßer, Alan L. Selman, Stephen Travers, and Liyu Zhang. Nonmitotic sets. Theoretical Computer Science, 410(21–23):2011–2023, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2009:PCM

[2086] Wenbin Chen, Matthew C. Schmidt, and Nagiza F. Samatova. On parameterized complexity of the Multi-MCS problem. Theoretical Computer Science, 410(21–23):2024–2032, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guignard:2009:CNK

[2087] Adrien Guignard and Éric Sopena. Compound Node–Kayles on paths. Theoretical Computer Science, 410(21– 23):2033–2044, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fleischner:2009:CGF

[2088] Herbert Fleischner, Egbert Mujuni, Daniël Paulusma, and Stefan Szeider. Covering graphs with few complete bipartite subgraphs. Theoretical Computer Science, 410(21–23):2045– 2053, May 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dantchev:2009:TRL

[2089] Stefan Dantchev, Barnaby Martin, and Mark Rhodes. Tight rank lower bounds for the Sherali–Adams proof system. Theoretical Computer Science, 410(21– 23):2054–2063, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nagoya:2009:CCC

[2090] Takayuki Nagoya and Seinosuke Toda. Computational complexity of computing a partial solution for the Graph Automorphism problems. Theoretical Computer Science, 410(21–23):2064– 2071, May 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alcon:2009:CCG

[2091] Liliana Alcón, Luerbio Faria, Celina M. H. de Figueiredo, and Marisa Gutierrez. The complexity of clique graph recognition. Theoretical Computer Science, 410(21–23):2072–2083, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goldstein:2009:ASC

[2092] Ilya Goldstein. Asymptotic subword complexity of fixed points of group substitutions. Theoretical Computer Science, 410(21–23):2084–2098, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2009:OST

[2093] Ming Liu, Yinfeng Xu, Chengbin Chu, and Feifeng Zheng. Online scheduling on two uniform machines to minimize the makespan. Theoretical Computer Science, 410(21–23):2099–2109, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baldoni:2009:IER

[2094] Roberto Baldoni, Silvia Bonomi, Leonardo Querzoni, and Sara Tucci Piergiovanni. Investigating the existence and the regularity of Logarithmic Harary Graphs. Theoretical Computer Science, 410(21–23):2110– 2121, May 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lando:2009:ISN

[2095] Yuval Lando and Zeev Nutov. Inapproximability of survivable networks. Theoretical Computer Science, 410(21– 23):2122–2125, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Col:2009:QLT

[2096] M.-A. Jacob-Da Col and P. Tellier. Quasi-linear transformations and discrete tilings. Theoretical Computer Science, 410(21–23):2126–2134, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cherubini:2009:CWP

[2097] A. Cherubini and A. Kisielewicz. Collapsing words, permutation conditions and coherent colorings of trees. Theoretical Computer Science, 410(21–23): 2135–2147, May 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reidenbach:2009:MPW

[2098] Daniel Reidenbach and Johannes C. Schneider. Morphically primitive words. Theoretical Computer Science, 410(21–23):2148–2161, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2009:CRT

[2099] Xingwu Liu, Zhiwei Xu, and Jianzhong Pan. Classifying rendezvous tasks of arbitrary dimension. Theoretical Computer Science, 410(21–23):2162– 2173, May 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beimel:2009:MCA

[2100] Amos Beimel, Boaz Ben-Moshe, Yehuda Ben-Shimol, Paz Carmi, Eldad Chai, Itzik Kitroser, and Eran Omri. Matrix columns allocation problems. Theoretical Computer Science, 410(21–23):2174–2183, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bourgeois:2009:EAM

[2101] N. Bourgeois, B. Escoffier, and V. Th.

Paschos. Efficient approximation of min set cover. Theoretical Computer Science, 410(21–23):2184–2195, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yu:2009:TMT

[2102] Changyuan Yu. Truthful mechanisms for two-range-values variant of unrelated scheduling. Theoretical Computer Science, 410(21–23):2196–2206, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galatolo:2009:CBC

[2103] Stefano Galatolo, Mathieu Hoyrup, and Cristóbal Rojas. A constructive Borel–Cantelli lemma. Constructing orbits with required statistical properties. Theoretical Computer Science, 410(21–23):2207–2222, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yi:2009:MSS

[2104] Chih-Wei Yi. Maximum scan statistics and channel assignment problems in homogeneous wireless networks. Theoretical Computer Science, 410(21–23): 2223–2233, May 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leveque:2009:CAG

[2105] Benjamin Lévêque, Frédéric Maffray, Bruce Reed, and Nicolas Trotignon. Coloring Artemis graphs. Theoretical Computer Science, 410(21–23):2234– 2240, May 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bockenhauer:2009:AHD

[2106] Hans-Joachim Böckenhauer, Joachim Kneis, and Joachim Kupke. Approximation hardness of deadline-TSP reoptimization. Theoretical Computer Science, 410(21–23):2241–2249, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vagvolgyi:2009:DBT

[2107] Sándor Vágvölgyi. Deterministic bottom-up tree transducers and ground term rewrite systems. Theoretical Computer Science, 410(21–23):2250– 2278, May 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martinez:2009:MDP

[2108] Conrado Martínez and Helmut Prodinger. Moves and displacements of particular elements in Quicksort. Theoretical Computer Science, 410(21–23):2279– 2284, May 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhao:2009:MAA

[2109] Shang-Wei Zhao and Xiao-Shan Gao. Minimal achievable approximation ratio for MAX-MQ in finite fields. Theoretical Computer Science, 410(21–23): 2285–2290, May 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tian:2009:BOA

[2110] Ji Tian, Ruyan Fu, and Jinjiang Yuan.

A best online algorithm for scheduling on two parallel batch machines. Theoretical Computer Science, 410(21–23): 2291–2294, May 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoffmann:2009:FTS

[2111] Jan Hoffmann. Finding a tree structure in a resolution proof is NP-complete. Theoretical Computer Science, 410(21– 23):2295–2300, May 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2009:Pd

[2112] Anonymous. Preface. Theoretical Computer Science, 410(24–25):2301– 2307, May 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alhazov:2009:PCS

[2113] Artiom Alhazov, Ion Petre, and Vladimir Rogojin. The parallel complexity of signed graphs: Decidability results and an improved algorithm. Theoretical Computer Science, 410(24– 25):2308–2315, May 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cai:2009:BSO

[2114] Ying Cai and Cunsheng Ding. Binary sequences with optimal autocorrelation. Theoretical Computer Science, 410(24–25):2316–2322, May 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calude:2009:TW

[2115] Cristian S. Calude, Helmut Jürgensen, and Ludwig Staiger. Topology on words. Theoretical Computer Science, 410(24–25):2323–2335, May 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Campeanu:2009:IRL

[2116] Cezar Câmpeanu and Nicolae Santean. On the intersection of regex languages with regular languages. Theoretical Computer Science, 410(24–25): 2336–2344, May 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cassaigne:2009:CFB

[2117] Julien Cassaigne, Juhani Karhumäki, and Petri Salmela. Conjugacy of finite biprefix codes. Theoretical Computer Science, 410(24–25):2345–2351, May 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cavaliere:2009:ASN

[2118] Matteo Cavaliere, Oscar H. Ibarra, Gheorghe Păun, Omer Egecioglu, Mihai Ionescu, and Sara Woodworth. Asynchronous spiking neural P systems. Theoretical Computer Science, 410(24–25):2352–2364, May 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2009:SMD

[2119] Shihyen Chen, Bin Ma, and Kaizhong Zhang. On the similarity metric and the distance metric. Theoretical Computer Science, 410(24–25):2365– 2376, May 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Domaratzki:2009:SCP

[2120] Michael Domaratzki and Alexander Okhotin. State complexity of power. Theoretical Computer Science, 410(24– 25):2377–2392, May 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kari:2009:TRW

[2121] Lila Kari, Kalpana Mahalingam, and Shinnosuke Seki. Twin-roots of words and their properties. Theoretical Computer Science, 410(24–25):2393–2400, May 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krieger:2009:DLS

[2122] Dalia Krieger, Avery Miller, Narad Rampersad, Bala Ravikumar, and Jeffrey Shallit. Decimations of languages and state complexity. Theoretical Computer Science, 410(24–25): 2401–2409, May 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2009:TOP

[2123] Shuai Cheng Li and Ming Li. On two open problems of 2-interval patterns. Theoretical Computer Science, 410(24–25):2410–2423, May 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paun:2009:HMT

[2124] Andrei Păun, Mihaela Păun, and Alfonso Rodríguez-Patón. On the Hopcroft’s minimization technique for DFA and DFCA. Theoretical Computer Science, 410(24–25):2424–2430, May 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rampersad:2009:SCU

[2125] Narad Rampersad, Nicolae Santean, Jeffrey Shallit, and Bala Ravikumar. State complexity of unique rational operations. Theoretical Computer Science, 410(24–25):2431–2441, May 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yen:2009:MEU

[2126] Hsu-Chun Yen and Chien-Liang Chen. On minimal elements of upward-closed sets. Theoretical Computer Science, 410(24–25):2442–2452, May 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2009:Pe

[2127] G. Rozenberg. Preface. Theoretical Computer Science, 410(26):2453– 2454, June 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Friedrich:2009:CSD

[2128] Tobias Friedrich, Nils Hebbinghaus, and Frank Neumann. Comparison of simple diversity mechanisms on plateau functions. Theoretical Computer Science, 410(26):2455–2462, June 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jain:2009:NBC

[2129] Rahul Jain and Shengyu Zhang. New bounds on classical and quantum oneway communication complexity. Theoretical Computer Science, 410(26): 2463–2477, June 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2009:LGA

[2130] Xingyi Zhang, Xiangxiang Zeng, and Linqiang Pan. On languages generated by asynchronous spiking neural P systems. Theoretical Computer Science, 410(26):2478–2488, June 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Broadbent:2009:PQC

[2131] Anne Broadbent and Elham Kashefi.

Parallelizing quantum circuits. Theoretical Computer Science, 410(26): 2489–2510, June 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sudholt:2009:IPM

[2132] Dirk Sudholt. The impact of parametrization in memetic evolutionary algorithms. Theoretical Computer Science, 410(26):2511–2528, June 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2009:NQS

[2133] Lvzhou Li and Daowen Qiu. A note on quantum sequential machines. Theoretical Computer Science, 410(26):2529– 2535, June 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2009:Pe

[2134] Anonymous. Preface. Theoretical Computer Science, 410(26):i–vi, June 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Han:2009:SCB

[2135] Yo-Sub Han and Kai Salomaa. State complexity of basic operations on suffix-free regular languages. Theoretical Computer Science, 410(27–29): 2537–2548, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berenbrink:2009:EER

[2136] Petra Berenbrink, Colin Cooper, and Zengjian Hu. Energy efficient randomised communication in unknown AdHoc networks. Theoretical Computer Science, 410(27–29):2549–2561, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jain:2009:OSL

[2137] Sanjay Jain and Efim Kinber. Oneshot learners using negative counterexamples and nearest positive examples. Theoretical Computer Science, 410(27– 29):2562–2580, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Su:2009:NHA

[2138] Chi-Shiang Su, Jason Chao-Hsien Pan, and Tsung-Shin Hsu. A new heuristic algorithm for the machine scheduling problem with job delivery coordination. Theoretical Computer Science, 410(27–29):2581–2591, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Thakur:2009:LCP

[2139] Mayur Thakur and Rahul Tripathi. Linear connectivity problems in directed hypergraphs. Theoretical Computer Science, 410(27–29):2592–2618, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wu:2009:OTS

[2140] Weiwei Wu, Minming Li, and Enhong Chen. Optimal tree structures for group key tree management considering insertion and deletion cost. Theoretical Computer Science, 410(27–29): 2619–2631, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Park:2009:CMP

[2141] Jung-Heum Park and Sang Hyuk Son. Conditional matching preclusion for hypercube-like interconnection networks. Theoretical Computer Science, 410(27–29):2632–2640, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2009:PET

[2142] Jianer Chen, Iyad A. Kanj, and Ge Xia.

On parameterized exponential time complexity. Theoretical Computer Science, 410(27–29):2641–2648, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harvey:2009:CFT

[2143] David Harvey. A cache-friendly truncated FFT. Theoretical Computer Science, 410(27–29):2649–2658, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garg:2009:IPG

[2144] Sanchit Garg and Éric Schost. Interpolation of polynomials given by straight-line programs. Theoretical Computer Science, 410(27–29):2659– 2662, June 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fu:2009:EIF

[2145] Bin Fu, Yumei Huo, and Hairong Zhao. Exponential inapproximability and FP-TAS for scheduling with availability constraints. Theoretical Computer Science, 410(27–29):2663–2674, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hong:2009:CRS

[2146] Soonjo Hong and Sujin Shin. Cyclic renewal systems. Theoretical Computer Science, 410(27–29):2675–2684, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lasserre:2009:PPA

[2147] Jean B. Lasserre, Monique Laurent, and Philipp Rostalski. A prolongation– projection algorithm for computing the finite real variety of an ideal. Theoretical Computer Science, 410(27–29): 2685–2700, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeCarli:2009:LLT

[2148] F. De Carli, A. Frosini, S. Rinaldi, and A. Sorbi. Lattices of local twodimensional languages. Theoretical Computer Science, 410(27–29):2701– 2713, June 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chang:2009:SM

[2149] Ching-Lueh Chang and Yuh-Dauh Lyuu. Spreading messages. Theoretical Computer Science, 410(27–29): 2714–2724, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaz:2009:BCA

[2150] Josep Diaz, Fabrizio Grandoni, and Alberto Marchetti Spaccamela. Balanced cut approximation in random geometric graphs. Theoretical Computer Science, 410(27–29):2725–2731, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cao:2009:PPB

[2151] Zhigang Cao and Xiaoguang Yang.

A PTAS for parallel batch scheduling with rejection and dynamic job arrivals. Theoretical Computer Science, 410(27–29):2732–2745, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2009:PCC

[2152] Hong Liu, Haodi Feng, Daming Zhu, and Junfeng Luan. Parameterized computational complexity of control problems in voting systems. Theoretical Computer Science, 410(27–29):2746– 2753, June 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kashan:2009:NMM

[2153] Ali Husseinzadeh Kashan, Behrooz Karimi, and S. M. T. Fatemi Ghomi. A note on minimizing makespan on a single batch processing machine with nonidentical job sizes. Theoretical Computer Science, 410(27–29):2754– 2758, June 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sakai:2009:CLT

[2154] Yoshifumi Sakai. Computing the longest topological common subsequence of a symbol-wise totally ordered directed acyclic graph and a sequence. Theoretical Computer Science, 410(27–29):2759–2766, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ollinger:2009:APV

[2155] N. Ollinger and G. Richard. Automata on the plane vs particles and collisions. Theoretical Computer Science, 410(27–29):2767–2773, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rhodes:2009:CRP

[2156] Mark Rhodes. On the Chvátal rank of the Pigeonhole Principle. Theoretical Computer Science, 410(27–29): 2774–2778, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balkova:2009:NSR

[2157] L’ubomíra Balková and Edita Pelantová. A note on symmetries in the Rauzy graph and factor frequencies. Theoretical Computer Science, 410(27– 29):2779–2783, June 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2009:EBb

[2158] Anonymous. Editorial board. Theoretical Computer Science, 410(27–29): i–vi, June 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2009:Pf

[2159] Anonymous. Preface. Theoretical Computer Science, 410(30–32):2785– 2794, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Allouche:2009:PRO

[2160] Jean-Paul Allouche, Narad Rampersad, and Jeffrey Shallit. Periodicity, repetitions, and orbits of an automatic sequence. Theoretical Computer Science, 410(30–32):2795–2803, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baturo:2009:CSM

[2161] Paweł Baturo and Wojciech Rytter.

Compressed string-matching in standard Sturmian words. Theoretical Computer Science, 410(30–32):2804– 2810, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berstel:2009:CPW

[2162] Jean Berstel, Luc Boasson, and Olivier Carton. Continuant polynomials and worst-case behavior of Hopcroft’s minimization algorithm. Theoretical Computer Science, 410(30–32):2811–2822, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blondel:2009:NPF

[2163] Vincent D. Blondel, Julien Cassaigne, and Raphaël M. Jungers. On the number of α-power-free binary words for 2 < α ≤ 7/3. Theoretical Computer Science, 410(30–32):2823– 2833, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bu:2009:FCS

[2164] Dongbo Bu, Ming Li, Shuai Cheng Li, Jianbo Qian, and Jinbo Xu. Finding compact structural motifs. Theoretical Computer Science, 410(30–32): 2834–2839, August 20, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bucci:2009:CMG

[2165] Michelangelo Bucci, Aldo de Luca, and Alessandro De Luca. Characteristic morphisms of generalized episturmian words. Theoretical Computer Science, 410(30–32):2840–2859, August 20, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bucci:2009:NCP

[2166] Michelangelo Bucci, Alessandro De Luca, Amy Glen, and Luca Q. Zamboni. A new characteristic property of rich words. Theoretical Computer Science, 410(30–32):2860–2863, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bugeaud:2009:SSR

[2167] Yann Bugeaud, Christophe Reutenauer, and Samir Siksek. A Sturmian sequence related to the uniqueness conjecture for Markoff numbers. Theoretical Computer Science, 410(30–32): 2864–2869, August 20, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choffrut:2009:ELL

[2168] Christian Choffrut and Serge Grigorieff. The “equal last letter” predicate for words on infinite alphabets and classes of multitape automata. Theoretical Computer Science, 410(30–32): 2870–2884, August 20, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Currie:2009:DCH

[2169] James Currie and Narad Rampersad.

Dejean’s conjecture holds for n ≥ 30. Theoretical Computer Science, 410 (30–32):2885–2888, August 20, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czeizler:2009:SWE

[2170] Elena Czeizler and Wojciech Plandowski.

On systems of word equations over three unknowns with at most six occurrences of one of the unknowns. Theoretical Computer Science, 410(30–32): 2889–2909, August 20, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dassow:2009:SOP

[2171] Jürgen Dassow, Gema M. Martín, and Francisco J. Vico. Some operations preserving primitivity of words. Theoretical Computer Science, 410(30–32): 2910–2919, August 20, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaz:2009:STF

[2172] J. Díaz, L. Kirousis, D. Mitsche, and X. Pérez-Giménez. On the satisfiability threshold of formulas with three literals per clause. Theoretical Computer Science, 410(30–32):2920– 2934, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diekert:2009:SRA

[2173] Volker Diekert and Dalia Krieger. Some remarks about stabilizers. Theoretical Computer Science, 410(30–32):2935– 2946, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frid:2009:SEB

[2174] A. E. Frid. Simple equations on binary factorial languages. Theoretical Computer Science, 410(30–32):2947– 2956, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Halava:2009:PCP

[2175] Vesa Halava, Jarkko Kari, and Yuri Matiyasevich. On post correspondence problem for letter monotonic languages. Theoretical Computer Science, 410(30–32):2957–2960, August 20, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Han:2009:NSC

[2176] Yo-Sub Han and Kai Salomaa. Nondeterministic state complexity of nested word automata. Theoretical Computer Science, 410(30–32):2961–2971, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hromkovic:2009:LCC

[2177] Juraj Hromkovič, Holger Petersen, and Georg Schnitger. On the limits of the communication complexity technique for proving lower bounds on the size of minimal NFA’s. Theoretical Computer Science, 410(30–32):2972– 2981, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2009:SSS

[2178] Oscar H. Ibarra, Andrei Păun, and Alfonso Rodríguez-Patón. Sequential SNP systems based on min/max spike number. Theoretical Computer Science, 410(30–32):2982–2991, August 20, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mikhailova:2009:PAP

[2179] I. A. Mikhailova and M. V. Volkov.

Pattern avoidance by palindromes. Theoretical Computer Science, 410 (30–32):2992–2998, August 20, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nicolas:2009:ECM

[2180] François Nicolas, Veli Mäkinen, and Esko Ukkonen. Efficient construction of maximal and minimal representations of motifs of a string. Theoretical Computer Science, 410(30–32):2999– 3005, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Qiu:2009:HEM

[2181] Daowen Qiu and Sheng Yu. Hierarchy and equivalence of multi-letter quantum finite automata. Theoretical Computer Science, 410(30–32):3006– 3017, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Restivo:2009:BWT

[2182] Antonio Restivo and Giovanna Rosone. Burrows–Wheeler transform and palindromic richness. Theoretical Computer Science, 410(30–32):3018–3026, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tijdeman:2009:FWW

[2183] R. Tijdeman and L. Q. Zamboni. Fine and Wilf words for any periods II. Theoretical Computer Science, 410(30–32): 3027–3034, August 20, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2009:EBc

[2184] Anonymous. Editorial board. Theoretical Computer Science, 410(30–32): i–vi, August 20, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2009:JKC

[2185] Anonymous. Juhani Karhumaki with a capercaillie. Theoretical Computer Science, 410(30–32):xi, August 20, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2009:Pf

[2186] G. Rozenberg. Preface. Theoretical Computer Science, 410(33–34):3035– 3036, August 21, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cannata:2009:P

[2187] Nicola Cannata and Emanuela Merelli.

Preface. Theoretical Computer Science, 410(33–34):3037–3038, August 21, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Versari:2009:SBM

[2188] Cristian Versari and Nadia Busi.

Stochastic biological modelling in the presence of multiple compartments. Theoretical Computer Science, 410 (33–34):3039–3064, August 21, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ciocchetta:2009:BPF

[2189] Federica Ciocchetta and Jane Hillston. Bio-PEPA: a framework for the modelling and analysis of biological systems. Theoretical Computer Science, 410(33–34):3065–3084, August 21, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barbuti:2009:ILS

[2190] Roberto Barbuti, Giulio Caravagna, Andrea Maggiolo-Schettini, and Paolo Milazzo. An intermediate language for the stochastic simulation of biological systems. Theoretical Computer Science, 410(33–34):3085–3109, August 21, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bodei:2009:CFA

[2191] Chiara Bodei. A Control Flow Analysis for Beta-binders with and without static compartments. Theoretical Computer Science, 410(33–34):3110– 3127, August 21, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barnat:2009:AAT

[2192] J. Barnat, L. Brim, I. Černá, S. Dražan, J. Fabriková, and D. Šafránek. On algorithmic analysis of transcriptional regulation by LTL model checking. Theoretical Computer Science, 410(33–34): 3128–3148, August 21, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bartocci:2009:MSC

[2193] E. Bartocci, F. Corradini, M. R. Di Berardini, E. Entcheva, S. A. Smolka, and R. Grosu. Modeling and simulation of cardiac tissue using hybrid I/O automata. Theoretical Computer Science, 410(33–34):3149–3165, August 21, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cardelli:2009:PMR

[2194] Luca Cardelli, Emmanuelle Caron, Philippa Gardner, Ozan Kahramanoḡulları, and Andrew Phillips. A process model of Rho GTP-binding proteins. Theoretical Computer Science, 410(33–34):3166–3185, August 21, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2009:Pg

[2195] Anonymous. Preface. Theoretical Computer Science, 410(33–34):i–vi, August 21, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Campeanu:2009:P

[2196] Cezar Câmpeanu and Giovanni Pighizzini.

Preface. Theoretical Computer Science, 410(35):3187, August 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alhazov:2009:SCC

[2197] Artiom Alhazov, Erzsébet Csuhaj-Varjú, Carlos Martín-Vide, and Yurii Rogozhin. On the size of computationally complete hybrid networks of evolutionary processors. Theoretical Computer Science, 410(35):3188– 3197, August 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biegler:2009:SDD

[2198] Franziska Biegler and Kai Salomaa.

On the synchronized derivation depth of context-free grammars. Theoretical Computer Science, 410(35):3198– 3208, August 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bordihn:2009:DFA

[2199] Henning Bordihn, Markus Holzer, and Martin Kutrib. Determination of finite automata accepting subregular languages. Theoretical Computer Science, 410(35):3209–3222, August 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brzozowski:2009:SCH

[2200] Janusz Brzozowski and Stavros Konstantinidis. State-complexity hierarchies of uniform languages of alphabetsize length. Theoretical Computer Science, 410(35):3223–3235, August 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brzozowski:2009:PS

[2201] Janusz Brzozowski and Nicolae Santean. Predictable semiautomata. Theoretical Computer Science, 410(35): 3236–3249, August 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czeizler:2009:DCW

[2202] Elena Czeizler, Eugen Czeizler, Lila Kari, and Kai Salomaa. On the descriptional complexity of Watson–Crick automata. Theoretical Computer Science, 410(35):3250–3260, August 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dassow:2009:TCH

[2203] Jürgen Dassow, Ralf Stiebe, and Bianca Truthe. Two collapsing hierarchies of subregularly tree controlled languages. Theoretical Computer Science, 410(35):3261–3271, August 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esik:2009:ESC

[2204] Zoltán Ésik, Yuan Gao, Guangwu Liu, and Sheng Yu. Estimation of state complexity of combined operations. Theoretical Computer Science, 410(35): 3272–3280, August 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gruber:2009:LOR

[2205] Hermann Gruber and Markus Holzer. Language operations with regular expressions of polynomial size. Theoretical Computer Science, 410(35):3281– 3289, August 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Piao:2009:OSC

[2206] Xiaoxue Piao and Kai Salomaa. Operational state complexity of nested word automata. Theoretical Computer Science, 410(35):3290–3302, August 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mavronicolas:2009:P

[2207] Marios Mavronicolas. Preface. Theoretical Computer Science, 410(36):3303– 3304, August 31, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fotakis:2009:SCN

[2208] Dimitris Fotakis, Spyros Kontogiannis, Elias Koutsoupias, Marios Mavronicolas, and Paul Spirakis. The structure and complexity of Nash equilibria for a selfish routing game. Theoretical Computer Science, 410(36):3305– 3326, August 31, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Christodoulou:2009:CM

[2209] George Christodoulou, Elias Koutsoupias, and Akash Nanavati. Coordination mechanisms. Theoretical Computer Science, 410(36):3327–3336, August 31, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Busch:2009:ARG

[2210] Costas Busch and Malik Magdon-Ismail. Atomic routing games on maximum congestion. Theoretical Computer Science, 410(36):3337–3347, August 31, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Auletta:2009:DTM

[2211] Vincenzo Auletta, Roberto De Prisco, Paolo Penna, and Giuseppe Persiano. On designing truthful mechanisms for online scheduling. Theoretical Computer Science, 410(36):3348–3356, August 31, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fischer:2009:ARS

[2212] Simon Fischer and Berthold Vöcking. Adaptive routing with stale information. Theoretical Computer Science, 410(36):3357–3371, August 31, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chitturi:2009:UBS

[2213] B. Chitturi, W. Fahle, Z. Meng, L. Morales, C. O. Shields, I. H. Sudborough, and W. Voit. An (18/11)n upper bound for sorting by prefix reversals. Theoretical Computer Science, 410(36):3372–3390, August 31, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kutylowski:2009:OSM

[2214] Jarosław Kutyłowski and Friedhelm Meyer auf der Heide. Optimal strategies for maintaining a chain of relays between an explorer and a base camp. Theoretical Computer Science, 410(36):3391–3405, August 31, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ausiello:2009:SSS

[2215] Giorgio Ausiello, Paolo G. Franciosa, and Giuseppe F. Italiano. Small stretch (α, β)-spanners in the streaming model. Theoretical Computer Science, 410(36): 3406–3413, August 31, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elsasser:2009:RRR

[2216] R. Elsässer and T. Sauerwald. On the runtime and robustness of randomized broadcasting. Theoretical Computer Science, 410(36):3414–3427, August 31, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bockenhauer:2009:RST

[2217] Hans-Joachim Böckenhauer, Juraj Hromkovič, Richard Královič, Tobias Mömke, and Peter Rossmanith. Reoptimization of Steiner trees: Changing the terminal set. Theoretical Computer Science, 410(36):3428–3435, August 31, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Holub:2009:P

[2218] Jan Holub. Preface. Theoretical Computer Science, 410(37):3437, September 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salomaa:2009:DDC

[2219] Kai Salomaa, Sheng Yu, and Jinfeng Zan. Deciding determinism of caterpillar expressions. Theoretical Computer Science, 410(37):3438–3446, September 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kutrib:2009:RNP

[2220] Martin Kutrib, Andreas Malcher, and Larissa Werlein. Regulated nondeterminism in pushdown automata. Theoretical Computer Science, 410(37): 3447–3460, September 1, 2009. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ackerman:2009:EEW

[2221] Margareta Ackerman and Jeffrey Shallit. Efficient enumeration of words in regular languages. Theoretical Computer Science, 410(37):3461–3470, September 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crochemore:2009:NCS

[2222] M. Crochemore, C. Epifanio, A. Gabriele, and F. Mignosi. From Nerode’s congruence to suffix automata with mismatches. Theoretical Computer Science, 410(37):3471–3480, September 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Droste:2009:WAW

[2223] Manfred Droste and George Rahonis. Weighted automata and weighted logics with discounting. Theoretical Computer Science, 410(37):3481–3494, September 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Steinby:2009:DSD

[2224] Magnus Steinby and Cătălin Ionuţ Tîrnăucă. Defining syntax-directed translations by tree bimorphisms. Theoretical Computer Science, 410(37):

3495–3503, September 1, 2009. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jonoska:2009:FSA

[2225] Nataša Jonoska and Joni B. Pirnot. Finite state automata representing twodimensional subshifts. Theoretical Computer Science, 410(37):3504–3512, September 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Volkov:2009:SAP

[2226] M. V. Volkov. Synchronizing automata preserving a chain of partial orders. Theoretical Computer Science, 410(37):3513–3519, September 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anselmo:2009:CMT

[2227] Marcella Anselmo, Dora Giammarresi, and Maria Madonia. A computational model for tiling recognizable two-dimensional languages. Theoretical Computer Science, 410(37):3520– 3529, September 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mraz:2009:DWE

[2228] F. Mráz, F. Otto, and M. Plátek. The degree of word-expansion of lexicalized RRWW-automata — a new measure for the degree of nondeterminism of (context-free) languages. Theoretical Computer Science, 410(37):3530– 3538, September 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hogberg:2009:BFB

[2229] Johanna Högberg, Andreas Maletti, and Jonathan May. Backward and forward bisimulation minimization of tree automata. Theoretical Computer Science, 410(37):3539–3552, September 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mohri:2009:GSA

[2230] Mehryar Mohri, Pedro Moreno, and Eugene Weinstein. General suffix automaton construction algorithm and space bounds. Theoretical Computer Science, 410(37):3553–3562, September 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klein:2009:ABM

[2231] Shmuel T. Klein and Miri Kopel Ben-Nissan. Accelerating Boyer–Moore searches on binary texts. Theoretical Computer Science, 410(37):3563– 3571, September 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moshe:2009:JSC

[2232] Yossi Moshe. On the joint subword complexity of automatic sequences. Theoretical Computer Science, 410(38– 40):3573–3588, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Masakova:2009:RBP

[2233] Z. Masáková and E. Pelantová. Relation between powers of factors and the recurrence function characterizing Sturmian words. Theoretical Computer Science, 410(38–40):3589–3596, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2009:OPM

[2234] An Zhang, Yiwei Jiang, and Zhiyi Tan. Online parallel machines scheduling with two hierarchies. Theoretical Computer Science, 410(38–40):3597– 3605, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muller:2009:CHL

[2235] Johannes Müller and Christoph Spandl. A Curtis–Hedlund–Lyndon theorem for Besicovitch and Weyl spaces. Theoretical Computer Science, 410(38–40):3606–3615, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mishna:2009:TNH

[2236] Marni Mishna and Andrew Rechnitzer.

Two non-holonomic lattice walks in the quarter plane. Theoretical Computer Science, 410(38–40):3616–3630, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bein:2009:OCL

[2237] Wolfgang Bein, Leah Epstein, Lawrence L.

Larmore, and John Noga. Optimally competitive list batching. Theoretical Computer Science, 410(38–40):3631– 3639, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Komusiewicz:2009:ICE

[2238] Christian Komusiewicz, Falk Hüffner, Hannes Moser, and Rolf Niedermeier.

Isolation concepts for efficiently enumerating dense subgraphs. Theoretical Computer Science, 410(38–40):3640– 3654, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Uscka-Wehlou:2009:TER

[2239] Hanna Uscka-Wehlou. Two equivalence relations on digital lines with irrational slopes. A continued fraction approach to upper mechanical words. Theoretical Computer Science, 410(38–40):3655– 3669, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jungers:2009:OFW

[2240] Raphaël M. Jungers, Vladimir Y. Protasov, and Vincent D. Blondel. Overlap-free words and spectra of matrices. Theoretical Computer Science, 410(38–40):3670–3684, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Acerbi:2009:CSD

[2241] Luigi Acerbi, Alberto Dennunzio, and Enrico Formenti. Conservation of some dynamical properties for operations on cellular automata. Theoretical Computer Science, 410(38–40):3685–3693, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dorrigiv:2009:RDP

[2242] Reza Dorrigiv, Alejandro López-Ortiz, and J. Ian Munro. On the relative dominance of paging algorithms. Theoretical Computer Science, 410(38– 40):3694–3701, September 6, 2009.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hasunuma:2009:LT

[2243] Toru Hasunuma, Toshimasa Ishii, Hirotaka Ono, and Yushi Uno. An O(n^{1.75}) L(2, 1)-labeling of trees. Theoretical Computer Science, 410(38–40):3702– 3710, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biegler:2009:USW

[2244] Franziska Biegler, Mark Daley, Markus Holzer, and Ian McQuillan. On the uniqueness of shuffle on words and finite languages. Theoretical Computer Science, 410(38–40):3711–3724, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Levi:2009:ACO

[2245] Tamir Levi and Ami Litman. Accelerating certain outputs of merging and sorting networks. Theoretical Computer Science, 410(38–40):3725–3732, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kowalik:2009:IEC

[2246] Łukasz Kowalik. Improved edgecoloring with three colors. Theoretical Computer Science, 410(38–40):3733– 3742, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Foata:2009:NPC

[2247] Dominique Foata and Guo-Niu Han.

New permutation coding and equidistribution of set-valued statistics. Theoretical Computer Science, 410(38– 40):3743–3750, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amini:2009:HAT

[2248] Omid Amini, Stéphane Pérennes, and Ignasi Sau. Hardness and approximation of traffic grooming. Theoretical Computer Science, 410(38–40):3751– 3760, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ji:2009:PMS

[2249] Min Ji and T. C. E. Cheng. Parallelmachine scheduling of simple linear deteriorating jobs. Theoretical Computer Science, 410(38–40):3761–3768, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jiang:2009:SNT

[2250] Tao Jiang, Zevi Miller, and Dan Pritikin. Separation numbers of trees. Theoretical Computer Science, 410(38– 40):3769–3781, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paquin:2009:GCW

[2251] Geneviève Paquin. On a generalization of Christoffel words: epichristoffel words. Theoretical Computer Science, 410(38–40):3782–3791, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Augustine:2009:SPP

[2252] John Augustine, Sudarshan Banerjee, and Sandy Irani. Strip packing with precedence constraints and strip packing with release times. Theoretical Computer Science, 410(38–40):3792– 3803, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2009:PAU

[2253] Zi-Long Liu, Fang Tian, and Jun-Ming Xu. Probabilistic analysis of upper bounds for 2-connected distance kdominating sets in graphs. Theoretical Computer Science, 410(38–40):3804– 3813, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hermann:2009:CCO

[2254] Miki Hermann and Reinhard Pichler.

Complexity of counting the optimal solutions. Theoretical Computer Science, 410(38–40):3814–3825, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wittmann:2009:RGB

[2255] Dominik M. Wittmann, Daniel Schmidl, Florian Blöchl, and Fabian J. Theis. Reconstruction of graphs based on random walks. Theoretical Computer Science, 410(38–40):3826–3838, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beyersdorff:2009:NFE

[2256] Olaf Beyersdorff, Johannes Köbler, and Jochen Messner. Nondeterministic functions and the existence of optimal proof systems. Theoretical Computer Science, 410(38–40):3839–3855, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jonsson:2009:HCS

[2257] Peter Jonsson, Andrei Krokhin, and Fredrik Kuivinen. Hard constraint satisfaction problems have hard gaps at location 1. Theoretical Computer Science, 410(38–40):3856–3874, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2009:OSU

[2258] Ming Liu, Chengbin Chu, Yinfeng Xu, and Feifeng Zheng. Online scheduling on m uniform machines to minimize total (weighted) completion time. Theoretical Computer Science, 410(38– 40):3875–3881, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hua:2009:SMC

[2259] Qiang-Sheng Hua, Yuexuan Wang, Dongxiao Yu, and Francis C. M. Lau. Set multi-covering via inclusion– exclusion. Theoretical Computer Science, 410(38–40):3882–3892, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Keranen:2009:PAS

[2260] Veikko Keränen. A powerful abelian square-free substitution over 4 letters. Theoretical Computer Science, 410(38– 40):3893–3900, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Greco:2009:CCN

[2261] Gianluigi Greco and Francesco Scarcello. On the complexity of constrained Nash equilibria in graphical games. Theoretical Computer Science, 410(38–40):3901–3924, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biskup:2009:SSS

[2262] Marek Tomasz Biskup and Wojciech Plandowski. Shortest synchronizing strings for Huffman codes. Theoretical Computer Science, 410(38–40):3925– 3941, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2009:FAF

[2263] J. J. Liu, G. S. Huang, and Y. L.

Wang. A fast algorithm for finding the positions of all squares in a run-length encoded string. Theoretical Computer Science, 410(38–40):3942– 3948, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bulatov:2009:CWB

[2264] Andrei Bulatov, Martin Dyer, Leslie Ann Goldberg, Markus Jalsenius, and David Richerby. The complexity of weighted Boolean #CSP with mixed signs. Theoretical Computer Science, 410(38– 40):3949–3961, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dennunzio:2009:SAC

[2265] Alberto Dennunzio, Pierre Guillon, and Benoît Masson. Sand automata as cellular automata. Theoretical Computer Science, 410(38–40):3962–3974, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lee:2009:OST

[2266] Kangbok Lee, Joseph Y.-T. Leung, and Michael L. Pinedo. Online scheduling on two uniform machines subject to eligibility constraints. Theoretical Computer Science, 410(38–40):3975–3981, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mathissen:2009:EMT

[2267] Christian Mathissen. Existential MSO over two successors is strictly weaker than over linear orders. Theoretical Computer Science, 410(38–40):3982– 3987, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Morizumi:2009:LNN

[2268] Hiroki Morizumi. Limiting negations in non-deterministic circuits. Theoretical Computer Science, 410(38–40):3988– 3994, September 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Erdelyi:2009:GJN

[2269] Gábor Erdélyi, Lane A. Hemaspaandra, Jörg Rothe, and Holger Spakowski. Generalized juntas and NPhard sets. Theoretical Computer Science, 410(38–40):3995–4000, September 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carbone:2009:FFM

[2270] Marco Carbone, Paweł Sobociński, and Frank D. Valencia. Foreword: Festschrift for Mogens Nielsen’s 60th birthday. Theoretical Computer Science, 410(41):4001–4005, September 15, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beauxis:2009:PNA

[2271] Romain Beauxis and Catuscia Palamidessi. Probabilistic and nondeterministic aspects of anonymity. Theoretical Computer Science, 410(41):4006–4025, September 15, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Benes:2009:DMT

[2272] N. Beneš, J. Křetínský, K. G. Larsen, and J. Srba. On determinism in modal transition systems. Theoretical Computer Science, 410(41):4026–4043, September 15, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonchi:2009:RSS

[2273] Filippo Bonchi and Ugo Montanari.

Reactive systems, (semi-)saturated semantics and coalgebras on presheaves. Theoretical Computer Science, 410 (41):4044–4066, September 15, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

ElSalamouny:2009:AED

[2274] Ehab ElSalamouny, Karl Tikjøb Krukow, and Vladimiro Sassone. An analysis of the exponential decay principle in probabilistic trust models.

Theoretical Computer Science, 410 (41):4067–4084, September 15, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frandsen:2009:DMR

[2275] Gudmund Skovbjerg Frandsen and Peter Frands Frandsen. Dynamic matrix rank. Theoretical Computer Science, 410(41):4085–4093, September 15, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gazagnaire:2009:CMS

[2276] Thomas Gazagnaire, Blaise Genest, Loïc Hélouët, P. S. Thiagarajan, and Shaofa Yang. Causal Message Sequence Charts. Theoretical Computer Science, 410(41):4094–4110, September 15, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanGlabbeek:2009:CSE

[2277] R. J. van Glabbeek and G. D. Plotkin. Configuration structures, event structures and Petri nets. Theoretical Computer Science, 410(41):4111–4159, September 15, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Winskel:2009:PA

[2278] Glynn Winskel. Prime algebraicity. Theoretical Computer Science, 410 (41):4160–4168, September 15, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chadha:2009:DBT

[2279] Rohit Chadha and Mahesh Viswanathan. Deciding branching time properties for asynchronous programs. Theoretical Computer Science, 410(42):4169–4179, September 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Niebert:2009:EMC

[2280] Peter Niebert and Doron Peled. Efficient model checking for LTL with partial order snapshots. Theoretical Computer Science, 410(42):4180–4189, September 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Colson:2009:PSO

[2281] Loïc Colson and David Michel. Pedagogical second-order λ-calculus. Theoretical Computer Science, 410(42): 4190–4203, September 28, 2009. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

David:2009:DPC

[2282] René David. A direct proof of the confluence of combinatory strong reduction. Theoretical Computer Science, 410(42):4204–4215, September 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Preoteasa:2009:FRM

[2283] Viorel Preoteasa. Frame rule for mutually recursive procedures manipulating pointers. Theoretical Computer Science, 410(42):4216–4233, September 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mao:2009:MCP

[2284] Xuxin Mao and Luoshan Xu. Meet continuity properties of posets. Theoretical Computer Science, 410(42):4234–4240, September 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hadjidj:2009:FTM

[2285] Rachid Hadjidj and Hanifa Boucheneb. On-the-fly TCTL model checking for time Petri nets. Theoretical Computer Science, 410(42):4241–4261, September 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fainekos:2009:RTL

[2286] Georgios E. Fainekos and George J. Pappas. Robustness of temporal logic specifications for continuous-time signals. Theoretical Computer Science, 410(42):4262–4291, September 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Iliopoulos:2009:FSI

[2287] Costas Iliopoulos and Wojciech Rytter. Foreword: Special issue in honor of the 60th birthday of Prof. Maxime Crochemore. Theoretical Computer Science, 410(43):4293–4294, October 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Smyth:2009:NAP

[2288] W. F. Smyth and Shu Wang. A new approach to the periodicity lemma on strings with holes. Theoretical Computer Science, 410(43):4295–4302, October 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mozes:2009:FAC

[2289] Shay Mozes, Dekel Tsur, Oren Weimann, and Michal Ziv-Ukelson. Fast algorithms for computing tree LCS. Theoretical Computer Science, 410(43):4303–4314, October 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kapah:2009:IRE

[2290] Oren Kapah, Gad M. Landau, Avivit Levy, and Nitsan Oz. Interchange rearrangement: The element-cost model. Theoretical Computer Science, 410(43): 4315–4326, October 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Battaglia:2009:MPS

[2291] Giovanni Battaglia, Davide Cangelosi, Roberto Grossi, and Nadia Pisanti. Masking patterns in sequences: a new class of motif discovery with don’t cares. Theoretical Computer Science, 410(43):4327–4340, October 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ukkonen:2009:MMR

[2292] Esko Ukkonen. Maximal and minimal representations of gapped and nongapped motifs of a string. Theoretical Computer Science, 410(43):4341– 4349, October 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salson:2009:FSA

[2293] M. Salson, T. Lecroq, M. Léonard, and L. Mouchard. A four-stage algorithm for updating a Burrows–Wheeler transform. Theoretical Computer Science, 410(43):4350–4359, October 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Apostolico:2009:SCS

[2294] Alberto Apostolico and Fabio Cunial. The subsequence composition of a string. Theoretical Computer Science, 410(43):4360–4371, October 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Castiglione:2009:CSW

[2295] G. Castiglione, A. Restivo, and M. Sciortino. Circular Sturmian words and Hopcroft’s algorithm. Theoretical Computer Science, 410(43):4372– 4381, October 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amir:2009:ECR

[2296] Amihood Amir, Yonatan Aumann, Piotr Indyk, Avivit Levy, and Ely Porat. Efficient computations of ℓ_{1} and ℓ_{∞} rearrangement distances. Theoretical Computer Science, 410(43):4382– 4390, October 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Federico:2009:STC

[2297] Maria Federico and Nadia Pisanti. Suffix tree characterization of maximal motifs in biological sequences. Theoretical Computer Science, 410(43):4391– 4401, October 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lee:2009:DRS

[2298] Sunho Lee and Kunsoo Park. Dynamic rank/select structures with applications to run-length encoded texts. Theoretical Computer Science, 410(43): 4402–4413, October 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gonzalez:2009:RSD

[2299] Rodrigo González and Gonzalo Navarro.

Rank/select on dynamic compressed sequences and applications. Theoretical Computer Science, 410(43):4414– 4422, October 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beal:2009:CCS

[2300] Marie-Pierre Béal and Dominique Perrin. Completing codes in a sofic shift. Theoretical Computer Science, 410(43): 4423–4431, October 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeza-Yates:2009:SBM

[2301] Ricardo Baeza-Yates, Véronique Bruyère, Olivier Delgrange, and Rodrigo Scheihing. On the size of Boyer–Moore automata. Theoretical Computer Science, 410(43):4432–4443, October 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gal:2009:PSI

[2302] Anna Gál. Preface: Special issue of ICALP 2006 — dedicated to the memory of Ingo Wegener. Theoretical Computer Science, 410(44):4445, October 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dietzfelbinger:2009:MPD

[2303] Martin Dietzfelbinger. In memoriam: Prof. Dr. math. Ingo Wegener, 1950– 2008. Theoretical Computer Science, 410(44):4446–4447, October 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2009:CDF

[2304] Xi Chen and Xiaotie Deng. On the complexity of 2D discrete fixed point problem. Theoretical Computer Science, 410(44):4448–4456, October 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Finocchi:2009:ORS

[2305] Irene Finocchi, Fabrizio Grandoni, and Giuseppe F. Italiano. Optimal resilient sorting and searching in the presence of memory faults. Theoretical Computer Science, 410(44):4457– 4470, October 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marx:2009:PVM

[2306] Dániel Marx. A parameterized view on matroid optimization problems. Theoretical Computer Science, 410(44): 4471–4479, October 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sankowski:2009:MWB

[2307] Piotr Sankowski. Maximum weight bipartite matching in matrix multiplication time. Theoretical Computer Science, 410(44):4480–4488, October 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chaudhuri:2009:PRA

[2308] Kamalika Chaudhuri, Satish Rao, Samantha Riesenfeld, and Kunal Talwar. A push–relabel approximation algorithm for approximating the minimum-degree MST problem and its generalization to matroids. Theoretical Computer Science, 410(44):4489– 4503, October 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harren:2009:AAO

[2309] Rolf Harren. Approximation algorithms for orthogonal packing problems for hypercubes. Theoretical Computer Science, 410(44):4504–4532, October 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fleischer:2009:F

[2310] Rudolf Fleischer and Jinhui Xu. Foreword. Theoretical Computer Science, 410(45):4533, October 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Angel:2009:MHS

[2311] Eric Angel, Evripidis Bampis, and Laurent Gourvès. On the minimum hitting set of bundles problem. Theoretical Computer Science, 410(45):4534– 4542, October 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2009:AME

[2312] Zhi-Zhong Chen and Ruka Tanahashi. Approximating maximum edge 2-coloring in simple graphs via local improvement. Theoretical Computer Science, 410(45):4543–4553, October 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Betzler:2009:FPA

[2313] Nadja Betzler, Michael R. Fellows, Jiong Guo, Rolf Niedermeier, and Frances A. Rosamond. Fixedparameter algorithms for Kemeny rankings. Theoretical Computer Science, 410(45):4554–4570, October 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gutin:2009:MLB

[2314] Gregory Gutin, Igor Razgon, and Eun Jung Kim. Minimum leaf outbranching and related problems. Theoretical Computer Science, 410(45): 4571–4579, October 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bansal:2009:SSS

[2315] Nikhil Bansal, Ho-Leung Chan, and Kirk Pruhs. Speed scaling with a solar cell. Theoretical Computer Science, 410(45):4580–4587, October 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Miyoshi:2009:SFI

[2316] Naoto Miyoshi, Takeya Shigezumi, Ryuhei Uehara, and Osamu Watanabe. Scale free interval graphs. Theoretical Computer Science, 410(45):4588– 4600, October 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Falaschi:2009:F

[2317] Moreno Falaschi, Maurizio Gabbrielli, and Catuscia Palamidessi. Foreword. Theoretical Computer Science, 410(46):

4601–4602, November 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barbuti:2009:GLP

[2318] Roberto Barbuti. Giorgio Levi in Pisa.

Theoretical Computer Science, 410(46): 4603–4604, November 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guglielmi:2009:PPG

[2319] Alessio Guglielmi. Personal portrait of Giorgio Levi. Theoretical Computer Science, 410(46):4605–4607, November 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alpuente:2009:TNR

[2320] María Alpuente, Santiago Escobar, and José Iborra. Termination of narrowing revisited. Theoretical Computer Science, 410(46):4608–4625, November 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amato:2009:ASD

[2321] Gianluca Amato, James Lipton, and Robert McGrail. On the algebraic structure of declarative programming languages. Theoretical Computer Science, 410(46):4626–4671, November 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bagnara:2009:APC

[2322] Roberto Bagnara, Patricia M. Hill, and Enea Zaffanella. Applications of polyhedral computations to the analysis and verification of hardware and software systems. Theoretical Computer Science, 410(46):4672–4691, November 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bossi:2009:SLP

[2323] Annalisa Bossi. S-semantics for logic programming: a retrospective look. Theoretical Computer Science, 410(46): 4692–4703, November 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gras:2009:NSI

[2324] Daniel Cabeza Gras and Manuel V. Hermenegildo. Non-strict independencebased program parallelization using sharing and freeness information. Theoretical Computer Science, 410(46): 4704–4723, November 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cousot:2009:AIR

[2325] Patrick Cousot, Radhia Cousot, and Roberto Giacobazzi. Abstract interpretation of resolution-based semantics. Theoretical Computer Science, 410(46):4724–4746, November 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liang:2009:FPL

[2326] Chuck Liang and Dale Miller. Focusing and polarization in linear, intuitionistic, and classical logics. Theoretical Computer Science, 410(46):4747– 4768, November 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maher:2009:LCE

[2327] Michael J. Maher. Local consistency for extended CSPs. Theoretical Computer Science, 410(46):4769–4783, November 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ueda:2009:LHL

[2328] Kazunori Ueda. LMNtal as a hierarchical logic programming language. Theoretical Computer Science, 410(46): 4784–4800, November 1, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Civril:2009:SMV

[2329] Ali Çivril and Malik Magdon-Ismail.

On selecting a maximum volume submatrix of a matrix and related problems. Theoretical Computer Science, 410(47–49):4801–4811, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bayer:2009:PTP

[2330] Daniel Bayer, Van Bang Le, and H. N.

de Ridder. Probe threshold and probe trivially perfect graphs. Theoretical Computer Science, 410(47–49):4812– 4822, November 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dennunzio:2009:DDA

[2331] A. Dennunzio, P. Di Lena, E. Formenti, and L. Margara. On the directional dynamics of additive cellular automata. Theoretical Computer Science, 410(47–49):4823–4833, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hof:2009:PGC

[2332] Pim van ’t Hof, Daniël Paulusma, and Gerhard J. Woeginger. Partitioning graphs into connected parts. Theoretical Computer Science, 410(47–49):

4834–4843, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Regnault:2009:PAS

[2333] Damien Regnault, Nicolas Schabanel, and Éric Thierry. Progresses in the analysis of stochastic 2D cellular automata: a study of asynchronous 2D minority. Theoretical Computer Science, 410(47–49):4844–4855, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2009:SFJ

[2334] Shisheng Li and Jinjiang Yuan. Scheduling with families of jobs and delivery coordination under job availability. Theoretical Computer Science, 410 (47–49):4856–4863, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kis:2009:SMU

[2335] Tamás Kis. Scheduling multiprocessor UET tasks of two sizes. Theoretical Computer Science, 410(47–49): 4864–4873, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Domosi:2009:SYT

[2336] Pál Dömösi, Géza Horváth, and Laurent Vuillon. On the Shyr–Yu theorem. Theoretical Computer Science, 410(47–49):4874–4877, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Simonsen:2009:CCL

[2337] Jakob Grue Simonsen. On the computational complexity of the languages of general symbolic dynamical systems and beta-shifts. Theoretical Computer Science, 410(47–49):4878–4891, November 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huang:2009:CWF

[2338] Yunbao Huang. The complexity of C^{bω}-words of the form w̃ × w. Theoretical Computer Science, 410(47–49): 4892–4904, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhao:2009:IGG

[2339] Yingchao Zhao, Wei Chen, and Shang hua Teng. The isolation game: a game of distances. Theoretical Computer Science, 410(47–49):4905–4919, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alon:2009:HEM

[2340] Noga Alon and Uri Stav. Hardness of edge-modification problems. Theoretical Computer Science, 410(47–49): 4920–4927, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arvind:2009:PLJ

[2341] V. Arvind, Johannes Köbler, and Wolfgang Lindner. Parameterized learnability of juntas. Theoretical Computer Science, 410(47–49):4928–4936, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeFelice:2009:CRC

[2342] Clelia De Felice, Gabriele Fici, and Rosalba Zizza. A characterization of regular circular languages generated by marked splicing systems. Theoretical Computer Science, 410(47–49):4937– 4960, November 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garcia:2009:LRL

[2343] Pedro García, Manuel Vázquez de Parga, Antonio Cano, and Damián López. On locally reversible languages. Theoretical Computer Science, 410(47–49): 4961–4974, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cordasco:2009:NSW

[2344] Gennaro Cordasco and Luisa Gargano. Navigable Small-World networks with few random bits. Theoretical Computer Science, 410(47–49):4975–4988, November 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gassner:2009:HHI

[2345] Elisabeth Gassner, Johannes Hatzl, Sven O. Krumke, Heike Sperber, and Gerhard J. Woeginger. How hard is it to find extreme Nash equilibria in network congestion games? Theoretical Computer Science, 410(47–49): 4989–4999, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kowalik:2009:DAM

[2346] Łukasz Kowalik and Marcin Mucha. Deterministic 7/8-approximation for the metric maximum TSP. Theoretical Computer Science, 410(47–49):5000– 5009, November 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kao:2009:NWA

[2347] Jui-Yi Kao, Narad Rampersad, and Jeffrey Shallit. On NFAs where all states are final, initial, or both. Theoretical Computer Science, 410(47–49): 5010–5021, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cain:2009:AS

[2348] Alan J. Cain. Automaton semigroups. Theoretical Computer Science, 410(47–49):5022–5038, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2009:OSM

[2349] Ming Liu, Yinfeng Xu, Chengbin Chu, and Feifeng Zheng. Online scheduling to minimize modified total tardiness with an availability constraint. Theoretical Computer Science, 410(47–49): 5039–5046, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2009:SOM

[2350] Xingyu Chen, Leah Epstein, and Zhiyi Tan. Semi-online machine covering for two uniform machines. Theoretical Computer Science, 410(47–49):5047– 5062, November 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2009:HRA

[2351] Lei Chen, Changhong Lu, and Zhenbing Zeng. Hardness results and approximation algorithms for (weighted) paired-domination in graphs. Theoretical Computer Science, 410(47–49): 5063–5071, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2009:DPD

[2352] Lei Chen, Changhong Lu, and Zhenbing Zeng. Distance paired-domination problems on subclasses of chordal graphs. Theoretical Computer Science, 410(47–49):5072–5081, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Batu:2009:STA

[2353] Tuḡkan Batu, Petra Berenbrink, and Christian Sohler. A sublinear-time approximation scheme for bin packing. Theoretical Computer Science, 410 (47–49):5082–5092, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kiltz:2009:DCC

[2354] Eike Kiltz and David Galindo. Direct chosen-ciphertext secure identitybased key encapsulation without random oracles. Theoretical Computer Science, 410(47–49):5093–5111, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sgall:2009:PSO

[2355] Jiří Sgall, Hadas Shachnai, and Tami Tamir. Periodic scheduling with obligatory vacations. Theoretical Computer Science, 410(47–49):5112–5121, November 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khamis:2009:RAD

[2356] Soheir M. Khamis, Sameh S. Daoud, and Hanaa A. E. Essa. A randomized algorithm for determining dominating sets in graphs of maximum degree five. Theoretical Computer Science, 410(47–49):5122–5127, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Spoerhase:2009:CPP

[2357] J. Spoerhase and H.-C. Wirth. (r, p)- centroid problems on paths and trees. Theoretical Computer Science, 410 (47–49):5128–5137, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bang-Jensen:2009:DDU

[2358] Jørgen Bang-Jensen and Matthias Kriesell. Disjoint directed and undirected paths and cycles in digraphs. Theoretical Computer Science, 410 (47–49):5138–5144, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tirnauca:2009:NSC

[2359] Cristina Tîrnăucă and Satoshi Kobayashi. Necessary and sufficient conditions for learning with correction queries. Theoretical Computer Science, 410(47–49):5145–5157, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DAlessandro:2009:PCF

[2360] Flavio D’Alessandro, Benedetto Intrigila, and Stefano Varricchio. The Parikh counting functions of sparse contextfree languages are quasi-polynomials. Theoretical Computer Science, 410 (47–49):5158–5181, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2009:OSU

[2361] Wenjie Li, Jinjiang Yuan, Jianfa Cao, and Hailin Bu. Online scheduling of unit length jobs on a batching machine to maximize the number of early jobs with lookahead. Theoretical Computer Science, 410(47–49):5182–5187, November 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Che:2009:NQA

[2362] Ada Che, Vladimir Kats, and Eugene Levner. A note on a quadratic algorithm for the 2-cyclic robotic scheduling problem. Theoretical Computer Science, 410(47–49):5188–5190, November 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dubickas:2009:BWG

[2363] Artūras Dubickas. Binary words with a given Diophantine exponent. Theoretical Computer Science, 410(47–49): 5191–5195, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yu:2009:AEC

[2364] Dongxiao Yu, Jianfeng Hou, Guizhen Liu, Bin Liu, and Lan Xu. Acyclic edge coloring of planar graphs with large girth. Theoretical Computer Science, 410(47–49):5196–5200, November 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kucera:2009:F

[2365] Luděk Kučera. Foreword. Theoretical Computer Science, 410(50):5201, November 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blaser:2009:SAA

[2366] Markus Bläser and Andreas Meyer de Voltaire. Semisimple algebras of almost minimal rank over the reals. Theoretical Computer Science, 410 (50):5202–5214, November 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonsma:2009:FPB

[2367] Paul Bonsma and Luis Cereceda.

Finding Paths between graph colourings: PSPACE-completeness and superpolynomial distances. Theoretical Computer Science, 410(50):5215–5226, November 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crochemore:2009:RSA

[2368] Maxime Crochemore, Lucian Ilie, and Wojciech Rytter. Repetitions in strings: Algorithms and combinatorics. Theoretical Computer Science, 410(50):5227–5235, November 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duckworth:2009:LIS

[2369] William Duckworth and Michele Zito.

Large independent sets in random regular graphs. Theoretical Computer Science, 410(50):5236–5243, November 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koiran:2009:VTT

[2370] Pascal Koiran and Sylvain Perifel.

VPSPACE and a transfer theorem over the complex field. Theoretical Computer Science, 410(50):5244–5251, November 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Longpre:2009:CS

[2371] Luc Longpré and Pierre McKenzie. The complexity of Solitaire. Theoretical Computer Science, 410(50):5252–5260, November 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nikoletseas:2009:EPC

[2372] S. Nikoletseas, C. Raptopoulos, and P. G. Spirakis. Expander properties and the cover time of random intersection graphs. Theoretical Computer Science, 410(50):5261–5272, November 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salamon:2009:AMI

[2373] Gábor Salamon. Approximating the Maximum Internal Spanning Tree problem. Theoretical Computer Science, 410(50):5273–5284, November 17, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tani:2009:CFA

[2374] Seiichiro Tani. Claw finding algorithms using quantum walk. Theoretical Computer Science, 410(50):5285– 5297, November 17, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferragina:2009:F

[2375] Paolo Ferragina and Gad M. Landau. Foreword. Theoretical Computer Science, 410(51):5299, November 28, 2009.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bergeron:2009:NLT

[2376] Anne Bergeron, Julia Mixtacki, and Jens Stoye. A new linear time algorithm to compute the genomic distance via the double cut and join distance. Theoretical Computer Science, 410(51):5300–5316, November 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hundt:2009:CGA

[2377] Christian Hundt, Maciej Liśkiewicz, and Ragnar Nevries. A combinatorial geometrical approach to twodimensional robust pattern matching with scaling and rotation. Theoretical Computer Science, 410(51):5317– 5333, November 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amir:2009:ASM

[2378] Amihood Amir, Yonatan Aumann, Oren Kapah, Avivit Levy, and Ely Porat. Approximate string matching with address bit errors. Theoretical Computer Science, 410(51):5334–5346, November 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Keller:2009:LCP

[2379] Orgad Keller, Tsvi Kopelowitz, and Moshe Lewenstein. On the longest common parameterized subsequence. Theoretical Computer Science, 410 (51):5347–5353, November 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fischer:2009:FEB

[2380] Johannes Fischer, Veli Mäkinen, and Gonzalo Navarro. Faster entropybounded compressed suffix trees. Theoretical Computer Science, 410(51): 5354–5364, November 28, 2009. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kolpakov:2009:SGP

[2381] Roman Kolpakov and Gregory Kucherov. Searching for gapped palindromes. Theoretical Computer Science, 410 (51):5365–5373, November 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ma:2009:WGW

[2382] Bin Ma. Why greed works for shortest common superstring problem. Theoretical Computer Science, 410(51):5374– 5381, November 28, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2009:P

[2383] Boting Yang and Cao An Wang. Preface. Theoretical Computer Science, 410 (52):5383, December 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huffner:2009:ICC

[2384] Falk Hüffner, Christian Komusiewicz, Hannes Moser, and Rolf Niedermeier. Isolation concepts for clique enumeration: Comparison and computational experiments. Theoretical Computer Science, 410(52):5384–5397, December 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2009:PCV

[2385] Zhao Zhang, Xiaofeng Gao, and Weili Wu. PTAS for connected vertex cover in unit disk graphs. Theoretical Computer Science, 410(52):5398– 5402, December 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Danziger:2009:CAA

[2386] Peter Danziger, Eric Mendelsohn, Lucia Moura, and Brett Stevens. Covering arrays avoiding forbidden edges. Theoretical Computer Science, 410(52): 5403–5414, December 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cai:2009:AAC

[2387] Zhipeng Cai, Zhi-Zhong Chen, and Guohui Lin. A 3.4713-approximation algorithm for the capacitated multicast tree routing problem. Theoretical Computer Science, 410(52):5415– 5424, December 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Betzler:2009:PCC

[2388] Nadja Betzler and Johannes Uhlmann.

Parameterized complexity of candidate control in elections and related digraph problems. Theoretical Computer Science, 410(52):5425–5442, December 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brandstadt:2009:SPG

[2389] Andreas Brandstädt and Van Bang Le.

Simplicial powers of graphs. Theoretical Computer Science, 410(52):5443– 5454, December 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marzban:2009:CSP

[2390] Marjan Marzban, Qian-Ping Gu, and Xiaohua Jia. Computational study on planar dominating set problem. Theoretical Computer Science, 410(52): 5455–5466, December 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bocker:2009:GWP

[2391] S. Böcker, S. Briesemeister, Q. B. A. Bui, and A. Truss. Going weighted: Parameterized algorithms for cluster editing. Theoretical Computer Science, 410(52):5467–5480, December 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shen:2009:SAS

[2392] Zhizhang Shen, Ke Qiu, and Eddie Cheng. On the surface area of the (n, k)-star graph. Theoretical Computer Science, 410(52):5481– 5490, December 6, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Janssen:2009:PGV

[2393] Jeannette Janssen and Paweł Prałat. Protean graphs with a variety of ranking schemes. Theoretical Computer Science, 410(52):5491–5504, December 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wagner:2009:CIS

[2394] Peter Wagner and Andreas Brandstädt. The complete inclusion structure of leaf power classes. Theoretical Computer Science, 410(52):5505–5514, December 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bhattacharya:2009:OMM

[2395] Binay Bhattacharya, Mike Burmester, Yuzhuang Hu, Evangelos Kranakis, Qiaosheng Shi, and Andreas Wiese. Optimal movement of mobile sensors for barrier coverage of a planar region. Theoretical Computer Science, 410(52):5515–5528, December 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Soloveichik:2011:ESC

[2396] David Soloveichik and Erik Winfree.

Erratum to “The computational power of Benenson automata” [Theoret. Comput. Sci. 344 (2005) 279–297]. Theoretical Computer Science, 412(43): 6112–6113, October 7, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511003537. See

[323].

Ren:2011:NAC

[2397] Wei Ren and Qing Zhao. A note on ‘Algorithms for connected set cover problem and fault-tolerant connected set cover problem’. Theoretical Computer Science, 412(45):6451– 6454, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006189. See

[1973].

Blumensath:2013:ESS

[2398] Achim Blumensath. Erratum to “On the structure of graphs in the Caucal hierarchy” [Theoret. Comput. Sci. 400 (2008) 19–45]. Theoretical Computer Science, 475(??): 126–127, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513000108. See

[1622].

vanBreugel:2013:ASD

[2399] Franck van Breugel, Claudio Hermida, Michael Makkai, and James Worrell. Addendum to “Recursively defined metric spaces without contraction” [Theoret. Comput. Sci. 380 (1/2) (2007) 143–163]. Theoretical Computer Science, 492(??): 117–122, June 24, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300340X. See

[1186].